A Complete Bibliography of Publications in the Australian Journal of Statistics and the Australian & New Zealand Journal of Statistics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

03 June 2023 Version 1.14

Title word cross-reference

(0, 1) [285]. (M, S) [1075]. $185.00 [2275]. 2 [1708, 1911, 1277, 563]. 21 [563, 563]. 2^{m} [2129, 987]. 2^{r} [1848]. 2 × 2 [285, 904, 392]. 3 [2139, 1911]. 3 × 3 × K [1626]. 4^{3} [247]. 50(3) [1174]. A [742, 1220, 801]. α_{n} [1595]. Ar(1) [1491, 1530]. Ar(P) [1540]. β [460, 1459, 1990]. χ [340, 200]. χ^{2} [66, 119, 120, 154, 955]. D [1387, 987, 944]. δ [1393]. E [1237]. ϵ [1530]. F [2070, 1501, 811, 1218, 2288, 128, 295, 516]. G [2373]. γ = k_{3}k_{2}^{−3/2} [50]. K [762, 774, 2257, 2352, 115, 390, 421, 1710, 878, 367]. L [2168, 1156]. L_{1} [2252]. l_{P} [1201]. λ [1437]. M [1236, 1708, 1394, 1109, 2154, 390, 421, 2129]. μ [1665]. n [61, 641]. P [1686, 1869, 2102]. π [428]. Q [1665]. q ≥ 2 [1174]. R [1049, 732]. r = k [684]. R^{2} [2223]. S [1593]. s^{2}G [1002]. S_{q} [1174]. \sqrt{n} [1292].

T [944, 569, 985, 1361, 1037, 1004, 1047, 1811, 369, 1093, 232, 1694]. t^{2} [985]. U [549, 1111, 2085, 677, 1016]. V [987]. W [1424]. X [756]. Z [515, 1096].

-Confidence [460]. -Consistent [1292]. -Corrected [1393]. -Designs [1595]. -Distribution [128]. -Estimates [2154]. -Estimation [1236]. -Estimator [1708, 1049, 1394, 1109]. -Estimators [1156]. -Events [944]. -Invariant [1665]. -Isotropic [1201]. -Latinized [1361]. -moments [2168]. -Optimal [1387, 1220, 987, 1075]. -Processes [1665]. -Repetitions [1530]. -Sample [115, 762, 2257]. -Statistics [549, 1111, 2085, 677, 1016]. -Sufficient [1237]. -Test [569, 369, 811, 1218, 2288]. -Tests [985]. -Trimmed [1459, 1990]. -Values [1869]. -variables [1694]. -Wishart [2373].

1 [2275, 165]. 11 [1535]. 12th [490]. 1975 [424]. 1976 [447, 445]. 1977 [491, 490]. 1980’s [599]. 1990 [1132]. 1995 [1280].

2001 [1589]. 2004 [1739, 1953]. 2010 [1985]. 2011 [2009]. 2012 [2057]. 2013 [2105, 2106, 2109]. 2018 [2315, 2308]. 20th [1752]. 21st [1417]. 27th [445]. 28th [424]. 2nd [1955, 1956]. 2^{T M} [630].

30 [447]. 311 [2289]. 37th [203]. 39 [2289].

5th [491].

8 [2276].

9 [2267]. 978 [2276, 2275, 2267]. 978-1-4398-7885-9 [2267]. 978-1-4822-3623-1 [2275]. 978-1-4987-1956-8 [2276].

A. [1168, 1599]. Aalen [2097]. Abraham [2064]. Abridged [32]. Absence [2012, 1720]. Absolute [1840]. Absorbing [302, 178]. Accelerated [1818, 1048]. Accelerating [2371]. accelerometry [2301]. Acceptability [1414, 1447]. Acceptance [1470]. Access [648]. Accounting [1897]. Accounts* [426]. Accuracy [456, 955, 798, 1624, 1883, 2230]. Accurate [2049]. Acknowledgement [1180, 1322]. Acknowledgment [1347, 1407, 1463, 1497, 1551, 1615]. Action [2275]. Activity [522]. Acyclic [1825]. adaptation [2371]. Adaptive [1732, 1442, 2191, 1213, 1793, 2125, 1523, 2149, 1693, 2284, 1469, 1557, 2371, 2280, 2355]. Added [1352]. Addendum [1599, 477, 405, 88]. adding [1571]. Additional [802]. Additive [1387, 1731, 1965, 2097, 1465, 1976, 1841, 1784, 1799, 2269]. Additivity [932]. Address [599, 1417]. Adequacy [590, 2397]. Adequate [514]. Adjacent [1880, 870]. Adjusted [1621, 908, 1895, 1299, 2182]. Adjusting [1084]. Adjustment [1195, 1521, 2293]. Adjustments [1902]. Admissibility [513, 743]. Adrian [1981, 1715, 2346]. Adult [29]. Advanced [2119, 1046]. Adventure [647]. Adversarial [2365]. Advising [567]. Affected [55].

Affine [1619, 1963, 2384]. Affine-Equivariant [1963]. Affine-Invariant [1619, 2384]. AFL [2261]. After [61, 1848, 811, 1764, 1892, 1138, 2070, 2310]. Against [1015, 678, 2152, 812, 855, 1039, 130, 582, 1613, 616, 2263]. Age [214, 61, 1753, 1905, 1660, 1746, 1926, 2360]. Age-Cluster [1926]. Age-Dependent [214]. Ages [29]. Aggregate [1754, 1905]. Aggregated [1671]. Aggregating [1269]. Agreement [856]. Agresti [1955]. Agricultural [1431]. Agriculture [1992]. Ahmad [1504]. Ahmed [1472]. Aid [1936]. Aided [987, 1075]. Aided-Construction [987]. AIDS [946, 1132]. Air [1893, 140, 1968, 1967]. Air-Borne [140]. Aitkin [1969]. Akaike [1819]. al [2406]. Alan [2035, 1955]. Alcohol [902]. Alfred [363]. Algebraic [177]. Algorithm [1793, 1244, 1978, 630, 73, 1366, 1205, 1230, 1277, 1083, 1645, 1760, 1423, 2240, 2212, 2266, 2398, 2372, 2371]. Algorithmic [1038, 2132]. Algorithms [2267, 1993, 1918, 2378]. Alias [450]. aliased [2212]. Aligning [389]. All-Bias [1772]. Allele [1250]. Allergic [1797]. Allocation [174, 250, 4, 878, 1151, 469, 1577, 2336]. Allowing [1260]. Almost [1394, 2088]. Along [1225, 287]. Also [913]. Alternating [991]. Alternative [1458, 389, 381, 2310, 2237]. Alternatives [1015, 1412, 812, 1936, 130, 582, 1613, 616, 1689]. Amemiya [1183]. Among [1098, 798, 1706, 2229]. Analogue [68]. Analogues [149, 151, 2243]. analyse [2382]. Analyses [1383, 1661, 7, 2214]. Analysing [2202, 1499, 284, 733, 2351, 2377, 1981]. Analysis [2070, 1012, 406, 1203, 96, 1852, 1480, 62, 2080, 712, 980, 1910, 2010, 1254, 2133, 1194, 1938, 1466, 1261, 1028, 2007, 1672, 2191, 2276, 757, 1202, 1721, 1349, 1339, 1529, 1512, 451, 2018, 485, 207, 994, 1452, 71, 456, 1995, 1018, 1972, 2131, 1702, 773, 2052, 1965, 962, 1290, 772, 2083, 767, 1099, 1172, 1820, 1272, 49, 525, 299, 971, 2059, 2180, 896, 11, 23, 47, 1135, 2100, 1590, 2115, 925, 1594, 904, 1103, 1169, 1444, 1606, 426, 1284, 1889, 581, 359, 1248, 2061, 2162, 2185, 1288, 960, 448, 1976, 1513, 533, 871, 652, 2111, 1146, 1117, 1195]. Analysis [557, 745, 1789, 783, 1340, 2151, 2109, 858, 640, 1297, 1511, 1945, 854, 830, 1431, 1026, 1897, 1408, 1784, 1900, 1917, 459, 839, 2167, 1633, 2409, 2222, 2365, 2396, 2252, 1556, 1561, 2237, 2404, 2269, 2292, 2329, 2280, 2017, 2146, 2092, 2045, 1955, 2186, 2135, 2175, 2134, 1045, 2174]. Analytic [699, 1645]. Analytical [353]. Ancestral [1643]. Ancillaries [814]. Ancillarity [720, 1524, 805, 918]. Anderson [1044]. Angle [1263]. Angular [1584, 2153, 1197]. Anna [2357]. Annealing [1191]. Annette [1981]. Announcement [1391, 1534, 1496, 1519, 1335, 1498]. announcements [1550]. Annual [144, 169, 194, 226, 256, 87]. Anomalies [1752, 299, 634]. ANOVA [1560, 2116, 582, 2076, 2151]. ANOVA-type [2116]. Answer [1978]. Antedependence [1954]. Antón [1954]. Any [1375]. AO [1589]. AOPE [2282]. APES [2311]. Applicable [1655]. Application [2070, 2058, 454, 1091, 1938, 1672, 1778, 1591, 602, 2116, 1978, 844, 1937, 1817, 285, 1936, 586, 1754, 1329, 2083, 655, 706, 1820, 1465, 2071, 2084, 2047, 37, 2115, 1103, 1791, 619, 1942, 2036, 1031, 854, 1134, 2332, 1693, 2237, 1636]. Applications [1671, 1668, 2221, 1466, 2239, 1829, 1672, 2006, 2079, 2191, 1797, 2024, 1392, 1979, 2211, 1484, 1989, 2238, 1716, 769, 930, 2140, 1520, 2023, 1535, 2212, 2116, 2205, 1978, 2139, 2108, 1753, 1617, 2025, 671, 1999, 452, 1965, 2027, 2001, 2157, 1235, 2177, 2000, 2097, 1465, 2114, 2206, 2066, 1673, 2158, 2026, 2028, 2148, 2022, 1418, 2115, 1630, 2087, 918, 1700, 1908, 1429, 1606, 2019, 1406, 2117, 1405, 1464, 1441, 1701, 2237, 2125, 622, 1683, 1953, 2138, 2190, 2036, 852, 1727, 2199, 2149, 2098, 376, 2081, 2178, 1564, 1511, 1453, 1642, 1926, 1945, 1234, 1605, 1914, 1618, 1552]. Applications [1966, 1536, 1927, 1682, 1917, 1379, 1576, 1577, 1430, 2150, 2167, 1521, 2192, 2276, 2020, 2065, 2051, 2145, 2064, 2164, 2091, 2043]. Applied [1410, 647, 985, 1218, 157, 2047, 371, 1701, 430, 1311, 2401, 2146, 2018, 2017, 2031, 1813]. Applies [1522]. Appointment [1927]. Approach [440, 1114, 691, 1270, 1466, 1449, 638, 1038, 1153, 1962, 1452, 1723, 1928, 1929, 553, 1433, 1780, 2162, 533, 1457, 1942, 1017, 2190, 2036, 1873, 1239, 1324, 1951, 2194, 2397, 1556, 2246, 2232, 2304, 2343, 2258, 2278, 2417, 1969, 2067]. Approaches [96, 1081]. Approximate [1515, 1867, 587, 2335, 229, 2291, 2123, 2311]. Approximating [1855, 1163, 1888, 798, 1694]. Approximation [954, 515, 698, 1049, 787, 340, 1163, 1353, 955, 1399, 1021, 293, 1800, 516]. Approximations [1778, 1675, 769, 791, 1157, 252, 361, 1561, 1240, 729, 1703, 1096, 1800]. Ar1 [1453]. Arbitrary [244, 303]. Arch [1540]. Archer [832]. Arcs [880, 274]. Area [457, 688, 1936, 3, 724, 2155, 2138, 2004, 1960, 1919, 2412, 2271]. Areas [994]. ARFIMA [1340]. Argument [471, 671]. Arguments [1414, 592]. ARIMA [636]. Arising [719, 963, 904]. Arm [2126]. ARMA [1255, 1164, 329, 2372, 897, 534, 1964]. Array [563, 2129]. Arrays [996, 761, 2215]. Arrival [2022]. Arsenic [2149]. Article [1344, 1339, 1342, 1341, 1337, 1338, 1336, 1340, 1343]. Articles [1403]. Artificial [2132, 9]. Aspects [1828, 434, 550, 525, 1025, 129, 425]. Assays [1070]. Assess [2152]. Assessing [1797, 792, 1758, 1018, 553, 1314, 888, 1706, 2242, 1687]. Assessment [1259, 1271]. assessments [2361]. Assisted [1750, 1992]. Associated [1065, 302, 742, 1770, 2124, 1849]. Association [679, 2222, 964, 2160]. Associations [1893, 1967, 1968]. assuming [2296]. Assumption [85, 1076, 2216, 2387]. Assumptions [569, 2075]. Asymmetric [2141, 1888, 1914, 2212]. Asymmetry [1528]. Asymptotic [1092, 1080, 606, 701, 711, 751, 1412, 1730, 485, 71, 985, 1218, 636, 618, 692, 735, 1175, 912, 205, 216, 362, 825, 300, 359, 493, 1053, 852, 1212, 837, 122, 798, 700, 869, 906, 1861, 1786, 1800, 1156, 1593, 1447]. Asymptotically [1086, 504, 189, 438]. Asymptotics [1659, 1085, 1324, 2250, 2355]. Atlantic [1476]. Attain [2162]. Attainment [475, 1027]. Attraction [1294, 1424]. attributable [2379]. AUCs [2002]. auctions [2365]. AUD$94.99 [2276]. augmented [1568]. Australia [118, 144, 169, 194, 226, 256, 491, 445, 27, 2395, 434, 2273, 1272, 5, 101, 333, 490, 685, 31, 1699, 424, 327, 1511, 355, 975, 2167, 468]. Australian [2289, 185, 195, 211, 227, 267, 2211, 90, 174, 147, 973, 559, 1723, 2325, 435, 1033, 1030, 426, 1280, 502, 76, 230, 156, 1031, 969, 1025, 1, 32, 968]. Australians [1032]. Auto [1178]. Auto-Regressive [1178]. Autocorrelated [993, 1155, 1355]. Autocorrelation [1157, 637]. Autocorrelations [416, 722, 488]. Autocovariance [440, 1341]. Autocovariances [1081, 488]. Automatic [1026]. Autoregression [1374, 1537]. Autoregressions [722, 967]. Autoregressive [440, 513, 768, 1365, 1835, 891, 2139, 1411, 767, 735, 1099, 1487, 1128, 1127, 1292, 487, 806, 857, 2411, 2264, 2330]. Auxiliary [1388, 1307, 1891, 1486, 889, 1443, 470, 1472]. Average [416, 440, 480, 767, 690, 738, 1645, 1635]. Averaging [1704, 1785, 1937, 2278]. Avoiding [2107]. Award [1534, 1985, 2009, 2057, 2105, 1604, 1652, 2308, 2188]. Awarded [612, 686, 1062, 1168, 1222, 1233, 1310, 1391, 1496, 1715, 1090, 1078, 1603]. Awards [2105, 2106, 2188, 2189]. Axial [1387, 1337].

B [2289, 1967, 2385]. B-splines [2385]. B.I.B. [199]. back [2235]. Background [1029]. Backlogs [454]. Backward [2162]. Baddeley [1715, 2346]. Bahadur [1242, 1106]. Balance [841, 494, 282, 1002, 374, 2312]. Balanced [847, 1611, 595, 905, 1155, 656, 896, 430, 1299, 783, 870, 761, 644, 563, 1622, 1256, 261]. Balancing [855]. Ballot [1104]. Banach [1834]. Band [1894]. Bands [2087, 391, 1177]. Bandwidth [1200, 1607]. Barnett [1981]. Barrier [2157]. Bartlett [603]. Base [1369]. Based [454, 907, 1480, 68, 751, 2141, 1676, 403, 1315, 1414, 1447, 1171, 1723, 1253, 2139, 2108, 1871, 1961, 109, 788, 1063, 1242, 2074, 2071, 1039, 1249, 1870, 1308, 1211, 1845, 2061, 464, 1773, 2004, 592, 1040, 869, 800, 1109, 1914, 660, 516, 583, 616, 1445, 1760, 1294, 1241, 1799, 1724, 1633, 2214, 2363, 2364, 2324, 1514, 2205, 661, 2333, 1424, 2265, 2397, 1686, 2340, 2252, 1554, 2386, 1558, 2332, 2117, 2232, 2404, 2366, 1559, 1380, 1687]. Basin [1476]. Basing [11]. Basis [1626]. Basu [918]. Batches [2026]. Bathtub [1827]. Bathtub-Type [1827]. Bayes [651, 1556, 1006, 1128, 2309, 389, 2294, 1166, 1976, 472, 897, 1488, 407, 1756, 1745, 2101, 2110]. Bayesian [2092, 1968, 2031, 2132, 2299, 1969, 1967, 2058, 2374, 2344, 1839, 1864, 1560, 2260, 1829, 2079, 1038, 1339, 1962, 1785, 2130, 1893, 2395, 1723, 1566, 1928, 1929, 1753, 2163, 2396, 1987, 1014, 1702, 1988, 2319, 1648, 1965, 553, 141, 2001, 2083, 775, 2114, 2142, 2158, 2011, 1780, 2367, 2261, 1543, 1889, 260, 1608, 2303, 1373, 2379, 2123, 2036, 2304, 1340, 1727, 481, 2030, 1782, 1784, 2278, 2213, 2401, 2194, 2247]. Bayesian/Likelihood [1969]. bbreg [2382]. Be [462, 563, 2291]. Becker [1045]. Been [1699]. Before [1618]. Behaviour [1860, 150, 122]. Behavioural [1307, 1986, 1432, 1555]. Behrens [2214, 315, 314, 1868]. Being [1872, 430, 1565]. Bell [699, 571]. Bellman [717]. Belz [410]. Benchmark [976]. Bent [2084]. Bent-Cable [2084]. Bernoulli [1095, 2003, 1659, 859]. Berry [748, 1009]. Bessel [2382]. Best [1097, 1397, 462, 1004, 1047, 391, 740]. Beta [1042, 1829, 930, 627, 383, 1806, 1142, 1743]. Beta-Binomial [930, 1743]. Beta-Optimal [1042]. Better [775, 1037]. Bettors [775]. Between [1975, 2038, 404, 2170, 1829, 1797, 607, 1893, 1032, 471, 1201, 2002, 14, 1263, 1590, 1237, 177, 935, 501, 1895, 438, 419, 229, 468, 1968, 2212, 1967]. Between-Household [1590]. Beyond [1827, 2253]. Bhattacharyya [475]. Bias [1708, 1675, 1610, 721, 2287, 268, 1353, 1579, 967, 1609, 534, 1808, 1445, 1772, 1607, 2312]. Biased [1316, 2221, 2323]. biases [1421]. Biasing [786, 1147]. Biassed [609]. Bib [408]. Bibliography [334, 546, 994, 101, 333, 685, 409, 541, 709]. Bicentennial [968]. bid [2365]. Bienaymé [548]. bifunctional [2378]. Bifurcating [1365, 1099, 1487]. Big [2178]. Bilateral [575]. Binary [2012, 1410, 1711, 2024, 486, 2152, 1806, 1999, 959, 899, 1949, 1791, 2166, 1040, 1256, 2274, 2272]. Bingham [1174]. Binomial [916, 1344, 480, 930, 2023, 285, 452, 1175, 1468, 1144, 587, 1927, 1849, 1755, 1743, 1423, 2287, 1418, 1558, 2321]. Bioinformatics [2030]. Biology [1992]. Biomass [1105]. Biometrics [972]. Biometry [973]. Biostatisticians [2173]. Biostatistics [1991]. Biplots [1997, 1980]. Bird [733]. Birth [1781, 1655, 336, 2260]. Birth-Death [336, 2260]. Birth-Order [1781]. Biserial [556]. Bisexual [1663, 1807]. Bivariate [275, 125, 236, 1365, 1478, 1979, 572, 358, 317, 507, 550, 1664, 285, 452, 1121, 1619, 2077, 1613, 306, 703, 1087, 1267, 1733, 2048, 385, 1434, 60, 13, 108, 1572, 2317, 1418, 2326]. Blaschke [477, 384]. Blest [1861]. Blest-Type [1861]. Block [199, 2152, 1070, 988, 92, 1415, 684, 1107, 1075, 1623, 1299, 783, 870, 893, 817, 1622, 1057, 1206, 261, 1784, 483, 1430, 2150, 961, 1008, 2255]. Blocking [1188, 2370]. Blocks [199, 2152, 1531, 641, 1256]. Blockwise [1834]. Blue [1975]. BNPdensity [2374]. Boca [2276, 2275, 2267]. Body [2084]. Bonferroni [945]. Bonferroni-Type [945]. Book [1971, 35, 45, 51, 38, 59, 64, 70, 77, 84, 94, 99, 104, 111, 116, 123, 136, 145, 153, 161, 170, 175, 183, 190, 201, 209, 217, 224, 235, 246, 254, 264, 272, 280, 290, 305, 319, 332, 352, 365, 388, 399, 378, 387, 398, 412, 423, 433, 444, 455, 467, 489, 466, 478, 509, 521, 497, 498, 508, 520, 544, 531, 543, 580, 598, 565, 579, 597, 632, 613, 631, 645, 659, 676, 687, 705, 704, 746, 771, 759, 770, 784, 808, 820, 794, 807, 819, 834]. Book [861, 833, 848, 860, 874, 887, 901, 873, 886, 915, 928, 941, 914, 927, 940, 999, 1011, 1023, 998, 1010, 1022, 1060, 1059, 1077, 1102, 1116, 1088, 1101, 1115, 1131, 1130, 1149, 1167, 1179, 1193, 1207, 1221, 1231, 1245, 1257, 1268, 1278, 1289, 1300, 1309, 1321, 1334, 1346, 1362, 1376, 1389, 1401, 1403, 1416, 1426, 1438, 1450, 1460, 1473, 1483, 1493, 1510, 1517, 1532, 1546, 1562, 1575, 1587, 1601, 1614, 1629, 1640, 1737, 1725, 2078, 2227, 2080, 1980, 1997, 2146, 2133, 2092, 2007, 2045, 2006, 2103, 2079, 2276, 2275, 2094]. Book [1046, 2018, 2267, 1814, 2298, 1995, 2017, 1972, 2131, 2163, 2065, 2052, 1981, 2029, 2031, 2066, 2112, 2093, 2132, 2033, 1954, 2051, 1955, 1956, 2186, 2005, 2020, 2145, 2313, 2053, 1993, 2135, 1696, 1697, 2032, 2019, 2183, 2121, 2185, 2120, 2064, 2299, 2111, 1982, 2175, 2164, 2044, 1841, 2054, 2119, 2209, 1813, 2134, 1045, 2081, 2234, 1970, 2030, 1044, 1994, 1991, 2174, 2219, 2091, 1943, 339, 2173, 2226, 2247, 2233, 1983, 2043, 2055, 2184, 1969, 2042, 2306, 2067, 1992, 1957, 1996, 2406]. Books [219, 900, 1061]. Bootstrap [1188, 1052, 957, 1187, 2071, 1399, 2019, 1166, 1887, 1779, 1177, 2194, 2316, 2285, 2255]. Bootstrap-Based [2071]. Bootstrapped [937]. Bootstrapping [2200]. Boris [1323]. Borne [1855, 140]. Both [1489, 275, 1999, 2202]. Bottom [1939]. Bottom-Up [1939]. Bottomless [577]. Bound [1009, 1057]. Boundary [1778, 2399]. Boundary-Mode [1778]. bounded [2382]. Bounds [480, 417, 1610, 1710, 844, 1356, 988, 340, 1892, 2197, 1337, 206, 1620, 1021, 1853, 475, 608, 1808, 945, 1331, 483, 1190]. Bowel [1700]. Box [2040]. brain [2134]. Branching [663, 691, 476, 1807, 655, 692, 706, 726, 664]. Break [1291]. Break-Point [1291]. Breakdown [1836, 240, 2263]. Breaker [745]. Breaking [1228, 2212]. breakpoint [2232]. Breast [2083, 2115, 522, 558, 1605]. Breeding [1989, 1430]. Brian [2007]. Bridge [2118]. Bridging [2111]. British [1591]. Brown [1948]. Bubble [948]. Buckley [2193]. Budget [1734]. Buehler [1691, 1620]. BUGS [2092, 2236]. Building [463, 468]. Bulk [80, 2253]. Bunching [1137]. Bureau [971, 1280]. Burr [1157].

C [1062, 2276, 2267, 1821, 1822, 1044]. C. [1062, 1822]. Cable [2084]. Calculated [7]. Calculating [1503, 2048]. Calculation [790, 1331, 776]. Calculations [1038, 1576]. Calibration [19, 1750, 1773, 1422, 779]. Campbell [2308, 1985, 2009, 2057, 2105, 2188]. Can [1037, 1406, 2291]. Canadian [2275]. Canberra [1132, 1280]. Cancer [2083, 2158, 2115, 1700, 522, 558, 1966]. Canonical [757, 1202, 1225, 262, 2100, 493, 605]. Cantoni [1991]. Capability [2114]. Capture [747, 1820, 1986, 1660, 1853, 2174, 1742, 1143, 1556, 1555]. Capture-Recapture [747, 1820, 1660, 1742, 1143, 1556, 1555]. Card [2018]. Cardiovascular [1893, 1968, 1967]. career [2392]. Carlo [1993, 2294, 752, 1712, 1918, 1041, 381]. Carrier [1855]. Carrier-Borne [1855]. Carry [1471]. Carry-overs [1471]. Carryover [2170]. CART [2305]. Cartesian [1197]. Carver [191]. Case [1065, 1781, 491, 485, 1794, 934, 291, 2158, 1210, 205, 216, 362, 1993, 1004, 1047, 790, 1579, 472, 1286, 2149, 610, 1343, 1215, 2395, 1688, 2270, 2379, 2361, 2348, 1687]. Case-Control [1579, 2379, 1687]. Case-Deletion [1343]. Cash [557]. Catastrophe [714]. Categorical [2146, 2016, 1955, 2186, 652, 1728, 1632, 1687]. Categories [1528]. Category [856, 1951]. CATI [1475]. Cauchy [540]. Causal [1982]. Causes [284]. Caution [79]. Caution-Level [79]. Cavitation [948]. Cell [1099, 1487, 756, 1136, 335]. cells [2349]. Censored [1091, 1515, 2062, 1748, 1478, 1730, 1586, 1272, 367, 1594, 2063, 581, 744, 171, 198, 583, 884, 1036, 2128, 2264, 2145]. Censoring [1199, 1019, 266, 912, 2245]. Censorship [906]. Census [2211, 447]. Censuses [795]. Central [1358, 292, 119, 2085, 128, 232, 295, 1690, 144, 169, 194, 226, 256]. Centred [1785, 2285]. Centres [174]. century [1417, 1552, 1752]. cerebral [2349]. Certain [281, 297, 663, 348, 417, 810, 1356, 383, 56, 1370, 130, 206, 177, 560, 824, 830, 126]. certificate [2361]. CES [252, 361]. Chain [1129, 1626, 1051, 302, 683, 1175, 920, 1918, 587]. Chains [107, 1358, 1702, 1005, 1035, 163, 178, 419]. Challenges [1132, 413]. Chan [2007]. Chance [1286]. Chang [1993]. Change [267, 768, 1400, 1244, 1155, 1329, 2088, 2178, 1915, 1863, 895, 2239, 1688]. Change-of-Support [2178]. Change-Over [768, 1244, 1155, 895]. Change-Point [1329, 2088, 1863]. Changes [793, 373, 2323]. Changing [1702]. Channel [1838]. Characterisation [1252]. Characterisations [1253]. Characteristic [417, 392, 1582, 206, 1249, 838]. Characteristics [823, 1381, 1572]. Characterization [454, 1763, 394, 348, 215, 375, 859, 1189, 1500, 815, 464, 865, 626, 1607, 530, 1445]. Characterizations [540, 1880, 828, 872, 1148, 1147, 1816, 750, 385, 1241]. Characterizing [703, 1317]. Charles [379]. Chart [800]. Charts [983, 148, 1453, 108]. Checking [1490]. Checks [630]. Chemical [494, 282, 374]. Cheng [1333]. Chernick [2019]. Chernoff [865]. Chernoff-Type [865]. Chi [1398, 1527, 953, 954, 43, 1123, 196, 340, 955, 1425]. Chi-Square [43, 1123, 196]. Chi-Squared [1398, 953, 954, 340, 955, 1527, 1425]. Child [1499]. Childs [1359]. Choice [1988, 788, 1369, 1248, 344, 2291]. Cholesky [925]. Choosing [1432, 1230]. Christoph [1996]. Circle [1514, 880, 738]. Circles [787, 1298]. Circular [1480, 1811, 1542, 1904]. Claim [1839]. Claim-Size [1839]. Class [1076, 1540, 1094, 1807, 938, 1806, 1175, 1435, 762, 1455, 628, 1073, 1504, 1311, 1443, 644, 629, 438, 2037, 1881, 2326]. Classes [780, 1585, 672, 2289]. Classical [1342, 149, 151, 2064, 745]. Classification [883, 1185, 513, 680, 743, 1170, 1153, 42, 1939, 525, 502, 693, 407, 1951, 1567, 2367]. Classifications [250]. Climate [2294, 2239]. Clinic [1349, 1520]. Clinical [552, 2078, 1972, 1806, 2052, 2126, 2032, 1134, 1638]. Clocking [453]. Cloning [1906]. Closed [1658, 1736, 2001, 1858]. Closed-Form [1858]. Closeness [627]. Closest [1538]. Cloud [205, 216, 362, 2178]. Cluster [1385, 1673, 2125, 2149, 1926, 2396, 1693, 1557]. Clustered [1410, 1529, 2153, 1308, 2063, 1146, 1747, 1068, 2393]. Clustering [951, 2205, 2397, 1419, 2394]. Clusters [1379, 2396, 2397, 2401]. Co [213]. Co-Spectra [213]. Cocaine [1210]. Cochran [2243]. Cochrane [1013]. Cocktails [2109]. Coefficient [921, 1236, 1042, 1484, 512, 188, 313, 1292, 867, 1434, 801, 50, 990, 2172, 2237]. Coefficients [718, 262, 324, 625, 1066, 48, 585, 774, 7, 25]. cohort [2379]. Coins [1661]. cointegration [2332]. cointegration-based [2332]. collected [2292]. Collector [349]. Collinearity [742]. Colonial [970]. Colony [335, 832]. Colours [2109]. Column [1611, 908, 1191, 1205, 430, 1256, 1206, 836, 1724]. columnarity [2349]. Combination [37, 2230]. Combinations [943, 944, 803, 945, 2099]. Combining [766, 2015, 1999, 1082, 1951, 2271]. Comment [1932, 1933, 1968]. commented [2406]. Comments [661, 662]. Commodity [31]. Common [2107, 181, 1121, 582, 1082, 1145]. Commonwealth [970]. Companion [2120]. company [1566]. Comparative [1025]. Compared [574]. Comparing [1392, 1318, 1683]. Comparison [898, 1442, 601, 1161, 414, 789, 1171, 358, 36, 843, 1585, 682, 1806, 1367, 1503, 1363, 172, 266, 2142, 2071, 2028, 18, 1499, 669, 1041, 1703, 381, 1786, 1256, 1805, 1917, 24, 2072, 1266, 2344, 2417]. Comparisons [82, 713, 1124, 1592, 973, 718, 1154, 1110, 2000, 1242, 260, 1302, 1379, 1838]. Compartmental [877]. Compartments [778]. Competing [1282, 445, 1184, 2131, 1135, 2053]. Competition [499, 228, 1195, 1953]. Competitions [1234]. Complete [92, 1299, 583]. Completely [1398, 729]. Completes [1519]. Completions [355]. Complex [2006, 934, 149, 151, 2011, 1798, 303, 2223]. Compliance [1897]. Component [831, 485, 499, 2162, 2409, 2282, 2404]. Components [267, 1467, 196, 1285, 1172, 1235, 1319, 1581, 472, 1741, 407, 167, 978, 1295]. Composite [189]. Compositional [2109, 2329]. Compositions [2038]. Compound [20, 471, 2039, 1446]. compression [2409]. Computation [1501, 1435, 1158, 2194, 2123]. Computational [2134]. Compute [1091]. Computer [987, 1075, 1992]. Computer-Aided [1075]. Computer-Assisted [1992]. Computing [2107, 2236]. Concatenation [1910]. Concave [1201, 2266]. Concavity [1320]. Concentration [1319, 1608, 1174]. Concentrations [979]. Concepts [2065, 1993]. Concerning [416, 417, 1822]. Concise [734]. Concomitant [1478, 997]. Concordance [697]. Concurrences [2152]. Condition [689, 1305, 563]. Conditional [1709, 1913, 956, 1349, 1847, 1108, 2351, 1491, 1000, 1435, 625, 1646, 1189, 1950, 1003, 1056, 990, 2381]. Conditioning [1319, 1141, 1486]. Conditions [1852, 301, 1635]. Cone [1366]. Conference [753]. Confidence [2070, 829, 695, 1765, 2141, 460, 1213, 2023, 2169, 322, 1354, 57, 2050, 1848, 2002, 1187, 1328, 1369, 1456, 1525, 1764, 1868, 1892, 2073, 2127, 2285, 206, 2246, 2087, 623, 54, 2135, 892, 1293, 1773, 639, 391, 518, 1914, 1849, 1177, 776, 1239, 2099, 1799, 1894, 1686, 1691, 2310, 2284, 2416, 2103]. Confidentiality [320]. Confirmatory [2191]. Confounded [294, 511]. Congenital [1211]. Congratulatory [2]. Congruence [1706]. Conjectures [1666, 2414]. Connected [1626]. Connections [177, 1552]. Consistency [651, 2255, 2154]. Consistent [1292]. Constant [726, 2201]. Constrained [1367]. Constraints [1910, 564, 44, 1734, 2197, 2302, 1564]. constructed [2310]. Constructing [2107, 2130, 1949, 1205, 2284, 1773, 1008]. Construction [847, 668, 294, 1611, 1043, 1244, 1191, 656, 1455, 987, 1075, 511, 870, 893, 1809, 2102, 1724]. constructions [2281]. Constuction [261]. Consulting [2199, 1618]. Consumption [1364, 1464, 327]. Context [1140]. Contiguity [653]. Contiguous [836]. Contingency [1626, 1398, 1527, 931, 1529, 42, 196, 1338, 756, 1433, 904, 1613, 493, 605, 1396, 2160, 1528, 2381, 1425, 2228]. Continuous [665, 1677, 1763, 1259, 747, 1160, 714, 1064, 692, 1555, 1063, 560, 664, 391, 1017, 1952, 1125, 1230, 1277, 150, 163, 1742, 1894, 2377, 472]. Continuous-Time [1677, 1742, 1555]. Continuously [1732, 1999, 49]. Contraception [395]. Contractions [1678, 1724]. Contracts [463]. Contrast [1942]. Contrasts [811, 623, 2150]. Contributed [1880, 1877, 1886, 1885, 1884, 1889, 1887, 1888, 1878, 1879, 1881].

Contribution [832]. Contributions [2181, 121]. Control [898, 1538, 983, 1617, 1318, 1579, 461, 2041, 148, 1152, 1271, 1301, 1302, 2379, 1687, 2121]. Controlled [164, 186, 795, 1656]. Controls [18]. Controversial [2032]. Controversy [1752]. Convergence [829, 717, 1358, 1049, 722, 966, 1824, 1844, 1216, 1048, 1744, 1657, 401, 673, 487, 441, 949, 1635]. Convergent [926]. Conversations [1150]. Convex [562, 1987, 73, 741, 212]. Copt [1991]. Copula [1961, 2179, 2231]. copulae [2258]. Copulas [1847, 1831, 2216, 2183]. copy [2206]. Core [2084]. Cornish [363, 1158]. Correct [1214, 1978]. Corrected [1393, 1342, 1695]. Correcting [1063, 1119]. Correction [862, 929, 1708, 2287, 2332]. Corrections [138]. Correlated [1202, 112, 1539, 36, 1659, 1107, 870, 893, 611, 1809, 1775, 1554]. Correlation [1709, 1236, 275, 1458, 125, 236, 1076, 986, 718, 262, 550, 415, 556, 2179, 328, 1163, 625, 10, 2100, 1950, 356, 554, 585, 1895, 1761, 197, 182, 824, 798, 1434, 610, 7, 801, 1637, 2237, 2352, 2225, 1821]. correlation.’ [1822]. Correlations [828, 792, 493, 605]. Correspondence [1910, 2010, 2185]. Corrigenda [95, 106, 132, 143, 162, 192, 204, 253, 334, 566, 1219, 442, 330, 361, 362, 1428]. Corrigendum [1208, 495, 532, 646, 1089, 1924, 421, 1990, 706, 1047, 1923, 2289]. cortex [2349]. Cosine [1647]. Cosine-Series [1647]. cost [2409]. cost-compression [2409]. Costa [1821, 1822]. Costs [2157, 1372]. Council [144, 169, 194, 226, 256]. Count [2146, 1539, 2179, 1747, 2203, 733]. counter [2208]. Counting [1129, 2218]. Counts [1797, 1979, 1352, 1754, 1235, 1825, 1727, 1729, 2343]. Coupon [349]. Coupon-Collector [349]. Course [1392, 2199]. Court [1927]. covariables [2330]. Covariance [2070, 1113, 1852, 1940, 107, 601, 602, 1108, 1064, 1002, 982, 206, 581, 448, 871, 1041, 1146, 783, 167, 7, 806, 857, 1296]. Covariances [1054, 30, 488, 2229]. Covariate [680, 743, 1748, 2124, 1895, 1960, 2412]. Covariate-Adjusted [1895]. Covariates [1160, 1133, 1818, 2202, 1870, 1742]. Coverage [2070, 1848, 1764, 1892, 1281, 2123]. Coverages [576]. Cox [2040]. Crab [2149]. Craig [2017]. Cramér [492]. crash [2395]. CRC [2276, 2275, 2267]. create [2211]. Credibility [734]. credible [2379]. Cricket [1520, 2376, 2140]. Criminal [2067]. Criteria [1941, 1676, 1218, 938, 1507, 1819, 127, 2101, 1567, 1684, 2353, 2286]. Criterion [312, 1704, 1867, 484, 934, 603, 1964, 1786, 2342]. Critical [791, 1157, 341, 538, 1251]. Criticality [1659]. Crop [1511]. Cross [1471, 1977, 57, 1754, 9, 346, 1215, 2317, 2211, 2415, 2379, 2271]. cross-dependence [2317]. Cross-Over [9, 1471]. Cross-Product [57]. Cross-Sectional [1754, 346, 2415, 2379, 2271]. cross-sections [2211]. Cross-Validation [1977, 1215]. Crossed [407]. Crossing [2144, 1952, 640]. Crossings [1670]. Crossover [1809]. CSIR [972]. CSIRO [903]. Cube [1336]. Cubic [1387, 1722, 2087, 949]. Cumulant [50]. Cumulants [731]. Cumulative [1913, 699, 391, 148, 1894, 2381]. Cure [2115, 2063]. Curve [1203, 1091, 735, 742, 724, 779, 1134, 1805]. Curved [2139]. Curves [1827, 2144, 1161, 1506, 1468, 33, 1382, 1558]. Customer [114]. Customers [165]. Cusum [1211, 1453, 1926]. Cut [1930, 47]. Cut-Off [1930]. Cutting [1531]. Cyclic [780, 684]. Cyclical [1677].

D [2007, 2276, 947, 2139]. Dairying [3]. Dale [1954]. Daley [1934, 1654, 1404]. Dam [577, 755]. Damage [1253]. Daryl [1654, 1404]. Data [1012, 1901, 2012, 1203, 916, 1480, 712, 1019, 1091, 2146, 1515, 1194, 1711, 1160, 1829, 1672, 2006, 2191, 2276, 2062, 1202, 1349, 1364, 1478, 1730, 1586, 1484, 1578, 737, 457, 1188, 90, 789, 1539, 1891, 976, 1828, 843, 1383, 1995, 2017, 1928, 1817, 1228, 1999, 864, 2153, 2179, 788, 1906, 1754, 1905, 1224, 772, 1172, 1363, 2114, 1272, 1316, 2066, 2074, 1328, 1456, 1525, 1524, 2011, 1210, 2084, 982, 1151, 2016, 2028, 1067, 1135, 2148, 321, 1590, 2172, 2202, 1830, 2186, 2096, 2063, 2145, 1720, 1429, 1169, 1136, 1950, 581, 1464, 1320, 1216, 2124, 1596, 1895, 652, 1146, 346, 2190]. Data [2175, 1311, 1747, 2098, 2109, 2178, 1040, 641, 1025, 1511, 1152, 1453, 2091, 2203, 60, 2161, 1897, 1943, 733, 1900, 884, 1036, 839, 1904, 2280, 2382, 2260, 1567, 1684, 1632, 2356, 2364, 2351, 2205, 2396, 2359, 2393, 2206, 2310, 1418, 1554, 2274, 1630, 2377, 2415, 2327, 2304, 2271, 2292, 2301, 2329, 2362, 1568, 1687, 1981, 1954, 1955, 1045, 2029, 2112, 2174]. Data-adaptive [2280]. Data-Based [788]. Data-Driven [1904]. Databases [1429]. Dataset [2211]. Datasets [1906, 2401]. Date [49]. Dates [174]. Davidson [965]. Davies [416]. day [1520]. Dealing [1596, 1720]. Death [1655, 336, 2260]. Debate [1027]. debiased [2380]. December [1378, 1589]. deciding [2397]. Decision [1318, 15, 2367]. Decisions [11]. Decomposition [158, 2114, 2074, 925, 2124, 2381]. Decompositions [1003]. Deducible [85]. Deep [755]. Deficiency [606, 701]. Defined [1173, 273]. Definite [215, 262, 1296]. Definition [270, 405]. Definitional [1025]. degradation [1630]. Degree [945]. degrees [1694]. Delay [702, 924]. Delete [1911]. Delete- [1911]. Deletion [1343]. Deliberately [786]. Demand [31, 2239, 2273]. Demands [87]. Demographic [395, 753, 354]. Demonstrably [1097]. Demonstration [1909]. Densities [1201, 628, 589]. Density [400, 431, 1658, 1315, 1891, 651, 721, 788, 787, 1072, 1200, 1441, 1759, 179, 1119, 1631, 2266, 2372, 2303, 2417]. Dependence [1852, 1677, 2038, 1867, 1753, 317, 2085, 1039, 2138, 87, 1669, 2317, 2410, 2279, 2415, 2183]. dependency [2360, 2208]. Dependent [1448, 680, 743, 214, 711, 1133, 1188, 564, 1886, 1807, 42, 802, 1063, 617, 1136, 1284, 1405, 2201, 2348, 2154, 1215, 1635, 2318, 2255, 2327]. Depends [222]. Depth [2350, 2363, 2206]. depth-based [2363]. Derivation [199, 119, 1158, 868, 1056]. Derivative [259, 360, 1790]. Derivatives [1759, 1597]. Derived [471]. Descent [1576]. description [1554]. Descriptive [635]. Design [552, 1711, 1867, 1989, 1276, 1171, 844, 785, 2108, 1837, 1126, 436, 247, 971, 1612, 263, 1308, 1606, 1845, 1579, 1623, 729, 1125, 1277, 563, 1057, 975, 1688, 2370, 2007, 1972, 2052, 2111]. Design-Based [1171, 2108]. Designed [772, 896, 1594]. Designing [186, 1622]. Designs [847, 406, 668, 780, 294, 312, 510, 1019, 1749, 763, 2170, 768, 1375, 1471, 1442, 841, 494, 1509, 1387, 1611, 199, 243, 282, 595, 1094, 1043, 1400, 908, 1244, 905, 1585, 1155, 2152, 1070, 1734, 827, 984, 988, 92, 172, 1191, 1361, 1415, 1595, 408, 656, 899, 2171, 1002, 1287, 684, 1107, 1790, 987, 1075, 1205, 1531, 374, 409, 430, 500, 541, 1299, 511, 783, 870, 893, 761, 644, 817, 473, 1622, 1071, 1256, 1206, 261, 1784, 459, 483, 836, 895, 1190, 1430, 1809, 2072, 2102, 2150, 961, 1772, 1724, 1008, 2014, 2338, 1638, 2297, 2212, 2281, 2229, 2257]. designs [2129, 2312, 2262, 1813]. Detect [2171]. Detectable [1608]. Detecting [1746, 1789, 779, 2323, 568, 2401]. Detection [799, 2025, 1702, 4, 2083, 1235, 1688, 2114, 2071, 1433, 1720, 2117, 535, 2399, 1605, 1569, 1026, 1863, 2206, 2294, 2241, 1553]. Determination [1507, 897, 1294]. determine [2360, 2301]. Determined [2013]. Determining [1053]. Deterministic [240]. Development [434, 622]. Deviance [1584]. Deviation [1351]. Deviations [1668, 636, 1759, 13]. Diagnostic [2045, 1490, 1285, 1581, 888, 1744, 1826, 2123, 1904]. Diagnostics [1355, 1306, 1343, 1036, 1386, 2264]. Dialling [1475]. Dice [1592]. Dichotomising [1063]. dictated [2291]. Difference [404, 112, 2002, 1288, 728, 86, 1694]. Differenced [866]. Differences [468]. Differencing [1291]. Different [799, 766, 2169, 370, 1900]. Differential [348, 1736, 962, 963]. Difficult [1400]. Difficulties [683]. Diffusion [1298]. Diggle [1994]. Digit [1475]. Digits [41]. Dimension [1512, 773, 997, 2250, 2393, 2386, 2208]. Dimensional [1626, 1940, 1126, 1939, 2011, 1879, 2091, 1431, 2161, 1408, 937, 2231, 2309, 2294, 2241, 2280]. Dimensions [409, 541, 2384, 2216]. Direct [902, 1081, 2004, 167]. Directed [1825, 1632]. Direction [742, 1728, 1566]. Directional [843, 1337]. Directions [1263, 2217]. Dirichlet [1929, 1712]. Disabilities [666]. Disability [1753]. Disclosure [2121]. discontinuities [1684]. Discordancy [822]. Discovery [1429]. Discrepancy [426]. Discrete [1012, 1129, 348, 287, 1328, 1456, 1525, 1436, 1869, 1173, 1320, 703, 393, 1240, 163, 380, 1945, 2192, 2303]. Discrete-Time [1945]. Discretely [1812]. Discriminant [1203, 1254, 1512, 1018, 300, 427, 359, 1633]. Discriminating [607]. Discrimination [574, 1153, 602, 2309]. Discussion [1489, 446, 1477, 1947, 1392, 1414, 1452, 1795, 1946, 1945, 1948, 1643, 1512, 2288, 1417]. Disease [1829, 553, 1590, 816, 1045, 1685]. Diseases [1893, 1931, 1968, 1967]. Disequilibrium [1103]. Disparities [1776]. Disparity [1836]. Dispersed [1785]. Dispersion [1084, 341, 1485, 758, 1306, 1083, 1902, 1301, 1302, 1265, 1569]. Dispersions [843]. Displaced [97]. Displays [888]. dissimilarity [2214]. dissimilarity-based [2214]. Distance [694, 1829, 751, 1153, 1397, 1384, 1706, 1173, 981, 2364, 1553]. distances [1633]. Distinct [1655]. Distinguishability [856]. Distributed [1846, 880, 1093, 355]. Distribution [281, 297, 1065, 454, 540, 1839, 1940, 53, 1182, 275, 66, 822, 1344, 134, 68, 342, 125, 1186, 809, 681, 716, 796, 828, 1112, 954, 357, 1478, 620, 414, 853, 715, 698, 699, 930, 1171, 1501, 976, 764, 61, 1356, 269, 330, 375, 507, 627, 383, 484, 938, 934, 285, 452, 360, 1304, 453, 1110, 251, 1768, 1465, 1272, 787, 238, 292, 340, 26, 188, 760, 119, 1831, 1039, 244, 551, 100, 241, 872, 1336, 371, 524, 1211, 727, 603, 306, 1779, 89, 373, 1293, 1679, 391, 1373, 274, 1120, 1252, 401, 852, 865, 1267, 765, 626, 131, 1488, 693]. Distribution [991, 1074, 481, 376, 867, 1542, 1096, 385, 530, 965, 798, 571, 97, 293, 449, 830, 128, 171, 232, 283, 295, 516, 616, 13, 838, 1656, 50, 1858, 517, 1174, 1894, 933, 1296, 1810, 1593, 2396, 2255, 2377, 2373, 2261, 2372, 1683, 2217, 2312]. Distribution-Free [681, 796, 1112, 1039, 524, 1293, 449]. Distribution-Preserving [1296]. Distributional [1723, 420, 1816, 1428, 1413, 2387]. Distributions [1114, 1763, 1668, 513, 394, 1880, 953, 236, 431, 348, 751, 1273, 1730, 1484, 1854, 187, 607, 572, 242, 302, 1253, 1929, 471, 317, 550, 1774, 1197, 238, 151, 222, 367, 1437, 1436, 1258, 1370, 859, 912, 672, 133, 723, 1263, 1189, 1824, 1173, 752, 260, 493, 1507, 703, 1811, 1240, 148, 964, 2004, 837, 965, 629, 60, 198, 307, 1445, 303, 732, 733, 584, 1317, 108, 1241, 1217, 2192, 1692, 2221, 2350, 2245, 1420, 1561, 2406, 2242, 2388, 2392, 2335, 2326, 2406]. District [1577]. Disturbances [1339, 791, 690]. Divergence [1676, 1337, 1920, 1657]. Divisible [1240]. Divorce [90]. DNA [2206]. Do [1176, 1037, 424, 885, 1454]. Dobson [1981]. Doctor [1783]. Dogit [1762]. Doksum [699, 571]. Domain [797, 1198, 533, 534, 1424]. Domains [384, 1294, 1788]. Domijan [1795]. Dominance [760]. Dominating [1120]. Domination [1786]. Dosage [479]. Dose [2087]. Double [1065, 1105, 1502, 1271, 1782, 2343]. Double-Marking [1271]. Doubly [1272, 1096, 128, 232, 295]. Doubt [855]. Douglas [1992, 596, 646]. Down [898]. Draw [1104]. Drawn [125, 67]. Drift [419]. Driven [1904, 2411]. Drives [1030]. Drug [2177]. Dual [2014]. Dual-Objective [2014]. Duality [44]. Duc [1967]. Due [499, 1644, 1579, 2323]. Duplicate [1247]. Durability [1606]. Duration [1033, 1031]. Durbin [689, 535]. Dust [140]. Dynamic [2058, 1290, 1612, 2261, 223, 1264, 393, 2203, 2344, 2238, 2272, 2029]. Dynamical [1719]. Dynamics [948, 65, 1717].

E. [1192, 1222, 1391, 1078]. Each [1852]. Ear [1079]. Early [1028, 1989, 969, 1552]. Early-Generation [1989]. earnings [1566]. Earthquake [2114, 425]. Easter [1521]. Eclecticism [491]. ECM [2372]. Ecological [1675, 2117, 1957]. Econometric [288, 2020, 1031]. Economic [4, 2028, 567, 347, 753]. Economic-Demographic [753]. Economics [2163]. ecotoxicological [2237]. edge [2253]. Edited [2276, 2275, 1972]. Editing [258, 354]. Edition [2299, 2133]. Editor [913, 1279, 1492, 1534, 1628, 396, 1333, 1968, 657, 1821, 1545, 110, 899, 1649, 1519, 1604, 1652, 1822, 885, 1345, 1404, 350, 1639, 1967, 1332]. Editorial [39, 78, 185, 195, 211, 227, 1534, 1550, 1335, 1404, 1498]. Edmund [363]. edn [1955, 1956]. Education [1027, 667, 248]. Educational [1027, 1029].

Edward [1209]. Edwin [1150]. Effect [353, 2103, 525, 2135, 284, 30, 637, 165, 1040, 1382, 1571, 1521]. Effective [1632, 1141, 2354]. Effects [267, 1349, 112, 1877, 1084, 328, 1110, 1235, 959, 2011, 205, 216, 362, 1818, 2126, 1441, 1701, 2124, 1395, 407, 798, 1966, 1838, 1560, 2239, 2212, 2282, 2288, 2362]. Efficacy [1482]. Efficiency [312, 595, 905, 1013, 988, 526, 656, 1106, 995, 1053, 868, 922, 167, 1861, 1057, 483]. Efficiency-Balanced [656]. Efficient [1435, 2172, 2341, 2385, 815, 1107, 1394, 774, 2207, 1760, 2225, 2102, 2212, 2378]. Effort [4, 1432, 1453]. Efron [1592, 1583]. EGD [406]. Ehrenfest [2003]. Eigenvalue [663]. eigenvalues [2253]. Eigenvector [689]. Eigenvectors [776]. Eighteen [647]. Eighth [50]. Elderly [1893, 1968, 1967]. elections [2325]. Electoral [559, 1576, 1577]. electricity [2239]. Electroencephalographic [1838]. Elements [181, 61, 1834, 1379]. Eliciting [1825]. Elimination [2162, 2291]. Elliptical [1041]. Elliptical-Theory [1041]. Embedded [1239]. Emeritus [1589]. Emerson [1862]. emphasis [2246]. Empirical [1091, 1315, 1213, 1006, 1128, 1582, 2039, 1870, 1303, 389, 1166, 472, 1942, 1773, 1959, 897, 2060, 401, 852, 1488, 561, 838, 441, 2333, 1556]. Empirically [1048]. employed [2239]. Employment [1032, 1945]. Encounters [1223]. Enders [2017]. Endpoints [1972]. Enhance [1406]. En’ko [947]. ensemble [2384]. Ensuring [343]. Enter [15]. Entropy [1675, 1446, 1974, 2410, 2293]. enumeration [1553]. Envelope [360]. envelopes [2357]. Environment [30, 519, 1656, 1204]. Environmental [1406, 2401, 1957]. Environments [726]. Epidemic [712, 947, 1931, 1648, 1098, 1100]. Epidemics [946, 947, 1855]. Epidemiology [816, 2044]. Epistemology [2042]. Equal [308, 774, 469]. Equality [222, 367, 130, 1506, 1041, 1146, 2217]. Equation [313, 728]. Equations [1480, 1466, 348, 1544, 1730, 2116, 1736, 962, 268, 622, 1325, 127, 1386, 2224, 1570, 1982, 2055]. Equi [1775]. Equi-Correlated [1775]. Equicorrelated [1067]. Equicorrelation [1042, 867, 990]. Equileverage [1043]. Equineighboured [827, 984]. Equispaced [949]. Equivalence [2012, 1449, 1364, 1988, 1956, 1949, 1406, 1464]. Equivalent [504]. Equivariant [1963]. Ergodic [1092]. Ergodicity [1086, 419]. Erich [1995]. Errata [88]. Erratum [2389]. Error [1254, 112, 1785, 635, 781, 1731, 1342, 811, 618, 1197, 1625, 1063, 1831, 1314, 1921, 427, 359, 1779, 2088, 182, 700, 869, 1093, 382, 1960, 2182, 1331, 2312, 2341, 2332, 1695, 2412, 2328]. Error-In-Variables [1342]. Errors [1448, 768, 841, 429, 505, 1540, 112, 2107, 993, 559, 22, 802, 1644, 735, 959, 1002, 300, 738, 2060, 2098, 7, 25, 1865, 2154, 1809, 1775, 2264]. Errors-in-Variables [2107]. Eruption [2022]. Esséen [748, 1009]. Essentially [726]. Essentials [2175]. Estimable [679]. Estimate [1214, 1778, 2015, 1999, 1905, 618, 91, 700, 838, 2239]. Estimated [675, 2169, 359, 25, 1686, 2287, 2286]. Estimates [504, 570, 1364, 112, 512, 1358, 651, 1154, 721, 1054, 141, 1072, 389, 1166, 2159, 826, 1212, 487, 1537, 2154, 2090, 1560, 1421]. Estimating [1480, 663, 747, 1896, 1924, 1783, 1466, 809, 615, 600, 1105, 841, 539, 1730, 1350, 1979, 879, 977, 902, 976, 2116, 1736, 1903, 1055, 1754, 1690, 2011, 1210, 2026, 1258, 1573, 1121, 1660, 1263, 1237, 1830, 1700, 1441, 622, 1325, 1597, 1856, 876, 1679, 1741, 274, 777, 1395, 336, 488, 824, 74, 869, 171, 198, 2055, 588, 1743, 1423, 2224, 1570, 2372]. Estimation [1199, 1572, 528, 911, 1735, 907, 1839, 1114, 1940, 2317, 1913, 1236, 298, 710, 372, 1182, 2040, 439, 1170, 1749, 107, 400, 431, 1658, 576, 831, 85, 536, 2144, 357, 2024, 494, 917, 797, 1315, 1478, 620, 1578, 715, 1591, 1769, 930, 1520, 1812, 323, 1891, 719, 476, 688, 1383, 1877, 1184, 1886, 2050, 624, 1084, 420, 1807, 1228, 285, 1000, 17, 802, 1936, 1304, 1424, 1930, 251, 1137, 1884, 324, 252, 361, 1197, 1164, 1235, 1705, 2177, 1081, 2097, 1341, 1033, 3, 1200, 1316, 313, 1002, 1058, 1138, 1128, 1831, 1437, 1582, 1067, 1384, 1921, 1127, 804, 100, 241]. Estimation [875, 1780, 2118, 724, 741, 1446, 1869, 1308, 1173, 2077, 1100, 725, 1974, 1759, 137, 30, 1198, 1757, 924, 2155, 179, 967, 329, 472, 1523, 1942, 654, 1122, 738, 2138, 346, 534, 1120, 2054, 868, 1007, 1119, 1074, 1607, 273, 2004, 774, 803, 1031, 1292, 1068, 1178, 2343, 469, 589, 1001, 1381, 1502, 1873, 381, 610, 1729, 1380, 60, 1580, 621, 382, 1960, 1919, 425, 1262, 1760, 1598, 1312, 1382, 1881, 997, 979, 1174, 1755, 818, 1728, 1624, 1909, 2101, 1775, 1810, 2344, 1634, 2316, 2318, 2238, 2302, 2273, 1631, 2231, 2380, 2255, 2274, 2341, 2266, 1558, 1570, 2385, 2232, 2327]. estimation [2304, 2412, 2271, 2207, 2225, 2417]. Estimator [701, 921, 253, 1273, 1708, 879, 288, 1049, 788, 578, 1625, 1183, 1009, 1844, 1368, 1394, 1396, 740, 344, 673, 1522, 587, 869, 1109, 2089, 239, 782, 231, 2182, 1445, 2193, 1482, 2252, 2304, 1695]. Estimators [2086, 1388, 1840, 606, 1065, 1254, 403, 1282, 1730, 2107, 1171, 1750, 1961, 310, 1367, 2027, 172, 526, 2142, 1337, 268, 2071, 1314, 1485, 475, 919, 995, 889, 1082, 939, 347, 777, 325, 765, 407, 1443, 1016, 1093, 906, 381, 1786, 438, 1238, 1156, 2245, 2282, 2285, 2229, 1472, 1557, 2352, 1422, 2355]. Ethically [2093]. Euro [1661]. Europe [1661]. Eva [1991]. Evaluating [2181, 1196]. Evaluation [1896, 1924, 1475, 2045, 2141, 512, 2025, 2083, 752, 851, 1511, 1914, 1917]. Evan [1223]. Evanescent [1165, 221]. even [2129]. evenness [2312]. Event [1972, 2066, 2145, 1945, 2263, 2080]. Events [1664, 697, 943, 944, 2036, 945, 1879]. Ever [1076]. Evidence [1213, 1414, 1447, 855, 1027, 897, 1878]. Ewens [1310]. Exact [1765, 460, 1610, 1108, 2169, 1885, 1366, 1525, 370, 2087, 1869, 1949, 2126, 603, 2291, 1142, 674, 571, 782, 830, 303]. Exactly [343, 1906]. Examination [1270, 689, 835, 1269, 314, 2216, 523, 2075]. Example [1092, 1430]. Examples [710, 1680, 2247]. Exceedances [1860, 978]. Exceeds [1903]. Excel [1995]. excess [2256]. Exchange [373]. Exchangeable [2003, 996]. Exclusion [2061]. Exclusive [1791]. exhaustive [2403, 2311]. Existence [460, 91, 758, 1665, 2281]. expanded [2240]. Expansion [671, 1158, 300, 359, 727]. Expansions [549, 985, 1218, 128]. Expectancies [1754]. Expectation [1847, 359, 2254, 2398]. expectation-maximisation [2398]. Expectations [404, 1710, 1435, 756, 501, 654, 1196, 621]. Expected [1020, 427, 506, 2304]. expectile [2380]. Expenditure [1829, 1364, 346]. Experience [1028, 850, 354]. Experiences [2199, 849]. Experiment [82, 630, 328, 971, 1594, 1606, 284, 1230, 1143]. Experimental [312, 844, 827, 1118, 1531, 1125, 644, 975, 2370]. Experimentation [1307]. Experiments [406, 510, 747, 1867, 1509, 1307, 1133, 1848, 1126, 772, 1820, 1986, 795, 896, 1432, 2061, 1405, 458, 409, 430, 541, 182, 1277, 1992, 2235, 2297, 1556, 2206]. Explanatory [1382, 2411]. Explicit [688, 1680, 1413, 1428]. Exploited [1719, 2291]. Exploring [2201]. Exponential [1185, 1065, 1901, 298, 394, 1776, 1097, 1124, 1836, 453, 655, 706, 462, 760, 367, 1138, 912, 1211, 995, 1516, 260, 461, 626, 837, 1083, 171, 588, 584, 1324, 517, 1420, 1573]. Exposure [1051, 1063]. Exposure-Response [1063]. Expression [1858]. Expressions [846, 926]. Extended [1194, 1672, 1487, 92, 346, 2076, 2151, 2037]. Extending [610]. Extension [762, 473]. Extensions [1721, 1670, 918]. Extent [1699]. External [805]. Extinction [1663, 664]. Extra [995]. Extrapolative [501]. Extrema [812, 1039]. Extremal [1357]. Extreme [810, 1762, 1664, 266, 1976, 964, 765, 1879, 2293, 2081]. extreme-value [2293]. Extremes [1871].

F [2007, 2006, 2275, 1822]. F.A.A. [545]. FAA [1589, 2035]. Factor [2191, 1962, 988, 1299, 1756, 1645, 1057, 2151]. Factorial [406, 294, 1509, 1848, 630, 2171, 1169, 458, 511, 182, 1277, 563, 2129]. Factorisation [514]. Factors [1400, 742, 1299, 1745, 483, 2262]. Failure [1012, 1470, 462, 1138, 1436, 1818, 1596]. Faithful [2022]. Falling [2013]. False [1700]. Familial [1383, 2029]. Families [1028, 655, 706, 1235, 1485, 995, 1122, 1007, 1239, 468, 2388]. Family [1776, 471, 1774, 1436, 1173, 1920, 1811, 1029, 1445, 1324, 354, 468]. Farm [1258, 2028]. Fashion [1055]. Fast [2378, 2049, 2311, 2266, 2403]. Fault [4]. Fay [2138]. Feature [1703, 2403, 2305]. February [1589]. federal [2325]. Fertility [1431]. Feser [1991]. Festschrift [1654, 2346, 2391]. Few [907, 303, 2356]. Fewster [2308]. Fibrous [2139]. Fiducial [20]. Fieguth [1983]. Field [1867, 902, 1126, 1976, 1431, 2072]. Fields [1802, 1637]. Fifth [2133]. Filter [1054, 586]. Filters [1535]. Finance [1027, 2081]. Financial [1846, 2025]. Finch [662]. Find [284]. Finding [1087, 1206]. Findings [615]. Fine [1844]. Fine-Tuned [1844]. Finetti [1201]. Finger [1235]. Finite [114, 936, 1186, 85, 539, 797, 1171, 578, 1140, 609, 1486, 1665, 1741, 594, 1381, 1873, 782, 1262, 2393, 2366, 2242]. Finite-Sample [782]. Finiteness [872]. Firm [15]. First [768, 1392, 947, 1411, 970, 1128, 1127, 1288, 1679, 738, 644, 2365, 2330]. First- [1679]. First-Difference [1288]. First-Order [1411, 1128, 1127, 2330]. first-price [2365]. First-Year [1392]. Fish [1660]. Fisher [2214, 515, 715, 698, 1181, 764, 315, 314, 1224, 1868, 1158, 1699, 2217, 1434, 801, 1174]. Fishery [1466]. Fit [1467, 822, 1344, 739, 1393, 764, 56, 109, 1906, 316, 1285, 1647, 1920, 800, 438, 2168, 583, 616, 1904, 1143, 1692, 2364, 1514, 2265, 2334, 2387, 2242, 1687]. fits [1688]. Fitted [1449, 2096]. Fitting [1019, 948, 950, 951, 1363, 1140, 1163, 1468, 1683, 1559, 1469, 952, 1805, 949]. Fixed [829, 1626, 1054, 1680, 155, 525, 205, 560, 384, 639, 906, 1624]. Fixed-Width [829, 639]. Flats [1509]. Flexible [2084, 2022, 2102]. Flies [1118]. Fligner [762]. Flip [19]. Florida [2276, 2275, 2267]. Flow [287]. Flowering [2036]. Flows [2058, 917, 1591, 557]. focus [1683]. Focused [1835, 2304, 1964]. focussed [2342]. Fodder [87]. folded [2329]. folding [2208]. Foldover [1094, 1276]. Follow [2016]. Follow-Up [2016]. Following [768, 2207]. Food [327]. Foot [1829]. Foot-And-Mouth [1829]. Force [1591, 1030, 1029]. Forecasting [2058, 1290, 2360, 313, 1922, 2261]. Forecasts [1722, 2067]. Foreman [1752, 1452]. Forerunners [154]. Foreword [1653]. Form [699, 2181, 292, 313, 1183, 758, 131, 381, 33, 1858]. Formal [407]. Formation [1428, 1413, 335]. Forming [500]. Forms [953, 504, 954, 1723, 1356, 238, 846, 955, 926, 1196, 1262]. Formula [477, 384, 1142]. Formulae [107, 1847, 1269, 1670, 249]. Formulation [757]. Forthcoming [1232, 1246, 1348, 1360, 1377, 1390, 1402, 1409, 1427, 1439, 1451, 1461, 1462, 1474, 1495, 1508, 1518, 1533, 1547, 1563, 1574, 1588, 1600, 1616, 1627, 1641, 1650, 1681, 1698, 1713, 1726, 1738, 1751, 1766, 1777, 1787, 1796, 1803, 1815, 1823, 1832, 1842, 1851, 1859, 1866, 1874, 1882, 1890, 1898, 1907, 1916, 1925, 1935, 1944, 1958, 1973, 1984, 1998, 2008, 2021, 2034, 2046, 2056, 2068, 2082, 2095, 2104, 2113, 2122, 2136, 2147, 2156, 2165, 2176, 2187, 2198, 2204, 2210, 2220]. Forward [599, 2251]. Foundations [2276, 491, 2267]. Four [1113, 1655, 1585, 2071, 241, 516]. Four-Moment [516]. Four-Parameter [241]. Fourier [1316]. Fourth [1123]. Fraction [247, 1700, 2063]. Fractional [454, 2016, 987, 450, 563, 129]. fractionally [2289]. Fragile [1235]. Frailty [1499, 1284, 1760, 1598]. Frames [1001]. Framework [1717]. Fraser [1520]. Fréchet [1304]. Fréchet-Type [1304]. Free [681, 716, 796, 1112, 1110, 1005, 1039, 524, 1293, 449]. freedom [1694]. Freight [299]. Freight-Rate [299]. Frequency [534]. Frequentist [2073, 2142]. Friends [2132]. Frost [416]. fruit [2408]. Fuel [147]. Fuentes [1994]. Fulfilled [1699]. Full [2191, 1034]. Full-Information [2191]. Full-Time [1034]. Fully [1801, 1394]. Function [418, 53, 480, 400, 1186, 2024, 2141, 620, 699, 1171, 671, 966, 1055, 252, 361, 141, 1341, 787, 1582, 1249, 741, 300, 427, 212, 179, 1856, 639, 391, 777, 1120, 852, 1074, 55, 1142, 824, 1914, 838, 1858, 2398]. Functional [1828, 962, 1986, 2386, 1325, 1960, 33]. Functionals [829, 962, 963, 129]. Functions [417, 976, 1342, 109, 1201, 1503, 1646, 37, 31, 1237, 608, 816, 1826, 1122, 1888, 401, 1007, 296, 1894, 949, 2352]. Fundamental [477, 384]. Fundamentals [1996]. Further [276, 2073, 108]. Future [903]. fuzzy [2386].

G [1192, 1150, 2276, 1981, 1090, 165, 1045, 1967]. G. [970]. Gâteaux [1506]. Gain [156]. Gains [25]. Galambos [1046]. Galton [1663, 1181, 726, 286, 234]. Gambling [1674, 634]. Games [944, 1247]. GAMLSS [2406]. Gamma [1114, 471, 625, 371, 1598, 2372, 2399, 1568]. Gammas [627]. gammSlice [2269]. Gang [1381]. Gani [942, 1233, 1934, 1932, 1933]. Gaps [474]. GARCH [2157, 1888, 1865]. Gardner [1997]. Gardner-Lubbe [1997]. Gauss [1623, 1083]. Gaussian [1015, 1113, 1186, 414, 977, 1637, 983, 1491, 484, 2179, 360, 2231, 2397, 2340, 2253, 2377, 1701, 1942, 1351, 2178]. GD [668]. Gelfand [1994]. Gene [1235, 2036]. Genealogy [2010]. General [440, 1752, 757, 1769, 435, 1585, 923, 938, 802, 896, 725, 1844, 1623, 197, 131, 830, 1312, 736, 1693, 2258, 2286, 2225]. Generalisation [930, 881, 2171]. Generalised [2316, 2024, 2116, 2001, 2228, 2269, 2168, 2182, 2272, 2218, 2224, 2254, 2342, 2403, 2311]. Generalization [1625, 1619, 212, 1359, 1504, 1695]. Generalizations [1393]. Generalized [1633, 675, 1735, 1480, 1776, 1880, 1466, 1160, 1501, 1723, 301, 1762, 1420, 1965, 1465, 1437, 888, 257, 725, 1066, 960, 1712, 1976, 1862, 1798, 1305, 869, 1902, 459, 1799, 1810, 2207, 2065, 1841, 2055]. generate [2258]. Generated [49]. Generating [418, 1906, 501, 608, 765]. Generation [1989, 5, 2206]. Genest [1822]. Genetic [1829, 7, 25, 2362]. Genetics [719, 950, 1643, 1996]. genomic [2400, 2401]. Genotype [1705, 30, 1204, 2400]. Genotypic [30]. Geoff [2391]. Geoffrey [1378]. Geometric [1765, 672, 1189, 376, 1239, 1324, 8, 2075, 2218]. Geometrical [384]. Geometry [781, 177, 384, 477]. Geostatistical [2179]. Germany [1661]. germination [1683]. Getting [1978]. Gibbs [1792]. Gil [737]. Gini [1484, 1959, 1593]. Gittins [1331]. Given [1903, 1000, 1665, 641]. Glass [1606]. Global [1320, 2398]. Globally [2352]. Gnedenko [1323]. Golf [1234]. Gompertz [1905]. Gompertz-Type [1905]. Goodness [1692, 822, 1344, 1393, 1514, 764, 56, 109, 316, 800, 438, 2168, 583, 616, 1143, 2364, 2334, 2387, 1687]. Goodness-of-Fit [822, 1393, 764, 1692, 1514, 2334, 1687]. Gower [1997]. GPS [2324]. Graduate [2199]. Graham [1496]. Graph [1825]. Graphical [2058, 601, 1826, 2196, 1876, 1900, 2231, 2340, 2291]. Graphics [1995]. Grassia [627]. Greater [1144]. Greedy [2251]. Greenacre [1980]. Grids [949]. Griffiths [679]. Gross [917, 1591]. groundwater [1564]. Group [1852, 763, 1730, 2169, 369, 1638, 1686, 2287]. Group-Censored [1730]. Grouped [1484, 1320, 60]. Grouping [796, 1112, 138]. Groups [31]. Growth [1203, 663, 463, 1721, 735, 2047, 742, 1660, 724, 726, 815, 347, 33, 1382, 1134, 2408, 2414, 2405]. Grundy [758]. GSS [435]. GTFS [2324]. GTFS-based [2324]. Guarantee [1138]. Guarded [1414, 1447]. Guide [2093]. Guttorp [1994].

H [970, 832, 545, 410, 1967]. H. [546, 612]. Haenszel [1396, 2243]. Half [1142]. Half-Integer [1142]. Hall [1090]. Handbook [2078, 1994]. Handicapping [1234]. Hannan [1209, 329, 1078]. Harbour [1716]. hardback [2276, 2275, 2267]. Harmonic [844, 8]. Harold [1536]. Harris [717]. Hastings [2371]. Hausdorff [1153]. Hausman [2310]. Having [677, 1528]. Haynes [1948]. Hazard [1578, 242, 1885, 816, 2201, 2252]. Hazards [1913, 1019, 2097]. Head [1965]. Health [2163, 1754, 1981, 2273]. Heat [2167]. Heavy [1839, 1846, 1679, 1865]. Heavy-Tailed [1839, 1679, 1865]. Heavy-Tailed-Distributed [1846]. Hedging [2157]. Heiberger [1995]. Henderson [1535]. Henry [379, 1589]. Hereditary [486]. Heritier [1991]. Herriot [2138]. Heterogeneity [2171, 1590, 1098, 182, 1143]. Heterogeneous [1839, 1905, 1743, 2415]. Heteroscedastic [675, 763, 1367, 1137, 1647, 2143]. Heteroscedasticity [1157, 923, 617, 1176]. Heterosis [568]. Heuristic [458]. Heyde [1062]. Hidden [1889, 1918, 1741, 1727, 1381, 2279, 2270, 2301, 1971]. Hierarchical [1748, 1753, 2031, 1212, 1782, 2401]. High [1940, 1939, 2011, 1879, 2091, 2161, 2250, 2231, 2384, 2309, 2241, 2362, 2280]. High-Dimensional [1940, 1939, 2011, 1879, 2091, 2161, 2231, 2309, 2241, 2280]. high-throughput [2362]. Higher [1471, 474, 1750, 1883, 2361]. Higher-Order [1750, 1883]. Highly [1749]. Hilbe [2005]. Hilbert [924, 2110]. Hildreth [921]. Hildreth-Houck [921]. Historical [1752, 947, 548]. History [2080, 968, 2137, 1670, 1945]. Hitting [1298]. HIV [946, 1091, 1272]. HIV/AIDS [946]. Hodges [504, 869]. Holes [2149]. Homogeneity [1449, 841, 812, 582, 1516, 1457, 670, 1003, 845]. Homogeneous [1358]. Hong [2177]. Honorary [102, 159, 168, 496, 658]. Horvitz [578, 325, 673]. Hospital [1893, 1968, 1967]. Houck [921]. Household [1475, 1829, 1364, 2015, 666, 648, 2148, 1590, 327, 320]. Households [1379]. Housing [447, 355]. Huang [1695]. Huber [1708]. Human [913, 950, 1705, 413, 395, 2349]. Hurricanes [1476]. Hurwitz [1768]. Hybrid [2274, 2255]. Hyper [1987, 672]. Hyper-Geometric [672]. Hyper-Lassos [1987]. Hypercubic [406]. Hypergeometric [671, 1370, 551, 1252]. Hypothermia [2084]. Hypotheses [20, 1014, 1956, 189, 1145]. Hypothesis [1052, 1848, 1385, 1127, 230, 2195, 481]. Hypothetical [248].

Ideal [547]. Identical [472]. Identifiability [1515, 127]. Identification [1213, 1793, 529, 2180]. Identify [1926]. Identifying [1838]. Identity [842, 1087]. ignorability [2296]. Ignorable [1591, 2016, 1900]. Ignoring [1579]. II [1199, 2128, 2297, 505, 1447, 1424, 912, 1428, 1413]. Illustrated [3]. Image [1151, 2134, 1983]. Imbedding [163]. Imcomplete [893]. Immigration [655, 692, 706, 286, 234]. Immunoassay [1608, 1856]. Impact [1466, 1248, 395, 2273]. Impatient [165]. Imperfect [1720]. Implementation [1645]. Implementing [1188]. Implication [2012]. Implications [648, 1027]. implicit [2398]. Implicitly [273]. importance [1565, 2403, 2370]. Imprecise [1806]. Improved [440, 1307, 955, 1368, 2217]. Improved-Accuracy [955]. Improving [740, 945, 2241]. Imputation [1254, 2015, 2016, 2148, 1900, 1576, 2296]. INAR [2317]. Inaugural [1534, 1604, 1652]. Incidence [2158]. Incidental [1342]. Including [2117]. Income [1364, 1253, 1258, 2359]. Incomplete [1160, 451, 1070, 772, 155, 684, 677, 2202, 2117, 1513, 1075, 1623, 783, 870, 1142, 817, 470, 1001, 1622, 261, 1553, 2274]. Incompletely [1788]. Incorporating [896, 1895]. Increases [1642]. Increasing [997]. Incremental [1541]. Increments [1710]. Incubation [1272]. Independence [1709, 572, 10, 756, 1189, 604, 571, 1861, 368, 229, 584, 696, 1217, 1296]. Independent [883, 297, 831, 2023, 1227, 1326, 1021, 385, 641, 1755, 2404]. Indeterminate [1336]. Index [1403, 1494, 1548, 1549, 1767, 879, 1871, 1921, 1773, 1959, 1728, 1593, 2302, 2380, 2385]. Indexes [1440, 1602, 1651, 1714, 1804, 1843, 147]. Indicators [503, 2392]. Indices [2012, 1331]. Indirect [167]. Individual [2181, 1999]. Individuals [1853]. Induced [1129, 981, 2193]. Inducing [890]. Industry [1936, 1993]. Inequalities [137, 865, 981]. Inequality [678, 242, 1367, 2197, 628, 1368, 492, 1238, 1251]. Infection [980, 1079]. Infectious [553, 1045]. Inference [691, 711, 877, 956, 1864, 1776, 1752, 1365, 1307, 1962, 491, 2116, 1828, 1877, 2139, 2333, 2108, 1871, 957, 1648, 958, 878, 655, 706, 1175, 1487, 341, 959, 2074, 1223, 2172, 2143, 1486, 960, 1712, 2159, 2049, 744, 1488, 592, 2178, 1782, 1643, 2203, 660, 208, 1969, 990, 933, 1883, 2260, 661, 2319, 2347, 1689, 2366, 2291, 2272, 2033, 2173]. Inferences [1870]. inferiority [2007]. Infill [2355]. Infinitely [755]. Inflated [1311, 2218, 1539]. inflated-parameter [2218]. Inflows [755]. Influence [888, 1303, 1306, 1029, 2264, 1036, 1386]. Influences [1027]. Influential [1789, 779, 1569]. Information [1388, 1065, 2248, 2259, 2268, 2277, 2283, 2290, 2300, 2307, 2314, 2322, 2331, 2337, 2345, 2358, 2368, 2375, 2383, 2390, 2402, 2407, 2413, 766, 2191, 2353, 451, 976, 1541, 1868, 2073, 2127, 2142, 889, 567, 1964, 1443, 470, 1434, 1645, 1819, 2101, 1472, 2342, 1559, 2286]. Informative [1543, 1733]. Inheritance [214]. Inhomogeneous [1864, 2353]. Initial [199]. Injuries [1965]. Innovations [1912, 1374]. Inputs [577, 240]. Inrefence [961]. Insect [622]. Insights [1752]. Inspection [959]. Instability [1922]. Institute [203]. Instructions [885]. Instruments [49]. Insurance [1839, 1668, 2273]. Integer [1142, 2411, 2330]. integer-valued [2411, 2330]. Integral [1467, 242, 477, 384]. Integrated [1718, 2197, 1969, 2289]. Intelligence [2132]. Intensity [1962, 1481, 2351, 2352, 2355]. intensity-reweighted [2352]. Inter [2022]. Inter-Arrival [2022]. Interacting [430]. Interaction [270, 405, 311, 458, 1243, 1026, 1204]. Interactions [2170, 486, 30, 1567]. Intercept [650]. Interchangeability [358]. Interest [1752]. Interface [1995]. Internal [600]. International [753, 203]. Internet [1673]. Interplot [1195]. Interpretation [635, 683]. Interrelationships [1201]. Interruptions [165]. Interval [1091, 1765, 1748, 2141, 460, 1054, 251, 2142, 2071, 1067, 244, 2063, 2232, 639, 1914, 2145]. Interval-Censored [1091, 1748, 2063, 2145]. Intervals [2070, 829, 2103, 1213, 2023, 2130, 2169, 322, 1354, 2050, 1848, 636, 2002, 1187, 1525, 1764, 1868, 1892, 2073, 2127, 625, 623, 2135, 892, 1773, 1849, 2099, 1799, 1686, 2310, 2285, 2246, 2284, 2416, 2379, 2335]. Interview [1223, 1372]. Intra [1076]. Intra-Class [1076]. Intraclass [1236]. Introduction [1971, 1024, 2092, 971, 2019, 1841, 2033, 2299, 1970, 1943]. Introductory [2184]. Invariance [649, 793]. Invariant [1619, 391, 1665, 740, 2196, 1196, 1241, 1217, 2384]. Inventory [157, 527, 237]. Inverse [414, 233, 983, 846, 1370, 1198, 1789, 2149, 926, 806, 2287, 2377, 1693, 1683, 2373]. Inverses [257, 725]. Inverting [440, 2284]. Investigation [2245, 682]. Invited [446, 2235, 1947, 1512, 2107, 1795, 1946, 2069, 2096, 2166, 2124, 2123, 1945, 1948, 1876, 1883]. Irreversible [778]. ISBN [2276, 2275, 2267]. Isolating [959]. Isolation [1118]. Isometry [793]. Isotropic [1201]. Issue [2248, 2259, 2268, 2277, 2283, 2290, 2300, 2307, 2314, 2322, 2331, 2337, 2345, 2358, 2368, 2375, 2383, 2390, 2402, 2407, 2413, 968]. Issues [1674, 904, 2032, 1281]. Item [2079]. Items [959, 462]. Iterated [561]. Iteration [384, 945]. Iterative [1896, 1924, 2040, 1068, 459]. IV [1858]. Ivan [2007].

J [1192, 1222, 1310, 1150, 2275, 1821, 1981, 2005, 1822, 708, 1078]. J-F [1822]. J. [942, 1233]. Jackknife [701, 923, 1911]. Jackknifing [652, 1396, 736]. Jalaludin [1967]. James [1209, 1168, 1093, 2193]. Jannach [2276]. Janos [1046]. January [1378, 1739]. Johansen [1912]. John [1997, 1654]. Johnson [1739]. Joint [1860, 187, 360, 2231, 238, 723, 1513, 89, 2190, 674, 2066]. Joint-regression [1513]. Jonckheere [1412]. Jonckheere-Type [1412]. Jones [1970]. Jorgensen [1795]. Journal [185, 195, 211, 227, 2289, 968, 1335, 1]. judgement [2366]. Judges [1927]. July [445, 490]. Jump [1812, 179, 213, 1918, 2295]. Jumps [1431]. June [447]. Justice [2067]. Justification [312].

Kairat [2020]. Kalman [1054, 586]. Karenina [2357]. Karl [1972]. Ken [1752]. Kendall [931, 1961, 965]. Kernel [1448, 1658, 1315, 2107, 1891, 788, 1072, 1200, 2047, 2309, 2256, 1759, 949, 2273, 1631, 2252, 2355]. kernel-based [2252]. Keyword [1549]. Kind [281]. Knibbs [1132, 1280, 491, 445, 970, 1699]. Knowledge [1429]. Known [1903, 26, 1120]. Kolmogorov [1393, 813, 1384]. Komlós [769]. Kong [2177]. Krawtchouk [2003, 215]. Krige [1489]. Kruskal [1592, 2404]. Kshirsagar [1333]. Kullback [1676]. Kurtosis [1621, 935, 1119].

L [1971, 2007, 2006, 661, 1954]. Label [2196]. Label- [2196]. Labels [1151]. laboratory [2297]. Labour [1028, 1591, 920, 1030, 1791, 1029]. Labour-Force [1029]. LaBudde [2019]. Lack [2265, 1647, 1904]. Lack-of-Fit [1647, 1904]. Lacking [429, 505]. Lags [355]. Laguerre [2348]. Laird [1996]. Lancaster [545, 2109, 1589, 546, 612, 1455]. Landfalling [1476]. Lange [1996]. Langevin [853, 2217]. Laplace [1778, 875, 2328, 1800].

Large [1668, 1475, 792, 2024, 288, 655, 1200, 581, 2162, 1759, 554, 1834, 1351, 1271, 354, 2417, 706]. Large-Scale [2024, 288, 1271, 354, 1475, 2417]. Largest [390, 421, 1871, 585, 293, 229]. Laser [1817]. Lasso [1877, 1625, 1695, 2182]. Lassos [1987]. Late [1917, 1752, 970]. Late-Stage [1917]. Latent [1962, 2001, 2022, 2138, 293, 1881, 2296, 2404, 2112]. Latent-State [2022]. later [2297]. Latin [311]. Latinized [1361, 1190]. Lattice [294, 575, 982, 511, 459]. Lattices [731]. Law [1834, 561, 2192]. Lawless [2275]. Laws [536, 1500, 1147]. Layout [1243, 449]. Layouts [1801]. Lead [237]. Leading [970]. league [1418]. Learning [2067, 2406]. Least [675, 268, 299, 1183, 2162, 1523, 1457, 2151, 869, 487, 2297]. Least-Squares [268]. Leave [1314]. Leave-One-Out [1314]. Leavers [1029]. Lecture [1132, 490, 1280, 424, 1417, 1552, 491, 445, 1452, 1699, 2109]. Lehmann [504, 869]. Lehmann-Type [504]. Length [1316, 2059, 609, 1147, 2221, 1554, 2323]. length-biased [2221, 2323]. Length-Biasing [1147]. Length-Biassed [609]. Letter [913, 1279, 1492, 1628, 396, 1968, 657, 1821, 1545, 110, 899, 1649, 1822, 885, 1345, 350, 1639, 1967]. Letters [1333, 1472, 1332]. Level [1769, 1734, 1670, 2155, 2196, 640, 79, 2338, 1509]. Level-Invariant [2196]. Levels [1400, 1878, 1564]. Levene [1457, 845]. Leverage [888, 1902, 1274]. leveraging [2400]. Lévy [1384]. Liang [1410]. Life [102, 159, 168, 496, 658, 68, 1754, 2052, 1138, 2026, 5, 371, 557, 1915, 32, 1692, 2245]. Life-Test [68]. Ligges [2267]. Likelihood [1114, 743, 372, 1182, 1270, 357, 917, 1315, 1730, 1802, 1793, 1812, 476, 938, 934, 285, 1137, 2074, 1163, 91, 912, 1106, 2039, 1870, 2061, 1701, 1320, 995, 1856, 1942, 1773, 1959, 738, 2060, 1212, 2054, 296, 587, 1729, 1380, 60, 1800, 1445, 857, 1056, 1969, 1266, 1482, 1775, 1810, 2194, 2318, 1421, 2333, 1686, 2274, 1570, 2272]. Likelihood-Based [1315, 1870, 2061, 1773, 1380]. Likelihoods [1161, 321, 2369]. Limit [1357, 2013, 1358, 2085, 234, 996, 221, 937]. Limitations [1620]. Limited [976, 617]. Limiting [620, 292]. Limits [638, 20, 57, 26, 1328, 1369, 1456, 1691]. Line [1514, 287, 1070, 800]. Lineage [1099, 1487]. Linear [1840, 1505, 1448, 799, 1091, 1896, 1924, 1910, 2040, 766, 695, 1160, 570, 1038, 1676, 429, 505, 1339, 1043, 791, 36, 564, 1018, 923, 317, 1491, 2153, 966, 22, 44, 1367, 802, 1906, 2065, 2001, 2137, 1722, 932, 2148, 888, 2100, 2118, 1870, 2020, 257, 725, 300, 1441, 960, 1325, 1166, 1264, 1220, 1887, 1712, 1779, 728, 393, 346, 777, 2060, 2196, 650, 1212, 868, 803, 2098, 65, 1093, 906, 129, 1756, 2037, 1265, 1238, 2154, 736, 1324, 1036, 2099, 2075, 2101, 1775, 2344, 2240, 2282, 2265, 2288, 2254, 1468, 2385, 2229, 2342, 2369, 2328, 2403, 2207, 2286, 2311, 1682]. Linear [1982]. Linear-By-Linear [1896, 1924]. Linearity [1317]. Linearly [306]. Lines [1449, 1989, 1298]. LINEX [1074]. Linkage [1103, 2211]. Linked [2098, 800]. Linking [2028]. links [2212]. Linnik [1500]. Lipton [1192]. List [333, 685]. Lists [2177, 1830]. Lloyd [1739]. Lluch [346]. Loadings [2162]. Local [1448, 1802, 1722, 795, 1952, 1688]. Localized [1704, 1684]. Locally [1237, 1950]. Location [1732, 1259, 1182, 1261, 324, 367, 1619, 1073, 740, 1963, 171, 517, 2239, 2363, 2406]. locations [2324]. Log [1091, 1896, 1924, 2040, 1441, 1320, 2196, 2160, 60, 1265, 2222, 2266]. log-concave [2266]. Log-Concavity [1320]. Log-Linear [1091, 1896, 1924, 2040, 1441, 2196, 1265]. Log-Normal [60]. Logarithm [561]. Logarithmic [486]. Logistic [702, 574, 1349, 1383, 1338, 1336, 1434, 2292, 198, 1566, 2005]. Logit [1210, 2272]. Loglinear [1941, 825]. Lognormal [357, 100, 241]. Long [1448, 1677, 1886, 2114, 851, 1351, 1669, 1863, 818, 2400]. Long- [1886]. Long-Memory [1863, 818, 1351]. Long-Range [1448, 1677, 1669]. Long-term [2114, 2400]. Longest [1833]. Longitudinal [1203, 1480, 1711, 2211, 1928, 1224, 1033, 2029, 2066, 2112, 1954, 2084, 2172, 2202, 1284, 1216, 2124, 1895, 2190, 1727, 1031, 1025, 2359, 2274, 2327, 1568]. look [2235, 2288]. Looking [599]. Lopsided [2000]. Loss [1065, 2141, 2152, 741, 1446, 156, 1074, 1914, 2263, 2340, 2304]. loss-based [2340]. Lotto [1740]. Low [1178, 2250]. Low-Order [1178]. Lower [2197, 1337, 1021, 1853, 475, 3]. Loynes [1044]. Lubbe [1997]. Luck [1247].

M [1233, 1995, 2005, 2019, 1472, 1044, 1994, 1996, 165]. M. [661]. M/G/1 [165]. Macdonald [1971]. Machine [2067]. Magazine [1051]. Magnitude [1520]. Mail [890]. Maillardet [1970]. Main [1395]. Maitland [596, 646]. Major [769, 1235]. Making [328]. Malawi [1499]. Malformations [1211]. Malt [1706]. Man [551]. Management [1406]. Managing [1519]. Mann [1504]. Manova [2161]. Manpower [1064]. Mantel [2243, 1396]. Mantoux [1000]. Manufacturing [147]. many [2237, 1568]. many-to-many [2237]. Maori [1576]. mapping [2339, 2395]. Marginal [1349, 715, 2153, 1830, 1003, 1598]. Marginals [1626]. Margins [1455]. Maria [1991]. Maria-Pia [1991]. Marihuana [615]. Marine [1441]. Mark [2007, 1993]. Marked [1667]. Markers [1951]. Market [1028, 920, 15, 634]. Marking [1271]. Markov [1971, 2112, 1129, 1626, 372, 691, 1864, 107, 158, 187, 1051, 1812, 1479, 302, 683, 1358, 1702, 692, 1465, 2279, 1151, 882, 1005, 1035, 920, 625, 1889, 1079, 664, 1623, 2270, 1918, 755, 1741, 1727, 178, 2301, 419, 1669]. Markov-Renewal [372]. Markovian [158, 1064, 655, 706, 244]. Marks [1913, 1270, 835, 1269, 2348]. Martingales [291, 1952]. Mass [2015, 2256]. match [2261]. Matched [751, 1949, 744]. Matched-Paris [751]. Mathematical [948, 395, 547]. Mathematics [1392, 1577]. Mating [1807, 1250]. Matric [626]. Matric-Variate [626]. Matrices [601, 602, 1702, 206, 1706, 223, 1041, 1146, 450, 303]. Matrix [440, 1940, 986, 1064, 671, 392, 188, 1002, 925, 557, 693, 293, 806, 1174, 437, 1217, 2392]. matrix-normal [2392]. Maurice [410]. Maxima [125, 236, 178, 1294]. Maximal [1872, 918]. maximisation [2398]. Maximization [1869]. Maximize [53]. Maximum [416, 1477, 1114, 372, 357, 917, 1730, 2318, 1229, 1675, 476, 285, 2074, 91, 29, 2293, 1701, 1856, 1327, 738, 1212, 587, 824, 1530, 1729, 60, 1445, 1056, 1775, 1421, 2054]. Maximum-Entropy [1675]. May [462]. Mayo [1520]. McLachlan [2391]. MCMC [2295, 1744]. Mean [573, 342, 539, 1273, 1544, 1459, 844, 251, 26, 2011, 1138, 1121, 1263, 741, 628, 1790, 1082, 1976, 89, 939, 2088, 1330, 798, 1915, 1873, 382, 419, 1265, 1145, 933, 1217, 2347, 2217, 2343, 2213, 1990]. Means [53, 414, 1979, 390, 421, 1610, 1014, 1154, 878, 726, 582, 623, 1486, 538, 1808, 74, 1919, 8, 13, 108, 2099, 1636]. Measure [2038, 1458, 1621, 679, 1350, 1333, 1287, 1761, 1665, 1528, 1821, 1822]. measured [2412]. Measurement [90, 1731, 1063, 1025, 1960, 2328]. Measurements [1375, 994, 1644, 1895, 1071, 1838]. Measures [1709, 694, 751, 2103, 1133, 925, 1103, 1920, 1608, 2160, 653, 2334, 2403]. Measuring [2376, 2410]. Mechanisms [2202, 765]. Mechanistic [1794]. Medal [2315, 542, 612, 686, 1062, 1168, 1222, 1233, 1310, 1391, 1496, 1603, 1715, 1850, 1875, 2035, 2189, 1090, 1078]. Median [570, 370, 200, 628, 826, 1963, 1330, 730, 700, 1502, 229, 1690]. Medical [322]. Meixner [1455, 551, 672, 723, 1252]. Melanoma [581]. Melbourne [1465, 354]. Member [102, 159, 168, 496]. Members [658]. Memoriam [1739, 1589]. Memory [1863, 818, 2333, 1351, 2020]. memoryless [1692]. Mendel [1181]. Mersmann [2267]. Message [2, 2373]. Meta [1938, 2018, 1248, 2135]. Meta-Analysis [1938, 2018, 1248, 2135]. Metapopulation [1663]. Metastatic [522]. Methadone [1349]. Method [1410, 748, 1748, 1988, 1313, 959, 1581, 1435, 1437, 1582, 1314, 875, 1780, 1830, 524, 1248, 924, 1773, 1053, 86, 1068, 869, 1001, 610, 518, 1008, 2294, 2416]. Methodology [1618, 2094]. Methods [2086, 1423, 668, 1692, 1388, 1572, 2012, 1489, 1633, 2214, 1448, 1677, 1934, 1626, 1735, 1709, 1635, 1940, 1763, 2128, 1732, 1975, 1162, 1410, 1671, 1638, 1467, 1941, 1398, 1425, 1527, 1896, 1924, 2010, 2040, 1634, 1749, 2038, 1515, 1718, 1560, 1711, 1458, 1621, 2170, 1449, 1538, 2218, 1471, 1567, 1684, 1704, 681, 1442, 1565, 1412, 1765, 2062, 2144, 1748, 2141, 1721, 1676, 1977, 1387, 1611, 1540, 1544, 1730, 1758, 1586, 2015, 1802, 1529, 1962, 1578, 1213, 1459, 1708, 2245, 1632, 1470, 1512, 891, 1421, 1675, 2222, 1932, 1610, 1710, 1188, 1637, 1584, 2107, 2013, 2003, 2240, 1539, 1414, 1447]. Methods [2130, 1400, 2169, 1893, 1801, 1750, 1393, 1514, 1383, 1723, 1490, 1736, 1762, 1397, 1731, 1931, 2050, 2181, 1961, 1566, 1585, 1928, 1929, 1385, 2152, 2163, 1987, 2025, 1702, 1491, 1988, 1411, 1420, 1939, 2153, 2179, 1424, 1648, 1631, 2244, 1930, 2231, 1734, 1754, 2065, 1644, 1686, 1137, 1526, 2083, 1503, 1625, 1690, 2002, 2224, 1363, 1487, 1556, 1705, 1688, 1419, 1454, 1555, 1986, 1722, 1680, 1674, 1541, 1415, 1581, 2074, 1435, 1456, 1525, 1524, 1691, 1764, 2073, 2127, 2142, 2197, 2171, 2011, 2228, 1685, 1612, 2059, 2180, 2216, 2016, 2085, 1437, 1436, 1582, 2028, 47, 1455, 1573, 1384, 1921]. Methods [1583, 1646, 1619, 1719, 1433, 1485, 1554, 2100, 2118, 1706, 1647, 2039, 2172, 2202, 1746, 2246, 1446, 1561, 2087, 1432, 1468, 1519, 1558, 1620, 1949, 2126, 2230, 1570, 2223, 1693, 1399, 2063, 2145, 1720, 1922, 1499, 2229, 1444, 389, 2077, 1543, 1136, 2166, 2019, 1920, 1950, 501, 1613, 2143, 2061, 2162, 1974, 1759, 1486, 1516, 2232, 1757, 2155, 1712, 1579, 2124, 1976, 1513, 1707, 1531, 1744, 1609, 1523, 1608, 1457, 1942, 1623, 1394, 1689, 2159, 2049, 2200, 1918, 1679, 1959, 2201, 1396, 1761, 2041, 1741, 2208, 2088, 2060, 1559, 1395, 2196, 1733, 2048, 2076, 2151, 2243, 1488, 1964, 2195, 1952, 1695, 1963]. Methods [2217, 1469, 1740, 1747, 1798, 1443, 1607, 1522, 1557, 1703, 2134, 2004, 2160, 2081, 2109, 1542, 1636, 1530, 469, 1381, 1422, 2089, 1434, 1622, 1271, 1729, 1991, 1413, 1643, 2091, 1380, 2242, 2203, 2129, 1569, 1431, 1580, 1756, 1800, 2037, 1537, 2168, 1717, 1645, 2207, 1528, 1960, 2182, 1445, 1919, 2161, 2241, 1408, 1760, 1745, 2043, 1177, 1382, 1742, 2154, 2193, 2225, 1900, 1917, 1379, 1553, 1933, 2090, 1678, 2072, 2102, 1571, 1694, 2099, 1755, 1799, 2075, 1724, 2215, 1568, 1728, 1624, 1743, 2101, 1687, 1951, 2014, 2213, 1386, 2194, 2110, 2239, 2274, 2361, 2271, 1968, 1967, 2031, 2119]. Metrics [1384, 547]. Metropolis [2371]. Michael [1980]. Mid [1686]. Mid- [1686]. Midzuno [881]. Migration [600, 156]. Mile [477]. Mine [140]. Mineral [1405]. minification [2411]. Minimal [1626, 1938, 720]. Minimal-Repair [1938]. Minimax [741, 1597, 777, 1156]. minimising [2229]. Minimum [710, 716, 2409, 474, 1892, 1384, 677, 1173, 1608, 1942, 249, 981, 803, 824, 589, 382, 1554, 2304]. Mining [1429]. Mis [1490, 1238, 1251]. Mis-specification [1490]. Mis-Specified [1238, 1251]. Misclassification [883, 958, 300, 1286]. Misclassified [1729]. Mises [2217, 607, 1421, 1488]. Misreported [1644]. Missing [1913, 1254, 1891, 1383, 2017, 2074, 2016, 2202, 1870, 1260, 2124, 1178, 1742, 1900, 696, 24, 2344, 2364, 2359]. Missing-at-Random [2124]. Mistakes [2107]. Mixed [440, 294, 1470, 2065, 2001, 1465, 2029, 2148, 1441, 1701, 1712, 1148, 511, 1212, 2037, 1966, 1265, 1784, 1805, 468, 2240, 2282, 2288, 2254, 2269, 2326, 1682, 1568, 2272]. Mixed-Effects [1441, 1966]. Mixing [2118, 2077, 274, 1661, 441]. Mixture [1159, 1349, 1387, 2108, 1419, 1140, 1673, 2084, 1135, 2115, 1791, 1606, 1516, 2014, 2374, 2205, 2397, 2359, 2393, 2399, 2242, 2335, 2417, 2401]. mixture-based [2205, 2397]. Mixtures [510, 1854, 750, 1741, 1755, 2392]. Modal [2394, 2328]. Mode [1778, 628, 1330]. Model [1012, 1203, 1159, 675, 702, 298, 214, 778, 1019, 1091, 1941, 695, 763, 2170, 72, 1365, 2062, 1748, 757, 1038, 1676, 429, 505, 1339, 1544, 1962, 1835, 1785, 977, 1051, 486, 174, 947, 1171, 1750, 2205, 564, 1877, 1490, 1762, 2108, 1655, 1937, 923, 1126, 1367, 802, 529, 1648, 1965, 1329, 578, 1224, 1290, 2027, 767, 1099, 1487, 2097, 1986, 617, 1892, 525, 1210, 2084, 1138, 1436, 1582, 742, 1135, 1921, 1308, 1791, 1922, 825, 1499, 2077, 1543, 725, 1284, 1079, 2117, 1441, 1516, 1264, 1887, 2155, 1976, 1623, 2366, 527, 393, 897, 395, 407, 273, 2004, 592]. Model [1292, 1068, 1530, 906, 1782, 1380, 1234, 660, 1003, 1966, 1819, 1238, 1251, 1742, 1784, 568, 736, 1036, 1134, 1805, 2090, 2075, 335, 1743, 1247, 2381, 2400, 2344, 1560, 661, 2408, 2359, 1419, 2380, 2279, 2320, 2229, 2372, 1559, 2399, 2292, 2329, 2286, 2278, 2330, 1568, 2272]. Model* [313]. Model-Assisted [1750]. Model-Based [1171, 1308, 2004, 592, 660, 2117, 2366]. Model-Unbiasedness [578]. Modeling [2079, 2100, 1771, 1982, 2190, 2030, 2406, 1983]. Modelled [920]. Modelling [946, 1718, 948, 1307, 1854, 2349, 1716, 1133, 835, 1539, 1794, 1032, 2324, 950, 1931, 951, 1929, 1753, 2137, 1465, 2114, 1033, 2158, 2022, 1418, 2183, 1216, 1953, 413, 2036, 1747, 1564, 1902, 978, 1879, 1661, 1580, 1717, 2392, 2362, 952, 753, 1521, 949, 2374, 2351, 2325, 2393, 1630, 1558, 1570, 2270, 2388]. Models [1015, 665, 1840, 1971, 2058, 916, 1505, 921, 1860, 1668, 510, 1941, 1896, 1924, 2040, 766, 739, 1776, 1186, 1160, 1349, 1387, 1478, 1540, 1846, 1835, 1591, 1108, 288, 688, 683, 1253, 1731, 1886, 1961, 1342, 1491, 2153, 2179, 586, 2065, 1644, 636, 1137, 2001, 1164, 1503, 2157, 735, 1175, 1172, 1235, 1363, 1820, 1465, 2029, 1140, 2066, 2112, 1673, 139, 2158, 1338, 1954, 2047, 982, 1151, 1163, 1355, 14, 1106, 932, 1127, 2148, 1719, 888, 2118, 1818, 1647, 2115, 2172, 1870, 1869, 2096, 2053, 815, 1098, 1100, 1889, 1701, 311, 1320, 960, 1712, 1779, 329, 1609, 2049, 1918, 1306]. Models [738, 2201, 248, 2138, 1741, 535, 534, 2060, 2196, 826, 1212, 1727, 1841, 1964, 1311, 587, 1031, 1178, 1093, 1434, 1083, 1902, 1945, 1643, 518, 2203, 854, 1580, 1756, 1295, 2037, 1645, 1139, 1003, 1265, 146, 952, 1408, 1760, 1598, 1876, 1312, 1863, 1881, 2154, 1239, 1324, 1799, 1772, 818, 1266, 1728, 2101, 1775, 1386, 2344, 2296, 1632, 1421, 2222, 2240, 2302, 2289, 2333, 2282, 2265, 2231, 2340, 2288, 1555, 2206, 2279, 2228, 2254, 1573, 2334, 2332, 2385, 2415, 1683, 2270, 2342, 2291, 2269, 2405, 2387, 2264, 2403, 1569, 2207, 2301, 2286, 2225, 2311, 1682, 1687, 2401, 2005]. Moderate [1759]. moderation [2361]. Modern [2065, 1996]. Modification [573, 755]. Modified [1567, 1003, 1134, 2333]. MODIS [2178]. Modulo [454, 1241]. Moivre [2064]. Moment [1350, 1269, 1807, 842, 356, 1147, 516, 50, 1810]. Moments [1113, 1515, 392, 846, 1974, 306, 1507, 1087, 926, 2168]. Monitoring [1642, 2221]. Monotone [242, 340, 2236, 623]. Monotonic [1481]. Monotonically [1110]. Monotonicity [1666]. Monte [2294, 752, 1712, 1918, 1041, 381, 1993]. Mood [2171]. Moran [686, 1359]. Morgan [1967]. Mortality [27, 958, 1905, 29, 490, 1966]. Mosaic [1105]. Most [1950, 2213]. Mouth [1829]. move [2384]. Moving [416, 440, 767, 690, 738, 1635]. MPS [2388]. MSE [803]. Mulaik [1982]. Multi [911, 649, 1962, 688, 250, 1014, 1734, 2126, 1689, 2201, 1299, 2151, 1963, 470, 1838, 937, 1951]. Multi-Arm [2126]. Multi-Auxiliary [470]. Multi-Category [1951]. Multi-Channel [1838]. Multi-Dimensional [937]. Multi-Factor [1299, 2151]. Multi-Level [1734]. Multi-Normal [1014]. Multi-Response [649]. Multi-Sample [1963, 1689]. Multi-Stage [911, 688, 2201]. Multi-State [1962]. Multi-Way [250]. Multicollinearity [977]. Multidimensional [1358, 989, 1455, 1983]. Multigrade [665]. Multilevel [1945]. Multimix [1419]. multimodality [2256]. Multinomial [66, 269, 330, 375, 864, 1673, 1370, 1609, 750, 481]. Multinomials [120]. Multiperiod [1290]. Multiphase [2235, 2297]. Multiple [1388, 82, 713, 1592, 1043, 973, 718, 2139, 802, 1906, 1110, 2177, 2000, 2026, 18, 1318, 1830, 1700, 669, 889, 871, 535, 637, 2195, 858, 730, 801, 79, 1881, 1838, 1951, 2231, 1472, 2417]. Multiple-Record [1881]. Multiplicative [1371, 1253, 270, 405]. Multiplicity [666]. Multisample [934]. Multiserial [415]. Multistage [823]. Multistate [1754, 2053]. Multistrata [772]. Multivariate [2058, 694, 53, 2133, 754, 1793, 564, 2139, 1961, 1928, 383, 484, 2153, 1224, 149, 151, 139, 1338, 1612, 1058, 1831, 1370, 932, 133, 741, 1974, 1759, 727, 752, 790, 2404, 967, 1623, 2270, 1146, 2138, 2175, 740, 1825, 852, 964, 1963, 867, 1093, 629, 1152, 1879, 830, 1301, 1302, 1343, 1139, 584, 884, 1838, 2214, 2128, 2239, 2363, 2289, 2257, 1636, 2258, 2280, 2131]. Multivariate- [1093]. Murray [3, 1969]. Murrell [1943]. Murthy [1522]. music [2238, 2276]. Mutation [568]. Mynbaev [2020].

N [2289, 1841]. Naive [879, 2309]. Nan [1996]. narrower [2284]. National [426]. Natural [568]. Nbu [1746]. Nbu-Ness [1746]. Near [1244, 1327, 1899, 1206]. Near-Optimal [1244]. Nearest [1611, 1117]. Nearness [1154]. Necessary [563]. need [885]. Needed [1076]. Negative [1901, 1776, 1836, 1774, 1700, 223, 750, 1262, 1423, 1418, 2321]. Negatively [1770, 2154]. Negatives [1700]. Neighbour [1611, 1126, 1117, 1195, 839, 2312]. Neighbourhoods [1485]. Neither [1489]. Nerve [146]. Ness [1746]. Nested [908, 1499, 960, 2098, 1256]. nesting [1691]. Nets [2130]. Network [1929, 1825, 2324, 2244, 2279, 2369]. Networks [2011, 2247]. Neuwirth [1995]. News [6, 12, 16, 21, 28, 34, 40, 46, 52, 58, 63, 69, 75, 83, 93, 98, 103, 117, 124, 135, 142, 152, 160, 166, 176, 184, 193, 202, 210, 218, 225, 245, 255, 265, 271, 279, 304, 318, 331, 338, 351, 364, 377, 386, 397, 411, 422, 432, 443, 465]. Newton [1083]. next [2206]. Neyman [301, 420]. NGINAR [2218]. Niël [1997]. Niels [1045]. Nine [31]. Nineteenth [1552]. No [1790, 277]. Noel [2018]. Noise [1113, 962, 1680, 1058]. noisy [2304, 2401]. Non [1092, 1901, 1638, 1896, 1924, 1910, 2040, 1864, 134, 1880, 236, 431, 1261, 2007, 649, 574, 2144, 841, 1339, 460, 1591, 977, 485, 2363, 810, 358, 1987, 1491, 1774, 44, 1705, 292, 119, 2016, 1037, 47, 91, 932, 1646, 1468, 223, 1701, 1325, 1597, 472, 1942, 527, 430, 65, 1031, 2178, 1873, 1281, 128, 232, 283, 295, 1256, 1897, 1262, 1900, 1324, 961, 1728, 949, 2317, 2395, 2341, 2328, 2292, 2401]. Non-Additivity [932]. Non-Adjacent [1880]. Non-Binary [1256]. Non-Central [292, 119, 128, 232, 295]. Non-Compliance [1897]. Non-Convex [1987]. Non-Crossing [2144]. Non-Equispaced [949]. Non-Ergodic [1092]. Non-Existence [460, 91]. Non-Gaussian [977, 1491, 1701, 1942, 2178]. Non-Homogeneity [841]. Non-Identical [472]. Non-Ignorable [1591, 2016, 1900]. Non-inferiority [2007]. Non-Interacting [430]. Non-Iterative [1896, 1924, 2040]. Non-Linear [44, 1325, 65, 1324, 1468, 2328]. Non-Negative [1901, 1774, 223, 1262]. Non-Normal [134, 236, 649, 574, 1339, 485, 283, 2401]. Non-Orthogonal [961]. Non-Parametric [1864, 431, 1261, 358, 1705, 47, 1646, 1597, 1031, 1728, 1638, 2363, 2395, 2341]. non-randomized [2292]. Non-Response [1591, 1873, 1281]. Non-Sequential [527]. Non-Similar [1037]. Non-Stationary [810, 2317]. Non-Symmetrical [1910]. Noncentral [392, 151, 693, 1096].

Noncompliance [2061]. Nonconforming [959]. Noninferiority [1956]. Nonlinear [2086, 1540, 1133, 1374, 1794, 1818, 1701, 1902, 518, 1085, 2405, 1569]. Nonlinearity [1758]. Nonnegative [1296]. Nonnegative-Definite [1296]. Nonnull [934]. Nonorthogonal [2297]. Nonparamentric [831]. Nonparametric [907, 1913, 877, 1938, 638, 620, 2094, 2238, 2107, 749, 1801, 1828, 1184, 651, 2273, 1228, 1411, 1055, 1197, 2177, 2097, 1341, 1313, 1176, 1824, 1830, 2077, 1073, 1779, 2060, 650, 1966, 1177, 2374, 2115, 2246, 2225, 2417]. Nonregular [1106]. Nonreplicated [458, 1243]. Nontransitive [760]. NOPE [2282]. Norm [450]. Normal [573, 53, 1182, 275, 134, 342, 125, 236, 537, 649, 954, 574, 1161, 1339, 390, 421, 698, 607, 485, 1356, 1014, 1154, 1049, 251, 878, 1224, 2001, 1235, 26, 341, 846, 149, 139, 188, 1067, 1121, 872, 955, 741, 1021, 623, 727, 752, 790, 1082, 306, 1286, 1087, 148, 131, 867, 1542, 926, 1196, 60, 438, 198, 283, 1445, 1145, 13, 1036, 50, 108, 1217, 2416, 1683, 2392, 2335, 2401]. Normality [1092, 1080, 636, 1048]. Norman [1739]. Note [1080, 1388, 590, 39, 78, 1735, 418, 1880, 337, 1282, 1459, 1990, 451, 90, 323, 27, 947, 476, 564, 1255, 586, 140, 408, 1338, 2059, 91, 1067, 206, 932, 1500, 1543, 1790, 374, 605, 866, 738, 1165, 637, 783, 395, 922, 670, 376, 369, 74, 450, 309, 402, 469, 470, 594, 1605, 171, 295, 806, 857, 1745, 8, 895, 240, 1637, 2263, 1472, 2286]. Notes [6, 12, 16, 21, 28, 34, 40, 46, 52, 58, 63, 69, 75, 83, 93, 98, 103, 117, 124, 135, 142, 152, 160, 166, 176, 184, 193, 202, 210, 218, 225, 245, 255, 265, 271, 279, 304, 318, 331, 338, 351, 364, 377, 386, 397, 411, 422, 432, 443, 465, 695, 252, 141, 502, 943, 361]. Notice [105, 1391, 1496, 1519]. Notices [35, 38, 388, 399, 467, 489, 509, 521, 544, 580, 598, 632, 705, 771, 808, 820, 834, 861, 874, 887, 901, 915, 928, 941, 999, 1011, 1023, 1060, 1102, 1116, 1131]. November [1132, 1739, 1280]. NSW [1392]. Nuisance [933]. Null [603, 2417]. Number [1783, 480, 1375, 302, 2177, 140, 244, 1327, 1909, 2396, 2397, 2206]. Numbers [1648, 667, 1899, 1834, 1656]. Numerical [243]. Núñez [1954]. NZSA [2105, 2106, 2308].

O [546, 612, 2267, 545, 1970]. Obituaries [1192]. Obituary [278, 289, 363, 596, 1378, 1209, 410]. Objective [2014]. Objects [2139]. observation [2263]. Observations [907, 711, 1529, 2013, 2139, 1155, 802, 878, 1197, 795, 2202, 1107, 1899, 393, 1789, 870, 893, 798, 779, 2090, 1215, 2128, 1690, 2237, 1569, 2241]. Observed [877, 986, 1812, 2260]. Observer [856]. Obtained [471, 1869]. Occasion [1654]. Occasions [593, 643, 308, 539, 876, 939, 197, 345, 309]. Occupancy [2071, 2320]. Occupational [502]. occurs [2364]. October [424]. Odd [551]. Odds [2381, 2160]. Odds-symmetry [2381]. Off [1930]. Office [557]. Official [1452, 970, 850, 849]. Offspring [726]. Old [1905, 2360, 2022]. Old-Age [1905, 2360]. Oliver [1589]. Olse [1975]. Omitted [535]. Oncology [2078]. One [2070, 297, 454, 1273, 569, 199, 1520, 1801, 642, 1837, 1126, 773, 525, 370, 1314, 582, 1073, 1122, 461, 430, 1007, 884, 1241, 1560, 2297]. One- [1122]. One-day [1520]. One-Dimensional [1126]. One-Parameter [461]. One-Sample [642, 1073]. One-Sided [1273, 884]. One-Truncation [1007]. One-Way [2070, 1801, 525, 582, 1560]. online [2398]. Only [1878, 1690]. onto [1366]. OPE [2282]. Open [600, 1830]. Operations [15]. Operator [1162, 909]. Operators [963]. Optimal [1185, 1259, 691, 711, 1019, 1042, 2170, 1387, 1989, 1244, 1871, 878, 1612, 205, 216, 362, 2229, 501, 1845, 1405, 622, 1220, 189, 987, 1075, 1205, 538, 393, 1144, 1395, 555, 1813, 870, 893, 1071, 854, 1206, 1251, 479, 237, 2215, 2262, 1008, 2014, 1634, 2336]. Optimality [1867, 879, 912, 1620, 837, 2150]. Optimizing [1734]. Optimum [841, 414, 250, 374, 74, 402, 469]. Options [1723, 2157]. Oration [1417, 1552]. Orbits [1736]. Orcutt [1013]. Order [1065, 1113, 1162, 1880, 400, 768, 1471, 1478, 1781, 474, 2013, 1750, 1411, 1524, 1128, 1127, 1899, 1679, 738, 1964, 803, 1178, 1196, 1786, 307, 229, 1085, 1324, 1317, 1678, 1883, 2128, 1421, 2330]. Ordered [907, 1398, 1527, 1412, 1586, 1762, 1154, 1110, 878, 1020, 582, 1613, 2190, 2151, 1528, 1425, 1632, 2265]. Ordering [1908, 237]. Orderings [1293]. Orders [1220]. Ordinal [1896, 1924, 2040, 2191, 2108, 1955, 2205, 2228]. Ordinary [1736]. Ordinates [534]. Orginal [1344, 1339, 1342, 1341, 1337, 1338, 1336, 1340, 1343]. Orientations [384]. Original [334]. Originating [562]. Origins [260]. Orthant [752]. Orthogonal [43, 549, 908, 811, 269, 1834, 961, 2281, 2215, 330]. Orthogonality [1299]. Orthonormal [37, 2048]. Orthoscheme [1359]. O’Sullivan [1857, 1923]. Other [2058, 501, 1041, 1025, 1379]. OTRIMLE [2397]. Outbreak [553, 1590]. Outbreaks [1926]. outcome [2318, 2327]. outcome-dependent [2318, 2327]. Outcomes [1383, 2166, 1029, 2303]. Outlier [799, 1433, 858]. Outliers [694, 1213, 1793, 1274, 1431, 1580, 2350]. Outlook [2275]. Outputs [240]. Overdispersed [916]. Overdispersion [1266]. Overflow [114]. Overlap [787]. overs [1471]. Oversampled [1379]. Oversampling [1379]. Overtaking [277]. Overview [1429].

P [686, 1994]. P. [686, 947, 1090]. Package [2336, 2382, 2295, 2388, 2321]. Packages [2267]. pages [2276, 2275, 2267]. Pair [1058, 2352]. Paired [1260]. Pairs [1763, 2003, 1949, 744, 1299]. pairwise [2274]. pairwise-likelihood [2274]. Palindromic [1003]. panel [2344, 2310, 2415]. Paper [2058, 416, 446, 1947, 1512, 1932, 2107, 661, 477, 2108, 1795, 1946, 2069, 2137, 2097, 2047, 2096, 2166, 2125, 654, 1953, 2036, 2098, 1945, 1948, 1876, 1917, 1933, 1883]. Papers [2086, 2070, 2012, 1934, 2198, 2204, 2210, 2220, 1940, 2128, 1975, 1941, 2010, 2040, 2038, 1938, 1880, 2062, 2144, 2024, 2141, 1977, 1979, 2015, 1962, 1989, 1932, 2140, 2023, 2013, 2003, 2130, 2116, 1978, 1877, 2139, 1931, 1886, 2050, 1937, 1961, 1928, 1929, 2152, 1987, 2025, 1988, 1999, 1939, 2153, 1885, 1936, 1930, 1965, 1884, 2027, 2001, 2083, 2002, 2157, 2000, 1986, 2114, 2074, 2073, 2127, 2142, 2158, 2071, 2011, 2084, 2059, 2026, 2016, 2085, 2028, 1921, 2148, 2022, 2100, 2118, 2039, 2115, 2087, 1949, 2126, 2063, 1922, 2077, 1920, 1950, 1889, 2117, 2143, 2061, 2162, 1974, 1887, 2155]. Papers [2124, 1976, 1942, 2159, 2049, 1918, 1959, 1888, 2041, 2138, 2123, 2088, 2060, 2048, 2076, 2151, 1964, 1952, 1963, 2149, 1878, 2004, 2160, 2109, 2089, 1926, 1879, 2129, 2037, 1966, 1960, 1919, 2161, 1927, 1881, 2154, 1933, 2090, 2072, 2102, 2150, 2099, 2075, 2101, 1951, 2014, 2110, 1232, 1246, 1348, 1360, 1377, 1390, 1402, 1409, 1427, 1439, 1451, 1461, 1462, 1474, 1495, 1508, 1518, 1533, 1547, 1563, 1574, 1588, 1600, 1616, 1627, 1641, 1650, 1681, 1698, 1713, 1726, 1738, 1751, 1766, 1777, 1787, 1796, 1803, 1815, 1823, 1832, 1842, 1851, 1859, 1866, 1874, 1882]. papers [1890, 1898, 1907, 1916, 1925, 1935, 1944, 1958, 1973, 1984, 1998, 2008, 2021, 2034, 2046, 2056, 2068, 2082, 2095, 2104, 2113, 2122, 2136, 2147, 2156, 2165, 2176, 2187]. Paradox [719]. Paradoxical [615]. Parallel [1509, 1070, 276, 1915, 2292]. Parallelism [1411]. parallelogram [1636]. Parameter [1114, 1896, 1924, 1124, 357, 460, 1478, 930, 688, 1342, 1903, 788, 1054, 2127, 222, 1138, 100, 241, 560, 1004, 1047, 2049, 2200, 1122, 461, 2201, 777, 765, 1007, 1074, 1142, 1178, 2089, 1729, 2090, 517, 933, 2218, 2240, 2359, 2229, 2262]. Parameterisation* [1255]. Parameterization [1583]. Parameters [372, 348, 357, 715, 20, 1354, 512, 1111, 420, 1999, 586, 251, 324, 222, 367, 100, 241, 724, 54, 1679, 639, 336, 273, 774, 1068, 487, 60, 171, 198, 838, 1174, 1423, 1573]. Parametre [1122]. Parametric [1012, 1852, 1864, 1776, 431, 1261, 358, 1886, 1705, 2016, 47, 1646, 1485, 1173, 1216, 1597, 1007, 1031, 1879, 2203, 1239, 1728, 1883, 1638, 2363, 2395, 1556, 2341, 1558, 1570, 1559, 2412, 2271, 2272]. Parasite [2047]. Pareto [1395, 1810]. Paris [751]. Parity [1781]. Parsimonious [1771, 2392]. Partial [1709, 2002, 1759, 996, 1271, 1800, 1482, 734, 1561]. Partially [1840, 877, 1989, 984, 1870, 1887, 2151, 261, 1425, 2260, 2385]. Participant [1999]. Participation [1030]. particle [2347]. Particles [140]. Particular [1261, 1033, 787]. Partisan [1878]. Partition [43, 1514, 1503, 731, 857, 2396, 2386]. Partition-based [1514]. Partitioning [1398, 1425, 1527, 2051, 461]. party [2325]. Pass [1926]. passing [2373]. Past [903]. Pate [416]. Path [48]. Paths [562, 1630]. Pattern [1817, 591, 197, 1428, 1413, 2400, 2309]. Patterned [909]. Patterns [1477, 824, 2369]. Paul [1943, 1983]. Pavements [1617]. Payments [463]. PBIB [780]. PBIBD [2263]. Peace [1972]. Peak [611]. Pearson [1398, 1425, 1527, 301, 556, 529, 340, 154, 2237, 1858]. Pedigree [1224, 1172, 1136]. Pedigrees [1705]. Pelaez [737]. Penalised [2405, 1923, 2369]. Penalization [1987]. Penalized [1819, 1857, 1799]. Penalties [1469, 2037]. Penalty [1840, 2027]. pension [2360]. People [1909]. Performance [1684, 1704, 2045, 1018, 813, 1368, 1238, 2400, 2245]. Performances [1392]. Period [1272, 970]. Periodic [1771]. Periodogram [1594, 534]. Periods [1375, 1476]. Permanents [943]. Permutation [1505, 649, 751, 894, 2410, 1260, 871, 729, 945, 1838, 2257, 1414, 2006]. Perplexed [2093]. Personal [599, 1323]. Perspective [1536, 208]. Perturbation [1433]. perturbations [2414]. Pesarin [2006]. Peter [662].

Pharmaceutical [1993]. Phase [1884, 1744, 121, 239, 961, 2297]. Phenomena [1662]. Phenotype [1705]. phenotyping [2362]. Phylogenetics [2069]. Pia [1991]. Piece [274]. Piece-Wise [274]. Pilot [1977]. Pinto [1821, 1822]. Pioneer [2137]. Pitman [542, 612, 686, 1062, 1192, 1168, 1222, 1233, 1310, 1391, 1496, 1603, 1715, 1850, 1875, 2035, 2189, 2315, 1150, 1154, 1090, 1078]. Pixel [1151]. Place [291, 2406]. Planar [1112, 894, 252, 361]. Planning [436]. Plant [1989, 228, 1430]. Plant-Breeding [1989]. Plante [1822]. Plants [2036]. Play [1711, 1806]. Play-The-Winner [1711, 1806]. players [2376]. Playing [2199]. Plot [763, 1352, 2215]. Plots [642, 1490]. Plug [1769, 2159]. Plug-In [1769, 2159]. Plutonium [979]. Point [1477, 1671, 1259, 1864, 1350, 301, 951, 415, 1329, 1067, 1291, 2088, 1915, 1667, 1863, 2353, 2351, 1688, 2369, 2354]. Points [1261, 2025, 1680, 611, 1274]. Poisson [297, 404, 624, 550, 2027, 1481, 1680, 963, 2047, 859, 1573, 2334, 2039, 2270, 865, 1267, 1074, 2343, 81, 385, 530, 97, 1729, 1782, 1265, 1927, 2330, 2099, 933]. Poisson-exponential [1573]. Policies [237]. Policy [78, 185, 195, 211, 227, 821, 2273]. Pollen [1797]. Polling [1878]. Pollution [1893, 2149, 1968, 1967]. Polyhedral [1366]. Polynomial [1802, 2087, 274, 1772, 1688]. Polynomials [43, 2003, 215, 269, 330, 2048, 1196, 2236]. Polytope [2244]. Pooling [538, 1685]. Population [665, 267, 1940, 1913, 747, 1718, 1186, 326, 539, 2015, 447, 1171, 719, 1397, 1807, 878, 578, 370, 1121, 609, 1830, 815, 1853, 1486, 1198, 876, 774, 74, 1381, 1873, 1643, 1717, 1262, 1624, 1743, 1690, 2379]. Population-Size [1807]. Populations [1185, 537, 1097, 1124, 936, 85, 600, 574, 797, 390, 421, 1905, 1719, 1004, 1047, 1082, 1286, 461, 413, 74, 380, 594, 1145, 2366, 2336]. Positive [536, 1374, 215, 262, 846, 1039, 1500, 1950, 610]. Positively [1273]. Positivity [1091]. possible [2372]. Possibly [766, 370]. post [2295, 2366, 2405]. post-pretest [2405]. post-processing [2295]. post-shrinkage [2405]. post-stratified [2366]. Posterior [1092, 1080, 1484, 2289, 1824]. Postponable [165]. Potential [1718, 1452, 553]. Poverty [1936, 2339]. Power [1412, 1836, 147, 322, 314, 1734, 525, 1037, 9, 1888, 1678, 2192]. Power-Law [2192]. Powerful [1949, 1950, 2351, 2212, 2213]. Powers [56, 1662, 669]. PPGMMGA [2394]. PPS [253, 180, 343, 231]. Practical [1477, 2092, 1828, 1044, 2133]. Practice [1454, 2185, 2235, 2361, 2370, 1980]. Pre [782, 1238, 1251]. Pre-Test [782, 1238, 1251]. Precedence [1707]. Precedences [1860]. Precise [1014]. Predator [1736]. Predator-Prey [1736]. predict [2408]. Predicted [25]. predicting [1566]. Prediction [590, 946, 1186, 1835, 891, 1877, 1664, 636, 618, 2001, 1625, 2132, 462, 1067, 371, 1543, 2117, 667, 65, 55, 854, 1682, 1788, 2400, 1695, 2278]. Prediction-Focused [1835]. Predictive [2041, 2178]. Predictor [1735, 936, 2208]. Predictors [1731, 2124, 866, 1728]. Preface [968, 2249]. preferred [2325]. Preliminary [173, 1848, 538]. Prescribed [55]. Presence [2012, 1471, 1930, 2063, 1720, 1922, 458, 654, 1873, 1274, 1431, 621, 1865, 696, 933, 1685, 2328, 2323]. Presence/Absence [2012, 1720]. Present [903]. Presentation [1654]. Preserving [1296]. Presidential [599, 1417]. Press [2276, 2275, 2267]. Pretest [1840, 2310, 2405]. Prevalence [902, 1753]. Previous [896]. Prey [1736]. Price [147, 113, 373, 2365]. Prices [230]. Pricing [1723]. Primary [522, 1605]. Principal [485, 1319, 1581, 2162, 2409, 2404]. principle [1554]. Principles [649, 1759, 2173, 2356]. Prior [451, 1785, 1929, 1868, 2073, 2127, 2142, 2395, 2396, 2340]. Priori [381]. Priors [775, 1543]. Privacy [320]. private [2273]. Probabilistic [119]. Probabilities [643, 1839, 164, 308, 1749, 67, 72, 403, 539, 1470, 343, 2071, 1005, 1035, 1455, 37, 1853, 697, 752, 790, 1359, 1144, 2088, 2149, 402, 2381, 1571]. Probability [2070, 883, 528, 573, 1214, 576, 1046, 323, 642, 647, 1848, 310, 1000, 2132, 1764, 1892, 804, 1872, 1845, 384, 608, 1507, 2064, 1053, 1240, 2098, 653, 589, 1782, 437, 547, 2406]. Probability-Based [1845]. Probability-Linked [2098]. Probable [1900]. Probings [1752]. Probit [68, 2077]. Problem [312, 115, 537, 1112, 749, 287, 301, 315, 624, 314, 4, 958, 157, 1868, 1619, 223, 1073, 1147, 500, 1963, 1740, 774, 858, 1542, 965, 1372, 2214, 2363, 2357, 2280, 349]. Problems [946, 796, 902, 950, 1655, 957, 599, 567, 871, 413, 740, 650, 1849, 997, 547, 2367, 2417]. Procedure [342, 428, 792, 2141, 881, 1013, 1283, 1318, 1856, 1293, 1798, 641, 1914, 588, 1624]. Procedures [186, 898, 649, 1111, 1275, 582, 1911, 461, 2076, 1271, 24, 1332]. Process [1129, 440, 663, 1671, 513, 877, 768, 85, 1350, 158, 717, 242, 559, 1929, 966, 259, 360, 1659, 1481, 692, 625, 726, 664, 1712, 286, 2041, 393, 2060, 234, 336, 991, 1230, 561, 1642, 1661, 806, 126, 937, 2317, 2218, 2260, 2351, 2411, 2354]. Processes [1092, 1080, 416, 1086, 1677, 1671, 372, 691, 1864, 463, 714, 187, 1812, 476, 2139, 951, 624, 575, 1807, 550, 1411, 1754, 1137, 655, 706, 992, 882, 1128, 1298, 2039, 2020, 560, 1405, 1790, 1942, 1017, 1165, 1665, 2164, 1340, 1351, 1125, 1277, 594, 1301, 1302, 221, 1667, 1669, 441, 1635, 2353, 2327]. Processing [694, 1983, 2295]. Produced [49]. Product [57, 310, 1280, 610, 937]. Production [287, 252, 361]. Productivities [3]. Products [1662, 619]. Profession [599]. Professor [1589, 2035, 424, 545, 708]. Profile [1187, 1456]. Profiles [1414, 1447]. Program [1699, 1419]. Programming [22, 44, 73, 1612, 223, 393, 1970, 129]. Progress [435, 249]. Progressive [463]. progressively [2128]. Projected [2153]. Projection [2338, 936, 1366, 413, 753, 2394]. Promiscuous [1663]. Proof [1522, 1695]. propagation [2254]. Proper [1306]. Properly [1285]. Properties [281, 297, 1410, 403, 1276, 618, 2384, 1187, 2073, 2118, 825, 909, 728, 1212, 273, 296, 906, 782, 952, 1156, 2338, 2317]. Property [453, 1625, 1692, 2123, 1695]. Proportion [1170, 155, 1755, 2417]. Proportional [1019, 602, 343, 2377, 922]. Proportions [2103, 2023, 2013, 2169, 1227, 2260, 1686, 2287]. Proposal [1708]. Proposal- [1708]. Prospects [435, 599]. Protein [2036]. Providing [2199]. Proximity [1521]. Pseudo [2223, 1580]. Pseudo- [2223]. Pseudolikelihood [1477]. Pseudolikelihoods [2159]. pswor [428]. Public [821]. Publications [366, 942]. Publishing [647]. Purposes [436, 413]. pyramidal [2349].

Qian [1957]. Quadrat [420]. Quadratic [953, 954, 1387, 1356, 2181, 324, 238, 846, 955, 2087, 1173, 427, 1790, 131, 824, 926, 1196, 1262]. Quadrature [2191]. qualitative [2262]. Quality [1259, 2052, 1152, 1271, 1281]. Quanta [146]. Quantification [2084]. Quantile [576, 2062, 2144, 1228, 966, 1703, 594, 2246]. Quantiles [2050, 1304, 1020, 1399, 1919, 840, 2255, 1571]. quantitative [2262]. quantity [1571]. Quarter [247]. Quarterly [426]. Quasi [1802, 1137, 1237, 1950, 1464, 995, 150, 1266, 2228, 2272]. Quasi-Likelihood [1802, 1137, 995, 1266, 2272]. Quasi-Locally [1950]. Quasi-Score [1237]. Quasi-Stationary [150]. quasi-symmetry [2228]. Quasi-unit [1464]. Queensland [971]. Queries [2130]. Questioning [902]. Questions [786]. Queue [80, 122]. Queueing [1668]. Queues [114, 276, 1666, 81, 640]. Quine [286]. Quota [1281]. Quotient [1357].

R [1599, 2267, 2137, 2019, 2120, 1970, 1044, 2184, 2374, 2295, 2336, 2388, 1971, 2019, 1841, 1970, 1957, 2080, 1995, 2066, 2186, 2053, 2406, 2247]. R. [477]. Rachel [2308]. radius [2348]. Rain [328, 9]. Rain-Making [328]. Rainfall [1465, 205, 216, 362]. Raj [325]. Random [921, 186, 1860, 1475, 2170, 41, 1658, 954, 1349, 173, 1802, 2013, 2003, 181, 302, 562, 1877, 1356, 2050, 113, 316, 813, 2011, 1770, 1908, 1021, 619, 560, 591, 1790, 2124, 728, 1657, 1087, 230, 519, 1834, 693, 611, 1311, 407, 55, 1292, 150, 1040, 303, 1678, 1775, 1250, 1560, 2350, 1637, 2364, 1688, 2288, 2371, 2258]. Randomisation [1104, 1288]. Randomised [1282, 902, 1283]. Randomization [1444, 1744, 500, 660, 662, 661]. Randomized [1711, 615, 1806, 328, 1210, 2061, 729, 1798, 1622, 1897, 1206, 1784, 1332, 1638, 2292]. Randomly [880]. Range [1448, 1677, 134, 1886, 222, 89, 1669, 2238]. Rank [1203, 1732, 1458, 754, 1619, 1707, 1761, 2076, 1645, 2182, 1863, 884, 1821, 1822]. Rank-Sum [1707]. Ranked [1194, 1672, 61, 1526, 437, 1572, 1689]. ranked-set [1689]. ranking [2403]. Ranks [744]. Rao [492, 1305, 530]. Rare [553, 1211, 2090]. Rate [829, 663, 980, 1254, 1721, 1578, 1358, 1049, 299, 920, 1314, 1844, 359, 249, 347, 1729, 1782, 2414, 2252]. Rates [242, 811, 966, 1824, 726, 427, 1216, 1597, 2201, 401, 2222]. Ratio [743, 573, 85, 1979, 57, 938, 934, 1058, 1009, 912, 1106, 939, 1053, 2160, 296, 469, 857, 2222, 1686, 2360]. Ratio-To-Size [939]. Ratios [1871, 206, 876, 1196]. Raton [2276, 2275, 2267]. Ravishanker [2289]. Raw [823, 2301]. Ray [2289]. Rayleigh [259]. RC [668]. Re [1821, 1967]. read [2206]. read-depth [2206]. readings [2239]. real [2324]. real-time [2324]. Realized [1058]. Rearrangements [1834]. Reasoning [2000]. Recapture [747, 1307, 1820, 1986, 1660, 1432, 2174, 1742, 1143, 1624, 1556, 1555]. Received [219, 366, 1061]. Recidivist [1169]. Recipient [2308].

Recognition [1817, 2309]. Recommendation [899]. Recommendations [2166]. Reconsideration [357]. Reconsidered [1314, 48]. Reconstruction [591]. Record [1763, 1710, 1228, 1464, 464, 506, 1881, 2211]. Record-Breaking [1228]. Recording [49]. Records [1327]. Recovering [1541]. Rectangular [668, 575]. Rectangular-RC [668]. Recurrence [1862, 2048]. Recurrent [1945]. Recursive [439, 512, 2025, 802, 2077, 725, 1066, 1139, 2051]. Recursively [1173]. reduce [2416]. Reduced [780, 1203, 694, 1350, 313, 2085, 381, 1645]. Reduced-Rank [1203]. Reduction [1512, 721, 1353, 919, 2393, 2386, 2208]. Reductions [1524]. Referee [614, 707, 1180, 1322, 1347, 1407, 1463, 1497, 1551, 1615]. Referees [913, 885]. Reference [90, 322, 1033, 787, 522, 156]. References [2267]. reflected [2305]. Reflections [1670]. Regenerative [1662]. regime [2294]. Region [562]. Regional [753]. Regions [695, 2013, 1664, 54, 518, 776, 1239]. Registers [795]. Regression [2086, 1388, 675, 921, 1448, 253, 1735, 1763, 1410, 1019, 2339, 1261, 1829, 809, 504, 570, 681, 796, 1112, 574, 2144, 1748, 1161, 1105, 1339, 620, 1802, 894, 891, 451, 1133, 977, 1043, 1794, 181, 791, 1828, 1354, 1383, 1490, 1999, 1906, 586, 1055, 863, 1137, 2027, 1820, 1140, 1176, 1764, 1892, 2073, 2127, 2142, 149, 139, 188, 1355, 1831, 690, 23, 47, 1506, 2118, 1647, 2172, 2087, 1468, 2005, 2096, 619, 54, 524, 1543, 2166, 1911, 257, 2143, 1066, 889, 1597, 1779, 1895, 871, 750, 738, 535, 637, 1789, 1252, 650, 2098, 730, 1109, 469, 1093, 1297, 1873, 906]. Regression [1274, 782, 1380, 25, 854, 1343, 1580, 1139, 231, 2182, 1265, 1408, 1760, 1251, 1857, 1876, 1923, 1177, 1742, 1085, 1324, 997, 1317, 523, 1799, 1772, 2101, 2382, 1634, 2316, 1567, 1684, 2251, 1421, 1566, 2224, 2334, 1418, 1554, 2341, 2246, 1558, 2232, 1472, 1513, 2378, 2342, 2303, 2387, 2264, 1564, 2328, 2292, 2321, 1687]. Regression-Type [1388, 231, 253, 1472]. Regressions [1880, 1449, 36, 1792]. Regressive [1383, 1178]. Regressors [23, 1791]. Regular [829, 1239, 838]. Regularisation [2200]. Regularised [2161]. Regularly [431]. Reid [1795]. Rejoinder [1472, 1822, 850]. Related [1065, 394, 2103, 994, 902, 1674, 238, 1106, 530, 965]. Relation [312, 418, 1862]. Relationaships [799]. Relations [14, 1816, 2201, 419]. Relationship [1797, 1063, 1237, 501, 522, 438]. Relationships [1705, 935, 1325]. Relative [29, 2166, 1441, 167]. Relaxing [569, 2061]. Release [1780, 146, 952]. Reliability [907, 1827, 141]. Remark [986, 656]. Remarks [691, 23, 654, 1330, 126]. REML [2240, 1212, 2286]. Remotely [1151]. Renewal [1129, 372, 187, 989, 991, 1669]. Renewals [244]. Rental [157]. rep [2102]. Repair [1938]. Repeated [1711, 1375, 1133, 994, 2177, 526, 925, 1830, 1895, 1071, 2203, 854, 2090, 2323]. Repertoire [291]. Repetitions [1530]. Replaced [462]. Replacement [253, 164, 308, 67, 180, 233, 343, 231]. Replacements [61]. Replicated [1989, 1117]. Replicates [836]. Replication [450, 2150]. Reply [1968]. Report [144, 169, 194, 226, 256]. reported [2343]. Reporting [1782, 1605]. Reporting-Probability [1782]. Representation [1086]. Representations [1201, 852]. Reproduction [395]. reproductive [1569]. Resampling [1081, 1242, 1275]. Rescaled [1285]. Rescaling [1667]. Research [2106, 457, 1699, 502, 15, 2018]. Reserve [1441]. Reserves [557, 87]. Residual [1490, 1915, 1056, 2362]. Residuals [1584, 1084, 1066, 644, 1530, 1139, 523, 2265]. Residue [780]. Resistance [1229]. Resolution [987, 563]. Resolvable [1191, 1415, 1205, 817, 1724]. Resource [3]. Respect [348, 461]. Response [1934, 615, 649, 2079, 1307, 1282, 1591, 1584, 1400, 902, 1000, 1986, 1063, 890, 1283, 1287, 2087, 1432, 2208, 826, 1798, 1873, 1622, 1281, 1948, 1897, 1332, 1555, 2292, 662]. Responses [1978, 1806, 1541, 2202, 1297, 1380, 2359]. Restricted [1015, 251, 1856, 1368, 1212, 2089, 1238, 1557]. Restriction [2061, 1251]. Restrictions [674, 381]. Result [717, 1359, 1695]. Results [85, 80, 1447, 71, 149, 139, 882, 896, 918, 868, 1351, 530, 965, 561, 1428, 1413, 2261]. Reticulation [2036]. Retirement [1404]. Return [1476]. Returns [1758]. Reversible [1918, 2295]. Review [1971, 1737, 2078, 2227, 2080, 1980, 1997, 2146, 2133, 2092, 2235, 2007, 2045, 2006, 2103, 2079, 2276, 2275, 2094, 1046, 2018, 2267, 1995, 2017, 1972, 2131, 1227, 2163, 2065, 2052, 1981, 2029, 2031, 988, 2066, 2112, 2093, 2132, 2033, 1954, 2051, 1955, 1956, 2186, 2005, 2020, 2145, 2313, 2053, 1993, 2135, 2032, 2019, 2183, 2121, 2185, 2120, 2064, 2299, 2111, 1982, 2175, 2164, 2044, 1841, 2054, 2119, 2209, 2134, 1045, 2081, 1970, 2030, 1044, 1994, 1991, 2174, 2219, 2091, 1943, 2173, 2226, 2247, 2233, 1983, 2043, 2055, 2184, 1969, 2042, 2306, 2067, 753, 1992, 1957, 1996, 2406, 1696, 1697]. Reviews [45, 51, 59, 64, 70, 77, 84, 94, 99, 104, 111, 116, 123, 136, 145, 153, 161, 170, 175, 183, 190, 201, 209, 217, 224, 235, 246, 254, 264, 272, 280, 290, 305, 319, 332, 352, 365, 378, 387, 398, 412, 423, 433, 444, 455, 466, 478, 497, 498, 508, 520, 531, 543, 565, 579, 597, 613, 631, 645, 659, 676, 687, 704, 746, 759, 770, 784, 794, 807, 819, 833, 848, 860, 900, 873, 886, 914, 927, 940, 998, 1010, 1022, 1059, 1077, 1088, 1101, 1115, 1130]. Reviews [1149, 1167, 1179, 1193, 1207, 1221, 1231, 1245, 1257, 1268, 1278, 1289, 1300, 1309, 1321, 1334, 1346, 1362, 1376, 1389, 1401, 1403, 1416, 1426, 1438, 1450, 1460, 1473, 1483, 1493, 1510, 1517, 1532, 1546, 1562, 1575, 1587, 1601, 1614, 1629, 1640, 1725, 1814, 2298, 1813, 2234, 339]. Revisited [1706, 325]. reweighted [2352]. Rhinoconjunctivitis [1797]. Ribosomal [2036]. Rice [1670]. Richard [1995]. Richards [2408]. Richness [1778]. Ricker [1780]. Ridge [1235, 1109, 2089, 782]. Rigby [2406]. Right [1586]. Right-Censored [1586]. Rise [2069]. Risk [651, 1367, 773, 2197, 2166, 1166, 2041, 425, 2067, 2409, 2395, 2365, 2379]. Risks [445, 1184, 1135, 2053, 919, 2131]. rjmcmc [2295]. RMIT [1520]. Road [2058, 277, 1617, 2324]. Robinson [1970]. Robust [1199, 1236, 298, 1776, 1365, 2094, 718, 2116, 1354, 843, 1886, 1817, 1055, 863, 1099, 1172, 1235, 1705, 47, 1127, 1249, 1303, 1100, 1136, 2327, 1394, 1109, 2403, 1274, 1756, 1919, 2397, 2262, 1991]. Robustness [678, 1459, 1990, 2152, 636, 1256, 2263]. ROC [2045, 1468, 1558, 1951].

Rochester [1520]. Role [663, 956, 1752, 1069, 2199, 15, 903, 660, 661]. Root [1123, 1055, 293, 1865]. Roots [392, 206, 738, 303]. Rosén [737]. Ross [1666]. Ross-Type [1666]. Rotatability [1333, 1287]. Rotational [732]. Rothmann [2007]. Rough [456]. Round [548]. Rounded [321]. rounding [2359]. Routes [1883]. Roux [1997]. Row [1611, 908, 1191, 1205, 430, 1256, 1206, 836, 1724]. Row-And-Column [430]. Row-Column [908, 1191, 1205, 1256, 1206, 1724]. Rows [430]. Royal [2]. rps11 [2036]. Rubin [1305, 530]. Rudolph [2276]. Rugby [1953, 1418]. Ruin [1839]. Rule [883, 1185, 158, 1806, 1093]. Rules [743, 2367]. Run [1833, 1220]. Runs [41].

S [2007, 1472, 1957, 2263]. S&P200 [1723]. S. [1192]. Saddle [301]. Saddlepoint [1399, 1177]. Sales [174]. Salmaso [2006]. Same [1906, 468]. Sample [1940, 1357, 480, 115, 353, 2024, 853, 1708, 569, 749, 2116, 642, 1111, 1227, 1988, 957, 1290, 655, 706, 1081, 1313, 140, 266, 370, 795, 1039, 762, 1249, 814, 1399, 427, 1073, 1579, 89, 1087, 488, 2054, 1963, 1798, 774, 1016, 1040, 798, 700, 470, 594, 239, 782, 171, 616, 229, 1312, 588, 1217, 1296, 2250, 1634, 2211, 2320, 2255, 1630, 1689, 2336, 2354, 2257]. Sampled [539]. sampler [2384]. Samplers [1792, 2244]. Samples [680, 743, 822, 134, 125, 67, 326, 843, 371, 554, 198, 583, 50, 1755, 2350, 2318, 1685, 1689, 2366]. Sampling [253, 593, 643, 911, 164, 186, 308, 82, 1259, 1749, 72, 428, 1752, 823, 1442, 1105, 403, 180, 173, 233, 1470, 343, 1095, 1479, 881, 559, 1837, 1385, 420, 1930, 1884, 578, 1526, 155, 172, 1612, 804, 609, 1432, 560, 591, 2125, 1198, 1790, 1368, 876, 939, 1122, 1373, 1017, 197, 344, 1007, 1443, 1522, 2149, 592, 1230, 167, 1068, 345, 121, 220, 309, 402, 1001, 1502, 239, 1782, 1281, 7, 25, 231, 1262, 1238, 437, 1624, 1572, 2289, 2223, 1693, 1557, 2343, 2323, 1553]. Samuel [1991]. Sarmanov [2326]. Satisfactory [1226]. Scale [1182, 2024, 460, 1854, 1708, 749, 288, 1111, 812, 324, 793, 762, 1485, 581, 1004, 1047, 892, 1271, 171, 997, 1156, 354, 1475, 2406, 2399, 2417]. Scales [766, 1364, 1464]. Scaling [1270, 1269, 1816]. Scheffé [314, 779]. Scheme [113, 1211, 1693]. Schemes [403, 1470, 815, 479]. School [1029, 2361]. School-Leavers [1029]. Schur [1201]. Schur-Concave [1201]. Science [2018, 1406, 502, 2356, 2042]. Sciences [633]. Scientific [1970]. Score [1342, 1237, 2320, 2232, 2404]. score-based [2232]. Scores [1506, 2416]. Scotch [1706]. Scrambled [1297, 1380]. Screen [1605]. Screening [1700, 558, 2338, 2251]. sealed [2365]. sealed-bid [2365]. Search [1094, 1585, 1118, 1206, 2129]. Searching [479]. Searle [2137]. Seasonal [27, 17, 1981, 1291, 577, 755, 2293]. Seasonality [2114, 1476]. Second [281, 1350, 1421, 333, 1679, 1786, 1085, 1324]. Second-Order [1679, 1421]. Secondary [667]. Secretaries [1226]. Secretary [1372]. Sectional [1754, 346, 2415, 2379, 2271]. sections [2211, 2347]. seed [1683]. Seeding [205, 216, 362]. Seemingly [1792]. Segmentation [1918]. segmented [2232, 1564]. seismicity [2270]. Seismological [2114]. Selected [994, 1526]. Selecting [1097, 1538, 326, 390, 1397, 1018, 1004, 1144, 1226, 421, 1047].

Selection [1941, 537, 713, 1124, 1214, 1225, 1676, 1835, 1785, 343, 1108, 1937, 1111, 1939, 529, 155, 1764, 1892, 2059, 1436, 875, 1887, 2200, 1964, 2195, 2149, 167, 1274, 1643, 2037, 1819, 1312, 568, 2205, 1566, 2380, 2415, 2369, 2403, 2286, 2311, 2330]. Selective [437]. Selectors [1977]. Self [1605]. Self-reporting [1605]. Selling [174]. Semi [1852, 158, 1794, 882, 2115, 1558, 1570, 1079, 1216, 2271, 2203, 1559, 2412, 2272]. Semi-Markov [158, 882, 1079]. Semi-Mechanistic [1794]. Semi-nonparametric [2115]. Semi-Parametric [1852, 1216, 2203, 1558, 1570, 2271, 1559, 2412, 2272]. Semiparametric [1748, 1161, 1730, 1871, 1928, 1929, 1820, 1986, 2047, 1831, 1135, 2077, 2143, 2049, 2190, 1598, 1857, 1876, 1923, 1742, 2278, 2378]. Sen [758]. Senate [1104, 559]. Seneta [1391, 286]. Sensed [1151]. sensitive [2292]. Sensitivities [1605]. Sensitivity [1721, 320, 1900]. September [491]. Sequence [682, 1294]. Sequence-Comparison [682]. Sequences [41, 215, 810, 262, 866, 1834, 2255]. sequencing [2206]. Sequential [829, 528, 573, 691, 2296, 480, 207, 243, 71, 985, 1939, 815, 138, 189, 654, 527, 461, 639, 1373, 910, 1053, 1522, 369, 621, 479, 1624, 1008, 1638, 1561, 2294]. Serial [1113, 328, 356, 87]. Series [1971, 96, 1050, 62, 739, 1846, 891, 207, 595, 2302, 1452, 1490, 1886, 905, 2025, 671, 966, 1352, 526, 1316, 49, 1646, 1647, 1594, 1771, 533, 967, 1657, 347, 1825, 258, 55, 858, 1915, 926, 1580, 128, 1863, 818, 2218, 2239, 2211, 2333, 2342, 2271, 2129, 2207]. Server [122]. Service [80, 81]. Services [614, 707, 1180, 1322, 1347, 1407, 1463, 1497, 1551, 1615]. Session [203]. Set [1792, 1526, 1446, 919, 585, 461, 430, 437, 1572, 1689]. Set-Valued [919]. Sets [1658, 792, 1769, 36, 604, 430, 2350, 2304]. Setting [2064]. settings [2241]. Settler [156]. Several [308, 236, 326, 61, 1648, 1318, 1286, 1755]. Severity [553]. Sex [468]. Sexual [1118]. Shannon [1974]. shape [2347, 2406]. Share [230]. Shared [1778, 1598, 1250, 2359]. Sharp [1710]. Sharpening [1578]. Shayle [2137]. Sheather [1519]. Shelf [2026]. Shelf-Life [2026]. Shift [1732, 1241, 2294]. Shift-Invariant [1241]. shifted [2388]. Shock [1478]. Short [35, 38, 1765, 2139, 1886, 1525, 47, 667, 1351, 2333, 2020]. Short- [1351]. Short-Cut [47]. short-memory [2333, 2020]. Short-Range-Dependent [1886]. Shot [1680]. Should [1214, 424, 2155]. Shrinkage [1840, 891, 2027, 2097, 2405]. Siblings [1028]. Sided [1273, 1849, 884, 2335]. Sign [992]. Signal [2180, 1058]. Signals [1058, 2238]. Signed [1619, 744]. Signed-Rank [1619]. Significance [716, 337, 986, 855, 555, 1703, 79, 1204]. Significant [811]. Signs [1414, 1447]. Silhouettes [969]. Silverstone [1536]. Similar [1037]. Similarity [2012, 1885]. Simon [1841, 1520, 1519]. Simple [186, 1671, 298, 936, 428, 570, 476, 1978, 955, 825, 604, 189, 910, 296, 65, 1096, 641, 146, 2317, 1689]. simple-tree [1689]. Simplified [186, 288, 2074]. Simplifying [755, 1408, 2216]. Simulated [1191, 1701]. Simulation [1962, 1435, 1993, 669, 1970, 2239]. Simulator [2140]. Simultaneous [638, 1544, 268, 313, 859, 130, 2087, 623, 1386]. Single [712, 326, 2026, 1921, 1706, 122, 1728, 1332, 2302, 2380, 2129]. Single-Index [1921, 1728, 2302, 2380]. Single-Malt [1706]. Single-Server [122]. Singly [1527]. Singular [841, 1478, 1002, 2059, 2180]. Sinha [212]. Sir [191]. SIROMATH [850, 849]. Situation [157]. Situations [283]. Size [267, 1839, 747, 1807, 1227, 1988, 1830, 814, 2126, 1853, 939, 522, 588, 1624, 1743, 2250, 2230, 2354, 2103]. Sizes [66, 343, 2169, 1415, 120, 1590, 2135, 1040, 229]. Skellam [2261]. Skew [2001]. Skewed [1273]. Skewness [1621, 628, 935, 1330, 50]. Skill [1247]. Skip [1005]. Skip-Free [1005]. sleep [2301]. Sliced [1789]. Slippage [115, 1542]. Slopes [809, 1242]. Small [853, 1708, 2116, 1936, 1290, 756, 2155, 1075, 2138, 2004, 1016, 2412, 1071, 1960, 1919, 2304, 2271, 1571]. Small-Area [1936, 2004, 1919, 2271]. Smallest [229]. Smirnov [1393, 813]. Smooth [1489, 1467, 56, 109, 1285, 2115, 1267, 2387, 2168, 1904, 1565]. Smoothed [1977, 1054, 2285, 2232]. Smoother [1264]. Smoothers [1647]. Smoothing [1901, 993, 1817, 864, 788, 1722, 1258, 1303, 2049, 1819, 2193, 949, 1215, 2246, 2043]. Snapper [1441]. Soares [1821]. Social [2018, 457, 1753, 503, 502]. Society [118, 144, 169, 194, 226, 256, 491, 445, 490, 974, 424, 2, 76]. Socioeconomic [1027]. Software [1224, 2006]. Soil [979]. Sojourn [991]. Solution [314]. solutions [2214]. Some [668, 2012, 96, 916, 2214, 540, 1410, 85, 615, 1364, 403, 187, 80, 243, 282, 595, 1276, 1400, 1253, 434, 1356, 1358, 905, 1014, 310, 452, 586, 1275, 266, 1187, 149, 151, 139, 882, 23, 904, 825, 427, 581, 943, 876, 654, 728, 703, 248, 413, 1053, 897, 401, 729, 407, 837, 376, 1305, 296, 1330, 345, 1096, 385, 530, 561, 121, 1196, 129, 1302, 1552, 368, 850, 849, 221, 303, 733, 1085, 1324, 126, 237, 2317, 638, 1561, 885]. Song [1957]. source [2404]. Sources [1025, 1559]. South [974, 2361]. Space [114, 714, 1716, 1452, 1701, 2305]. Spaces [1834, 2110]. Spacings [1048, 583, 616]. Sparse [864, 2332, 2162, 1396, 1645, 2378, 2399]. Spatial [1477, 1867, 1350, 2325, 2001, 2083, 767, 1164, 1503, 1163, 1258, 2028, 2138, 1963, 2178, 1994, 2072, 2353, 2395, 2279, 2354, 2225, 2355]. Spatially [1202, 1469]. Spatially-adaptive [1469]. Spatially-Correlated [1202]. Spatio [2158, 2362]. Spatio-Temporal [2158, 2362]. Special [1512, 90, 322, 490, 723, 424, 790, 522, 156]. Species [1778, 2071, 1720, 519, 1656]. Specific [1746, 2124]. Specification [781, 1490]. Specified [1238, 1251, 1788]. Spectra [213]. Spectral [1086, 1940, 1658, 356, 179]. Spectrum [2059, 2180, 179]. Speed [1603]. Sphere [822, 764]. Spherical [789, 1263]. Sphericity [1159, 484, 934]. Spline [993, 1264, 1469, 1819, 1772]. Splines [2058, 1722, 1857, 1923, 2043, 1805, 1799, 949, 1215, 2246, 2385]. Split [763, 2215]. Split-Plot [763, 2215]. Splitting [1704, 1567, 1684]. Sports [1674]. Spread [809]. Spreadsheet [1995]. Spring [841, 1002]. Spurious [456]. Spying [2396]. Square [43, 1123, 196, 1613, 311, 382, 1528, 2381]. Squared [1398, 953, 954, 340, 955, 1425, 1527]. Squares [675, 268, 299, 1183, 2162, 1523, 1457, 869, 487]. SSAI [1654]. Stabilisation [2090]. Stability [1066, 2403]. Stable [536, 1912, 572, 813]. Stage [593, 911, 537, 1544, 688, 2108, 1385, 1283, 268, 582, 1757, 876, 939, 2064, 2201, 1798, 1443, 220, 588, 1917, 2125, 1332]. Stage-Wise [2201]. stages [2301]. Stan [2236]. Standard [635, 1936, 867, 700, 869, 2182, 13]. Standardized [996, 1561]. Stanley [191, 1982]. Starship [1437]. State [714, 1962, 1452, 2022, 1701, 2167]. State-Space [714]. States [1032, 32]. Stationary [1086, 1671, 1658, 1350, 810, 1702, 1942, 150, 2317, 2164]. Stations [971]. Statistic [1833, 1398, 1527, 699, 1369, 779, 616, 801, 229, 1425, 1424, 2257]. Statistical [96, 203, 946, 956, 1132, 2, 1854, 491, 2267, 486, 1893, 950, 2139, 1617, 855, 957, 958, 599, 633, 2069, 1503, 2206, 959, 2093, 2033, 971, 14, 1223, 1719, 1956, 2236, 2032, 426, 76, 960, 2155, 2088, 2119, 1661, 1618, 425, 208, 733, 2042, 2167, 961, 1996, 1247, 2221, 2356, 1968, 2406, 1967, 118, 144, 169, 194, 226, 256, 491, 445, 490, 974, 424, 2045, 2121, 2044, 2134, 1983, 1969]. Statistician [291, 970]. Statisticians [903, 1454]. Statistics [1065, 1050, 2078, 1162, 1467, 1880, 400, 751, 1478, 1781, 2013, 2289, 1452, 549, 1995, 434, 1111, 682, 1999, 958, 970, 832, 503, 1454, 2085, 101, 333, 685, 677, 177, 1326, 2135, 356, 669, 1048, 1504, 1899, 535, 637, 156, 401, 729, 981, 2199, 803, 1351, 1016, 969, 514, 830, 1343, 307, 583, 850, 849, 1863, 1317, 1678, 1992, 1957, 1217, 547, 2128, 340, 2223, 1552, 185, 195, 211, 227, 1280, 424, 1417, 1, 2299, 1044, 1994, 2275, 2094, 968, 2091, 2184]. Steele [1947, 1946, 1948]. Stefan [1956]. Stein [748, 1368, 1093]. Step [898]. Stephane [1991]. Stepwise [2141, 1326, 1914, 568]. Stereological [2347]. Stereology [1979]. Stigmatized [1381]. Stimulation [9]. Stirling [281, 1992]. Stochastic [1092, 1080, 665, 1677, 702, 778, 85, 1721, 1846, 486, 242, 22, 1648, 962, 992, 963, 1582, 1353, 1908, 393, 2164, 674, 1234, 854, 146, 237, 2414]. Stochastically [1120]. Stock [1758, 373]. Stopping [1006, 1373]. Storage [287, 240]. Straight [1449]. Strata [960]. Strategies [876, 1206, 2405, 2185]. Strategy [760]. Stratification [173, 1837, 922, 74, 402, 469, 2336]. Stratified [1749, 2108, 1612, 1308, 1198, 2366]. Stratifying [1837]. stratifyR [2336]. Streams [1259]. Strength [563]. Stress [952]. stretch [2384]. Strong [1371, 1834, 2154]. Structural [799, 288, 47, 448, 1982]. Structure [267, 353, 556, 2179, 2011, 964, 2160, 965, 1335, 857, 961, 1637]. Structures [1886, 1296, 2225, 2370]. Stuart [1378]. Student [2120, 667, 2392, 1037, 283]. Studentizing [1188]. Students [1392]. Studies [2052, 1099, 1487, 276, 795, 14, 1660, 1579, 2149, 1742, 1134, 1838, 2379, 1993, 2111]. Study [463, 486, 243, 1794, 1034, 2158, 2087, 581, 669, 1025, 2395, 2237, 2270, 2361, 2348, 457]. Sub [1039]. Sub-Sample [1039]. Subexponential [1668]. Subject [1019, 1731, 2197, 2124, 240]. Subject-Specific [2124]. Subjects [2170]. Subsample [812]. Subsampling [1912]. Subsequent [15]. Subset [1785, 1690]. Subsets [1395]. subspace [2394]. subtractive [2403]. Subunit [939]. Success [528]. Successive [308, 197, 345, 220]. Sufficiency [590, 337, 720, 1524]. Sufficient [2393, 1237, 814, 563]. Sugar [971]. Sugnet [1997]. Sum [297, 2023, 1021, 1707, 148]. Summary [1999]. Summation [454, 1241].

Sums [181, 1770, 1020, 1148, 996, 840, 2265, 2253, 1561]. Super [1953]. Superadditive [2154]. Supercritical [717]. Superimposed [500]. Superpopulation [1186]. Superposition [242]. Superpositions [1671]. Supplementary [541]. Supply [1791, 527]. Support [113, 1757, 1878, 2178]. Supports [212]. Supreme [1927]. Supremum [475]. Surely [2088]. Surface [1400, 562, 1287, 826, 1951]. Surrogate [1535]. Surveillance [1211]. Survey [1475, 1364, 1591, 902, 976, 786, 1978, 435, 785, 3, 666, 2148, 490, 2096, 1146, 344, 592, 1960, 975, 354, 734, 2250, 1033, 1031, 1025, 354]. Surveys [353, 2024, 1282, 2015, 436, 890, 526, 648, 795, 821, 2028, 1308, 2117, 1798, 470, 1622, 239, 610, 1281, 854, 320, 2090, 1634, 2319, 2292, 1553, 2054]. Survival [1091, 2062, 1586, 2131, 1201, 2063, 1499, 1284, 581, 2190]. Susceptibles [1648, 1098]. sustainable [2360]. Swings [27]. Switching [1678, 2323]. Sydney [1968, 1967, 1716, 457, 1893, 373]. Symmetric [1676, 572, 1070, 316, 867, 616, 2208]. Symmetrical [1910, 120, 1142]. Symmetrically [611]. Symmetries [732]. Symmetry [1833, 1275, 133, 1613, 1733, 380, 1003, 2381, 2228]. Symptoms [1797]. System [778, 181, 346, 165, 65, 127, 667]. Systematic [72]. Systems [1938, 435, 1736, 1064, 248, 519, 1740, 1915, 1881, 237].

T [1168, 2020]. Table [1398, 2160, 1425]. Tables [1626, 1527, 931, 1529, 42, 196, 285, 1754, 1338, 684, 5, 756, 1433, 904, 1613, 1513, 493, 605, 1396, 1071, 1528, 32, 2381, 2228]. Tagging [1382]. Tail [879, 1871, 2253, 1773, 2385]. Tailed [1839, 1846, 1679, 1865]. Talk [1050]. task [2323]. task-switching [2323]. Tau [931, 1961]. tax [2273]. Taylor [207]. Teaching [1050, 1044, 1618]. Team [1234]. Technical [1404]. Technique [458]. Techniques [688, 890, 1258, 724, 448]. Technologies [1943]. Technology [1466]. Teenage [1030]. Telephone [648, 1281]. Temperature [2084, 2239]. Temporal [2158, 1258, 2410, 2362]. Tennis [745]. Term [1519, 667, 851, 2400, 2114]. Terminal [146]. Terminating [562]. Ternary [847]. Terry [1603]. Tertiary [1027, 248]. tessellations [2348]. Test [1732, 1467, 573, 66, 739, 68, 1042, 480, 41, 689, 1912, 1989, 894, 569, 1108, 572, 749, 718, 1801, 2116, 1393, 811, 812, 1218, 1411, 484, 1806, 109, 1885, 1000, 316, 992, 617, 2171, 1037, 1506, 120, 1619, 756, 1647, 1326, 1700, 604, 1260, 1950, 9, 1744, 1457, 538, 910, 1144, 2088, 729, 481, 670, 369, 1040, 641, 800, 571, 782, 449, 830, 1343, 857, 1238, 1251, 884, 517, 1250, 2288, 2320, 2256, 2217, 2257, 2280]. Tester [19]. Testing [1015, 1159, 1833, 799, 573, 342, 986, 1758, 2169, 1052, 1397, 1848, 1385, 934, 2288, 1685, 222, 367, 690, 1039, 1646, 371, 2063, 582, 228, 1066, 458, 585, 697, 1516, 179, 213, 1146, 1733, 981, 2195, 1952, 1031, 616, 1865, 584, 1145, 1204, 1692, 1686, 2287, 1956]. Tests [2070, 1505, 1467, 1091, 822, 716, 754, 2006, 1412, 414, 1586, 1229, 1836, 789, 791, 1157, 1514, 358, 764, 923, 985, 1014, 56, 1285, 1481, 813, 692, 735, 2000, 1242, 266, 1581, 370, 1039, 762, 880, 912, 1106, 130, 205, 216, 362, 1127, 1249, 1746, 1326, 1949, 2126, 2096, 1920, 1406, 138, 1073, 1613, 189, 1779, 1291, 1707, 871, 1041, 1243, 1053, 650, 1267, 555, 729, 837, 867, 380, 1861, 438, 2168, 368, 79, 1849, 696, 1838, 1904, 1143, 845, 1632, 2363, 2364, 2265, 2230, 2404, 2284, 2243, 2387, 2213]. Textbooks [1992]. th [1710]. Theil [451, 524]. Their [348, 769, 1018, 2085, 697, 33, 273]. them [2107]. Theodorescu [654]. Theorem [215, 1358, 989, 44, 619, 212, 286, 1017, 234, 514, 748]. Theorems [2013, 2085, 996, 221, 2398]. Theory [2086, 1423, 1692, 1388, 1572, 2012, 1489, 1633, 2214, 1448, 1677, 1934, 1626, 1735, 1709, 1635, 1940, 1763, 2128, 1732, 1975, 1410, 1638, 1467, 1941, 1398, 1425, 1527, 2010, 2040, 1634, 1749, 2038, 1515, 1718, 1560, 1711, 948, 1458, 1621, 2170, 1449, 1538, 2218, 1471, 1567, 1684, 1704, 638, 1442, 1565, 2006, 2079, 1412, 1765, 2062, 2144, 1748, 2141, 1721, 1676, 1977, 1387, 1611, 1540, 1544, 1730, 1758, 1586, 2015, 1802, 853, 1529, 1962, 1578, 1459, 1708, 2245, 1632, 1470, 1512, 1421, 1675, 2222, 1932, 1610, 1710, 445, 485, 1637, 1584, 2013, 2003, 2240, 1539, 1414, 1447, 2130, 1400, 2169]. Theory [1801, 719, 1750, 810, 1393, 1514, 1383, 1723, 1490, 1736, 1762, 1397, 1731, 1931, 2050, 2181, 1961, 1566, 1585, 1928, 1929, 1385, 2152, 1987, 1702, 1491, 1988, 1411, 938, 1420, 1939, 2153, 2179, 1424, 1648, 1631, 2244, 1930, 2231, 1734, 1754, 1644, 1686, 1526, 1625, 1690, 2002, 2224, 1487, 1556, 1705, 1688, 1419, 1454, 1555, 1986, 1722, 1680, 1674, 1541, 1415, 1581, 2074, 1435, 1456, 1525, 1524, 1691, 1764, 2033, 2073, 2127, 2142, 2197, 149, 2171, 2011, 2228, 1685, 1612, 2059, 2180, 2216, 2016, 2085, 1437, 1436, 1582, 1455, 1573, 1384, 1921, 1583, 1646, 1619, 1719, 1433, 1485, 1554, 2100]. Theory [2118, 1706, 1647, 2039, 2172, 2202, 1746, 2246, 1446, 1561, 2087, 1432, 1468, 1558, 1620, 1949, 2126, 2230, 1570, 2223, 1693, 1399, 2063, 1720, 1922, 2229, 223, 1444, 2077, 1543, 1920, 1950, 1613, 2143, 2061, 2162, 2185, 1974, 1759, 1486, 1516, 2232, 1757, 2155, 1712, 1579, 1976, 1513, 1707, 1531, 1744, 1609, 1523, 1608, 1457, 1942, 1623, 1394, 1689, 2159, 2049, 2200, 1918, 1679, 1959, 1041, 2201, 1396, 1761, 2041, 1741, 197, 2208, 2088, 2060, 1559, 2164, 1395, 2196, 1733, 2048, 2076, 2151, 2243, 1488, 693, 1964, 2195, 1952, 1695, 1963, 2217, 1469, 1740, 1747, 1798, 1443, 1607, 1522, 1557, 1703, 2004]. Theory [2160, 2109, 345, 1542, 1636, 1530, 1381, 1422, 2089, 1434, 1622, 1729, 1413, 1643, 2091, 1380, 2242, 2203, 2129, 1569, 1431, 1580, 1756, 1800, 2037, 1537, 2168, 1717, 1645, 2207, 1528, 1960, 2182, 1445, 1919, 2161, 2241, 1408, 1760, 1745, 1382, 1742, 2154, 2193, 2225, 1553, 1933, 2090, 1678, 2072, 2102, 1571, 1694, 2099, 1755, 1799, 2075, 1724, 437, 2215, 1568, 1728, 1624, 1743, 2101, 734, 1687, 1951, 2014, 2213, 1386, 2194, 547, 2110, 1519, 1046]. Therapeutic [2084]. There [1720]. Thermodynamics [337]. Thinking [2044]. Third [685, 2299]. Thomas [420]. Thompson [578, 325, 673]. Thornett [661]. Those [2117]. Three [1114, 1398, 357, 1721, 1509, 358, 1837, 247, 268, 100, 1318, 669, 409, 541, 2338, 1425, 2216]. Three-Decision [1318]. three-level [2338]. Three-Parameter [1114, 357, 100]. Three-Quarter [247]. Three-Stage [268]. Three-Way [1398, 1837, 1425]. Threshold [1860, 1758, 1846, 1127, 1311, 978]. throughput [2362]. Tie [745]. Tie-Breaker [745]. Ties [1672, 1596]. Tight [1328, 1369]. Tilting [1313].

Time [1012, 665, 1971, 96, 440, 1050, 1677, 702, 747, 739, 1846, 891, 1716, 1133, 1452, 1034, 1490, 1886, 1972, 1064, 1837, 2025, 989, 966, 586, 1352, 692, 526, 2066, 49, 1138, 1436, 1646, 1818, 1594, 1771, 2145, 1284, 1405, 664, 349, 1596, 533, 967, 1373, 2201, 393, 1017, 347, 1825, 991, 258, 55, 858, 150, 163, 81, 1945, 1580, 1742, 1863, 818, 2239, 2245, 2324, 2333, 1555, 1630, 2342, 2271, 2207]. Time-Dependent [1405]. Time-Series [1863, 2271]. Time-to-Event [1972, 2066, 2145]. Times [462, 1006, 2022, 622, 1230, 237]. Timing [821]. Tolerance [638, 26, 2335]. tomography [2244]. Tool [502, 1904, 2351, 2377]. tools [2123]. Top [1196]. Top-Order [1196]. Topics [2045]. Total [841, 494, 1002, 1788]. Totals [2015]. Tournaments [1247]. Toxicological [2141, 1914]. trading [2332]. Traffic [2058, 1673, 1825, 2395]. trans [2294]. trans-dimensional [2294]. Transaction [2157]. Transform [875, 924, 2076]. Transformation [2040, 737, 1123, 471, 875, 1760, 2386, 2416]. Transformations [694, 1295]. Transformed [627, 1637]. Transforms [1680]. transit [2324]. Transition [323, 1034, 951, 1702, 2084, 1005, 1035]. Transitions [1032, 1945]. Translated [260]. Translation [963, 1217]. Translation-Invariant [1217]. Transmission [1590]. Transmitter [146]. Transportation [457]. travel [2324]. Trawl [1466]. Treatment [2152, 1110, 2126, 284, 1299, 1040, 674, 1838, 2150, 2370]. Treatments [2170, 1318, 430, 500]. Tree [1704, 1554, 1952, 1430, 1567, 1684, 1689]. Tree-based [1554]. trees [2339]. treewidth [2291]. Trend [754, 2114, 1220, 851, 654, 621]. Trends [711, 633, 1564]. Trial [1900]. Trials [552, 62, 1989, 322, 1972, 1988, 1806, 2126, 1117, 1431, 1917, 1430, 2072, 1638, 2007, 2052, 2032]. Triangular [507, 1806]. triaxial [2301]. Tribute [1378, 1323]. Tributes [1378]. Trimmed [1833, 1182, 1273, 1544, 1459, 1793, 1990, 1610, 2245, 1636]. Trimming [1213]. Truncated [297, 306, 148, 1747, 376, 60, 588, 13, 108, 2192, 1694]. Truncation [53, 275, 1122, 1007]. Trust [647]. Try [863]. Tuberculosis [1000]. Tumor [522]. Tumours [1605]. Tuned [1844]. Turning [2025]. Tusnády [769]. twentieth [1552]. Twenty20 [2140]. Twin [1363]. Two [2070, 1626, 593, 643, 53, 1527, 2038, 404, 537, 1124, 1449, 67, 1442, 1509, 539, 1544, 414, 1269, 2023, 749, 71, 2108, 1111, 1385, 1227, 1411, 1154, 957, 1884, 1329, 1503, 2002, 276, 1313, 1283, 370, 1138, 2028, 1433, 1780, 1263, 726, 582, 2125, 1757, 1082, 1513, 876, 939, 1122, 461, 1243, 1144, 519, 430, 1007, 1798, 1443, 2160, 592, 121, 220, 309, 2089, 1926, 239, 449, 1431, 662, 1849, 588, 1145, 961, 1332, 2289, 2325, 2414, 2357, 682, 2320, 1559, 2335, 1694]. Two-Dimensional [1626, 1431]. Two-level [1509]. Two-Moment [1269]. Two-Parameter [1124, 1138]. two-party [2325]. Two-Pass [1926]. Two-Phase [1884, 121, 239, 961]. Two-Release [1780]. Two-Sample [1111, 957, 1313, 370, 2320]. Two-Sided [1849, 2335]. Two-Species [519]. Two-Stage [593, 1544, 1385, 1283, 582, 1757, 876, 939, 1798, 1443, 220, 2125, 1332]. Two-Truncation [1122]. Two-Way [1329, 1433, 1243, 449, 1527, 1513].

Type [1388, 1827, 504, 1412, 1153, 310, 1304, 1905, 49, 1384, 1666, 889, 1707, 865, 945, 561, 779, 1861, 231, 1858, 253, 2116, 1472, 1199, 2128, 420, 1424, 912]. Type-II [2128]. Types [799, 1648].

U [2267]. U.S. [1927]. UK£ [2267]. Ultimate [1663]. Umbrella [1293]. Unavailability [1256]. Unbiased [1082, 1122, 1007, 1557, 670, 589, 382, 1262]. Unbiasedness [578]. unbounded [2372]. Uncertain [1648, 2073, 2127, 2142]. Uncertainty [2130]. Unconditional [2243]. Unconstrained [1977]. Uncontrolled [14, 23]. under-reported [2343]. Underlying [85]. Understanding [1997, 2135, 2119]. Unemployed [1026]. Unemployment [1033, 920, 1025]. Unequal [308, 1749, 537, 67, 72, 793, 1415, 804, 1444, 2149, 2150]. Unequally [1117]. Unified [638, 1717]. Uniform [454, 1076, 828, 985]. Uniformity [789, 880, 839]. Uniformizable [882]. Unimodality [1371, 1035]. Union [1953]. Unique [1507]. Uniqueness [1665]. Unit [738, 1865, 1464]. Units [1375, 67]. Univariate [197, 2266, 2336]. Universal [1786]. Universities [249]. University [1520]. Unknown [2141, 390, 421, 26, 582, 1291, 774]. Unmasking [1911]. Unrelated [1792, 1283]. Unreplicated [1989]. unusual [2241]. Unvarying [586]. Upcrossings [259]. Upon [222, 740]. Upper [844, 1328, 1369, 1892, 1057, 483, 1190, 1456]. Urns [2003]. Usage [615]. Use [1091, 43, 451, 1675, 207, 1094, 2152, 1905, 666, 1210, 756, 1670, 554, 2155, 1963, 470, 1622, 733, 838, 2221]. Used [68, 14, 725]. Useful [693]. Usefulness [2085]. Users [2177]. Using [1388, 1971, 2058, 1833, 680, 743, 1475, 1776, 400, 1449, 2239, 615, 2191, 1105, 1349, 1153, 2015, 1484, 1213, 1785, 457, 1893, 1929, 1111, 812, 1965, 324, 2177, 1722, 1191, 1673, 2279, 2158, 1612, 2084, 1151, 1831, 1258, 1384, 1506, 2148, 1830, 1468, 1284, 1066, 1441, 1974, 889, 1712, 1523, 816, 744, 2041, 2190, 1964, 1952, 1074, 1443, 2098, 167, 1970, 1297, 1502, 1271, 1782, 1800, 2037, 1960, 2301, 1408, 1819, 1881, 1755, 1572, 2260, 2211, 1968, 2324, 2295, 2287, 1690, 1419, 2206, 1554, 1353, 2385, 2236, 2261, 2406, 1472, 2416, 2366, 2123, 2369, 2286, 1694, 1967, 2184]. Utilising [2168]. Utilize [2073, 2127, 2142]. Utilizing [1868].

Vaccine [1482]. Valid [1288]. Validation [1977, 1215]. Valley [3]. Value [791, 1157, 810, 1762, 1903, 370, 1976, 964, 765, 2081, 1251, 24, 1788, 2293]. Valued [1999, 919, 2411, 2330]. Values [1763, 1254, 341, 266, 1020, 1869, 1260, 427, 464, 506, 538, 1178, 696, 2344, 2376]. Variability [284]. Variable [1225, 1939, 529, 1352, 1631, 1063, 617, 2380, 1072, 1764, 216, 362, 875, 2415, 1486, 527, 2369, 2195, 1274, 2037, 2330, 355, 2251, 1566, 2291, 2311]. Variables [297, 2058, 275, 954, 2024, 2107, 2003, 1891, 302, 2108, 1018, 1356, 2181, 1837, 1342, 1999, 316, 2001, 813, 846, 14, 723, 1770, 1908, 1021, 619, 604, 1087, 535, 385, 926, 1196, 1382, 441, 2296, 2411, 1694]. Variance [883, 911, 1735, 710, 1749, 107, 2024, 429, 505, 1282, 112, 1785, 1750, 499, 1064, 2282, 196, 1884, 618, 1172, 1235, 1081, 26, 525, 11, 1921, 804, 1646, 677, 925, 1308, 1444, 582, 1950, 300, 1597, 1856, 1457, 1826, 182, 407, 167, 1016, 589, 978, 1642, 7, 1295, 1645, 952, 845, 1296, 2255, 2341, 2372, 1422, 2312, 1571, 2280]. Variance-Covariance [107, 1064, 167]. Variances [1015, 537, 678, 2141, 841, 390, 421, 2171, 130, 872, 1444, 475, 488, 670]. Variants [1305, 2410]. Variate [757, 1202, 112, 86, 626]. Variate-Difference [112]. Variates [53, 36, 131, 2258]. Variation [1753, 17, 960, 995, 774, 838, 1555, 2206]. Variational [1889, 2143, 1976, 2101, 2373, 2294]. Variations [267, 980, 22, 648]. Varietal [62]. Variety [1511, 1917, 2072]. Variograms [1069]. Various [1503]. Vary [1853]. Varying [643, 164, 431, 1375, 403, 539, 512, 310, 586, 2172, 726, 402]. Varying-coefficient [2172]. Vatolkin [2276]. Vech [909]. Vector [564, 1965, 2127, 1942, 1340, 2224, 2332, 2372]. Vectors [1225]. Vegetation [1105]. Vehicle [1825, 2324]. vehicles [2324]. Verified [1700]. Versatile [1655]. Versions [679]. Versus [795, 560, 1927]. vertical [2347]. VGAM [2321]. Via [1763, 1161, 1736, 1817, 1976, 2200, 1317, 2316, 2393, 1418, 1647, 2269, 2138, 2392]. Vicente [1954]. Victoria [2395, 832, 1991, 2167]. Victoria-Feser [1991]. Victorian [1392, 667]. Victory [1520]. View [599]. Views [662]. vine [2216]. Viruses [1829]. visit [2327]. Visits [1783, 1893, 1968, 1967]. visualisation [2216]. Visualising [2400]. Visualization [1512, 2160]. Vital [832, 101, 333, 685]. Vladimirovich [1323]. Volatilities [1888]. Volatility [1846, 1582, 1922]. Volume [1403, 1494, 1548, 1602, 1651, 1804]. Volumes [1440, 1549]. vote [2325]. Vs [902].

W [1971, 1044, 1992]. W. [1310, 832]. Waiting [114, 349]. Wald [312]. Wales [974, 2361]. Walk [230, 611, 150, 2371]. Walks [113]. Wallis [1592, 2404]. Walsh [109]. Warranty [1138]. Was [1720]. Watson [1378, 822, 689, 1663, 726, 286, 535, 234]. wavelet [1554]. Waves [2167]. Way [2070, 1398, 1801, 1837, 250, 1329, 525, 1433, 582, 1243, 2160, 449, 1425, 1527, 1560, 1513]. Weak [1357, 1692]. Weakly [996, 2255]. Weekday [1878]. Weekday-Only [1878]. Weibull [1114, 68]. Weighing [312, 841, 494, 282, 1002, 263, 374, 761]. Weight [841, 494, 1176, 1002, 355, 2408]. Weighted [2086, 1467, 2023, 688, 1937, 2047, 1183, 724, 1770, 1523, 1457, 1761, 826, 730, 1537, 2222, 1821, 2319, 1822]. Weighting [2050, 1379]. Weights [1414, 1447, 1937, 1960]. Weihs [2276, 2267]. Wellek [1956]. Welsh [2035, 1934, 1544]. Western [1272]. Wheat [62]. Where [1400, 1142]. Which [1369, 273, 1522]. Whiskies [1706]. White [1113, 962]. Whither [1553]. Whitney [1504]. Whole [2152, 127]. Whose [883]. Wickens [379]. Width [829, 639, 2416]. Wiens [2007]. Wilcoxon [2086, 1049, 1619, 1707, 1504, 910, 1537]. Wilcoxon-Type [1707]. Wildlife [1717]. Wilk [742]. Wilkinson [1496]. William [596, 646]. Williams [1222, 1223, 708]. Window [2059]. Winds [1716]. Winner [1711, 1806]. Winsorized [1808]. Wise [2201, 274]. Wishart [292, 392, 2253, 2373, 965, 293]. Within [1852, 562, 893]. Without [164, 308, 67, 180, 233, 343]. Wold [625]. Wolff [654]. Wood [1841]. Wool [823, 174]. Work [1034, 970, 548]. Working [850, 849]. World [490]. Worst [1097]. worth [1453]. Wrapped [607, 1811]. wrist [2301].

X [1535, 1235, 30]. X-11 [1535].

Yates [758, 630]. Year [1392, 1029]. Years [647]. Young [1032, 29]. Youth [1025, 1026].

Zealand [1364, 101, 1464, 1878, 1552, 1576, 1577]. Zeger [1410]. Zero [297, 1539, 10, 2088, 1311, 1607, 2359, 1568]. zero-augmented [1568]. Zero-bias [1607]. Zero-inflated [1539]. Zero-One-Truncated [297]. zeros [1568]. Zeta [1768]. Zimmerman [1954]. Zucchini [1971].

References

Wickens:1959:AJS

[1] P. C. Wickens. The Australian Journal of Statistics. Australian Journal of Statistics, 1(1):1–2, April 1959. CODEN AUJSA3. ISSN 0004-9581.

Campion:1959:CMR

[2] H. Campion. A congratulatory message from the Royal Statistical Society. Australian Journal of Statistics, 1(1):2, April 1959. CODEN AU-JSA3. ISSN 0004-9581.

Jarrett:1959:ERP

[3] F. G. Jarrett. Estimation of resource productivities as illustrated by a survey of the Lower Murray Valley dairying area. Australian Journal of Statistics, 1(1):3–11, April 1959. CODEN AUJSA3. ISSN 0004-9581.

Hanson:1959:PEA

[4] M. A. Hanson. A problem in economic allocation of effort in fault detection. Australian Journal of Statistics, 1(1):12–18, April 1959. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1959:GLT

[5] H. O. Lancaster. Generation life tables for Australia. Australian Journal of Statistics, 1(1):19–33, April 1959. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1959:NNa

[6] Anonymous. News and notes. Australian Journal of Statistics, 1(1):34, April 1959. CODEN AUJSA3. ISSN 0004-9581.

Tallis:1959:SEG

[7] G. M. Tallis. Sampling errors of genetic correlation coefficients calculated from analyses of variance and covariance. Australian Journal of Statistics, 1(2):35–43, August 1959. CODEN AUJSA3. ISSN 0004-9581.

Weiler:1959:NHG

[8] H. Weiler. Note on harmonic and geometric means. Australian Journal of Statistics, 1(2):44–46, August 1959. CODEN AUJSA3. ISSN 0004-9581.

Moran:1959:PCT

[9] P. A. P. Moran. The power of a cross-over test for the artificial stimulation of rain. Australian Journal of Statistics, 1(2):47–52, August 1959. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1959:ZCI

[10] H. O. Lancaster. Zero correlation and independence. Australian Journal of Statistics, 1(2):53–56, August 1959. CODEN AUJSA3. ISSN 0004-9581.

Konijn:1959:BDA

[11] H. S. Konijn. Basing decisions on an analysis of variance. Australian Journal of Statistics, 1(2):57–68, August 1959. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1959:NNb

[12] Anonymous. News and notes. Australian Journal of Statistics, 1(2): 69–72, August 1959. CODEN AUJSA3. ISSN 0004-9581.

Weiller:1959:MSD

[13] H. Weiller. Means and standard deviations of a truncated normal bivariate distribution. Australian Journal of Statistics, 1(3):73–81, November 1959. CODEN AUJSA3. ISSN 0004-9581.

Konijn:1959:MUS

[14] H. S. Konijn. Models used in statistical studies of relations between uncontrolled variables. Australian Journal of Statistics, 1(3):82–93, November 1959. CODEN AUJSA3. ISSN 0004-9581.

Sharp:1959:RRF

[15] D. P. Sharp. The role of research in a firm’s decision to enter a market and in its subsequent operations. Australian Journal of Statistics, 1(3): 94–99, November 1959. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1959:NNc

[16] Anonymous. News and notes. Australian Journal of Statistics, 1(3): 100–101, November 1959. CODEN AUJSA3. ISSN 0004-9581.

Hannan:1960:ESV

[17] E. J. Hannan. The estimation of seasonal variation. Australian Journal of Statistics, 2(1):1–15, April 1960. CODEN AUJSA3. ISSN 0004-9581.

Konjin:1960:MCC

[18] H. S. Konjin. Multiple comparison with controls. Australian Journal of Statistics, 2(1):16–18, April 1960. CODEN AUJSA3. ISSN 0004-9581.

Elliott:1960:CFT

[19] A. G. L. Elliott. Calibration of a flip tester. Australian Journal of Statistics, 2(1):19–31, April 1960. CODEN AUJSA3. ISSN 0004-9581.

Cornish:1960:FLP

[20] E. A. Cornish. Fiducial limits for parameters in compound hypotheses. Australian Journal of Statistics, 2(1):32–40, April 1960. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1960:NNa

[21] Anonymous. News and notes. Australian Journal of Statistics, 2(1):40, April 1960. CODEN AUJSA3. ISSN 0004-9581.

Hanson:1960:ESV

[22] M. A. Hanson. Errors and stochastic variations in linear programming. Australian Journal of Statistics, 2(2):41–46, August 1960. CODEN AU-JSA3. ISSN 0004-9581.

Konijn:1960:SRR

[23] H. S. Konijn. Some remarks on regression analysis with uncontrolled regressors. Australian Journal of Statistics, 2(2):47–52, August 1960. CODEN AUJSA3. ISSN 0004-9581.

Wilkinson:1960:CMV

[24] G. N. Wilkinson. Comparison of missing value procedures. Australian Journal of Statistics, 2(2):53–65, August 1960. CODEN AUJSA3. ISSN 0004-9581.

Tallis:1960:SEE

[25] G. M. Tallis. The sampling errors of estimated genetic regression coefficients and the errors of predicted genetic gains. Australian Journal of Statistics, 2(2):66–77, August 1960. CODEN AUJSA3. ISSN 0004-9581.

Jilek:1960:TLN

[26] Miloš Jílek and Otakar Líkař. Tolerance limits of the normal distribution with known variance and unknown mean. Australian Journal of Statistics, 2(2):78–83, August 1960. CODEN AUJSA3. ISSN 0004-9581.

Dewdney:1960:NSS

[27] J. C. H. Dewdney. A note on seasonal swings in mortality in Australia, 1861–1955. Australian Journal of Statistics, 2(2):84–90, August 1960. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1960:NNb

[28] Anonymous. News and notes. Australian Journal of Statistics, 2(2): 91–92, August 1960. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1960:MYA

[29] H. O. Lancaster. Mortality at young adult ages — a relative maximum. Australian Journal of Statistics, 2(3):93–96, November 1960. CODEN AUJSA3. ISSN 0004-9581.

Morley:1960:EGX

[30] F. H. W. Morley. An effect of genotype X environment interactions on estimation of genotypic covariances. Australian Journal of Statistics, 2 (3):97–101, November 1960. CODEN AUJSA3. ISSN 0004-9581.

Leser:1960:DFN

[31] C. E. V. Leser. Demand functions for nine commodity groups in Australia. Australian Journal of Statistics, 2(3):102–113, November 1960. CODEN AUJSA3. ISSN 0004-9581.

Wickens:1960:ALT

[32] P. C. Wickens. Abridged life tables for the Australian states, 1953–1955. Australian Journal of Statistics, 2(3):114–121, November 1960. CODEN AUJSA3. ISSN 0004-9581.

Verhagen:1960:GCT

[33] A. M. W. Verhagen. Growth curves and their functional form. Australian Journal of Statistics, 2(3):122–127, November 1960. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1960:NNc

[34] Anonymous. News and notes. Australian Journal of Statistics, 2(3): 128–130, November 1960. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1960:SBN

[35] Anonymous. Short book notices. Australian Journal of Statistics, 2(3): 130, November 1960. CODEN AUJSA3. ISSN 0004-9581.

Finney:1961:CLR

[36] D. J. Finney and M. R. Sampford. The comparison of linear regressions for sets of correlated variates. Australian Journal of Statistics, 3(1):1–19, April 1961. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1961:CPA

[37] H. O. Lancaster. The combination of probabilities: an application of orthonormal functions. Australian Journal of Statistics, 3(1):20–33, April 1961. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1961:SBN

[38] Anonymous. Short book notices. Australian Journal of Statistics, 3(1): 34, April 1961. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1961:EN

[39] Anonymous. Editorial note. Australian Journal of Statistics, 3(1):34, April 1961. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1961:NNa

[40] Anonymous. News and notes. Australian Journal of Statistics, 3(1): 35–36, April 1961. CODEN AUJSA3. ISSN 0004-9581.

Bofinger:1961:RTS

[41] Eve Bofinger and V. J. Bofinger. A runs test for sequences of random digits. Australian Journal of Statistics, 3(2):37–41, August 1961. CODEN AUJSA3. ISSN 0004-9581.

Gregory:1961:CTD

[42] G. Gregory. Contingency tables with a dependent classification. Australian Journal of Statistics, 3(2):42–47, August 1961. CODEN AUJSA3. ISSN 0004-9581.

Claringbold:1961:UOP

[43] P. J. Claringbold. The use of orthogonal polynomials in the partition of chi-square. Australian Journal of Statistics, 3(2):48–63, August 1961. CODEN AUJSA3. ISSN 0004-9581.

Hanson:1961:DTN

[44] M. A. Hanson. A duality theorem in non-linear programming with nonlinear constraints. Australian Journal of Statistics, 3(2):64–72, August 1961. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1961:BRa

[45] Anonymous. Book reviews. Australian Journal of Statistics, 3(2):72–75, August 1961. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1961:NNb

[46] Anonymous. News and notes. Australian Journal of Statistics, 3(2): 75–76, August 1961. CODEN AUJSA3. ISSN 0004-9581.

Konijn:1961:NPR

[47] H. S. Konijn. Non-parametric, robust and short-cut methods in regression and structural analysis. Australian Journal of Statistics, 3(3):77–87, November 1961. CODEN AUJSA3. ISSN 0004-9581.

Moran:1961:PCR

[48] P. A. P. Moran. Path coefficients reconsidered. Australian Journal of Statistics, 3(3):87–93, November 1961. CODEN AUJSA3. ISSN 0004-9581.

Jowett:1961:ACG

[49] G. H. Jowett. The analysis of continuously generated time series date of the type produced by recording instruments. Australian Journal of Statistics, 3(3):94–107, November 1961. CODEN AUJSA3. ISSN 0004-9581.

Wilkinson:1961:EMC

[50] G. N. Wilkinson. The eighth moment and cumulant of the distribution for normal samples of the skewness coefficient γ = k_{3}k_{2}^{−3/2}. Australian Journal of Statistics, 3(3):108–109, November 1961. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1961:BRb

[51] Anonymous. Book reviews. Australian Journal of Statistics, 3(3):110– 114, November 1961. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1961:NNc

[52] Anonymous. News and notes. Australian Journal of Statistics, 3(3): 115–117, November 1961. CODEN AUJSA3. ISSN 0004-9581.

Beattie:1962:TTV

[53] A. W. Beattie. Truncation in two variates to maximize a function of the means of a normal multivariate distribution. Australian Journal of Statistics, 4(1):1–3, April 1962. CODEN AUJSA3. ISSN 0004-9581.

Maritz:1962:CRR

[54] J. S. Maritz. Confidence regions for regression parameters. Australian Journal of Statistics, 4(1):4–10, April 1962. CODEN AUJSA3. ISSN 0004-9581.

Saunders:1962:PRT

[55] L. R. Saunders. The prediction of a random time series affected by a prescribed time function. Australian Journal of Statistics, 4(1):11–24, April 1962. CODEN AUJSA3. ISSN 0004-9581.

Hamdan:1962:PCS

[56] M. A. Hamdan. The powers of certain smooth tests of goodness of fit. Australian Journal of Statistics, 4(1):25–40, April 1962. CODEN AU-JSA3. ISSN 0004-9581.

Fisher:1962:CLC

[57] Sir Ronald A. Fisher. Confidence limits for a cross-product ratio. Australian Journal of Statistics, 4(1):41, April 1962. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1962:NNa

[58] Anonymous. News and notes. Australian Journal of Statistics, 4(1): 42–44, April 1962. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1962:BRa

Tallis:1962:MLE

[60] G. M. Tallis and S. S. Y. Young. Maximum likelihood estimation of parameters of the normal, log-normal, truncated normal and bivariate normal distributions from grouped data. Australian Journal of Statistics, 4(2):49–54, August 1962. CODEN AUJSA3. ISSN 0004-9581.

Gani:1962:ADR

[61] J. Gani and G. F. Yeo. On the age distribution of n ranked elements after several replacements. Australian Journal of Statistics, 4(2):55–60, August 1962. CODEN AUJSA3. ISSN 0004-9581.

Balaam:1962:ASW

[62] L. N. Balaam and R. D. Hunter. The analysis of a series of wheat varietal trials. Australian Journal of Statistics, 4(2):61–70, August 1962. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1962:NNb

[63] Anonymous. News and notes. Australian Journal of Statistics, 4(2): 70–71, August 1962. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1962:BRb

[64] Anonymous. Book reviews. Australian Journal of Statistics, 4(2):72–76, August 1962. CODEN AUJSA3. ISSN 0004-9581.

Saunders:1962:DSN

[65] L. R. Saunders. The dynamics of a simple non-linear prediction system. Australian Journal of Statistics, 4(3):77–85, November 1962. CODEN AUJSA3. ISSN 0004-9581.

Bennett:1962:STM

[66] R. W. Bennett. Sizes of the χ^{2} test in the multinomial distribution. Australian Journal of Statistics, 4(3):86–88, November 1962. CODEN AUJSA3. ISSN 0004-9581.

Brewer:1962:STU

[67] K. R. W. Brewer and G. C. Udny. Samples of two units drawn with unequal probabilities without replacement. Australian Journal of Statistics, 4(3):89–100, November 1962. CODEN AUJSA3. ISSN 0004-9581.

Bhattacharya:1962:PAU

[68] S. K. Bhattacharya. On a probit analogue used in a life-test based on the Weibull distribution. Australian Journal of Statistics, 4(3):101–105, November 1962. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1962:NNc

[69] Anonymous. News and notes. Australian Journal of Statistics, 4(3):105, November 1962. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1962:BRc

[70] Anonymous. Book reviews. Australian Journal of Statistics, 4(3):106– 110, November 1962. CODEN AUJSA3. ISSN 0004-9581.

Ewens:1963:TAR

[71] W. J. Ewens. Two asymptotic results in sequential analysis. Australian Journal of Statistics, 5(1):1–4, April 1963. CODEN AUJSA3. ISSN 0004-9581.

Brewer:1963:MSS

[72] K. R. W. Brewer. A model of systematic sampling with unequal probabilities. Australian Journal of Statistics, 5(1):5–13, April 1963. CODEN AUJSA3. ISSN 0004-9581.

Hanson:1963:ACP

[73] M. A. Hanson. An algorithm for convex programming. Australian Journal of Statistics, 5(1):14–19, April 1963. CODEN AUJSA3. ISSN 0004-9581.

Sethi:1963:NOS

[74] V. K. Sethi. A note on optimum stratification of populations for estimating the population means. Australian Journal of Statistics, 5(1):20–33, April 1963. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1963:NNa

[75] Anonymous. News and notes. Australian Journal of Statistics, 5(1): 34–35, April 1963. CODEN AUJSA3. ISSN 0004-9581.

Moran:1963:ASS

[76] P. A. P. Moran. The Australian Statistical Society. Australian Journal of Statistics, 5(1):36, April 1963. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1963:BRa

Anonymous:1963:NEP

[78] Anonymous. Note on editorial policy: Note on editorial policy. Australian Journal of Statistics, 5(2):i, August 1963. CODEN AUJSA3. ISSN 0004-9581.

Verhagen:1963:CLM

[79] A. M. W. Verhagen. The “caution-level” in multiple tests of significance. Australian Journal of Statistics, 5(2):41–48, August 1963. CODEN AU-JSA3. ISSN 0004-9581.

Craven:1963:SRB

[80] B. D. Craven. Some results for the bulk service queue. Australian Journal of Statistics, 5(2):49–56, August 1963. CODEN AUJSA3. ISSN 0004-9581.

Shanbhag:1963:QPS

[81] D. N. Shanbhag. On queues with Poisson service time. Australian Journal of Statistics, 5(2):57–61, August 1963. CODEN AUJSA3. ISSN 0004-9581.

Balaam:1963:MCS

[82] L. N. Balaam. Multiple comparisons — a sampling experiment. Australian Journal of Statistics, 5(2):62–84, August 1963. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1963:NNb

[83] Anonymous. News and notes. Australian Journal of Statistics, 5(2):85, August 1963. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1963:BRb

[84] Anonymous. Book reviews. Australian Journal of Statistics, 5(2):86–91, August 1963. CODEN AUJSA3. ISSN 0004-9581.

Brewer:1963:REF

[85] K. R. W. Brewer. Ratio estimation and finite populations: Some results deducible from the assumption of an underlying stochastic process. Australian Journal of Statistics, 5(3):93–105, November 1963. CODEN AUJSA3. ISSN 0004-9581.

Rao:1963:VDM

[86] J. N. K. Rao and G. Tintner. On the variate difference method. Australian Journal of Statistics, 5(3):106–116, November 1963. CODEN AUJSA3. ISSN 0004-9581.

Powell:1963:SDA

[87] Alan Powell. Serial dependence in annual demands for fodder reserves. Australian Journal of Statistics, 5(3):117–124, November 1963. CODEN AUJSA3. ISSN 0004-9581.

Stenhouse:1963:EA

[88] N. S. Stenhouse. Errata and addendum. Australian Journal of Statistics, 5(3):125–126, November 1963. CODEN AUJSA3. ISSN 0004-9581.

Niven:1963:JDS

[89] B. S. Niven. The joint distribution of the sample mean and the sample range. Australian Journal of Statistics, 5(3):127–132, November 1963. CODEN AUJSA3. ISSN 0004-9581.

Day:1963:NMD

[90] L. H. Day. A note on the measurement of divorce, with special reference to Australian data. Australian Journal of Statistics, 5(3):133–142, November 1963. CODEN AUJSA3. ISSN 0004-9581.

Konijn:1963:NNE

[91] H. S. Konijn. Note on the non-existence of a maximum likelihood estimate. Australian Journal of Statistics, 5(3):143–146, November 1963. CODEN AUJSA3. ISSN 0004-9581.

John:1963:ECB

[92] Peter W. M. John. Extended complete block designs. Australian Journal of Statistics, 5(3):147–152, November 1963. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1963:NNc

[93] Anonymous. News and notes. Australian Journal of Statistics, 5(3): 153–154, November 1963. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1963:BRc

[94] Anonymous. Book reviews. Australian Journal of Statistics, 5(3):155– 159, November 1963. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1963:C

[95] Anonymous. Corrigenda. Australian Journal of Statistics, 5(3):159, November 1963. CODEN AUJSA3. ISSN 0004-9581.

Anderson:1964:SAS

[96] T. W. Anderson. Some approaches to the statistical analysis of time series. Australian Journal of Statistics, 6(1):1–11, April 1964. CODEN AUJSA3. ISSN 0004-9581.

Staff:1964:DPD

[97] P. J. Staff. The displaced Poisson distribution. Australian Journal of Statistics, 6(1):12–20, April 1964. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1964:NNa

[98] Anonymous. News and notes. Australian Journal of Statistics, 6(1):21, April 1964. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1964:BRa

[99] Anonymous. Book reviews. Australian Journal of Statistics, 6(1):22–28, April 1964. CODEN AUJSA3. ISSN 0004-9581.

Lambert:1964:EPT

[100] J. A. Lambert. Estimation of parameters in the three-parameter lognormal distribution. Australian Journal of Statistics, 6(2):29–32, August 1964. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1964:BVS

[101] H. O. Lancaster. Bibliography of vital statistics in Australia and New Zealand. Australian Journal of Statistics, 6(2):33–99, August 1964. CO-DEN AUJSA3. ISSN 0004-9581. See also second and third lists [333, 685].

Anonymous:1964:HLM

[102] Anonymous. Honorary life member. Australian Journal of Statistics, 6 (2):100, August 1964. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1964:NNb

[103] Anonymous. News and notes. Australian Journal of Statistics, 6(2):101, August 1964. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1964:BRb

[104] Anonymous. Book reviews. Australian Journal of Statistics, 6(2):102– 107, August 1964. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1964:N

[105] Anonymous. Notice. Australian Journal of Statistics, 6(2):108, August 1964. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1964:C

[106] Anonymous. Corrigenda. Australian Journal of Statistics, 6(2):108, August 1964. CODEN AUJSA3. ISSN 0004-9581.

Bofinger:1964:VCF

[107] Eve Bofinger. Variance-covariance formulae and estimation for Markov chains. Australian Journal of Statistics, 6(3):109–116, November 1964. CODEN AUJSA3. ISSN 0004-9581.

Williams:1964:FCM

[108] Jean M. Williams and H. Weiler. Further charts for the means of truncated normal bivariate distributions. Australian Journal of Statistics, 6 (3):117–129, November 1964. CODEN AUJSA3. ISSN 0004-9581.

Hamdan:1964:STG

[109] M. A. Hamdan. A smooth test of goodness of fit based on the Walsh functions. Australian Journal of Statistics, 6(3):130–136, November 1964. CODEN AUJSA3. ISSN 0004-9581.

Haig:1964:LEL

[110] B. D. Haig. Letter to the Editor: Letter to the Editor. Australian Journal of Statistics, 6(3):137–139, November 1964. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1964:BRc

[111] Anonymous. Book reviews. Australian Journal of Statistics, 6(3):143– 145, November 1964. CODEN AUJSA3. ISSN 0004-9581.

Chew:1965:ECE

[112] Victor Chew. Effects of correlated errors on variate-difference estimates of error variance. Australian Journal of Statistics, 7(1):1–6, April 1965. CODEN AUJSA3. ISSN 0004-9581.

Heathcote:1965:RWP

[113] C. R. Heathcote. Random walks and a price support scheme. Australian Journal of Statistics, 7(1):7–14, April 1965. CODEN AUJSA3. ISSN 0004-9581.

Bhat:1965:COQ

[114] U. N. Bhat. Customer overflow in queues with finite waiting space. Australian Journal of Statistics, 7(1):15–19, April 1965. CODEN AUJSA3. ISSN 0004-9581.

Bofinger:1965:SSP

[115] V. J. Bofinger. The k-sample slippage problem. Australian Journal of Statistics, 7(1):20–31, April 1965. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1965:BRa

[116] Anonymous. Book reviews. Australian Journal of Statistics, 7(1):32, April 1965. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1965:NNa

[117] Anonymous. News and notes. Australian Journal of Statistics, 7(1):32, April 1965. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1965:SSA

[118] Anonymous. The Statistical Society of Australia: The Statistical Society of Australia. Australian Journal of Statistics, 7(1):33–36, April 1965. CODEN AUJSA3. ISSN 0004-9581.

Kerridge:1965:PDN

[119] D. Kerridge. A probabilistic derivation of the non-central χ^{2} distribution. Australian Journal of Statistics, 7(2):37–39, August 1965. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1965:STS

[120] H. O. Lancaster and T. A. I. Brown. Sizes of the χ^{2} test in the symmetrical multinomials. Australian Journal of Statistics, 7(2):40–44, August 1965. CODEN AUJSA3. ISSN 0004-9581.

Singh:1965:SCT

[121] D. Singh and B. D. Singh. Some contributions to two-phase sampling. Australian Journal of Statistics, 7(2):45–47, August 1965. CODEN AU-JSA3. ISSN 0004-9581.

Shanbhag:1965:ABS

[122] D. N. Shanbhag. On the asymptotic behaviour of a single-server queue. Australian Journal of Statistics, 7(2):48–53, August 1965. CODEN AU-JSA3. ISSN 0004-9581.

Anonymous:1965:BRb

[123] Anonymous. Book reviews. Australian Journal of Statistics, 7(2):54–56, August 1965. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1965:NNb

[124] Anonymous. News and notes. Australian Journal of Statistics, 7(2):54, August 1965. CODEN AUJSA3. ISSN 0004-9581.

Bofinger:1965:CMS

[125] Eve Bofinger and V. J. Bofinger. The correlation of maxima in samples drawn from a bivariate normal distribution. Australian Journal of Statistics, 7(3):57–61, November 1965. CODEN AUJSA3. ISSN 0004-9581.

Wilkins:1965:SRC

[126] C. A. Wilkins. Some remarks on a certain process. Australian Journal of Statistics, 7(3):62–66, November 1965. CODEN AUJSA3. ISSN 0004-9581.

Wegge:1965:ICS

[127] Leon L. Wegge. Identifiability criteria for a system of equations as a whole. Australian Journal of Statistics, 7(3):67–77, November 1965. CO-DEN AUJSA3. ISSN 0004-9581.

Tiku:1965:SED

[128] M. L. Tiku. Series expansions for the doubly non-central F -distribution. Australian Journal of Statistics, 7(3):78–89, November 1965. CODEN AUJSA3. ISSN 0004-9581.

Swarup:1965:SAL

[129] Kanti Swarup. Some aspects of linear fractional functionals programming. Australian Journal of Statistics, 7(3):90–104, November 1965. CODEN AUJSA3. ISSN 0004-9581.

Krishnaiah:1965:STE

[130] P. R. Krishnaiah. Simultaneous tests for the equality of variances against certain alternatives. Australian Journal of Statistics, 7(3):105– 109, November 1965. CODEN AUJSA3. ISSN 0004-9581.

Robinson:1965:DGQ

[131] J. Robinson. The distribution of a general quadratic form in normal variates. Australian Journal of Statistics, 7(3):110–114, November 1965. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1965:C

[132] Anonymous. Corrigenda. Australian Journal of Statistics, 7(3):114, November 1965. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1965:SMD

[133] H. O. Lancaster. Symmetry in multivariate distributions. Australian Journal of Statistics, 7(3):115–126, November 1965. CODEN AUJSA3. ISSN 0004-9581.

Bhattacharjee:1965:DRN

[134] G. P. B. Bhattacharjee. Distribution of range in non-normal samples. Australian Journal of Statistics, 7(3):127–141, November 1965. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1965:NNc

[135] Anonymous. News and notes. Australian Journal of Statistics, 7(3): 142–143, November 1965. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1965:BRc

[136] Anonymous. Book reviews. Australian Journal of Statistics, 7(3):144– 150, November 1965. CODEN AUJSA3. ISSN 0004-9581.

Moran:1966:EI

[137] P. A. P. Moran. Estimation from inequalities. Australian Journal of Statistics, 8(1):1–8, April 1966. CODEN AUJSA3. ISSN 0004-9581.

McNeil:1966:GCS

[138] D. R. McNeil. Grouping corrections for sequential tests. Australian Journal of Statistics, 8(1):9–21, April 1966. CODEN AUJSA3. ISSN 0004-9581.

Kabe:1966:SRN

[139] D. G. Kabe. Some results for the normal multivariate regression models. Australian Journal of Statistics, 8(1):22–27, April 1966. CODEN AUJSA3. ISSN 0004-9581.

Joffe:1966:NND

[140] A. D. Joffe. A note on the number of dust particles in a sample of airborne mine dust. Australian Journal of Statistics, 8(1):28–31, April 1966. CODEN AUJSA3. ISSN 0004-9581.

Holla:1966:NBE

[141] M. S. Holla. Notes: Bayesian estimates of the reliability function. Australian Journal of Statistics, 8(1):32–35, April 1966. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1966:NNa

[142] Anonymous. News and notes. Australian Journal of Statistics, 8(1): 36–37, April 1966. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1966:Ca

[143] Anonymous. Corrigenda. Australian Journal of Statistics, 8(1):37, April 1966. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1966:SSA

[144] Anonymous. The Statistical Society of Australia: Annual report of the Central Council for 1965. Australian Journal of Statistics, 8(1):38–41, April 1966. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1966:BRa

[145] Anonymous. Book reviews. Australian Journal of Statistics, 8(1):42–51, April 1966. CODEN AUJSA3. ISSN 0004-9581.

Vere-Jones:1966:SSM

[146] D. Vere-Jones. Simple stochastic models for the release of Quanta of transmitter from a nerve terminal. Australian Journal of Statistics, 8(2): 53–63, August 1966. CODEN AUJSA3. ISSN 0004-9581.

Drane:1966:FPP

[147] N. T. Drane. Fuel and power price indexes, Australian manufacturing, 1908–9 to 1960–61. Australian Journal of Statistics, 8(2):64–79, August 1966. CODEN AUJSA3. ISSN 0004-9581.

Rai:1966:CSC

[148] G. Rai. Cumulative sum control charts for truncated normal distributions. Australian Journal of Statistics, 8(2):80–86, August 1966. CODEN AUJSA3. ISSN 0004-9581.

Kabe:1966:CASa

[149] D. G. Kabe. Complex analogues of some results in classical normal multivariate regression theory. Australian Journal of Statistics, 8(2): 87–91, August 1966. CODEN AUJSA3. ISSN 0004-9581.

Seneta:1966:QSB

[150] E. Seneta. Quasi-stationary behaviour in the random walk with continuous time. Australian Journal of Statistics, 8(2):92–98, August 1966. CODEN AUJSA3. ISSN 0004-9581.

Kabe:1966:CASb

[151] D. G. Kabe. Complex analogues of some classical noncentral multivariate distributions. Australian Journal of Statistics, 8(2):99–103, August 1966. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1966:NNb

[152] Anonymous. News and notes. Australian Journal of Statistics, 8(2):104, August 1966. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1966:BRb

[153] Anonymous. Book reviews. Australian Journal of Statistics, 8(2):105– 115, August 1966. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1966:FP

[154] H. O. Lancaster. Forerunners of the Pearson χ^{2}. Australian Journal of Statistics, 8(3):117–126, November 1966. CODEN AUJSA3. ISSN 0004-9581.

James:1966:ISS

[155] J. W. James. Incomplete sampling for selection of a fixed proportion. Australian Journal of Statistics, 8(3):127–137, November 1966. CODEN AUJSA3. ISSN 0004-9581.

Price:1966:AMS

[156] Charles Price. Australian migration statistics, 1959–65 (with special reference to settler gain and loss). Australian Journal of Statistics, 8(3): 138–153, November 1966. CODEN AUJSA3. ISSN 0004-9581.

Jain:1966:IPA

[157] H. C. Jain. An inventory problem applied to a rental situation. Australian Journal of Statistics, 8(3):154–162, November 1966. CODEN AUJSA3. ISSN 0004-9581.

Cinlar:1966:DSM

[158] E. Cinlar. Decomposition of a semi-Markov process under a Markovian rule. Australian Journal of Statistics, 8(3):163–170, November 1966. CO-DEN AUJSA3. ISSN 0004-9581.

Anonymous:1966:HLM

[159] Anonymous. Honorary life member. Australian Journal of Statistics, 8 (3):171–172, November 1966. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1966:NNc

[160] Anonymous. News and notes. Australian Journal of Statistics, 8(3): 173–174, November 1966. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1966:BRc

[161] Anonymous. Book reviews. Australian Journal of Statistics, 8(3):175– 179, November 1966. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1966:Cb

[162] Anonymous. Corrigenda. Australian Journal of Statistics, 8(3):180, November 1966. CODEN AUJSA3. ISSN 0004-9581.

Seneta:1967:IDC

[163] E. Seneta. On imbedding discrete chains in continuous time. Australian Journal of Statistics, 9(1):1–7, April 1967. CODEN AUJSA3. ISSN 0004-9581.

Avadhani:1967:CSV

[164] M. S. Avadhani and B. V. Sukhatme. Controlled sampling with varying probabilities with and without replacement. Australian Journal of Statistics, 9(1):8–15, April 1967. CODEN AUJSA3. ISSN 0004-9581.

Rao:1967:EPI

[165] S. Subba Rao. The effect of postponable interruptions on a M/G/1 system with impatient customers. Australian Journal of Statistics, 9(1): 16–34, April 1967. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1967:NNa

[166] Anonymous. News and notes. Australian Journal of Statistics, 9(1):34, April 1967. CODEN AUJSA3. ISSN 0004-9581.

Scheinberg:1967:SVR

[167] Eliyahu Scheinberg. The sampling variance of the relative efficiency of indirect to direct selection when using variance-covariance components. Australian Journal of Statistics, 9(1):35–40, April 1967. CODEN AU-JSA3. ISSN 0004-9581.

Anonymous:1967:HLM

[168] Anonymous. Honorary life member. Australian Journal of Statistics, 9 (1):41, April 1967. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1967:SSA

[169] Anonymous. The Statistical Society of Australia: Annual report of the Central Council for 1966. Australian Journal of Statistics, 9(1):42–45, April 1967. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1967:BRa

Tiku:1967:NEL

[171] M. L. Tiku. A note on estimating the location and scale parameters of the exponential distribution from a censored sample. Australian Journal of Statistics, 9(2):49–54, August 1967. CODEN AUJSA3. ISSN 0004-9581.

John:1967:ECS

[172] S. John. Estimators for the comparison of sampling designs. Australian Journal of Statistics, 9(2):55–56, August 1967. CODEN AUJSA3. ISSN 0004-9581.

Chikkagoudar:1967:SPR

[173] M. S. Chikkagoudar. Sampling with preliminary random stratification. Australian Journal of Statistics, 9(2):57–70, August 1967. CODEN AU-JSA3. ISSN 0004-9581.

Dent:1967:MAS

[174] W. Dent. A model for the allocation of sales dates to Australian wool selling centres. Australian Journal of Statistics, 9(2):71–81, August 1967. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1967:BRb

[175] Anonymous. Book reviews. Australian Journal of Statistics, 9(2):82–88, August 1967. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1967:NNb

[176] Anonymous. News and notes. Australian Journal of Statistics, 9(2):88, August 1967. CODEN AUJSA3. ISSN 0004-9581.

Linnik:1967:CCB

[177] Yu. V. Linnik. On certain connections between algebraic geometry and statistics. Australian Journal of Statistics, 9(3):89–92, November 1967. CODEN AUJSA3. ISSN 0004-9581.

Seneta:1967:MAM

[178] E. Seneta. On the maxima of absorbing Markov chains. Australian Journal of Statistics, 9(3):93–102, November 1967. CODEN AUJSA3. ISSN 0004-9581.

Nicholls:1967:ESD

[179] D. F. Nicholls. Estimation of the spectral density function when testing for a jump in the spectrum. Australian Journal of Statistics, 9(3):103– 108, November 1967. CODEN AUJSA3. ISSN 0004-9581.

Chikkagoudar:1967:PSR

[180] M. S. Chikkagoudar. On PPS sampling with and without replacement. Australian Journal of Statistics, 9(3):109–118, November 1967. CODEN AUJSA3. ISSN 0004-9581.

Eagleson:1967:RSS

[181] G. K. Eagleson and H. O. Lancaster. The regression system of sums with random elements in common. Australian Journal of Statistics, 9(3): 120–125, November 1967. CODEN AUJSA3. ISSN 0004-9581.

Robinson:1967:VHE

[182] J. Robinson and L. N. Balaam. Variance heterogeneity and error correlation in factorial experiments. Australian Journal of Statistics, 9(3): 126–130, November 1967. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1967:BRc

[183] Anonymous. Book reviews. Australian Journal of Statistics, 9(3):131– 132, November 1967. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1967:NNc

[184] Anonymous. News and notes. Australian Journal of Statistics, 9(3):132, November 1967. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1967:AJS

[185] Anonymous. The Australian Journal of Statistics editorial policy. Australian Journal of Statistics, 9(3):132–133, November 1967. CODEN AUJSA3. ISSN 0004-9581.

Avadhani:1968:SPD

[186] M. S. Avadhani and B. V. Sukhatme. Simplified procedures for designing controlled simple random sampling. Australian Journal of Statistics, 10 (1):1–7, April 1968. CODEN AUJSA3. ISSN 0004-9581.

Cinlar:1968:SJD

[187] Erhan Çinlar. Some joint distributions for Markov renewal processes. Australian Journal of Statistics, 10(1):8–20, April 1968. CODEN AU-JSA3. ISSN 0004-9581.

Kabe:1968:DRC

[188] D. G. Kabe. On the distribution of the regression coefficient matrix of a normal distribution. Australian Journal of Statistics, 10(1):21–23, April 1968. CODEN AUJSA3. ISSN 0004-9581.

Nagnur:1968:AOS

[189] B. N. Nagnur. Asymptotically optimal sequential tests for simple and composite hypotheses. Australian Journal of Statistics, 10(1):24–30, April 1968. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1968:BRa

[190] Anonymous. Book reviews. Australian Journal of Statistics, 10(1):31–40, April 1968. CODEN AUJSA3. ISSN 0004-9581.

Maitland:1968:SSC

[191] D. W. Maitland. Sir Stanley Carver 1897–1967. Australian Journal of Statistics, 10(1):41–43, April 1968. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1968:Ca

[192] Anonymous. Corrigenda. Australian Journal of Statistics, 10(1):43, April 1968. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1968:NNa

[193] Anonymous. News and notes. Australian Journal of Statistics, 10(1):43, April 1968. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1968:SSA

[194] Anonymous. The Statistical Society of Australia: Annual report of the Central Council for 1967. Australian Journal of Statistics, 10(1):44–47, April 1968. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1968:AJSa

[195] Anonymous. The Australian Journal of Statistics editorial policy. Australian Journal of Statistics, 10(1):48, April 1968. CODEN AUJSA3. ISSN 0004-9581.

Hamdan:1968:VCC

[196] M. A. Hamdan. The variance of the components of chi-square in contingency tables. Australian Journal of Statistics, 10(2):49–55, August 1968. CODEN AUJSA3. ISSN 0004-9581.

Prabhu-Ajgaonkar:1968:TUS

[197] S. G. Prabhu-Ajgaonkar. The theory of univariate sampling on successive occasions under the general correlation pattern. Australian Journal of Statistics, 10(2):56–63, August 1968. CODEN AUJSA3. ISSN 0004-9581.

Tiku:1968:EPN

[198] M. L. Tiku. Estimating the parameters of normal and logistic distributions from censored samples. Australian Journal of Statistics, 10(2): 64–74, August 1968. CODEN AUJSA3. ISSN 0004-9581.

Das:1968:DIB

[199] M. N. Das and A. C. Kulshreshtha. On derivation of initial blocks of B.I.B. designs with more than one initial block. Australian Journal of Statistics, 10(2):75–82, August 1968. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1968:M

[200] H. O. Lancaster. The median of χ. Australian Journal of Statistics, 10 (2):83, August 1968. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1968:BRb

[201] Anonymous. Book reviews. Australian Journal of Statistics, 10(2):84– 102, August 1968. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1968:NNb

[202] Anonymous. News and notes. Australian Journal of Statistics, 10(2): 102–103, August 1968. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1968:SIS

[203] Anonymous. 37th session of the International Statistical Institute. Australian Journal of Statistics, 10(2):103–104, August 1968. CODEN AU-JSA3. ISSN 0004-9581.

Anonymous:1968:Cb

[204] Anonymous. Corrigenda. Australian Journal of Statistics, 10(2):103–104, August 1968. CODEN AUJSA3. ISSN 0004-9581.

Kulkarni:1968:OAT

[205] S. R. Kulkarni. On the optimal asymptotic tests for the effects of cloud seeding on rainfall (1). The case of fixed effects. Australian Journal of Statistics, 10(3):105–115, November 1968. CODEN AUJSA3. ISSN 0004-9581.

Krishnaiah:1968:NCB

[206] P. R. Krishnaiah and P. K. Pathak. A note on confidence bounds for certain ratios of characteristic roots of covariance matrices. Australian Journal of Statistics, 10(3):116–119, November 1968. CODEN AUJSA3. ISSN 0004-9581.

Day:1968:UTS

[207] N. E. Day. The use of Taylor series in sequential analysis. Australian Journal of Statistics, 10(3):120–124, November 1968. CODEN AUJSA3. ISSN 0004-9581.

Verhagen:1968:PSI

[208] A. M. W. Verhagen. A perspective in statistical inference. Australian Journal of Statistics, 10(3):125–137, November 1968. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1968:BRc

[209] Anonymous. Book reviews. Australian Journal of Statistics, 10(3):138– 143, November 1968. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1968:NNc

[210] Anonymous. News and notes. Australian Journal of Statistics, 10(3): 143, November 1968. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1968:AJSb

[211] Anonymous. The Australian Journal of Statistics editorial policy. Australian Journal of Statistics, 10(3):144, November 1968. CODEN AU-JSA3. ISSN 0004-9581.

Mehndiratta:1969:GTS

[212] S. L. Mehndiratta. A generalization of a theorem of Sinha on supports of a convex function. Australian Journal of Statistics, 11(1):1–6, April 1969. CODEN AUJSA3. ISSN 0004-9581.

Nicholls:1969:TJC

[213] D. F. Nicholls. Testing for a jump in co-spectra. Australian Journal of Statistics, 11(1):7–13, April 1969. CODEN AUJSA3. ISSN 0004-9581.

Bartoszynski:1969:MAD

[214] Robert Bartoszynski and Howard D’Abrera. A model of age-dependent inheritance. Australian Journal of Statistics, 11(1):14–28, April 1969. CODEN AUJSA3. ISSN 0004-9581.

Eagleson:1969:CTP

[215] G. K. Eagleson. A characterization theorem for positive definite sequences on the Krawtchouk polynomials. Australian Journal of Statistics, 11(1):29–38, April 1969. CODEN AUJSA3. ISSN 0004-9581.

Kulkarni:1969:OAT

[216] S. R. Kulkarni. On the optimal asymptotic tests for the effects of cloud seeding on rainfall (2). The case of variable effects. Australian Journal of Statistics, 11(1):39–51, April 1969. CODEN AUJSA3. ISSN 0004-9581. See corrigenda [362].

Anonymous:1969:BRa

[217] Anonymous. Book reviews. Australian Journal of Statistics, 11(1):52–57, April 1969. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1969:NNa

[218] Anonymous. News and notes. Australian Journal of Statistics, 11(1):58, April 1969. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1969:BRB

[219] Anonymous. Books received. Australian Journal of Statistics, 11(1):58, April 1969. CODEN AUJSA3. ISSN 0004-9581.

Singh:1969:TSS

[220] D. Singh and O. P. Kathuria. On two-stage successive sampling. Australian Journal of Statistics, 11(2):59–66, August 1969. CODEN AU-JSA3. ISSN 0004-9581.

Vere-Jones:1969:SLT

[221] D. Vere-Jones. Some limit theorems for evanescent processes. Australian Journal of Statistics, 11(2):67–78, August 1969. CODEN AUJSA3. ISSN 0004-9581.

Khatri:1969:TEP

[222] C. G. Khatri and M. C. Jaiswal. Testing the equality of parameters in the distributions when the range depends upon the parameter. Australian Journal of Statistics, 11(2):79–84, August 1969. CODEN AUJSA3. ISSN 0004-9581.

Mandl:1969:TNN

[223] P. Mandl and E. Seneta. The theory of non-negative matrices in a dynamic programming problem. Australian Journal of Statistics, 11(2): 85–96, August 1969. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1969:BRc

[224] Anonymous. Book reviews. Australian Journal of Statistics, 11(2):97– 103, August 1969. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1969:NNb

[225] Anonymous. News and notes. Australian Journal of Statistics, 11(2): 104–105, August 1969. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1969:SSA

[226] Anonymous. The Statistical Society of Australia: Annual report of the Central Council for 1968. Australian Journal of Statistics, 11(2):106–109, August 1969. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1969:AJS

[227] Anonymous. The Australian Journal of Statistics editorial policy. Australian Journal of Statistics, 11(2):110, August 1969. CODEN AUJSA3. ISSN 0004-9581.

McGilchrist:1969:TPC

[228] C. A. McGilchrist. Testing for plant competition. Australian Journal of Statistics, 11(3):111–119, November 1969. CODEN AUJSA3. ISSN 0004-9581.

Walsh:1969:SSA

[229] John E. Walsh. Sample sizes for approximate independence between sample median and largest (or smallest) order statistic. Australian Journal of Statistics, 11(3):120–122, November 1969. CODEN AUJSA3. ISSN 0004-9581.

Praetz:1969:ASP

[230] P. D. Praetz. Australian share prices and the random walk hypothesis. Australian Journal of Statistics, 11(3):123–139, November 1969. CODEN AUJSA3. ISSN 0004-9581.

Tripathi:1969:RTE

[231] T. P. Tripathi. A regression-type estimator in sampling with PPS and with replacement. Australian Journal of Statistics, 11(3):140–148, November 1969. CODEN AUJSA3. ISSN 0004-9581. See corrigendum [253].

Tiku:1969:DNC

[232] M. L. Tiku. Doubly non-central t distribution. Australian Journal of Statistics, 11(3):149–154, November 1969. CODEN AUJSA3. ISSN 0004-

Chikkagoudar:1969:ISR

[233] M. S. Chikkagoudar. Inverse sampling without replacement. Australian Journal of Statistics, 11(3):155–165, November 1969. CODEN AUJSA3. ISSN 0004-9581.

Quine:1969:LTG

[234] M. P. Quine and E. Seneta. A limit theorem for the Galton–Watson process with immigration. Australian Journal of Statistics, 11(3):166– 173, November 1969. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1969:BRd

[235] Anonymous. Book reviews. Australian Journal of Statistics, 11(3):174– 178, November 1969. CODEN AUJSA3. ISSN 0004-9581.

Bofinger:1970:CMS

[236] V. J. Bofinger. The correlation of maxima in several bivariate non-normal distributions. Australian Journal of Statistics, 12(1):1–7, April 1970. CODEN AUJSA3. ISSN 0004-9581.

Wright:1970:SOO

[237] Gordon P. Wright. Some optimal ordering policies for inventory systems with stochastic lead times. Australian Journal of Statistics, 12(1):8–12, April 1970. CODEN AUJSA3. ISSN 0004-9581.

Jensen:1970:JDQ

[238] D. R. Jensen. The joint distribution of quadratic forms and related distributions. Australian Journal of Statistics, 12(1):13–22, April 1970. CODEN AUJSA3. ISSN 0004-9581.

Srivastava:1970:TPS

[239] Surendra K. Srivastava. A two-phase sampling estimator in sample surveys. Australian Journal of Statistics, 12(1):23–27, April 1970. CODEN AUJSA3. ISSN 0004-9581.

Wu:1970:NSD

[240] P. M. Wu. A note on storage with deterministic outputs and inputs subject to breakdown. Australian Journal of Statistics, 12(1):28–32, April 1970. CODEN AUJSA3. ISSN 0004-9581.

Lambert:1970:EPF

[241] J. A. Lambert. Estimation of parameters in the four-parameter lognormal distribution. Australian Journal of Statistics, 12(1):33–43, April 1970. CODEN AUJSA3. ISSN 0004-9581.

Enns:1970:SSP

[242] E. G. Enns. A stochastic superposition process and an integral inequality for distributions with monotone hazard rates. Australian Journal of Statistics, 12(1):44–49, April 1970. CODEN AUJSA3. ISSN 0004-9581.

Day:1970:NSS

[243] N. E. Day. A numerical study of some sequential designs. Australian Journal of Statistics, 12(1):50–57, April 1970. CODEN AUJSA3. ISSN 0004-9581.

Kshirsagar:1970:DNM

[244] A. M. Kshirsagar and Y. P. Gupta. Distribution of the number of Markovian renewals in an arbitrary interval. Australian Journal of Statistics, 12(1):58–63, April 1970. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1970:NNa

[245] Anonymous. News and notes. Australian Journal of Statistics, 12(1): 64–65, April 1970. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1970:BRa

[246] Anonymous. Book reviews. Australian Journal of Statistics, 12(1):66–71, April 1970. CODEN AUJSA3. ISSN 0004-9581.

John:1970:TQF

[247] Peter W. M. John. A three-quarter fraction of the 4^{3} design. Australian Journal of Statistics, 12(2):73–77, August 1970. CODEN AUJSA3. ISSN 0004-9581.

Pollard:1970:SHM

[248] A. H. Pollard. Some hypothetical models in systems of tertiary education. Australian Journal of Statistics, 12(2):78–91, August 1970. CODEN AUJSA3. ISSN 0004-9581.

Pollard:1970:MRP

[249] A. H. Pollard. Minimum rate of progress formulae at universities. Australian Journal of Statistics, 12(2):92–103, August 1970. CODEN AU-JSA3. ISSN 0004-9581.

Grassia:1970:OAM

[250] A. Grassia. Optimum allocation in multi-way classifications. Australian Journal of Statistics, 12(2):104–111, August 1970. CODEN AUJSA3. ISSN 0004-9581.

Heiny:1970:EPN

[251] Robert L. Heiny and M. M. Siddiqui. Estimation of the parameters of a normal distribution when the mean is restricted to an interval. Australian Journal of Statistics, 12(2):112–117, August 1970. CODEN AUJSA3. ISSN 0004-9581.

Hoa:1970:NEC

[252] Tran Van Hoa. Notes on the estimation of the CES production function by planar approximations. Australian Journal of Statistics, 12(2):118– 123, August 1970. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1970:CRT

[253] Anonymous. Corrigenda: A regression-type estimator in sampling with PPS and with replacement. Australian Journal of Statistics, 12(2):124, August 1970. CODEN AUJSA3. ISSN 0004-9581. See [231].

Anonymous:1970:BRb

[254] Anonymous. Book reviews. Australian Journal of Statistics, 12(2):125– 130, August 1970. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1970:NNb

[255] Anonymous. News and notes. Australian Journal of Statistics, 12(2): 131, August 1970. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1970:SSA

[256] Anonymous. The Statistical Society of Australia: Annual report of the Central Council for 1969. Australian Journal of Statistics, 12(2):132–135, August 1970. CODEN AUJSA3. ISSN 0004-9581.

McGilchrist:1970:GIL

[257] C. A. McGilchrist and A. G. L. Elliott. Generalized inverses in linear regression. Australian Journal of Statistics, 12(3):137–140, November 1970. CODEN AUJSA3. ISSN 0004-9581.

Roughan:1970:ETS

[258] J. L. Roughan and H. H. Evans. Editing time series. Australian Journal of Statistics, 12(3):141–149, November 1970. CODEN AUJSA3. ISSN 0004-9581.

Hasofer:1970:DUR

[259] A. M. Hasofer. On the derivative and the upcrossings of the Rayleigh process. Australian Journal of Statistics, 12(3):150–151, November 1970. CODEN AUJSA3. ISSN 0004-9581.

Naik:1970:BCO

[260] Umesh D. Naik. Bayesian comparisons of the origins of translated exponential distributions. Australian Journal of Statistics, 12(3):152–159, November 1970. CODEN AUJSA3. ISSN 0004-9581.

Wallis:1970:CPB

[261] W. D. Wallis. A constuction for partially balanced incomplete block designs. Australian Journal of Statistics, 12(3):160–161, November 1970. CODEN AUJSA3. ISSN 0004-9581.

Griffiths:1970:PDS

[262] R. C. Griffiths. Positive definite sequences and canonical correlation coefficients. Australian Journal of Statistics, 12(3):162–165, November 1970. CODEN AUJSA3. ISSN 0004-9581.

Kulshreshtha:1970:NWD

[263] A. C. Kulshreshtha and A. Dey. A new weighing design. Australian Journal of Statistics, 12(3):166–168, November 1970. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1970:BRc

[264] Anonymous. Book reviews. Australian Journal of Statistics, 12(3):169– 176, November 1970. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1970:NNc

[265] Anonymous. News and notes. Australian Journal of Statistics, 12(3): 176, November 1970. CODEN AUJSA3. ISSN 0004-9581.

Johnson:1971:CST

[266] N. L. Johnson. Comparison of some tests of sample censoring of extreme values. Australian Journal of Statistics, 13(1):1–6, April 1971. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1971:EVC

[267] Anonymous. The effects of variations in the components of change on the size and structure of the Australian population: 1911–66. Australian Journal of Statistics, 13(1):7–18, April 1971. CODEN AUJSA3. ISSN 0004-9581.

Kakwani:1971:BTS

[268] K. C. Kakwani. The bias of the three-stage least-squares estimators in simultaneous equations. Australian Journal of Statistics, 13(1):19–26, April 1971. CODEN AUJSA3. ISSN 0004-9581.

Griffiths:1971:OPM

[269] B. C. Griffiths. Orthogonal polynomials on the multinomial distribution. Australian Journal of Statistics, 13(1):27–35, April 1971. CODEN AUJSA3. ISSN 0004-9581. See corrigenda [330].

Lancaster:1971:MDI

[270] H. O. Lancaster. The multiplicative definition of interaction. Australian Journal of Statistics, 13(1):36–44, April 1971. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1971:NNa

[271] Anonymous. News and notes. Australian Journal of Statistics, 13(1): 45–46, April 1971. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1971:BRa

[272] Anonymous. Book reviews. Australian Journal of Statistics, 13(1):47–56, April 1971. CODEN AUJSA3. ISSN 0004-9581.

Salmond:1971:EPM

[273] C. E. Salmond. The estimation of parameters of a model which are implicitly defined from their properties. Australian Journal of Statistics, 13(2):57–63, August 1971. CODEN AUJSA3. ISSN 0004-9581.

Preston:1971:EMD

[274] P. P. Preston. Estimating the mixing distribution by piece-wise polynomial arcs. Australian Journal of Statistics, 13(2):64–76, August 1971. CODEN AUJSA3. ISSN 0004-9581.

Begier:1971:CBN

[275] Mary H. Begier and M. A. Hamdan. Correlation in a bivariate normal distribution with truncation in both variables. Australian Journal of Statistics, 13(2):77–82, August 1971. CODEN AUJSA3. ISSN 0004-9581.

Hunter:1971:FST

[276] Jeffrey J. Hunter. Further studies on two queues in parallel. Australian Journal of Statistics, 13(2):83–93, August 1971. CODEN AUJSA3. ISSN 0004-9581.

Cowan:1971:RNO

[277] Richard J. Cowan. A road with no overtaking. Australian Journal of Statistics, 13(2):94–116, August 1971. CODEN AUJSA3. ISSN 0004-

Anonymous:1971:Oa

[278] Anonymous. Obituary. Australian Journal of Statistics, 13(2):117, August 1971. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1971:NNb

[279] Anonymous. News and notes. Australian Journal of Statistics, 13(2): 118, August 1971. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1971:BRb

[280] Anonymous. Book reviews. Australian Journal of Statistics, 13(2):119– 130, August 1971. CODEN AUJSA3. ISSN 0004-9581.

Ahuja:1971:CPS

[281] J. C. Ahuja. Certain properties of the Stirling distribution of the second kind. Australian Journal of Statistics, 13(3):133–136, November 1971. CODEN AUJSA3. ISSN 0004-9581.

Dey:1971:SCB

[282] A. Dey. On some chemical balance weighing designs. Australian Journal of Statistics, 13(3):137–141, November 1971. CODEN AUJSA3. ISSN 0004-9581.

Tiku:1971:SDU

[283] M. L. Tiku. Student’s distribution under non-normal situations. Australian Journal of Statistics, 13(3):142–148, November 1971. CODEN AUJSA3. ISSN 0004-9581.

Moean:1971:AEF

[284] P. A. P. Moean. Analysing an experiment to find the causes of variability of a treatment effect. Australian Journal of Statistics, 13(3):149–153, November 1971. CODEN AUJSA3. ISSN 0004-9581.

Hamdan:1971:MLE

[285] M. A. Hamdan and E. O. Martinson. Maximum likelihood estimation in the bivariate binomial (0, 1) distribution:application to 2 × 2 tables. Australian Journal of Statistics, 13(3):154–158, November 1971. CODEN AUJSA3. ISSN 0004-9581.

Pakes:1971:TQS

[286] A. G. Pakes. On a theorem of Quine and Seneta for the Galton–Watson process with immigration. Australian Journal of Statistics, 13(3):159– 164, November 1971. CODEN AUJSA3. ISSN 0004-9581.

Finch:1971:SPA

[287] P. D. Finch. A storage problem along a production line of discrete flow. Australian Journal of Statistics, 13(3):165–167, November 1971. CODEN AUJSA3. ISSN 0004-9581.

Dhrymes:1971:SSE

[288] Phoebus J. Dhrymes. A simplified structural estimator for large-scale econometric models. Australian Journal of Statistics, 13(3):168–175, November 1971. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1971:Ob

[289] Anonymous. Obituary. Australian Journal of Statistics, 13(3):176, November 1971. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1971:BRc

[290] Anonymous. Book reviews. Australian Journal of Statistics, 13(3):177– 187, November 1971. CODEN AUJSA3. ISSN 0004-9581.

Heyde:1972:MCP

[291] C. C. Heyde. Martingales: a case for a place in the statistician’s repertoire. Australian Journal of Statistics, 14(1):1–9, April 1972. CODEN AUJSA3. ISSN 0004-9581.

Jensen:1972:LFN

[292] D. B. Jensen. The limiting form of the non-central Wishart distribution. Australian Journal of Statistics, 14(1):10–16, April 1972. CODEN AUJSA3. ISSN 0004-9581.

Sugiyama:1972:ADL

[293] T. Sugiyama. Approximation for the distribution of the largest latent root of a Wishart matrix. Australian Journal of Statistics, 14(1):17–24, April 1972. CODEN AUJSA3. ISSN 0004-9581.

Baktoe:1972:CCM

[294] B. L. Baktoe and W. T. Federer. Construction of confounded mixed factorial and mixed lattice designs. Australian Journal of Statistics, 14 (1):25–36, April 1972. CODEN AUJSA3. ISSN 0004-9581.

Tiku:1972:NDD

[295] M. L. Tiku. A note on the distribution of the doubly non-central F distribution. Australian Journal of Statistics, 14(1):37–40, April 1972. CODEN AUJSA3. ISSN 0004-9581.

Sarndal:1972:PSS

[296] Carl-Erik Särndal. Properties of some simple functions of the likelihood ratio. Australian Journal of Statistics, 14(1):41–49, April 1972. CODEN AUJSA3. ISSN 0004-9581.

Ahuja:1972:CPD

[297] J. C. Ahuja and E. Enneking. Certain properties of the distribution of the sum of independent zero-one-truncated Poisson variables. Australian Journal of Statistics, 14(1):50–53, April 1972. CODEN AUJSA3. ISSN 0004-9581.

Bao:1972:RES

[298] P. V. Bao. Robust estimation for a simple exponential model. Australian Journal of Statistics, 14(1):54–62, April 1972. CODEN AUJSA3. ISSN 0004-9581.

Kerr:1972:LSA

[299] J. D. Kerr. Least squares analysis of freight-rate anomalies. Australian Journal of Statistics, 14(1):63–67, April 1972. CODEN AUJSA3. ISSN 0004-9581.

McLachlan:1972:AEV

[300] G. J. McLachlan. An asymptotic expansion for the variance of the errors of misclassification of the linear discriminant function. Australian Journal of Statistics, 14(1):68–72, April 1972. CODEN AUJSA3. ISSN 0004-9581.

Francis:1972:SPC

[301] Richard L. Francis and Howard D. Meeks. On saddle point conditions and the generalized Neyman–Pearson problem. Australian Journal of Statistics, 14(1):73–78, April 1972. CODEN AUJSA3. ISSN 0004-9581.

Enns:1972:DNR

[302] E. G. Enns. The distributions of a number of random variables associated with the absorbing Markov chain. Australian Journal of Statistics, 14 (1):79–83, April 1972. CODEN AUJSA3. ISSN 0004-9581.

Waikar:1972:EDF

[303] V. B. Waikar, T. C. Chang, and P. R. Krishnaiah. Exact distributions of a few arbitrary roots of some complex random matrices. Australian Journal of Statistics, 14(1):84–88, April 1972. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1972:NNa

[304] Anonymous. News and notes. Australian Journal of Statistics, 14(1): 89–91, April 1972. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1972:BRa

[305] Anonymous. Book reviews. Australian Journal of Statistics, 14(1):92–96, April 1972. CODEN AUJSA3. ISSN 0004-9581.

Nath:1972:MLT

[306] G. Baikunth Nath. Moments of a linearly truncated bivariate normal distribution. Australian Journal of Statistics, 14(2):97–102, August 1972. CODEN AUJSA3. ISSN 0004-9581.

Tiku:1972:DOS

[307] M. L. Tiku and H. J. Malik. On the distributions of order statistics. Australian Journal of Statistics, 14(2):103–108, August 1972. CODEN AUJSA3. ISSN 0004-9581.

Avadhani:1972:SSS

[308] M. S. Avadhani and B. V. Sukhatme. Sampling on several successive occasions with equal and unequal probabilities and without replacement. Australian Journal of Statistics, 14(2):109–119, August 1972. CODEN AUJSA3. ISSN 0004-9581.

Singh:1972:NST

[309] Ravindra Singh. A note on sampling over two occasions. Australian Journal of Statistics, 14(2):120–122, August 1972. CODEN AUJSA3. ISSN 0004-9581.

Gupta:1972:SPT

[310] P. C. Gupta. On some product type estimators in varying probability. Australian Journal of Statistics, 14(2):123–128, August 1972. CODEN AUJSA3. ISSN 0004-9581.

Milliken:1972:IML

[311] George A. Milliken and Franklin A. Graybill. Interaction models for the Latin square. Australian Journal of Statistics, 14(2):129–138, August 1972. CODEN AUJSA3. ISSN 0004-9581.

Banerjee:1972:JWC

[312] K. S. Banerjee. In justification of Wald’s criterion of efficiency of experimental designs in relation to the weighing problem. Australian Journal of Statistics, 14(2):139–142, August 1972. CODEN AUJSA3. ISSN 0004-

Kakwani:1972:RFC

[313] N. C. Kakwani and R. H. Court. Reduced form coefficient estimation and forecasting from a simultaneous equation model*. Australian Journal of Statistics, 14(2):143–160, August 1972. CODEN AUJSA3. ISSN 0004-9581.

Gipps:1972:EPS

[314] P. G. Gipps and Susan M. Smart. An examination of the power of Scheffé’s solution to the Fisher–Behrens problem. Australian Journal of Statistics, 14(2):161–166, August 1972. CODEN AUJSA3. ISSN 0004-9581.

Fraser:1972:BFP

[315] D. A. S. Fraser and F. Streit. On the Behrens–Fisher problem. Australian Journal of Statistics, 14(2):167–171, August 1972. CODEN AUJSA3. ISSN 0004-9581.

Heathcote:1972:TGF

[316] C. E. Heathcote. A test of goodness of fit for symmetric random variables. Australian Journal of Statistics, 14(2):172–181, August 1972. CODEN AUJSA3. ISSN 0004-9581.

Griffiths:1972:LDB

[317] E. C. Griffiths. Linear dependence in bivariate distributions. Australian Journal of Statistics, 14(2):182–187, August 1972. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1972:NNb

[318] Anonymous. News and notes. Australian Journal of Statistics, 14(2): 188–189, August 1972. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1972:BRb

[319] Anonymous. Book reviews. Australian Journal of Statistics, 14(2):190– 195, August 1972. CODEN AUJSA3. ISSN 0004-9581.

Ware:1972:CPS

[320] H. Ware and J. C. Caldwell. Confidentiality, privacy and sensitivity in household surveys. Australian Journal of Statistics, 14(3):197–203, November 1972. CODEN AUJSA3. ISSN 0004-9581.

Lambert:1972:LRD

[321] J. A. Lambert. Likelihoods with rounded data. Australian Journal of Statistics, 14(3):204–210, November 1972. CODEN AUJSA3. ISSN 0004-

Dudewicz:1972:CIP

[322] Edward J. Dudewicz. Confidence intervals for power, with special reference to medical trials. Australian Journal of Statistics, 14(3):211–216, November 1972. CODEN AUJSA3. ISSN 0004-9581.

Dent:1972:NTP

[323] Warren Dent. A note on transition probability estimation. Australian Journal of Statistics, 14(3):217–221, November 1972. CODEN AUJSA3. ISSN 0004-9581.

Hirai:1972:ESL

[324] Abdul Samad Hirai. Estimation of scale and location parameters, using quadratic coefficients. Australian Journal of Statistics, 14(3):222–226, November 1972. CODEN AUJSA3. ISSN 0004-9581.

Rao:1972:HTR

[325] T. J. Rao. Horvitz–Thompson and Des Raj estimators revisited. Australian Journal of Statistics, 14(3):227–230, November 1972. CODEN AUJSA3. ISSN 0004-9581.

Brewer:1972:SSS

[326] K. R. W. Brewer, L. J. Early, and S. F. Joyce. Selecting several samples from a single population. Australian Journal of Statistics, 14(3):231–239, November 1972. CODEN AUJSA3. ISSN 0004-9581.

Podder:1972:HCF

[327] N. Podder. Household consumption of food in Australia. Australian Journal of Statistics, 14(3):240–255, November 1972. CODEN AUJSA3. ISSN 0004-9581.

Hannan:1972:ESC

[328] E. J. Hannan and P. A. P. Moran. The effects of serial correlation on a randomized rain-making experiment. Australian Journal of Statistics, 14(3):256–261, November 1972. CODEN AUJSA3. ISSN 0004-9581.

Nioholls:1972:HEA

[329] D. F. Nioholls. On Hannan’s estimation of ARMA models. Australian Journal of Statistics, 14(3):262–269, November 1972. CODEN AUJSA3. ISSN 0004-9581.

Griffiths:1972:COP

[330] R. C. Griffiths. Corrigenda: Orthogonal polynomials on the multinomial distribution. Australian Journal of Statistics, 14(3):270, November 1972. CODEN AUJSA3. ISSN 0004-9581. See [269].

Anonymous:1972:NNc

[331] Anonymous. News and notes. Australian Journal of Statistics, 14(3): 271, November 1972. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1972:BRc

[332] Anonymous. Book reviews. Australian Journal of Statistics, 14(3):272– 276, November 1972. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1973:BVS

[333] H. O. Lancaster. Bibliography of vital statistics in Australia: a second list. Australian Journal of Statistics, 15(1):1–25, August 1973. CODEN AUJSA3. ISSN 0004-9581. See also first and third lists [101, 685].

Anonymous:1973:COB

[334] Anonymous. Corrigenda to original bibliography. Australian Journal of Statistics, 15(1):26, August 1973. CODEN AUJSA3. ISSN 0004-9581.

Yeo:1973:MCC

[335] G. F. Yeo. On a model of cell colony formation. Australian Journal of Statistics, 15(1):27–34, August 1973. CODEN AUJSA3. ISSN 0004-9581.

Reynolds:1973:EPB

[336] John F. Reynolds. On estimating the parameters of a birth-death process. Australian Journal of Statistics, 15(1):35–43, August 1973. CODEN AUJSA3. ISSN 0004-9581.

Chakrabarti:1973:NSS

[337] C. G. Chakrabarti. A note on the significance of sufficiency in thermodynamics. Australian Journal of Statistics, 15(1):44–48, August 1973. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1973:NNa

[338] Anonymous. News and notes. Australian Journal of Statistics, 15(1): 49–50, August 1973. CODEN AUJSA3. ISSN 0004-9581.

Vere-Jones:1973:BR

[339] D. Vere-Jones. Book reviews. Australian Journal of Statistics, 15(1): 51–64, August 1973. CODEN AUJSA3. ISSN 0004-9581.

Jensen:1973:MBC

[340] D. R. Jensen. Monotone bounds on the chi-squared approximation to the distribution of Pearson’s χ statistics. Australian Journal of Statistics, 15 (2):65–70, November 1973. CODEN AUJSA3. ISSN 0004-9581.

John:1973:CVI

[341] S. John. Critical values for inference about normal dispersion. Australian Journal of Statistics, 15(2):71–79, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Billard:1973:PTM

[342] L. Billard. A procedure for testing the mean of a normal distribution. Australian Journal of Statistics, 15(2):80–86, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Das:1973:PSR

[343] M. N. Das and S. Mohanty. On PPS sampling without replacement ensuring selection probabilities exactly proportional to sizes. Australian Journal of Statistics, 15(2):87–94, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Rao:1973:CES

[344] J. N. K. Rao and M. P. Singh. On the choice of estimator in survey sampling. Australian Journal of Statistics, 15(2):95–104, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Sen:1973:STS

[345] A. R. Sen. Some theory of sampling on successive occasions. Australian Journal of Statistics, 15(2):105–110, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Powell:1973:ELE

[346] Alan Powell. Estimation of Lluch’s extended linear expenditure system from cross-sectional data. Australian Journal of Statistics, 15(2):111– 117, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Praetz:1973:GRE

[347] P. D. Praetz. Growth rate estimators for economic time series. Australian Journal of Statistics, 15(2):118–127, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Boswell:1973:CCD

[348] M. T. Boswell and G. P. Patil. Characterization of certain discrete distributions by differential equations with respect to their parameters. Australian Journal of Statistics, 15(2):128–131, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Nath:1973:WTC

[349] Harmindar B. Nath. Waiting time in the Coupon-Collector’s Problem. Australian Journal of Statistics, 15(2):132–135, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Sherman:1973:LE

[350] Douglas Sherman. Letter to the Editor. Australian Journal of Statistics, 15(2):136, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1973:NNb

[351] Anonymous. News and notes. Australian Journal of Statistics, 15(2): 137, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1973:BRa

[352] Anonymous. Book reviews. Australian Journal of Statistics, 15(2):137– 144, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Brewer:1973:ESS

[353] K. R. W. Brewer and R. W. Mellor. The effect of sample structure on analytical surveys. Australian Journal of Statistics, 15(3):145–152, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Young:1973:ELS

[354] C. M. Young. Editing a large-scale demographic survey — the experience of the Melbourne Family Survey. Australian Journal of Statistics, 15(3): 153–166, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Williams:1973:VWD

[355] Ross A. Williams. Variable weight distributed lags: Housing completions in Australia. Australian Journal of Statistics, 15(3):167–179, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Miller:1973:SCS

[356] Kenneth S. Miller. Serial correlation and spectral moment statistics. Australian Journal of Statistics, 15(3):180–184, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Calitz:1973:MLE

[357] Fred Calitz. Maximum likelihood estimation of the parameters of the three-parameter lognormal distribution — a reconsideration. Australian Journal of Statistics, 15(3):185–190, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Ferguson:1973:CTN

[358] Nancy L. Ferguson. Comparison of three non-parametric tests for bivariate interchangeability. Australian Journal of Statistics, 15(3):191–209, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Mclachlan:1973:AEE

[359] G. J. Mclachlan. An asymptotic expansion of the expectation of the estimated error rate in discriminant analysis. Australian Journal of Statistics, 15(3):210–214, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Hasofer:1973:JDE

[360] A. M. Hasofer. The joint distribution of the envelope of a Gaussian process and its derivative. Australian Journal of Statistics, 15(3):215– 216, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Hoa:1973:CNE

[361] Tran Van Hoa. Corrigenda: Notes on the estimation of the CES production function by planar approximations. Australian Journal of Statistics, 15(3):217, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Kulkarni:1973:COA

[362] S. R. Kulkarni. Corrigenda: On the optimal asymptotic tests for the effects of cloud seeding on rainfall. (2) The case of variable effects. Australian Journal of Statistics, 15(3):217, November 1973. CODEN AU-JSA3. ISSN 0004-9581. See [216].

Anonymous:1973:OEA

[363] Anonymous. Obituary: Edmund Alfred Cornish (1909–1973). Australian Journal of Statistics, 15(3):218–220, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1973:NNc

[364] Anonymous. News and notes. Australian Journal of Statistics, 15(3): 221–222, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1973:BRb

[365] Anonymous. Book reviews. Australian Journal of Statistics, 15(3):223– 231, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1973:PR

[366] Anonymous. Publications received. Australian Journal of Statistics, 15 (3):232, November 1973. CODEN AUJSA3. ISSN 0004-9581.

Khatri:1974:TEL

[367] C. G. Khatri. On testing the equality of location parameters in k censored exponential distributions. Australian Journal of Statistics, 16(1):1–10, April 1974. CODEN AUJSA3. ISSN 0004-9581.

Thomas:1974:STI

[368] David Thomas and Nancy Ferguson. Some tests for independence. Australian Journal of Statistics, 16(1):11–19, April 1974. CODEN AUJSA3. ISSN 0004-9581.

Searle:1974:NSG

[369] G. C. L. Searle. A note on a sequential group t-test. Australian Journal of Statistics, 16(1):20–25, April 1974. CODEN AUJSA3. ISSN 0004-9581.

Kelleher:1974:ETS

[370] Grace J. Kelleher. Exact two-sample median tests when one “sample” value possibly from a different population. Australian Journal of Statistics, 16(1):26–29, April 1974. CODEN AUJSA3. ISSN 0004-9581.

Lingappaiah:1974:PSG

[371] G. S. Lingappaiah. Prediction in samples from the gamma distribution as applied to life testing. Australian Journal of Statistics, 16(1):30–32, April 1974. CODEN AUJSA3. ISSN 0004-9581.

Basawa:1974:MLE

[372] I. V. Basawa. Maximum likelihood estimation of parameters in renewal and Markov-renewal processes. Australian Journal of Statistics, 16(1): 33–43, April 1974. CODEN AUJSA3. ISSN 0004-9581.

Osborn:1974:DPC

[373] Denise R. Osborn. The distribution of price changes on the Sydney Stock Exchange. Australian Journal of Statistics, 16(1):44–49, April 1974. CO-DEN AUJSA3. ISSN 0004-9581.

Nigam:1974:NOC

[374] A. K. Nigam. A note on optimum chemical balance weighing designs. Australian Journal of Statistics, 16(1):50–52, April 1974. CODEN AU-JSA3. ISSN 0004-9581.

Griffiths:1974:CMD

[375] R. C. Griffiths. A characterization of the multinomial distribution. Australian Journal of Statistics, 16(1):53–56, April 1974. CODEN AUJSA3. ISSN 0004-9581.

Sandland:1974:NSA

[376] R. L. Sandland. A note on some applications of the truncated geometric distribution. Australian Journal of Statistics, 16(1):57–58, April 1974. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1974:NNa

[377] Anonymous. News and notes. Australian Journal of Statistics, 16(1):59, April 1974. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1974:BRa

[378] Anonymous. Book reviews. Australian Journal of Statistics, 16(1):60–70, April 1974. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1974:CHW

[379] H. O. Lancaster. Charles Henry Wickens: 1872–1939. Australian Journal of Statistics, 16(2):71–82, August 1974. CODEN AUJSA3. ISSN 0004-9581.

Shirahata:1974:TSD

[380] Shingo Shirahata. On tests of symmetry for discrete populations. Australian Journal of Statistics, 16(2):83–90, August 1974. CODEN AU-JSA3. ISSN 0004-9581.

Sowey:1974:PRR

[381] E. R. Sowey. A priori restrictions in reduced form estimation: a Monte Carlo comparison of alternative estimators. Australian Journal of Statistics, 16(2):91–105, August 1974. CODEN AUJSA3. ISSN 0004-9581.

Thornett:1974:MMS

[382] M. L. Thornett. Minimum mean square error unbiased estimation. Australian Journal of Statistics, 16(2):106–107, August 1974. CODEN AU-JSA3. ISSN 0004-9581.

Gupta:1974:CMB

[383] R. D. Gupta. On a certain multivariate beta distribution. Australian Journal of Statistics, 16(2):108–110, August 1974. CODEN AUJSA3. ISSN 0004-9581.

Miles:1974:FFB

[384] R. E. Miles. The fundamental formula of Blaschke in integral geometry and geometrical probability, and its iteration, for domains with fixed orientations. Australian Journal of Statistics, 16(2):111–118, August 1974. CODEN AUJSA3. ISSN 0004-9581. See addendum [477].

Shanbhag:1974:SCB

[385] D. N. Shanbhag and G. Rajamannar. Some characterizations of the bivariate distribution of independent Poisson variables. Australian Journal of Statistics, 16(2):119–125, August 1974. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1974:NNb

[386] Anonymous. News and notes. Australian Journal of Statistics, 16(2): 126, August 1974. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1974:BRb

[387] Anonymous. Book reviews. Australian Journal of Statistics, 16(2):127– 132, August 1974. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1974:BNa

[388] Anonymous. Book notices. Australian Journal of Statistics, 16(2):132– 134, August 1974. CODEN AUJSA3. ISSN 0004-9581.

Maritz:1974:AEA

[389] J. S. Maritz. Aligning of estimates: an alternative to empirical Bayes methods. Australian Journal of Statistics, 16(3):135–143, November 1974. CODEN AUJSA3. ISSN 0004-9581.

Chiu:1974:SPL

[390] W. K. Chiu. Selecting the m populations with largest means from k normal populations with unknown variances. Australian Journal of Statistics, 16(3):144–147, November 1974. CODEN AUJSA3. ISSN 0004-9581. See corrigendum [421].

Phadia:1974:BIC

[391] E. G. Phadia. Best invariant confidence bands for a continuous cumulative distribution function. Australian Journal of Statistics, 16(3):148– 152, November 1974. CODEN AUJSA3. ISSN 0004-9581.

Johnson:1974:MCR

[392] Dallas E. Johnson. On the moments of the characteristic roots of a noncentral (2 × 2) Wishart matrix. Australian Journal of Statistics, 16 (3):153–155, November 1974. CODEN AUJSA3. ISSN 0004-9581.

Pliska:1974:DPM

[393] Stanley R. Pliska. A dynamic programming model for optimal observations of a discrete time linear stochastic process. Australian Journal of Statistics, 16(3):156–162, November 1974. CODEN AUJSA3. ISSN 0004-9581.

Beg:1974:CER

[394] M. I. Beg and S. N. U. A. Kirmani. On a characterization of exponential and related distributions. Australian Journal of Statistics, 16(3):163–166, November 1974. CODEN AUJSA3. ISSN 0004-9581.

Rives:1974:NDI

[395] Norfleet W. Rives, Jr. A note on the demographic impact of contraception in a mathematical model of human reproduction. Australian Journal of Statistics, 16(3):167–173, November 1974. CODEN AUJSA3. ISSN 0004-9581.

Binet:1974:LE

[396] F. E. Binet. Letter to the Editor. Australian Journal of Statistics, 16(3): 174, November 1974. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1974:NNc

[397] Anonymous. News and notes. Australian Journal of Statistics, 16(3): 175, November 1974. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1974:BRc

[398] Anonymous. Book reviews. Australian Journal of Statistics, 16(3):176– 186, November 1974. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1974:BNb

[399] Anonymous. Book notices. Australian Journal of Statistics, 16(3):187, November 1974. CODEN AUJSA3. ISSN 0004-9581.

Bofinger:1975:EDF

[400] Eve Bofinger. Estimation of a density function using order statistics. Australian Journal of Statistics, 17(1):1–7, March 1975. CODEN AU-JSA3. ISSN 0004-9581.

Quine:1975:RCS

[401] M. P. Quine and J. Robinson. Rates of convergence for some statistics from empirical distribution functions. Australian Journal of Statistics, 17(1):8–11, March 1975. CODEN AUJSA3. ISSN 0004-9581.

Singh:1975:NOS

[402] Ravindra Singh. A note on optimum stratification in sampling with varying probabilities. Australian Journal of Statistics, 17(1):12–21, March 1975. CODEN AUJSA3. ISSN 0004-9581.

Chaudhubi:1975:SPE

[403] Arijit Chaudhubi. Some properties of estimators based on sampling schemes with varying probabilities. Australian Journal of Statistics, 17 (1):22–28, March 1975. CODEN AUJSA3. ISSN 0004-9581.

Best:1975:DBT

[404] D. J. Best. The difference between two Poisson expectations. Australian Journal of Statistics, 17(1):29–33, March 1975. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1975:MDI

[405] H. O. Lancaster. The multiplicative definition of interaction an addendum. Australian Journal of Statistics, 17(1):34–36, March 1975. CODEN AUJSA3. ISSN 0004-9581.

Aggaewal:1975:AEH

[406] K. E. Aggaewal. Analysis of EGD and hypercubic designs as factorial experiments. Australian Journal of Statistics, 17(1):36–42, March 1975. CODEN AUJSA3. ISSN 0004-9581.

Sahai:1975:SFB

[407] Hardeo Sahai and Alcedo A. Ramírez Martínez. Some formal Bayes estimators of variance components in a crossed classification random effects model. Australian Journal of Statistics, 17(1):43–48, March 1975. CO-DEN AUJSA3. ISSN 0004-9581.

Kageyama:1975:NBD

[408] Sanpei Kageyama and Goeo Ishu. A note on bib designs. Australian Journal of Statistics, 17(1):49–50, March 1975. CODEN AUJSA3. ISSN 0004-9581.

Preece:1975:BDE

[409] D. A. Preece. Bibliography of designs for experiments in three dimensions. Australian Journal of Statistics, 17(1):51–55, March 1975. CODEN AUJSA3. ISSN 0004-9581.

Williams:1975:OMH

[410] B. J. Williams. Obituary: Maurice H. Belz (1897–1975). Australian Journal of Statistics, 17(1):56–57, March 1975. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1975:NNa

[411] Anonymous. News and notes. Australian Journal of Statistics, 17(1):58, March 1975. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1975:BRa

[412] Anonymous. Book reviews. Australian Journal of Statistics, 17(1):59–62,

March 1975. CODEN AUJSA3. ISSN 0004-9581.

Pollard:1975:MHP

[413] J. H. Pollard. Modelling human populations for projection purposes — some of the problems and challenges. Australian Journal of Statistics, 17(2):63–76, June 1975. CODEN AUJSA3. ISSN 0004-9581.

Chhikara:1975:OTC

[414] Raj S. Chhikara. Optimum tests for comparison of two inverse Gaussian distribution means. Australian Journal of Statistics, 17(2):77–83, June 1975. CODEN AUJSA3. ISSN 0004-9581.

Hamdan:1975:PMC

[415] M. A. Hamdan and R. S. Schulman. On point multiserial correlation. Australian Journal of Statistics, 17(2):84–86, June 1975. CODEN AU-JSA3. ISSN 0004-9581.

Anderson:1975:PDP

[416] O. D. Anderson. On a paper by Davies, Pate and Frost concerning maximum autocorrelations for moving average processes. Australian Journal of Statistics, 17(2):87–90, June 1975. CODEN AUJSA3. ISSN 0004-9581.

Daley:1975:CBC

[417] D. J. Daley. Certain bounds concerning characteristic functions. Australian Journal of Statistics, 17(2):91–93, June 1975. CODEN AUJSA3. ISSN 0004-9581.

Barnett:1975:NGF

[418] N. S. Barnett. A note on a generating function relation. Australian Journal of Statistics, 17(2):94–95, June 1975. CODEN AUJSA3. ISSN 0004-9581.

Tweedie:1975:RBE

[419] Richard L. Tweedie. Relations between ergodicity and mean drift for Markov chains. Australian Journal of Statistics, 17(2):96–102, June 1975. CODEN AUJSA3. ISSN 0004-9581.

Gleeson:1975:QSE

[420] A. C. Gleeson and J. B. Douglas. Quadrat sampling and the estimation of Neyman Type A and Thomas distributional parameters. Australian Journal of Statistics, 17(2):103–113, June 1975. CODEN AUJSA3. ISSN 0004-9581.

Chiu:1975:CSP

[421] W. K. Chiu. Corrigendum: Selecting the m populations with largest means from k normal populations with unknown variances. Australian Journal of Statistics, 17(2):114, June 1975. CODEN AUJSA3. ISSN 0004-9581. See [390].

Anonymous:1975:NNb

[422] Anonymous. News and notes. Australian Journal of Statistics, 17(2): 115, June 1975. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1975:BRb

[423] Anonymous. Book reviews. Australian Journal of Statistics, 17(2):116– 120, June 1975. CODEN AUJSA3. ISSN 0004-9581.

Moran:1975:SSA

[424] P. A. P. Moran. Statistical Society of Australia special lecture (October 28th, 1975): What should a Professor of Statistics do? Australian Journal of Statistics, 17(3):121–133, September 1975. CODEN AUJSA3. ISSN 0004-9581.

Veee-Jones:1975:SAE

[425] D. Veee-Jones. Statistical aspects of the estimation of earthquake risk. Australian Journal of Statistics, 17(3):134–147, September 1975. CO-DEN AUJSA3. ISSN 0004-9581.

McDonald:1975:ASD

[426] John McDonald and Peter Monk. An analysis of the statistical discrepancy in the Australian quarterly national accounts*. Australian Journal of Statistics, 17(3):148–160, September 1975. CODEN AUJSA3. ISSN 0004-9581.

McLachlan:1975:SEV

[427] G. J. McLachlan. Some expected values for the error rates of the sample quadratic discriminant function. Australian Journal of Statistics, 17(3): 161–165, September 1975. CODEN AUJSA3. ISSN 0004-9581.

Brewer:1975:SPS

[428] Kenneth E. W. Brewer. A simple procedure for sampling π pswor. Australian Journal of Statistics, 17(3):166–172, September 1975. CODEN AUJSA3. ISSN 0004-9581.

Chambers:1975:LME

[429] E. L. Chambers and C. E. Heathcote. A linear model with errors lacking a variance. Australian Journal of Statistics, 17(3):173–185, September 1975. CODEN AUJSA3. ISSN 0004-9581. See also Part II [505].

Preece:1975:BDR

[430] D. A. Preece and W. B. Hall. Balanced designs for row-and-column experiments with two non-interacting sets of treatments, one set not being applied to all the rows. Australian Journal of Statistics, 17(3): 186–191, September 1975. CODEN AUJSA3. ISSN 0004-9581.

Bofinger:1975:NPE

[431] Eve Bofinger. Non-parametric estimation of density for regularly varying distributions. Australian Journal of Statistics, 17(3):192–195, September 1975. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1975:NNc

[432] Anonymous. News and notes. Australian Journal of Statistics, 17(3): 196–198, September 1975. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1975:BRc

[433] Anonymous. Book reviews. Australian Journal of Statistics, 17(3):199– 202, September 1975. CODEN AUJSA3. ISSN 0004-9581.

Gani:1976:SAD

[434] J. Gani. Some aspects in the development of statistics in Australia. Australian Journal of Statistics, 18(1–2):1–20, March 1976. CODEN AUJSA3. ISSN 0004-9581.

Foreman:1976:WGS

[435] E. K. Foreman, W. McLennan, D. G. Leaver, and M. W. Dickinson. Why general survey systems — Australian GSS progress and prospects. Australian Journal of Statistics, 18(1–2):21–36, March 1976. CODEN AUJSA3. ISSN 0004-9581.

Holt:1976:DSP

[436] D. Holt and T. M. F. Smith. The design of surveys for planning purposes. Australian Journal of Statistics, 18(1–2):37–44, March 1976. CODEN AUJSA3. ISSN 0004-9581.

Yanagawa:1976:RSS

[437] T. Yanagawa and S. Shirahata. Ranked set sampling theory with selective probability matrix. Australian Journal of Statistics, 18(1–2):45–52, March 1976. CODEN AUJSA3. ISSN 0004-9581.

Talus:1976:RBC

[438] G. M. Talus and P. L. Chesson. The relationship between a class of asymptotically normal estimators and goodness of fit tests. Australian Journal of Statistics, 18(1–2):53–61, March 1976. CODEN AUJSA3. ISSN 0004-9581.

Beinmuth:1976:RE

[439] James E. Beinmuth. Recursive estimation. Australian Journal of Statistics, 18(1–2):62–72, March 1976. CODEN AUJSA3. ISSN 0004-9581.

Anderson:1976:IAI

[440] O. D. Anderson. An improved approach to inverting the autocovariance matrix of a general mixed autoregressive moving average time process. Australian Journal of Statistics, 18(1–2):73–75, March 1976. CODEN AUJSA3. ISSN 0004-9581.

Wuhees:1976:CEP

[441] C. S. Wuhees. On the convergence of empirical processes of mixing variables. Australian Journal of Statistics, 18(1–2):76–83, March 1976. CO-DEN AUJSA3. ISSN 0004-9581.

Dent:1976:C

[442] Warren Dent. Corrigenda. Australian Journal of Statistics, 18(1–2): 84–85, March 1976. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1976:NN

[443] Anonymous. News and notes. Australian Journal of Statistics, 18(1–2): 87, March 1976. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1976:BRa

[444] Anonymous. Book reviews. Australian Journal of Statistics, 18(1–2): 88–100, March 1976. CODEN AUJSA3. ISSN 0004-9581.

David:1976:SSA

[445] H. A. David. Statistical Society of Australia the Knibbs Lecture for 1976 (July 27th, 1976): The theory of competing risks. Australian Journal of Statistics, 18(3):101–110, September 1976. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1976:IDP

[446] Anonymous. Invited discussion of the paper. Australian Journal of Statistics, 18(3):110–114, September 1976. CODEN AUJSA3. ISSN 0004-9581.

Cole:1976:CPH

[447] R. W. Cole. The census of population and housing, 30 June 1976. Australian Journal of Statistics, 18(3):115–130, September 1976. CODEN AUJSA3. ISSN 0004-9581.

Mukherjee:1976:TCS

[448] Bishwa Nath Mukherjee. Techniques of covariance structural analysis. Australian Journal of Statistics, 18(3):131–150, September 1976. CO-DEN AUJSA3. ISSN 0004-9581.

Tan:1976:DFT

[449] K. C. Tan and A. P. Gore. A distribution-free test for two-way layout. Australian Journal of Statistics, 18(3):151–157, September 1976. CODEN AUJSA3. ISSN 0004-9581.

Shirakura:1976:NNA

[450] Teruhiro Shirakura. A note on the norm of alias matrices in fractional replication. Australian Journal of Statistics, 18(3):158–160, September 1976. CODEN AUJSA3. ISSN 0004-9581.

Court:1976:NTU

[451] E. H. Court. A note on Theil’s use of incomplete prior information in regression analysis. Australian Journal of Statistics, 18(3):161–162, September 1976. CODEN AUJSA3. ISSN 0004-9581.

Hamdan:1976:BBD

[452] M. A. Hamdan and D. R. Jensen. A bivariate binomial distribution and some applications. Australian Journal of Statistics, 18(3):163–169, September 1976. CODEN AUJSA3. ISSN 0004-9581.

Hawkins:1976:CPE

[453] Douglas M. Hawkins and Samuel Kotz. A clocking property of the exponential distribution. Australian Journal of Statistics, 18(3):170–172, September 1976. CODEN AUJSA3. ISSN 0004-9581.

Arnold:1976:CUD

[454] Babry C. Arnold and Glen Meeden. A characterization of the uniform distribution based on summation modulo one, with application to fractional backlogs. Australian Journal of Statistics, 18(3):173–175, September 1976. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1976:BRb

[455] Anonymous. Book reviews. Australian Journal of Statistics, 18(3):176– 188, September 1976. CODEN AUJSA3. ISSN 0004-9581.

Finch:1977:RAS

[456] Peter D. Finch. Rough analysis and spurious accuracy. Australian Journal of Statistics, 19(1):1–21, March 1977. CODEN AUJSA3. ISSN 0004-9581.

Davis:1977:USA

[457] J. R. Davis. Using the Sydney Area Transportation Study data for social research. Australian Journal of Statistics, 19(1):22–31, March 1977. CODEN AUJSA3. ISSN 0004-9581.

Milliken:1977:HTT

[458] George A. Milliken and Dale Rasmuson. A heuristic technique for testing for the presence of interaction in nonreplicated factorial experiments. Australian Journal of Statistics, 19(1):32–38, March 1977. CODEN AU-JSA3. ISSN 0004-9581.

Williams:1977:IAG

[459] E. R. Williams. Iterative analysis of generalized lattice designs. Australian Journal of Statistics, 19(1):39–42, March 1977. CODEN AUJSA3. ISSN 0004-9581.

Chen:1977:NEE

[460] Hubert J. Chen. Non-existence of an exact β-confidence interval for a scale parameter. Australian Journal of Statistics, 19(1):43–55, March 1977. CODEN AUJSA3. ISSN 0004-9581.

Perng:1977:TSP

[461] S. K. Perng and John D. Groves. Two sequential procedures for partitioning a set of one-parameter exponential populations with respect to a control. Australian Journal of Statistics, 19(1):56–60, March 1977. CODEN AUJSA3. ISSN 0004-9581.

Kaminsky:1977:BPE

[462] Kenneth S. Kaminsky. Best prediction of exponential failure times when items may be replaced. Australian Journal of Statistics, 19(1):61–62, March 1977. CODEN AUJSA3. ISSN 0004-9581.

Breth:1977:GPS

[463] M. Breth. Growth processes: a study of progressive payments on building contracts. Australian Journal of Statistics, 19(1):63–69, March 1977. CODEN AUJSA3. ISSN 0004-9581.

Nagaraja:1977:CBR

[464] H. N. Nagaraja. On a characterization based on record values. Australian Journal of Statistics, 19(1):70–73, March 1977. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1977:NN

[465] Anonymous. News and notes. Australian Journal of Statistics, 19(1):74, March 1977. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1977:BRa

[466] Anonymous. Book reviews. Australian Journal of Statistics, 19(1):75–80, March 1977. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1977:BNa

[467] Anonymous. Book notices. Australian Journal of Statistics, 19(1):80–81, March 1977. CODEN AUJSA3. ISSN 0004-9581.

Young:1977:FBD

[468] Christabel M. Young. Family building differences between same sex and mixed sex families in Australia. Australian Journal of Statistics, 19(2): 83–95, June 1977. CODEN AUJSA3. ISSN 0004-9581. Check author name??

Singh:1977:NOS

[469] Ravindra Singh. A note on optimum stratification for equal allocation with ratio and regression methods of estimation. Australian Journal of Statistics, 19(2):96–104, June 1977. CODEN AUJSA3. ISSN 0004-9581.

Singh:1977:NUI

[470] Randhir Singh. A note on the use of incomplete multi-auxiliary information in sample surveys. Australian Journal of Statistics, 19(2):105–107, June 1977. CODEN AUJSA3. ISSN 0004-9581.

Grassia:1977:FDA

[471] A. Grassia. On a family of distributions with argument between 0 and 1 obtained by transformation of the gamma and derived compound distributions. Australian Journal of Statistics, 19(2):108–114, June 1977. CODEN AUJSA3. ISSN 0004-9581.

OBryan:1977:EBE

[472] Thomas E. O’Bryan and V. Susarla. Empirical Bayes estimation with non-identical components. Continuous case. Australian Journal of Statistics, 19(2):115–125, June 1977. CODEN AUJSA3. ISSN 0004-9581.

Singla:1977:ED

[473] S. L. Singla. An extension of & designs. Australian Journal of Statistics, 19(2):126–131, June 1977. CODEN AUJSA3. ISSN 0004-9581.

Cressie:1977:MHO

[474] Noel Cressie. The minimum of higher order gaps. Australian Journal of Statistics, 19(2):132–143, June 1977. CODEN AUJSA3. ISSN 0004-9581.

Mase:1977:ASB

[475] Shigeru Mase. On the attainment of the supremum of the Bhattacharyya lower bounds for variances of estimators. Australian Journal of Statistics, 19(2):144–151, June 1977. CODEN AUJSA3. ISSN 0004-9581.

Feigin:1977:NML

[476] Paul D. Feigin. A note on maximum likelihood estimation for simple branching processes. Australian Journal of Statistics, 19(2):152–154, June 1977. CODEN AUJSA3. ISSN 0004-9581.

Flrey:1977:ARM

[477] Wm. J. Flrey. Addendum to R. E. Miles’ Paper on the Fundamental Formula of Blaschke in Integral Geometry. Australian Journal of Statistics, 19(2):155–156, June 1977. CODEN AUJSA3. ISSN 0004-9581. See [384].

Anonymous:1977:BRb

[478] Anonymous. Book reviews. Australian Journal of Statistics, 19(2):157– 162, June 1977. CODEN AUJSA3. ISSN 0004-9581.

Wei:1977:SSS

[479] Lee-Jen Wei. Sequential searching schemes for an optimal dosage. Australian Journal of Statistics, 19(3):163–171, September 1977. CODEN AUJSA3. ISSN 0004-9581.

Billard:1977:BAS

[480] L. Billard. Bounds for the average sample number function for a binomial sequential test. Australian Journal of Statistics, 19(3):172–178, September 1977. CODEN AUJSA3. ISSN 0004-9581.

Samiuddin:1977:BTH

[481] M. Samiuddin and Muhammad Hanif. Bayesian test of hypothesis in the multinomial distribution. Australian Journal of Statistics, 19(3):179–184, September 1977. CODEN AUJSA3. ISSN 0004-9581.

Wood:1977:A

[482] J. T. Wood. 466 and all that. Australian Journal of Statistics, 19(3): 185–193, September 1977. CODEN AUJSA3. ISSN 0004-9581.

Williams:1977:UBE

[483] E. R. Williams and H. D. Patterson. Upper bounds for efficiency factors in block designs. Australian Journal of Statistics, 19(3):194–201, September 1977. CODEN AUJSA3. ISSN 0004-9581.

Gupta:1977:DST

[484] A. K. Gupta. On the distribution of sphericity test criterion in the multivariate Gaussian distribution. Australian Journal of Statistics, 19 (3):202–205, September 1977. CODEN AUJSA3. ISSN 0004-9581.

Davis:1977:ATP

[485] A. W. Davis. Asymptotic theory for principal component analysis: Nonnormal case. Australian Journal of Statistics, 19(3):206–212, September 1977. CODEN AUJSA3. ISSN 0004-9581.

Davis:1977:SSS

[486] A. W. Davis. Statistical study of a stochastic logarithmic model with hereditary binary interactions. Australian Journal of Statistics, 19(3): 213–225, September 1977. CODEN AUJSA3. ISSN 0004-9581.

Shibata:1977:CLS

[487] Ritei Shibata. Convergence of least squares estimates of autoregressive parameters. Australian Journal of Statistics, 19(3):226–235, September 1977. CODEN AUJSA3. ISSN 0004-9581.

Robinson:1977:EVC

[488] P. M. Robinson. Estimating variances and covariances of sample autocorrelations and autocovariances. Australian Journal of Statistics, 19(3): 236–240, September 1977. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1977:BNb

[489] Anonymous. Book notices. Australian Journal of Statistics, 19(3):241, September 1977. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1978:SSA

[490] H. O. Lancaster. Statistical Society of Australia special lecture: (July 12th, 1977): World mortality survey. Australian Journal of Statistics, 20 (1):1–42, March 1978. CODEN AUJSA3. ISSN 0004-9581.

Cox:1978:SSA

[491] D. R. Cox. Statistical Society of Australia the Knibbs Lecture for 1977: (September 5th, 1977): Foundations of statistical inference: the case for eclecticism. Australian Journal of Statistics, 20(1):43–59, March 1978. CODEN AUJSA3. ISSN 0004-9581.

Pitman:1978:CRI

[492] E. J. G. Pitman. The Cramér–Rao inequality. Australian Journal of Statistics, 20(1):60–74, March 1978. CODEN AUJSA3. ISSN 0004-9581.

ONeill:1978:ADC

[493] M. E. O’Neill. Asymptotic distributions of the canonical correlations from contingency tables. Australian Journal of Statistics, 20(1):75–82, March 1978. CODEN AUJSA3. ISSN 0004-9581.

Chacko:1978:ETW

[494] Annie Chacko and A. Dey. On the estimation of total weight in chemical balance weighing designs. Australian Journal of Statistics, 20(1):83–86, March 1978. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1978:C

[495] Anonymous. Corrigendum. Australian Journal of Statistics, 20(1):87, March 1978. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1978:HLM

[496] Anonymous. Honorary life member. Australian Journal of Statistics, 20 (1):88, March 1978. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1978:BRa

[497] Anonymous. Book reviews. Australian Journal of Statistics, 20(1):89–97, March 1978. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1978:BRb

[498] Anonymous. Book reviews. Australian Journal of Statistics, 20(1):95–97, March 1978. CODEN AUJSA3. ISSN 0004-9581.

Federer:1978:CVD

[499] W. T. Federer and O. O. Ladipo. A component of variance due to competition. Australian Journal of Statistics, 20(2):99–110, June 1978. CODEN AUJSA3. ISSN 0004-9581.

Preece:1978:RPF

[500] D. A. Preece, R. A. Bailey, and H. D. Patterson. A randomization problem in forming designs with superimposed treatments. Australian Journal of Statistics, 20(2):111–125, June 1978. CODEN AUJSA3. ISSN 0004-9581.

McDonald:1978:RBO

[501] John McDonald. On the relationship between the optimal extrapolative and other expectations generating methods. Australian Journal of Statistics, 20(2):126–135, June 1978. CODEN AUJSA3. ISSN 0004-9581.

Mcdonnell:1978:NAO

[502] Patrick Mcdonnell, F. Lancaster Jones, and Paul Duncan-Jones. Notes on the Australian occupational classification as a social science research tool. Australian Journal of Statistics, 20(2):136–142, June 1978. CODEN AUJSA3. ISSN 0004-9581.

Horn:1978:SSS

[503] Robert V. Horn. From social statistics to social indicators. Australian Journal of Statistics, 20(2):143–152, June 1978. CODEN AUJSA3. ISSN 0004-9581.

Brown:1978:AEF

[504] B. M. Brown and D. G. Kildea. Asymptotically equivalent forms of Hodges–Lehmann-type estimates in regression. Australian Journal of Statistics, 20(2):153–160, June 1978. CODEN AUJSA3. ISSN 0004-9581.

Chambers:1978:LME

[505] R. L. Chambers and C. R. Heathcote. A linear model with errors lacking a variance II. Australian Journal of Statistics, 20(2):161–175, June 1978. CODEN AUJSA3. ISSN 0004-9581. See also Part I [429].

Nagaraja:1978:EVR

[506] H. N. Nagaraja. On the expected values of record values. Australian Journal of Statistics, 20(2):176–182, June 1978. CODEN AUJSA3. ISSN 0004-9581.

Griffiths:1978:BTD

[507] R. C. Griffiths. On a bivariate triangular distribution. Australian Journal of Statistics, 20(2):183–185, June 1978. CODEN AUJSA3. ISSN 0004-

Anonymous:1978:BRc

[508] Anonymous. Book reviews. Australian Journal of Statistics, 20(2):186– 190, June 1978. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1978:BNa

[509] Anonymous. Book notices. Australian Journal of Statistics, 20(2):190– 193, June 1978. CODEN AUJSA3. ISSN 0004-9581.

Becker:1978:MDE

[510] Niels Becker. Models and designs for experiments with mixtures. Australian Journal of Statistics, 20(3):195–208, September 1978. CODEN AUJSA3. ISSN 0004-9581.

Raktoe:1978:CCM

[511] B. L. Raktoe, A. A. Rayner, and D. O. Chalton. On construction of confounded mixed factorial and lattice designs. Australian Journal of Statistics, 20(3):209–218, September 1978. CODEN AUJSA3. ISSN 0004-9581.

Freebairn:1978:RCE

[512] J. W. Freebairn. Recursive coefficient estimates for the evaluation of varying parameters. Australian Journal of Statistics, 20(3):219–228, September 1978. CODEN AUJSA3. ISSN 0004-9581.

Bandyopadhyay:1978:CDA

[513] Shibdas Bandyopadhyay. Classification, distributions and admissibility in autoregressive process. Australian Journal of Statistics, 20(3):229–239, September 1978. CODEN AUJSA3. ISSN 0004-9581.

Speed:1978:FTA

[514] T. P. Speed. A factorisation theorem for adequate statistics. Australian Journal of Statistics, 20(3):240–249, September 1978. CODEN AUJSA3. ISSN 0004-9581.

Chaubey:1978:NAF

[515] Yogendra P. Chaubey and Govind S. Mudholkar. A new approximation for Fisher’s Z. Australian Journal of Statistics, 20(3):250–256, September 1978. CODEN AUJSA3. ISSN 0004-9581.

Tiku:1978:FMA

[516] M. L. Tiku and D. Y. N. Yip. A four-moment approximation based on the F distribution. Australian Journal of Statistics, 20(3):257–261, September 1978. CODEN AUJSA3. ISSN 0004-9581.

Wong:1978:TLP

[517] P. G. Wong. A test of the location parameter of the exponential distribution. Australian Journal of Statistics, 20(3):262–269, September 1978. CODEN AUJSA3. ISSN 0004-9581.

Sundaraij:1978:MCR

[518] N. Sundaraij. A method for confidence regions for nonlinear models. Australian Journal of Statistics, 20(3):270–274, September 1978. CO-DEN AUJSA3. ISSN 0004-9581.

Prajneshu:1978:TSS

[519] Prajneshu. Two-species systems in random environment. Australian Journal of Statistics, 20(3):275–281, September 1978. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1978:BRd

[520] Anonymous. Book reviews. Australian Journal of Statistics, 20(3):282– 287, September 1978. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1978:BNb

[521] Anonymous. Book notices. Australian Journal of Statistics, 20(3):288– 289, September 1978. CODEN AUJSA3. ISSN 0004-9581.

Tallis:1979:RMA

[522] G. M. Tallis, P. Leppard, T. J. O’Neill, and G. Sarfaty. The relationship of metastatic activity to primary tumor size with special reference to breast cancer. Australian Journal of Statistics, 21(1):1–17, March 1979. CODEN AUJSA3. ISSN 0004-9581.

Wilson:1979:ERR

[523] S. R. Wilson. Examination of regression residuals. Australian Journal of Statistics, 21(1):18–29, March 1979. CODEN AUJSA3. ISSN 0004-9581.

Maritz:1979:TMD

[524] J. S. Maritz. On Theil’s method in distribution-free regression. Australian Journal of Statistics, 21(1):30–35, March 1979. CODEN AUJSA3. ISSN 0004-9581.

Kanji:1979:PAA

[525] G. K. Kanji. Power aspects of analysis of variance in fixed effect model one-way classification. Australian Journal of Statistics, 21(1):36–44, March 1979. CODEN AUJSA3. ISSN 0004-9581.

Jones:1979:ETS

[526] Roger G. Jones. The efficiency of time series estimators for repeated surveys. Australian Journal of Statistics, 21(1):45–56, March 1979. CODEN AUJSA3. ISSN 0004-9581.

Panda:1979:NSI

[527] Radha Madhab Panda. A non-sequential inventory model with variable supply. Australian Journal of Statistics, 21(1):57–65, March 1979. CO-DEN AUJSA3. ISSN 0004-9581.

Alvo:1979:SEP

[528] Mayer Alvo and Paul Cabilio. On the sequential estimation of the probability of success. Australian Journal of Statistics, 21(1):66–70, March 1979. CODEN AUJSA3. ISSN 0004-9581.

Hasofer:1979:ISV

[529] A. M. Hasofer and R. B. Davis. On the identification of the selection variable in Pearson’s selection model. Australian Journal of Statistics, 21(1):71–77, March 1979. CODEN AUJSA3. ISSN 0004-9581.

Shanbhag:1979:SRR

[530] D. N. Shanbhag and J. Panaretos. Some results related to the Rao– Rubin characterization of the Poisson distribution. Australian Journal of Statistics, 21(1):78–83, March 1979. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1979:BRa

[531] Anonymous. Book reviews. Australian Journal of Statistics, 21(1):84–92, March 1979. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1979:C

[532] Anonymous. Corrigendum. Australian Journal of Statistics, 21(1):92, March 1979. CODEN AUJSA3. ISSN 0004-9581.

Nicholls:1979:ATS

[533] D. F. Nicholls. The analysis of time series — the time domain approach. Australian Journal of Statistics, 21(2):93–120, June 1979. CODEN AU-JSA3. ISSN 0004-9581.

Pukkila:1979:BPO

[534] T. Pukkila. The bias in periodogram ordinates and the estimation of ARMA models in the frequency domain. Australian Journal of Statistics, 21(2):121–128, June 1979. CODEN AUJSA3. ISSN 0004-9581.

Praetz:1979:DOV

[535] Peter D. Praetz. The detection of omitted variables by Durbin–Watson statistics in multiple regression models. Australian Journal of Statistics, 21(2):129–138, June 1979. CODEN AUJSA3. ISSN 0004-9581.

Brockwell:1979:EPS

[536] P. J. Brockwell and B. M. Brown. Estimation for the positive stable laws, I. Australian Journal of Statistics, 21(2):139–148, June 1979. CODEN AUJSA3. ISSN 0004-9581.

Bofinger:1979:TSS

[537] Eve Bofinger. Two stage selection problem for normal populations with unequal variances. Australian Journal of Statistics, 21(2):149–156, June 1979. CODEN AUJSA3. ISSN 0004-9581.

Ohtani:1979:OCV

[538] Kazuhiro Ohtani and Toshihisa Toyoda. Optimal critical values for a preliminary test in pooling means. Australian Journal of Statistics, 21 (2):157–161, June 1979. CODEN AUJSA3. ISSN 0004-9581.

Chaudhuri:1979:EMF

[539] A. Chaudhuri and R. Arnab. On estimating the mean of a finite population sampled on two occasions with varying probabilities. Australian Journal of Statistics, 21(2):162–165, June 1979. CODEN AUJSA3. ISSN 0004-9581.

Arnold:1979:SCC

[540] B. C. Arnold. Some characterizations of the Cauchy distribution. Australian Journal of Statistics, 21(2):166–169, June 1979. CODEN AU-JSA3. ISSN 0004-9581.

Preece:1979:SBD

[541] D. A. Preece. Supplementary bibliography of designs for experiments in three dimensions. Australian Journal of Statistics, 21(2):170–172, June 1979. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1979:PM

[542] Anonymous. Pitman Medal. Australian Journal of Statistics, 21(2):173– 174, June 1979. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1979:BRb

Anonymous:1979:BN

[544] Anonymous. Book notices. Australian Journal of Statistics, 21(2):184– 185, June 1979. CODEN AUJSA3. ISSN 0004-9581.

Moran:1979:PHL

[545] P. A. P. Moran. Professor H. O. Lancaster F.A.A. Australian Journal of Statistics, 21(3):187, September 1979. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1979:BHL

[546] Anonymous. Bibliography of H. O. Lancaster. Australian Journal of Statistics, 21(3):188–192, September 1979. CODEN AUJSA3. ISSN 0004-9581.

Zolotarev:1979:IMP

[547] V. M. Zolotarev. Ideal metrics in the problems of probability theory and mathematical statistics. Australian Journal of Statistics, 21(3):193–208, September 1979. CODEN AUJSA3. ISSN 0004-9581.

Seneta:1979:RHW

[548] E. Seneta. Round the historical work on Bienaymé. Australian Journal of Statistics, 21(3):209–220, September 1979. CODEN AUJSA3. ISSN 0004-9581.

Eagleson:1979:OES

[549] G. K. Eagleson. Orthogonal expansions and U-statistics. Australian Journal of Statistics, 21(3):221–237, September 1979. CODEN AUJSA3. ISSN 0004-9581.

Griffiths:1979:ACB

[550] R. C. Griffiths, R. K. Milne, and R. Wood. Aspects of correlation in bivariate Poisson distributions and processes. Australian Journal of Statistics, 21(3):238–255, September 1979. CODEN AUJSA3. ISSN 0004-9581.

Lai:1979:OMM

[551] C. D. Lai and D. Vere-Jones. Odd man out — the Meixner hypergeometric distribution. Australian Journal of Statistics, 21(3):256–265, September 1979. CODEN AUJSA3. ISSN 0004-9581.

Armitage:1979:DCT

[552] P. Armitage. The design of clinical trials. Australian Journal of Statistics, 21(3):266–281, September 1979. CODEN AUJSA3. ISSN 0004-9581.

Heyde:1979:APS

[553] C. C. Heyde. On assessing the potential severity of an outbreak of a rare infectious disease: a Bayesian approach. Australian Journal of Statistics, 21(3):282–292, September 1979. CODEN AUJSA3. ISSN 0004-9581.

Moran:1979:UCL

[554] P. A. P. Moran. The use of correlation in large samples. Australian Journal of Statistics, 21(3):293–300, September 1979. CODEN AUJSA3. ISSN 0004-9581.

Robinson:1979:OTS

[555] J. Robinson. Optimal tests of significance. Australian Journal of Statistics, 21(3):301–310, September 1979. CODEN AUJSA3. ISSN 0004-9581.

Hamdan:1979:SPB

[556] M. A. Hamdan. On the structure of Pearson’s biserial correlation. Australian Journal of Statistics, 21(3):311–314, September 1979. CODEN AUJSA3. ISSN 0004-9581.

Pollard:1979:MAC

[557] J. H. Pollard. Matrix analysis of the cash flows and reserves of a life Office. Australian Journal of Statistics, 21(3):315–324, September 1979. CODEN AUJSA3. ISSN 0004-9581.

Tallis:1980:SBC

[558] G. M. Tallis, T. J. O’Neu, P. Leppard, and G. Sarfaty. Screening for breast cancer. Australian Journal of Statistics, 22(1):1–23, March 1980. CODEN AUJSA3. ISSN 0004-9581.

Fischer:1980:SEE

[559] A. J. Fischer. Sampling errors in the electoral process for the Australian Senate. Australian Journal of Statistics, 22(1):24–39, March 1980. CO-DEN AUJSA3. ISSN 0004-9581.

McDunnough:1980:FVR

[560] Philip McDunnough and David B. Wolfson. Fixed versus random sampling of certain continuous parameter processes. Australian Journal of Statistics, 22(1):40–49, March 1980. CODEN AUJSA3. ISSN 0004-9581.

Shorack:1980:SLI

[561] Galen R. Shorack. Some law of the iterated logarithm type results for the empirical process. Australian Journal of Statistics, 22(1):50–59, March 1980. CODEN AUJSA3. ISSN 0004-9581.

Enns:1980:RPO

[562] E. G. Enns and P. F. Ehlers. Random paths originating within a convex region and terminating on its surface. Australian Journal of Statistics, 22(1):60–68, March 1980. CODEN AUJSA3. ISSN 0004-9581.

Shirakura:1980:NSC

[563] Teruhiro Shirakura. Necessary and sufficient condition for a balanced array of strength 21 to be a balanced fractional 2 factorial design of resolution 21. Australian Journal of Statistics, 22(1):69–74, March 1980. CODEN AUJSA3. ISSN 0004-9581.

Fisher:1980:NML

[564] N. I. Fisher and H. M. Hudson. A note on the multivariate linear model with constraints on the dependent vector. Australian Journal of Statistics, 22(1):75–78, March 1980. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1980:BRa

[565] Anonymous. Book reviews. Australian Journal of Statistics, 22(1):79–96, March 1980. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1980:C

[566] Anonymous. Corrigenda. Australian Journal of Statistics, 22(1):97–98, March 1980. CODEN AUJSA3. ISSN 0004-9581.

Norton:1980:IPE

[567] W. E. Norton. Information problems and economic advising. Australian Journal of Statistics, 22(2):101–124, June 1980. CODEN AUJSA3. ISSN 0004-9581.

Watterson:1980:DNS

[568] G. A. Watterson and R. Anderson. Detecting natural selection: the stepwise mutation model with heterosis. Australian Journal of Statistics, 22(2):125–142, June 1980. CODEN AUJSA3. ISSN 0004-9581.

Cressie:1980:RAO

[569] Noel Cressie. Relaxing assumptions in the one sample t-test. Australian Journal of Statistics, 22(2):143–153, June 1980. CODEN AUJSA3. ISSN 0004-9581.

Brown:1980:MES

[570] B. M. Brown. Median estimates in simple linear regression. Australian Journal of Statistics, 22(2):154–165, June 1980. CODEN AUJSA3. ISSN 0004-9581.

Spurr:1980:EDB

[571] Barrie D. Spurr. The exact distribution of the Bell–Doksum test of independence. Australian Journal of Statistics, 22(2):166–171, June 1980. CODEN AUJSA3. ISSN 0004-9581.

DeSilva:1980:TIB

[572] B. M. DeSilva and R. C. Griffiths. A test of independence for bivariate symmetric stable distributions. Australian Journal of Statistics, 22(2): 172–177, June 1980. CODEN AUJSA3. ISSN 0004-9581.

Baruah:1980:MSP

[573] H. K. Baruah and G. P. Bhattacharjee. A modification of the sequential probability ratio test for testing a normal mean. Australian Journal of Statistics, 22(2):178–187, June 1980. CODEN AUJSA3. ISSN 0004-9581.

Byth:1980:LRC

[574] K. Byth and G. J. McLachlan. Logistic regression compared to normal discrimination for non-normal populations. Australian Journal of Statistics, 22(2):188–196, June 1980. CODEN AUJSA3. ISSN 0004-9581.

Gleeson:1980:BPR

[575] A. C. Gleeson and C. A. McGilchrist. Bilateral processes on a rectangular lattice. Australian Journal of Statistics, 22(2):197–206, June 1980. CODEN AUJSA3. ISSN 0004-9581.

Breth:1980:QEP

[576] M. Breth. Quantile estimation and probability coverages. Australian Journal of Statistics, 22(2):207–211, June 1980. CODEN AUJSA3. ISSN 0004-9581.

Phatarfod:1980:BDS

[577] Ravindra M. Phatarfod. The bottomless dam with seasonal inputs. Australian Journal of Statistics, 22(2):212–217, June 1980. CODEN AU-JSA3. ISSN 0004-9581.

Ho:1980:MUH

[578] E. W. H. Ho. Model-unbiasedness and the Horvitz–Thompson estimator in finite population sampling. Australian Journal of Statistics, 22(2): 218–225, June 1980. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1980:BRb

Anonymous:1980:BNa

[580] Anonymous. Book notices. Australian Journal of Statistics, 22(2):233– 235, June 1980. CODEN AUJSA3. ISSN 0004-9581.

McLachlan:1980:CAS

[581] G. J. McLachlan and J. N. Holt. The covariance analysis of some censored survival data from a large scale study of melanoma. Australian Journal of Statistics, 22(3):237–249, September 1980. CODEN AUJSA3. ISSN 0004-9581.

Marcus:1980:TSP

[582] Ruth Marcus. Two-stage procedures for testing homogeneity of means against ordered alternatives in one-way ANOVA with common unknown variance. Australian Journal of Statistics, 22(3):250–259, September 1980. CODEN AUJSA3. ISSN 0004-9581.

Tiku:1980:GFS

[583] M. L. Tiku. Goodness of fit statistics based on the spacings of complete or censored samples. Australian Journal of Statistics, 22(3):260–275, September 1980. CODEN AUJSA3. ISSN 0004-9581.

Weer:1980:TIM

[584] D. R. Weer and A. P. Basu. Testing for independence in multivariate exponential distributions. Australian Journal of Statistics, 22(3):276– 288, September 1980. CODEN AUJSA3. ISSN 0004-9581.

Moran:1980:TLS

[585] P. A. P. Moran. Testing the largest of a set of correlation coefficients. Australian Journal of Statistics, 22(3):289–297, September 1980. CO-DEN AUJSA3. ISSN 0004-9581.

Hatanaka:1980:NNA

[586] Michio Hatanaka. Note: a note on the application of the Kalman filter to regression models with some parameters varying over time and others unvarying. Australian Journal of Statistics, 22(3):298–306, September 1980. CODEN AUJSA3. ISSN 0004-9581.

Saunders:1980:AML

[587] I. W. Saunders. An approximate maximum likelihood estimator for chain binomial models. Australian Journal of Statistics, 22(3):307–316, September 1980. CODEN AUJSA3. ISSN 0004-9581.

Watson:1980:TSP

[588] Dan A. Watson and Saul Blumenthal. A two stage procedure for estimating the size of a truncated exponential sample. Australian Journal of Statistics, 22(3):317–327, September 1980. CODEN AUJSA3. ISSN 0004-9581.

Singh:1980:MVU

[589] Jagbir Singh. Minimum variance unbiased estimation of probability densities. Australian Journal of Statistics, 22(3):328–331, September 1980. CODEN AUJSA3. ISSN 0004-9581.

Akahira:1980:NNP

[590] Masafumi Akahira and Kei Takeuchi. Note: a note on prediction sufficiency (adequacy) and sufficiency. Australian Journal of Statistics, 22 (3):332–335, September 1980. CODEN AUJSA3. ISSN 0004-9581.

Moore:1980:RSP

[591] Marc Moore. Random sampling in pattern reconstruction. Australian Journal of Statistics, 22(3):336–340, September 1980. CODEN AUJSA3. ISSN 0004-9581.

Sarndal:1980:TMB

[592] C. E. Särndal. Two model-based inference arguments in survey sampling. Australian Journal of Statistics, 22(3):341–348, September 1980. CODEN AUJSA3. ISSN 0004-9581.

Arnab:1980:TSS

[593] Raghunath Arnab. Two-stage sampling over two occasions. Australian Journal of Statistics, 22(3):349–357, September 1980. CODEN AUJSA3. ISSN 0004-9581.

Singh:1980:NNS

[594] Kesar Singh. Note: a note on sample quantile processes of finite populations. Australian Journal of Statistics, 22(3):358–363, September 1980. CODEN AUJSA3. ISSN 0004-9581.

Dey:1980:SSE

[595] A. Dey and M. Singh. Some series of efficiency balanced designs. Australian Journal of Statistics, 22(3):364–367, September 1980. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1980:ODW

[596] Anonymous. Obituary: Douglas William Maitland (1925–1980). Australian Journal of Statistics, 22(3):368–369, September 1980. CODEN AUJSA3. ISSN 0004-9581. See corrigendum [646].

Anonymous:1980:BRc

[597] Anonymous. Book reviews. Australian Journal of Statistics, 22(3):370– 376, September 1980. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1980:BNb

[598] Anonymous. Book notices. Australian Journal of Statistics, 22(3):376– 378, September 1980. CODEN AUJSA3. ISSN 0004-9581.

Heyde:1981:PAL

[599] C. C. Heyde. Presidential address: Looking forward into the 1980’s: a personal view of the problems and prospects for the statistical profession. Australian Journal of Statistics, 23(1):1–14, March 1981. CODEN AUJSA3. ISSN 0004-9581.

Brosnan:1981:EIM

[600] Peter Brosnan. On estimating internal migration for open populations. Australian Journal of Statistics, 23(1):15–20, March 1981. CODEN AU-JSA3. ISSN 0004-9581.

Campbell:1981:GCC

[601] N. A. Campbell. Graphical comparison of covariance matrices. Australian Journal of Statistics, 23(1):21–37, March 1981. CODEN AUJSA3. ISSN 0004-9581.

Dargahi-Noubary:1981:ADW

[602] G. R. Dargahi-Noubary. An application of discrimination when covariance matrices are proportional. Australian Journal of Statistics, 23(1): 38–44, March 1981. CODEN AUJSA3. ISSN 0004-9581.

Nandi:1981:END

[603] S. B. Nandi. The exact null distribution of Bartlett’s criterion. Australian Journal of Statistics, 23(1):45–51, March 1981. CODEN AUJSA3. ISSN 0004-9581.

Maritz:1981:STI

[604] A. Maritz. A simple test for the independence of sets of variables. Australian Journal of Statistics, 23(1):52–57, March 1981. CODEN AUJSA3. ISSN 0004-9581.

ONeill:1981:NCC

[605] M. E. O’Neill. A note on the canonical correlations from contingency tables. Australian Journal of Statistics, 23(1):58–66, March 1981. CODEN AUJSA3. ISSN 0004-9581.

Akahira:1981:ADE

[606] Masafumi Akahira. On asymptotic deficiency of estimators. Australian Journal of Statistics, 23(1):67–72, March 1981. CODEN AUJSA3. ISSN 0004-9581.

Collett:1981:DBM

[607] D. Collett and T. Lewis. Discriminating between the von Mises and wrapped normal distributions. Australian Journal of Statistics, 23(1): 73–79, March 1981. CODEN AUJSA3. ISSN 0004-9581.

Narayan:1981:BPG

[608] P. Narayan. On bounds for probability generating functions. Australian Journal of Statistics, 23(1):80–90, March 1981. CODEN AUJSA3. ISSN 0004-9581.

Littlejohn:1981:LBS

[609] R. P. Littlejohn. Length-biassed sampling from a finite population. Australian Journal of Statistics, 23(1):91–94, March 1981. CODEN AUJSA3. ISSN 0004-9581.

Srivenkataramana:1981:EPM

[610] T. Srivenkataramana and D. S. Tracy. Extending product method of estimation to positive correlation case in surveys. Australian Journal of Statistics, 23(1):95–100, March 1981. CODEN AUJSA3. ISSN 0004-9581.

Sabharwal:1981:PPS

[611] Y. P. Sabharwal and B. R. Gupta. Peak points in a symmetrically correlated random walk. Australian Journal of Statistics, 23(1):101–108, March 1981. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1981:PMA

[612] Anonymous. Pitman Medal awarded to H. O. Lancaster. Australian Journal of Statistics, 23(1):109, March 1981. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1981:BRa

[613] Anonymous. Book reviews. Australian Journal of Statistics, 23(1):110– 136, March 1981. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1981:RS

[614] Anonymous. Referees’ services 1980. Australian Journal of Statistics, 23 (1):137, March 1981. CODEN AUJSA3. ISSN 0004-9581.

Brewer:1981:EMU

[615] K. R. W. Brewer. Estimating marihuana usage using randomized response — some paradoxical findings. Australian Journal of Statistics, 23 (2):139–148, June 1981. CODEN AUJSA3. ISSN 0004-9581.

Tiku:1981:GFS

[616] M. L. Tiku. A goodness of fit statistic based on the sample spacings for testing a symmetric distribution against symmetric alternatives. Australian Journal of Statistics, 23(2):149–158, June 1981. CODEN AU-JSA3. ISSN 0004-9581.

Jarque:1981:THL

[617] Carlos M. Jarque. A test for heteroscedasticity in a limited dependent variable model. Australian Journal of Statistics, 23(2):159–163, June 1981. CODEN AUJSA3. ISSN 0004-9581.

Hong-Zhi:1981:APE

[618] An Hong-Zhi. Asymptotic properties of an estimate of the prediction error variance. Australian Journal of Statistics, 23(2):164–176, June 1981. CODEN AUJSA3. ISSN 0004-9581.

Maller:1981:TPR

[619] R. A. Maller. A theorem on products of random variables, with application to regression. Australian Journal of Statistics, 23(2):177–185, June 1981. CODEN AUJSA3. ISSN 0004-9581.

Cheng:1981:NER

[620] Kuang-Fu Cheng and Pi-Erh Lin. Nonparametric estimation of a regression function: Limiting distribution. Australian Journal of Statistics, 23 (2):186–195, June 1981. CODEN AUJSA3. ISSN 0004-9581.

Theodorescu:1981:SEE

[621] Radu Theodorescu and Hans Wolff. Sequential estimation of expectations in the presence of trend. Australian Journal of Statistics, 23(2):196–203, June 1981. CODEN AUJSA3. ISSN 0004-9581.

Morton:1981:OEE

[622] R. Morton. Optimal estimating equations with applications to insect development times. Australian Journal of Statistics, 23(2):204–213, June 1981. CODEN AUJSA3. ISSN 0004-9581.

Marcus:1981:SCI

[623] Ruth Marcus. Simultaneous confidence intervals for monotone contrasts of normal means. Australian Journal of Statistics, 23(2):214–223, June 1981. CODEN AUJSA3. ISSN 0004-9581.

Gelfand:1981:EPP

[624] Alan E. Gelfand. An estimation problem with Poisson processes. Australian Journal of Statistics, 23(2):224–231, June 1981. CODEN AU-JSA3. ISSN 0004-9581.

Lai:1981:CCC

[625] C. D. Lai. On conditional correlation coefficients of a Wold Markov process of gamma intervals. Australian Journal of Statistics, 23(2):232– 237, June 1981. CODEN AUJSA3. ISSN 0004-9581.

Richards:1981:CMV

[626] Donald St. P. Richards. A characterization of a matric-variate exponential distribution. Australian Journal of Statistics, 23(2):238–239, June 1981. CODEN AUJSA3. ISSN 0004-9581.

Griffiths:1981:CGT

[627] David Griffiths and Christine Schafer. Closeness of Grassia’s transformed gammas and the beta distribution. Australian Journal of Statistics, 23 (2):240–246, June 1981. CODEN AUJSA3. ISSN 0004-9581.

MacGillivray:1981:MMM

[628] H. L. MacGillivray. The mean, median, mode inequality and skewness for a class of densities. Australian Journal of Statistics, 23(2):247–250, June 1981. CODEN AUJSA3. ISSN 0004-9581.

Sivazlian:1981:CMD

[629] B. D. Sivazlian. A class of multivariate distributions. Australian Journal of Statistics, 23(2):251–255, June 1981. CODEN AUJSA3. ISSN 0004-9581.

Gupta:1981:CYA

[630] B. C. Gupta. Checks on Yates’s algorithm for a 2^{T M} factorial experiment. Australian Journal of Statistics, 23(2):256–258, June 1981. CODEN AU-JSA3. ISSN 0004-9581.

Anonymous:1981:BRb

Anonymous:1981:BN

[632] Anonymous. Book notices. Australian Journal of Statistics, 23(2):272,

June 1981. CODEN AUJSA3. ISSN 0004-9581.

Heyde:1981:TSS

[633] C. C. Heyde. Trends in the statistical sciences. Australian Journal of Statistics, 23(3):273–286, September 1981. CODEN AUJSA3. ISSN 0004-9581.

Tuckwell:1981:AGM

[634] R. H. Tuckwell. Anomalies in the gambling market. Australian Journal of Statistics, 23(3):287–295, September 1981. CODEN AUJSA3. ISSN 0004-9581.

Finch:1981:DIS

[635] Peter D. Finch. A descriptive interpretation of standard error. Australian Journal of Statistics, 23(3):296–299, September 1981. CODEN AUJSA3. ISSN 0004-9581.

Heuts:1981:ARP

[636] R. M. J. Heuts. Asymptotic robustness of prediction intervals of ARIMA models to deviations from normality. Australian Journal of Statistics, 23 (3):300–308, September 1981. CODEN AUJSA3. ISSN 0004-9581.

Praetz:1981:NEA

[637] Peter Praetz. A note on the effect of autocorrelation on multiple regression statistics. Australian Journal of Statistics, 23(3):309–313, September 1981. CODEN AUJSA3. ISSN 0004-9581.

Breth:1981:STL

[638] M. Breth. Simultaneous tolerance limits and a unified approach to some nonparametric theory. Australian Journal of Statistics, 23(3):314–327, September 1981. CODEN AUJSA3. ISSN 0004-9581.

Perng:1981:SFW

[639] S. K. Perng, David P. Hasza, and C. Paik. Sequential fixed-width confidence interval for a function of the parameters. Australian Journal of Statistics, 23(3):328–336, September 1981. CODEN AUJSA3. ISSN 0004-9581.

Shanthikumar:1981:LCA

[640] J. G. Shanthikumar. On level crossing analysis of queues. Australian Journal of Statistics, 23(3):337–342, September 1981. CODEN AUJSA3. ISSN 0004-9581.

Shirahata:1981:STP

[641] Shingo Shirahata. A simple test procedure when data are given by n independent blocks. Australian Journal of Statistics, 23(3):343–351, September 1981. CODEN AUJSA3. ISSN 0004-9581.

Fisher:1981:OSP

[642] N. I. Fisher. One-sample probability plots. Australian Journal of Statistics, 23(3):352–359, September 1981. CODEN AUJSA3. ISSN 0004-9581.

Arnab:1981:STO

[643] R. Arnab. Sampling on two occasions with varying probabilities. Australian Journal of Statistics, 23(3):360–364, September 1981. CODEN AUJSA3. ISSN 0004-9581.

Sharma:1981:CED

[644] V. K. Sharma. A class of experimental designs balanced for first residuals. Australian Journal of Statistics, 23(3):365–370, September 1981. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1981:BRc

[645] Anonymous. Book reviews. Australian Journal of Statistics, 23(3):371– 381, September 1981. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1981:CDW

[646] Anonymous. Corrigendum: Douglas William Maitland. Australian Journal of Statistics, 23(3):381, September 1981. CODEN AUJSA3. ISSN 0004-9581. See [596].

Gani:1982:APE

[647] J. Gani. An adventure in publishing: Eighteen years of the applied probability trust. Australian Journal of Statistics, 24(1):1–17, March 1982. CODEN AUJSA3. ISSN 0004-9581.

Jones:1982:VHT

[648] Roger Jones. Variations in household telephone access: Implications for telephone surveys. Australian Journal of Statistics, 24(1):18–32, March 1982. CODEN AUJSA3. ISSN 0004-9581.

Brockwell:1982:NNI

[649] Peter J. Brockwell, Paul W. Mielke, Jr., and John Robinson. On nonnormal invariance principles for multi-response permutation procedures. Australian Journal of Statistics, 24(1):33–41, March 1982. CODEN AU-JSA3. ISSN 0004-9581.

Rao:1982:NTI

[650] K. S. Madhava Rao and A. P. Gore. Nonparametric tests for intercept in linear regression problems. Australian Journal of Statistics, 24(1):42–50, March 1982. CODEN AUJSA3. ISSN 0004-9581.

Ghorai:1982:BRC

[651] J. K. Ghorai, H. Rubin, and H. Rubin. Bayes risk consistency of nonparametric Bayes density estimates. Australian Journal of Statistics, 24 (1):51–66, March 1982. CODEN AUJSA3. ISSN 0004-9581.

Parr:1982:JCD

[652] William C. Parr and H. Dennis Tolley. Jackknifing in categorical data analysis. Australian Journal of Statistics, 24(1):67–79, March 1982. CO-DEN AUJSA3. ISSN 0004-9581.

Scott:1982:CPM

[653] David J. Scott. Contiguity of probability measures. Australian Journal of Statistics, 24(1):80–88, March 1982. CODEN AUJSA3. ISSN 0004-9581.

Pakes:1982:SRP

[654] Anthony G. Pakes. Some remarks on the paper by Theodorescu and Wolff: “Sequential Estimation of Expectations in the Presence of Trend”. Australian Journal of Statistics, 24(1):89–97, March 1982. CODEN AU-JSA3. ISSN 0004-9581.

Hudson:1982:LSI

[655] Irene L. Hudson. Large sample inference for Markovian exponential families with application to branching processes with immigration. Australian Journal of Statistics, 24(1):98–112, March 1982. CODEN AUJSA3. ISSN 0004-9581. See corrigendum [706].

Kageyama:1982:RCE

[656] Sanpei Kageyama. Remark on construction of efficiency-balanced designs. Australian Journal of Statistics, 24(1):113–115, March 1982. CO-DEN AUJSA3. ISSN 0004-9581.

Finch:1982:LE

[657] P. D. Finch. Letter to the Editor. Australian Journal of Statistics, 24 (1):116–118, March 1982. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1982:HLM

[658] Anonymous. Honorary life members. Australian Journal of Statistics, 24 (1):119–121, March 1982. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1982:BRa

[659] Anonymous. Book reviews. Australian Journal of Statistics, 24(1):122– 134, March 1982. CODEN AUJSA3. ISSN 0004-9581.

Thornett:1982:RRM

[660] M. L. Thornett. The role of randomization in model-based inference. Australian Journal of Statistics, 24(2):137–145, June 1982. CODEN AU-JSA3. ISSN 0004-9581.

Finch:1982:CPM

[661] Peter D. Finch. Comments on paper by M. L. Thornett: The role of randomization in model based inference. Australian Journal of Statistics, 24(2):146–147, June 1982. CODEN AUJSA3. ISSN 0004-9581.

Thornett:1982:TVR

[662] M. L. Thornett. Two views of randomization?: Response to Peter Finch’s comments. Australian Journal of Statistics, 24(2):148–150, June 1982. CODEN AUJSA3. ISSN 0004-9581.

Asmussen:1982:RCE

[663] Søren Asmussen. On the role of a certain eigenvalue in estimating the growth rate of a branching process. Australian Journal of Statistics, 24 (2):151–159, June 1982. CODEN AUJSA3. ISSN 0004-9581.

Narayan:1982:ETC

[664] Prakash Narayan. On the extinction time of a continuous time Markov branching process. Australian Journal of Statistics, 24(2):160–164, June 1982. CODEN AUJSA3. ISSN 0004-9581.

Agrafiotis:1982:SMP

[665] G. K. Agrafiotis. Stochastic multigrade population models in continuous time. Australian Journal of Statistics, 24(2):165–170, June 1982. CODEN AUJSA3. ISSN 0004-9581.

Jones:1982:UMH

[666] Roger G. Jones. The use of multiplicity in a household survey of disabilities. Australian Journal of Statistics, 24(2):171–178, June 1982. CODEN AUJSA3. ISSN 0004-9581.

Nicholls:1982:STP

[667] Miles G. Nicholls. Short term prediction of student numbers in the Victorian Secondary Education System. Australian Journal of Statistics, 24 (2):179–190, June 1982. CODEN AUJSA3. ISSN 0004-9581.

Agrawal:1982:SMC

[668] H. L. Agrawal and Jagdish Prasad. Some methods of construction of GD–RC and rectangular-RC designs. Australian Journal of Statistics, 24(2):191–200, June 1982. CODEN AUJSA3. ISSN 0004-9581.

Morley:1982:SSP

[669] Clive L. Morley. A simulation study of the powers of three multiple comparison statistics. Australian Journal of Statistics, 24(2):201–210, June 1982. CODEN AUJSA3. ISSN 0004-9581.

Samiuddin:1982:NUT

[670] M. Samiuddin. A note on an unbiased test of homogeneity of variances. Australian Journal of Statistics, 24(2):211–215, June 1982. CODEN AU-JSA3. ISSN 0004-9581.

Gupta:1982:SEH

[671] Rameshwar D. Gupta and Donald St P. Richards. Series expansion for a hypergeometric function of matrix argument with applications. Australian Journal of Statistics, 24(2):216–220, June 1982. CODEN AU-JSA3. ISSN 0004-9581.

Lai:1982:MCM

[672] C. D. Lai. Meixner classes and Meixner hyper-geometric distributions. Australian Journal of Statistics, 24(2):221–233, June 1982. CODEN AU-JSA3. ISSN 0004-9581.

Robinson:1982:CHT

[673] P. M. Robinson. On the convergence of the Horvitz–Thompson estimator. Australian Journal of Statistics, 24(2):234–238, June 1982. CODEN AUJSA3. ISSN 0004-9581.

Shonkwiler:1982:JTE

[674] J. S. Shonkwiler. The joint treatment of exact and stochastic restrictions. Australian Journal of Statistics, 24(2):239–244, June 1982. CODEN AU-JSA3. ISSN 0004-9581.

Anh:1982:EGL

[675] V. V. Anh. Estimated generalized least squares for a heteroscedastic regression model. Australian Journal of Statistics, 24(2):245–254, June 1982. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1982:BRb

Lee:1982:ISH

[677] Alan J. Lee. On incomplete U-statistics having minimum variance. Australian Journal of Statistics, 24(3):275–282, September 1982. CODEN AUJSA3. ISSN 0004-9581.

Brown:1982:RAI

[678] B. M. Brown. Robustness against inequality of variances. Australian Journal of Statistics, 24(3):283–295, September 1982. CODEN AUJSA3. ISSN 0004-9581.

Bolyai:1982:EVG

[679] Sándor Csörgö Bolyai and Peter Hall. Estimable versions of Griffiths’ measure of association. Australian Journal of Statistics, 24(3):296–308, September 1982. CODEN AUJSA3. ISSN 0004-9581.

Bandyopadhyay:1982:CCU

[680] Shibdas Bandyopadhyay. Covariate classification using dependent samples. Australian Journal of Statistics, 24(3):309–317, September 1982. CODEN AUJSA3. ISSN 0004-9581.

Brown:1982:DFM

[681] B. M. Brown and J. S. Maritz. Distribution-free methods in regression. Australian Journal of Statistics, 24(3):318–331, September 1982. CO-DEN AUJSA3. ISSN 0004-9581.

Gordon:1982:ITS

[682] A. D. Gordon. An investigation of two sequence-comparison statistics. Australian Journal of Statistics, 24(3):332–342, September 1982. CO-DEN AUJSA3. ISSN 0004-9581.

Finch:1982:DIM

[683] Peter D. Finch. Difficulties of interpretation with Markov chain models. Australian Journal of Statistics, 24(3):343–349, September 1982. CO-DEN AUJSA3. ISSN 0004-9581.

Lamacraft:1982:TCI

[684] Robin R. Lamacraft and W. B. Hall. Tables of cyclic incomplete block designs: r = k. Australian Journal of Statistics, 24(3):350–360, September 1982. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1982:BVS

[685] H. O. Lancaster. Bibliography of vital statistics in Australia: a third list. Australian Journal of Statistics, 24(3):361–380, September 1982.

CODEN AUJSA3. ISSN 0004-9581. See also first and second lists [101, 333].

Anonymous:1982:PMA

[686] Anonymous. Pitman Medal awarded to P. A. P. Moran. Australian Journal of Statistics, 24(3):383, September 1982. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1982:BRc

[687] Anonymous. Book reviews. Australian Journal of Statistics, 24(3):384– 400, September 1982. CODEN AUJSA3. ISSN 0004-9581.

Feigin:1983:WAT

[688] P. D. Feigin, R. L. Tweedie, and C. Belyea. Weighted area techniques for explicit parameter estimation in multi-stage models. Australian Journal of Statistics, 25(1):1–16, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Cassing:1983:EEC

[689] Shirley A. Cassing and Kenneth J. White. An examination of the eigenvector condition in the Durbin–Watson test. Australian Journal of Statistics, 25(1):17–22, April 1983. CODEN AUJSA3. ISSN 0004-9581.

King:1983:TMA

[690] Maxwell L. King. Testing for moving average regression disturbances. Australian Journal of Statistics, 25(1):23–34, April 1983. CODEN AU-JSA3. ISSN 0004-9581.

Basawa:1983:ROI

[691] I. V. Basawa and Niels Becker. Remarks on optimal inference for Markov branching processes: a sequential approach. Australian Journal of Statistics, 25(1):35–46, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Hudson:1983:ATC

[692] Irene L. Hudson. Asymptotic tests for the continuous time Markov branching process with immigration. Australian Journal of Statistics, 25(1):47–57, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Rogers:1983:NDR

[693] G. S. Rogers and D. G. Kabe. On the noncentral distribution of a random matrix useful in classification theory. Australian Journal of Statistics, 25 (1):58–63, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Barnett:1983:RDM

[694] Vic Barnett. Reduced distance measures and transformations in processing multivariate outliers. Australian Journal of Statistics, 25(1):64–75, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Beth:1983:NNC

[695] M. Beth. Notes: Notes on confidence regions for a linear model. Australian Journal of Statistics, 25(1):76–84, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Wei:1983:TIP

[696] L. J. Wei. Tests for independence in the presence of missing values. Australian Journal of Statistics, 25(1):85–90, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Moran:1983:TCE

[697] P. A. P. Moran. Testing the concordance of events with their probabilities. Australian Journal of Statistics, 25(1):91–95, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Clark:1983:NAF

[698] R. M. Clark and B. J. Morrison. A normal approximation to the Fisher distribution. Australian Journal of Statistics, 25(1):96–104, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Clarke:1983:AFC

[699] A. P. Clarke and W. Marwood. An analytic form for the cumulative distribution function of the Bell–Doksum statistic. Australian Journal of Statistics, 25(1):105–108, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Sheather:1983:EAS

[700] S. J. Sheather and J. S. Maritz. An estimate of the asymptotic standard error of the sample median. Australian Journal of Statistics, 25(1):109– 122, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Akahira:1983:ADJ

[701] Masafumi Akahira. Asymptotic deficiency of the jackknife estimator. Australian Journal of Statistics, 25(1):123–129, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1983:SLM

[702] Anonymous. A stochastic logistic model with time delay. Australian Journal of Statistics, 25(1):130–135, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Papageorgiou:1983:CSB

[703] H. Papageorgiou. On characterizing some bivariate discrete distributions. Australian Journal of Statistics, 25(1):136–144, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1983:BRa

[704] Anonymous. Book reviews. Australian Journal of Statistics, 25(1):145– 162, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1983:BN

[705] Anonymous. Book notices. Australian Journal of Statistics, 25(1):162– 164, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Hudson:1983:CLS

[706] Irene L. Hudson. Corrigendum: Large sample inference for Markovian exponential families with application to branching processes with immigration. Australian Journal of Statistics, 25(1):165, April 1983. CODEN AUJSA3. ISSN 0004-9581. See [655].

Anonymous:1983:RS

[707] Anonymous. Referees’ services 1982. Australian Journal of Statistics, 25 (1):167, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Pitman:1983:PJW

[708] E. J. G. Pitman. Professor E. J. Williams. Australian Journal of Statistics, 25(1):169–170, April 1983. CODEN AUJSA3. ISSN 0004-9581.

Williams:1983:B

[709] E. J. Williams. Bibliography. Australian Journal of Statistics, 25(2): 171–173, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Bartlett:1983:EMV

[710] M. S. Bartlett. Examples of minimum variance estimation. Australian Journal of Statistics, 25(2):174–181, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Basawa:1983:RTA

[711] I. V. Basawa. Recent trends in asymptotic optimal inference for dependent observations. Australian Journal of Statistics, 25(2):182–190, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Becker:1983:ADS

[712] Niels G. Becker. Analysis of data from a single epidemic. Australian Journal of Statistics, 25(2):191–197, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Bofinger:1983:MCS

[713] Eve Bofinger. Multiple comparisons and selection. Australian Journal of Statistics, 25(2):198–207, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Brockwell:1983:CPC

[714] P. J. Brockwell, J. M. Gani, and S. I. Resnick. Catastrophe processes with continuous state-space. Australian Journal of Statistics, 25(2):208–226, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Clark:1983:EPM

[715] R. M. Clark. Estimation of parameters in the marginal Fisher distribution. Australian Journal of Statistics, 25(2):227–237, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Brown:1983:MSD

[716] B. M. Brown. Minimum significance and distribution free tests. Australian Journal of Statistics, 25(2):238–248, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Cohn:1983:CRS

[717] Harry Cohn. On the convergence result for the supercritical Bellman– Harris process. Australian Journal of Statistics, 25(2):249–255, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Eagleson:1983:RTM

[718] G. K. Eagleson. A robust test for multiple comparisons of correlation coefficients. Australian Journal of Statistics, 25(2):256–263, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Ewens:1983:EPA

[719] W. J. Ewens. An estimation paradox arising in population genetics theory. Australian Journal of Statistics, 25(2):264–272, February 1983. CO-DEN AUJSA3. ISSN 0004-9581.

Gordon:1983:AMS

[720] Ian Gordon. Ancillarity and minimal sufficiency. Australian Journal of Statistics, 25(2):273–777, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Hall:1983:RBD

[721] Peter Hall. On the reduction of bias in density estimates. Australian Journal of Statistics, 25(2):278–286, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Hannan:1983:CAA

[722] E. J. Hannan and L. Kavalieris. The convergence of autocorrelations and autoregressions. Australian Journal of Statistics, 25(2):287–297, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1983:SJD

[723] H. O. Lancaster. Special joint distributions of Meixner variables. Australian Journal of Statistics, 25(2):298–309, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Leedow:1983:WAT

[724] M. I. Leedow and R. L. Tweedie. Weighted area techniques for the estimation of the parameters of a growth curve. Australian Journal of Statistics, 25(2):310–320, February 1983. CODEN AUJSA3. ISSN 0004-9581.

McGilchrist:1983:GIU

[725] C. A. McGilchrist, R. L. Sandland, and J. L. Hennessy. Generalized inverses used in recursive estimation of the general linear model. Australian Journal of Statistics, 25(2):321–328, February 1983. CODEN AUJSA3. ISSN 0004-9581.

MacPhee:1983:GWB

[726] I. M. MacPhee and H.-J. Schuh. A Galton–Watson branching process in varying environments with essentially constant offspring means and two rates of growth. Australian Journal of Statistics, 25(2):329–338, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Moran:1983:NEM

[727] P. A. P. Moran. A new expansion for the multivariate normal distribution. Australian Journal of Statistics, 25(2):339–344, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Pakes:1983:SPR

[728] Anthony G. Pakes. Some properties of a random linear difference equation. Australian Journal of Statistics, 25(2):345–357, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Robinson:1983:AST

[729] J. Robinson. Approximations to some test statistics for permutation tests in a completely randomized design. Australian Journal of Statistics, 25 (2):358–369, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Seneta:1983:WMM

[730] E. Seneta. The weighted median and multiple regression. Australian Journal of Statistics, 25(2):370–377, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Speed:1983:CPL

[731] T. P. Speed. Cumulants and partition lattices. Australian Journal of Statistics, 25(2):378–388, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Watson:1983:DRS

[732] G. S. Watson. Distributions in R with rotational symmetries. Australian Journal of Statistics, 25(2):389–394, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Watterson:1983:SSD

[733] G. A. Watterson. Some statistical distributions of use in analysing bird count data. Australian Journal of Statistics, 25(2):395–401, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Zehnwirth:1983:CTC

[734] Ben Zehnwirth. Credibility theory: a concise partial survey. Australian Journal of Statistics, 25(2):402–411, February 1983. CODEN AUJSA3. ISSN 0004-9581.

Hudson:1983:ATG

[735] Irene L. Hudson. Asymptotic tests for growth curve models with autoregressive errors. Australian Journal of Statistics, 25(3):413–424, September 1983. CODEN AUJSA3. ISSN 0004-9581.

Weber:1983:JGL

[736] N. C. Weber and A. H. Welsh. Jackknifing the general linear model. Australian Journal of Statistics, 25(3):425–436, September 1983. CO-DEN AUJSA3. ISSN 0004-9581.

Csorgo:1983:GPR

[737] Sándor Csörgö. Gil–Pelaez–Rosén transformation of data. Australian Journal of Statistics, 25(3):437–441, September 1983. CODEN AUJSA3. ISSN 0004-9581.

Pesaran:1983:NNM

[738] M. H. Pesaran. Note: a note on the maximum likelihood estimation of regression models with first order moving average errors with roots on the unit circle. Australian Journal of Statistics, 25(3):442–448, September 1983. CODEN AUJSA3. ISSN 0004-9581.

Beveridge:1983:TFT

[739] Stephen Beveridge. On a test of fit in time series models. Australian Journal of Statistics, 25(3):449–452, September 1983. CODEN AUJSA3. ISSN 0004-9581.

Puri:1983:IUB

[740] Madan L. Puri and Dan A. Ralescu. Improving upon the best invariant estimator in multivariate location problems. Australian Journal of Statistics, 25(3):453–462, September 1983. CODEN AUJSA3. ISSN 0004-9581.

Lin:1983:MEM

[741] Pi-Erh Lin and Amany Mousa. Minimax estimation of a multivariate normal mean under a convex loss function. Australian Journal of Statistics, 25(3):463–466, September 1983. CODEN AUJSA3. ISSN 0004-9581.

Kshirsagar:1983:DCF

[742] Anant M. Kshirsagar and Thomas M. Davis. Direction and collinearity factors of Wilk’s A associated with the growth curve model. Australian Journal of Statistics, 25(3):467–481, September 1983. CODEN AUJSA3. ISSN 0004-9581.

Bandyopadhyay:1983:ALR

[743] Shibdas Bandyopadhyay. Admissibility of likelihood ratio rules in covariate classification using dependent samples. Australian Journal of Statistics, 25(3):482–486, September 1983. CODEN AUJSA3. ISSN 0004-9581.

Pettitt:1983:ICM

[744] A. N. Pettitt. Inference for censored matched pairs using signed ranks. Australian Journal of Statistics, 25(3):487–495, September 1983. CO-DEN AUJSA3. ISSN 0004-9581.

Pollard:1983:ACT

[745] G. H. Pollard. An analysis of classical and tie-breaker tennis. Australian Journal of Statistics, 25(3):496–505, September 1983. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1983:BRb

[746] Anonymous. Book reviews. Australian Journal of Statistics, 25(3):506– 516, September 1983. CODEN AUJSA3. ISSN 0004-9581.

Becker:1984:EPS

[747] Niels G. Becker. Estimating population size from capture-recapture experiments in continuous time. Australian Journal of Statistics, 26(1): 1–7, April 1984. CODEN AUJSA3. ISSN 0004-9581.

Barbour:1984:SMB

[748] A. D. Barbour and Peter Hall. Stein’s method and the Berry–Esséen Theorem. Australian Journal of Statistics, 26(1):8–15, April 1984. CO-DEN AUJSA3. ISSN 0004-9581.

Deshpande:1984:TNT

[749] Jayant V. Deshpandé and Kalpana Kusum. A test for the nonparametric two sample scale problem. Australian Journal of Statistics, 26(1):16–24, April 1984. CODEN AUJSA3. ISSN 0004-9581.

Papageorgiou:1984:CMN

[750] H. Papageorgiou. Characterizations of multinomial and negative multinomial mixtures by regression. Australian Journal of Statistics, 26(1): 25–29, April 1984. CODEN AUJSA3. ISSN 0004-9581.

Brockwell:1984:ADM

[751] Peter J. Brockwell and Paul W. Mielke. Asymptotic distributions of matched-Paris permutation statistics based on distance measures. Australian Journal of Statistics, 26(1):30–38, April 1984. CODEN AUJSA3. ISSN 0004-9581.

Moran:1984:MCE

[752] P. A. P. Moran. The Monte Carlo evaluation of orthant probabilities for multivariate normal distributions. Australian Journal of Statistics, 26 (1):39–44, April 1984. CODEN AUJSA3. ISSN 0004-9581.

Williams:1984:RED

[753] Pamela Williams. Regional economic-demographic modelling and projection: Review of an international conference. Australian Journal of Statistics, 26(1):45–57, April 1984. CODEN AUJSA3. ISSN 0004-9581.

Brown:1984:RTM

[754] B. M. Brown and S. I. Resnick. Rank tests for multivariate trend. Australian Journal of Statistics, 26(1):58–67, April 1984. CODEN AUJSA3. ISSN 0004-9581.

Phatarfod:1984:IDD

[755] Ravindra M. Phatarfod. The infinitely deep dam with seasonal Markov inflows — a simplifying modification. Australian Journal of Statistics, 26(1):68–74, April 1984. CODEN AUJSA3. ISSN 0004-9581.

Lawal:1984:UTI

[756] H. Bayo Lawal and Graham J. G. Upton. On the use of X as a test of independence in contingency tables with small cell expectations. Australian Journal of Statistics, 26(1):75–85, April 1984. CODEN AUJSA3. ISSN 0004-9581.

Campbell:1984:CVA

[757] N. A. Campbell. Canonical variate analysis — a general model formulation. Australian Journal of Statistics, 26(1):86–96, April 1984. CODEN AUJSA3. ISSN 0004-9581.

Liu:1984:ESY

[758] Tzen-Ping Liu. Existence of a Sen–Yates–Grundy form for a dispersion. Australian Journal of Statistics, 26(1):97–104, April 1984. CODEN AU-JSA3. ISSN 0004-9581.

Anonymous:1984:BRa

[759] Anonymous. Book reviews. Australian Journal of Statistics, 26(1):105– 109, April 1984. CODEN AUJSA3. ISSN 0004-9581.

Kaminsky:1984:SND

[760] K. S. Kaminsky, E. M. Luks, and P. I. Nelson. Strategy, nontransitive dominance and the exponential distribution. Australian Journal of Statistics, 26(2):111–118, June 1984. CODEN AUJSA3. ISSN 0004-9581.

Saha:1984:BAW

[761] G. M. Saha and Sanpei Kageyama. Balanced arrays and weighing designs. Australian Journal of Statistics, 26(2):119–124, June 1984. CO-DEN AUJSA3. ISSN 0004-9581.

Koffler:1984:SEF

[762] Stephen L. Koffler and Douglas A. Penfield. A K-sample extension to Fligner’s class of tests for scale. Australian Journal of Statistics, 26(2): 125–132, June 1984. CODEN AUJSA3. ISSN 0004-9581.

Bhuyan:1984:GSP

[763] K. C. Bhuyan. A group of split-plot designs with a heteroscedastic model. Australian Journal of Statistics, 26(2):133–141, June 1984. CODEN AU-JSA3. ISSN 0004-9581.

Fisher:1984:GFT

[764] N. I. Fisher and D. J. Best. Goodness-of-fit tests for Fisher’s distribution on the sphere. Australian Journal of Statistics, 26(2):142–150, June 1984. CODEN AUJSA3. ISSN 0004-9581.

Revfeim:1984:GMP

[765] K. J. A. Revfeim. Generating mechanisms of, and parameter estimators for, the extreme value distribution. Australian Journal of Statistics, 26 (2):151–159, June 1984. CODEN AUJSA3. ISSN 0004-9581.

Berman:1984:CIL

[766] Mark Berman. Combining information from linear models on possibly different scales. Australian Journal of Statistics, 26(2):160–168, June 1984. CODEN AUJSA3. ISSN 0004-9581.

Huang:1984:AMA

[767] Jun S. Huang. The autoregressive moving average model for spatial analysis. Australian Journal of Statistics, 26(2):169–178, June 1984. CODEN AUJSA3. ISSN 0004-9581.

Bora:1984:CDE

[768] A. C. Bora. Change-over designs with errors following a first order autoregressive process. Australian Journal of Statistics, 26(2):179–188, June 1984. CODEN AUJSA3. ISSN 0004-9581.

Csorgo:1984:KMT

[769] Sándor Csörgö and Peter Hall. The Komlós–Major–Tusnády approximations and their applications. Australian Journal of Statistics, 26(2): 189–218, June 1984. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1984:BRb

[770] Anonymous. Book reviews. Australian Journal of Statistics, 26(2):219– 225, June 1984. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1984:BN

[771] Anonymous. Book notices. Australian Journal of Statistics, 26(2):225, June 1984. CODEN AUJSA3. ISSN 0004-9581.

Houtman:1984:AMD

[772] A. Houtman, T. P. Speed, and A. C. Nielsen. The analysis of multistrata designed experiments with incomplete data. Australian Journal of Statistics, 26(3):227–246, September 1984. CODEN AUJSA3. ISSN 0004-9581.

Hasofer:1984:RAM

[773] A. M. Hasofer, D. Veneziano, and H. Banon. Risk analysis in more than one dimension. Australian Journal of Statistics, 26(3):247–254, September 1984. CODEN AUJSA3. ISSN 0004-9581.

Samanta:1984:EEP

[774] M. Samanta. Efficient estimation of parameters in the K sample problem with equal but unknown population coefficients of variation. Australian Journal of Statistics, 26(3):255–262, September 1984. CODEN AUJSA3. ISSN 0004-9581.

Isaacson:1984:BPB

[775] Dean Isaacson and Glen Meeden. Better priors for Bayesian bettors. Australian Journal of Statistics, 26(3):263–271, September 1984. CO-DEN AUJSA3. ISSN 0004-9581.

Watson:1984:CCR

[776] G. S. Watson. The calculation of confidence regions for eigenvectors. Australian Journal of Statistics, 26(3):272–276, September 1984. CO-DEN AUJSA3. ISSN 0004-9581.

Puri:1984:LME

[777] Madan L. Puri, Dan A. Ralescu, and Stefan S. Ralescu. Linear minimax estimators for estimating a function of the parameter. Australian Journal of Statistics, 26(3):277–283, September 1984. CODEN AUJSA3. ISSN 0004-9581.

Becker:1984:SMI

[778] Niels Becker and Caroline Finch. On a stochastic model for an irreversible system of compartments. Australian Journal of Statistics, 26(3):284–289, September 1984. CODEN AUJSA3. ISSN 0004-9581.

Spiegelman:1984:NSD

[779] Clifford H. Spiegelman. A new statistic for detecting influential observations in a Scheffé type calibration curve. Australian Journal of Statistics, 26(3):290–297, September 1984. CODEN AUJSA3. ISSN 0004-9581.

Agrawal:1984:RRC

[780] H. C. Agrawal and C. R. Nair. Reduced residue classes cyclic PBIB designs. Australian Journal of Statistics, 26(3):298–309, September 1984. CODEN AUJSA3. ISSN 0004-9581.

Fisher:1984:GSE

[781] Gordon Fisher and Michael McAleer. The geometry of specification error. Australian Journal of Statistics, 26(3):310–322, September 1984. CODEN AUJSA3. ISSN 0004-9581.

Srivastava:1984:EFS

[782] V. K. Srivastava and D. E. A. Giles. Exact finite-sample properties of a pre-test estimator in ridge regression. Australian Journal of Statistics, 26(3):323–336, September 1984. CODEN AUJSA3. ISSN 0004-9581.

Ratcliff:1984:NAC

[783] D. Ratcliff, E. R. Williams, and T. P. Speed. A note on the analysis of covariance in balanced incomplete block designs. Australian Journal of Statistics, 26(3):337–341, September 1984. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1984:BRc

[784] Anonymous. Book reviews. Australian Journal of Statistics, 26(3):342– 347, September 1984. CODEN AUJSA3. ISSN 0004-9581.

Foreman:1985:SD

[785] E. K. Foreman. Survey design. Australian Journal of Statistics, 27(1): 1–19, April 1985. CODEN AUJSA3. ISSN 0004-9581.

Fischer:1985:DBS

[786] A. J. Fischer and R. D. Lewis. Deliberately biasing survey questions. Australian Journal of Statistics, 27(1):20–35, April 1985. CODEN AU-JSA3. ISSN 0004-9581.

Jarrett:1985:DFA

[787] R. G. Jarrett, S. Liow, and B. J. T. Morgan. Density function approximation with particular reference to the distribution of overlap of circles. Australian Journal of Statistics, 27(1):36–43, April 1985. CODEN AU-JSA3. ISSN 0004-9581.

Hart:1985:DBC

[788] Jeffrey D. Hart. Data-based choice of the smoothing parameter for a kernel density estimator. Australian Journal of Statistics, 27(1):44–52, April 1985. CODEN AUJSA3. ISSN 0004-9581.

Diggle:1985:CTU

[789] P. J. Diggle, N. I. Fisher, and A. J. Lee. A comparison of tests of uniformity for spherical data. Australian Journal of Statistics, 27(1): 53–59, April 1985. CODEN AUJSA3. ISSN 0004-9581.

Moran:1985:CMN

[790] P. A. P. Moran. Calculation of multivariate normal probabilities — another special case. Australian Journal of Statistics, 27(1):60–67, April 1985. CODEN AUJSA3. ISSN 0004-9581.

Evans:1985:CVA

[791] Merran A. Evans and Maxwell L. King. Critical value approximations for tests of linear regression disturbances. Australian Journal of Statistics, 27(1):68–83, April 1985. CODEN AUJSA3. ISSN 0004-9581.

Cameron:1985:NPA

[792] M. A. Cameron and G. K. Eagleson. A new procedure for assessing large sets of correlations. Australian Journal of Statistics, 27(1):84–95, April 1985. CODEN AUJSA3. ISSN 0004-9581.

James:1985:IIU

[793] Ian R. James. On the invariance of isometry under unequal scale changes. Australian Journal of Statistics, 27(1):96–101, April 1985. CODEN AU-JSA3. ISSN 0004-9581.

Anonymous:1985:BRa

[794] Anonymous. Book reviews. Australian Journal of Statistics, 27(1):102– 110, April 1985. CODEN AUJSA3. ISSN 0004-9581.

Kish:1985:SSV

[795] Leslie Kish. Sample surveys versus experiments, controlled observations, censuses, registers, and local studies. Australian Journal of Statistics, 27 (2):111–122, June 1985. CODEN AUJSA3. ISSN 0004-9581.

Brown:1985:GPD

[796] B. M. Brown. Grouping problems in distribution-free regression. Australian Journal of Statistics, 27(2):123–134, June 1985. CODEN AU-JSA3. ISSN 0004-9581.

Chaudhuri:1985:DEF

[797] Arijit Chaudhuri and Bahul Mukerjee. Domain estimation in finite populations. Australian Journal of Statistics, 27(2):135–137, June 1985. CO-DEN AUJSA3. ISSN 0004-9581.

Sharma:1985:ECA

[798] Subhash C. Sharma. The effects of correlation among observations on the accuracy of approximating the distribution of sample mean by its asymptotic distribution. Australian Journal of Statistics, 27(2):138–150, June 1985. CODEN AUJSA3. ISSN 0004-9581.

Barnett:1985:DTD

[799] Vic Barnett. Detection and testing of different types of outlier in linear structural relationaships. Australian Journal of Statistics, 27(2):151–162, June 1985. CODEN AUJSA3. ISSN 0004-9581.

Shirahata:1985:GFT

[800] S. Shirahata, K. Wakimoto, and T. Tarumi. A goodness of fit test based on the linked line chart. Australian Journal of Statistics, 27(2):163–171, June 1985. CODEN AUJSA3. ISSN 0004-9581.

Venable:1985:MCC

[801] W. N. Venable. The multiple correlation coefficient and Fisher’s A statistic. Australian Journal of Statistics, 27(2):172–182, June 1985. CODEN AUJSA3. ISSN 0004-9581.

Haslett:1985:REG

[802] Stephen Haslett. Recursive estimation of the general linear model with dependent errors and multiple additional observations. Australian Journal of Statistics, 27(2):183–188, June 1985. CODEN AUJSA3. ISSN 0004-9581.

Samanta:1985:MME

[803] M. Samanta. Minimum MSE estimation by linear combinations of order statistics. Australian Journal of Statistics, 27(2):189–194, June 1985. CODEN AUJSA3. ISSN 0004-9581.

Kumar:1985:VEU

[804] Pranesh Kumar, V. K. Gupta, and S. K. Agarwal. On variance estimation in unequal probability sampling. Australian Journal of Statistics, 27(2): 195–201, June 1985. CODEN AUJSA3. ISSN 0004-9581.

Lloyd:1985:EA

[805] C. J. Lloyd. On external ancillarity. Australian Journal of Statistics, 27 (2):202–220, June 1985. CODEN AUJSA3. ISSN 0004-9581.

Verbyla:1985:NIC

[806] A. P. Verbyla. A note on the inverse covariance matrix of the autoregressive process. Australian Journal of Statistics, 27(2):221–224, June 1985. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1985:BRb

[807] Anonymous. Book reviews. Australian Journal of Statistics, 27(2):225– 229, June 1985. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1985:BNa

[808] Anonymous. Book notices. Australian Journal of Statistics, 27(2):229– 230, June 1985. CODEN AUJSA3. ISSN 0004-9581.

Breth:1985:ESD

[809] M. Breth. On estimating the spread of a distribution about its regression slopes. Australian Journal of Statistics, 27(3):231–235, September 1985. CODEN AUJSA3. ISSN 0004-9581.

Feeney:1985:EVT

[810] G. A. Feeney and P. K. Sen. Extreme value theory for certain nonstationary sequences. Australian Journal of Statistics, 27(3):236–245, September 1985. CODEN AUJSA3. ISSN 0004-9581.

Gordon:1985:ERO

[811] Ian Gordon. Error rates for orthogonal contrasts after a significant F - test. Australian Journal of Statistics, 27(3):246–251, September 1985. CODEN AUJSA3. ISSN 0004-9581.

Gore:1985:THA

[812] A. P. Gore and Ashok Shanubhogue. A test of homogeneity against scale alternatives using subsample extrema. Australian Journal of Statistics, 27(3):252–258, September 1985. CODEN AUJSA3. ISSN 0004-9581.

Hua:1985:PKS

[813] Tsushung A. Hua. On the performance of Kolmogorov–Smirnov tests on stable random variables. Australian Journal of Statistics, 27(3):259–263, September 1985. CODEN AUJSA3. ISSN 0004-9581.

Lloyd:1985:ASS

[814] C. J. Lloyd. Ancillaries sufficient for the sample size. Australian Journal of Statistics, 27(3):264–272, September 1985. CODEN AUJSA3. ISSN 0004-9581.

Manjunath:1985:CES

[815] S. M. Manjunath. Characterization of efficient sequential schemes for population growth models. Australian Journal of Statistics, 27(3):273– 282, September 1985. CODEN AUJSA3. ISSN 0004-9581.

ONeill:1985:EDU

[816] Terence J. O’Neill, G. M. Tallis, and P. Leppard. The epidemiology of a disease using hazard functions. Australian Journal of Statistics, 27(3): 283–297, September 1985. CODEN AUJSA3. ISSN 0004-9581.

Sharma:1985:RIB

[817] V. K. Sharma and M. N. Das. On resolvable incomplete block designs. Australian Journal of Statistics, 27(3):298–302, September 1985. CO-DEN AUJSA3. ISSN 0004-9581.

Yajima:1985:ELM

[818] Yoshihiro Yajima. On estimation of long-memory time series models. Australian Journal of Statistics, 27(3):303–320, September 1985. CO-DEN AUJSA3. ISSN 0004-9581.

Anonymous:1985:BRc

[819] Anonymous. Book reviews. Australian Journal of Statistics, 27(3):321– 324, September 1985. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1985:BNb

[820] Anonymous. Book notices. Australian Journal of Statistics, 27(3):324– 325, September 1985. CODEN AUJSA3. ISSN 0004-9581.

Kish:1986:TSP

[821] Leslie Kish. Timing of surveys for public policy. Australian Journal of Statistics, 28(1):1–12, April 1986. CODEN AUJSA3. ISSN 0004-9581.

Best:1986:GFD

[822] D. J. Best and N. I. Fisher. Goodness-of-fit and discordancy tests for samples from the Watson distribution on the sphere. Australian Journal of Statistics, 28(1):13–31, April 1986. CODEN AUJSA3. ISSN 0004-9581.

Brown:1986:MSR

[823] G. H. Brown. Multistage sampling for raw wool characteristics. Australian Journal of Statistics, 28(1):32–45, April 1986. CODEN AUJSA3. ISSN 0004-9581.

Scariano:1986:EMM

[824] Stephen M. Scariano. Estimating the maximum (minimum) of a quadratic function under certain correlation patterns. Australian Journal of Statistics, 28(1):46–51, April 1986. CODEN AUJSA3. ISSN 0004-9581.

Maller:1986:SAP

[825] R. A. Maller. Some asymptotic properties of the simple loglinear model. Australian Journal of Statistics, 28(1):52–71, April 1986. CODEN AU-JSA3. ISSN 0004-9581.

Rao:1986:WME

[826] P. V. Rao and J. D. Lees. Weighted median estimates for response surface models. Australian Journal of Statistics, 28(1):72–78, April 1986. CODEN AUJSA3. ISSN 0004-9581.

Ipinyomi:1986:EED

[827] R. A. Ipinyomi. Equineighboured experimental designs. Australian Journal of Statistics, 28(1):79–88, April 1986. CODEN AUJSA3. ISSN 0004-9581.

Brown:1986:CCU

[828] Timothy C. Brown, Donald I. Cartwright, and G. K. Eagleson. Correlations and characterizations of the uniform distribution. Australian Journal of Statistics, 28(1):89–96, April 1986. CODEN AUJSA3. ISSN 0004-9581.

Aerts:1986:CRS

[829] M. Aerts and H. Callaert. The convergence rate of sequential fixedwidth confidence intervals for regular functionals. Australian Journal of Statistics, 28(1):97–106, April 1986. CODEN AUJSA3. ISSN 0004-9581.

Tang:1986:EDC

[830] J. Tang and A. K. Gupta. Exact distribution of certain general test statistics in multivariate analysis. Australian Journal of Statistics, 28 (1):107–114, April 1986. CODEN AUJSA3. ISSN 0004-9581.

Breth:1986:ICN

[831] M. Breth. Independent component nonparamentric estimation. Australian Journal of Statistics, 28(1):115–123, April 1986. CODEN AU-JSA3. ISSN 0004-9581.

Hopper:1986:CWH

[832] John L. Hopper. The contribution of W. H. Archer to vital statistics in the Colony of Victoria. Australian Journal of Statistics, 28(1):124–137, April 1986. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1986:BRa

[833] Anonymous. Book reviews. Australian Journal of Statistics, 28(1):138– 141, April 1986. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1986:BNa

[834] Anonymous. Book notices. Australian Journal of Statistics, 28(1):142, April 1986. CODEN AUJSA3. ISSN 0004-9581.

Daley:1986:MEM

[835] D. J. Daley and E. Seneta. Modelling examination marks. Australian Journal of Statistics, 28(2):143–153, June 1986. CODEN AUJSA3. ISSN 0004-9581.

Williams:1986:RCD

[836] E. R. Williams. Row and column designs with contiguous replicates. Australian Journal of Statistics, 28(2):154–163, June 1986. CODEN AU-JSA3. ISSN 0004-9581.

Samanta:1986:AOS

[837] M. Samanta. On asymptotic optimality of some tests for exponential distributions. Australian Journal of Statistics, 28(2):164–172, June 1986. CODEN AUJSA3. ISSN 0004-9581.

Welsh:1986:UED

[838] A. H. Welsh. On the use of the empirical distribution and characteristic function to estimate parameters of regular variation. Australian Journal of Statistics, 28(2):173–181, June 1986. CODEN AUJSA3. ISSN 0004-

Williams:1986:NAU

[839] E. R. Williams. Neighbour analysis of uniformity data. Australian Journal of Statistics, 28(2):182–191, June 1986. CODEN AUJSA3. ISSN 0004-9581.

Watson:1986:QS

[840] Ray Watson and Ian Gordon. On quantiles of sums. Australian Journal of Statistics, 28(2):192–199, June 1986. CODEN AUJSA3. ISSN 0004-9581.

Ceranka:1986:OSS

[841] B. Ceranka and K. Katulska. Optimum singular spring balance weighing designs with non-homogeneity of the variances of errors for estimating the total weight. Australian Journal of Statistics, 28(2):200–205, June 1986. CODEN AUJSA3. ISSN 0004-9581.

Jones:1986:MI

[842] M. C. Jones. A moment identity. Australian Journal of Statistics, 28(2): 206–212, June 1986. CODEN AUJSA3. ISSN 0004-9581.

Fisher:1986:RCD

[843] N. I. Fisher. Robust comparison of dispersions for samples of directional data. Australian Journal of Statistics, 28(2):213–219, June 1986. CO-DEN AUJSA3. ISSN 0004-9581.

Fitzpatrick:1986:UBH

[844] Simon Fitzpatrick and Richard G. Jarrett. Upper bounds for the harmonic mean, with an application to experimental design. Australian Journal of Statistics, 28(2):220–229, June 1986. CODEN AUJSA3. ISSN 0004-9581.

Yitnosumarto:1986:LTV

[845] S. Yitnosumarto and M. E. O’Neill. On Levene’s tests of variance homogeneity. Australian Journal of Statistics, 28(2):230–241, June 1986. CODEN AUJSA3. ISSN 0004-9581.

Jones:1986:EIM

[846] M. C. Jones. Expressions for inverse moments of positive quadratic forms in normal variables. Australian Journal of Statistics, 28(2):242–250, June 1986. CODEN AUJSA3. ISSN 0004-9581.

Agarwal:1986:CBT

[847] G. G. Agarwal and Sanjay Kumar. Construction of balanced ternary designs. Australian Journal of Statistics, 28(2):251–255, June 1986. CO-DEN AUJSA3. ISSN 0004-9581.

Anonymous:1986:BRb

[848] Anonymous. Book reviews. Australian Journal of Statistics, 28(2):256– 264, June 1986. CODEN AUJSA3. ISSN 0004-9581.

Tweedie:1986:WOS

[849] R. L. Tweedie. Working with official statistics: Some SIROMATH experiences. Australian Journal of Statistics, 28(3):265–286, September 1986. CODEN AUJSA3. ISSN 0004-9581. See rejoinder [850].

Trewin:1986:WOS

[850] Dennis Trewin and David Leaver. Working with official statistics: Some SIROMATH experience — a rejoinder. Australian Journal of Statistics, 28(3):287–293, September 1986. CODEN AUJSA3. ISSN 0004-9581. See [849].

Nicholls:1986:ELT

[851] D. F. Nicholls, C. R. Heathcote, and R. B. Cunningham. The evaluation of long term trend I. Australian Journal of Statistics, 28(3):294–313, September 1986. CODEN AUJSA3. ISSN 0004-9581.

Ralescu:1986:ARM

[852] Stefan Ralescu. Asymptotic representations of the multivariate empirical distribution function and applications. Australian Journal of Statistics, 28(3):314–334, September 1986. CODEN AUJSA3. ISSN 0004-9581.

Chou:1986:SST

[853] Rouh jane Chou. Small sample theory of the Langevin distribution. Australian Journal of Statistics, 28(3):335–344, September 1986. CODEN AUJSA3. ISSN 0004-9581.

Tam:1986:OPS

[854] S. M. Tam. Optimal prediction in stochastic regression models with application to the analysis of repeated surveys. Australian Journal of Statistics, 28(3):345–353, September 1986. CODEN AUJSA3. ISSN 0004-9581.

Hall:1986:SSB

[855] Peter Hall and Ben Selinger. Statistical significance: Balancing evidence against doubt. Australian Journal of Statistics, 28(3):354–370, September 1986. CODEN AUJSA3. ISSN 0004-9581.

Darroch:1986:CDO

[856] J. N. Darroch and P. I. McCloud. Category distinguishability and observer agreement. Australian Journal of Statistics, 28(3):371–388, September 1986. CODEN AUJSA3. ISSN 0004-9581.

Verbyla:1986:NPL

[857] A. P. Verbyla. A note on a partition of the likelihood ratio test for autoregressive covariance structure. Australian Journal of Statistics, 28 (3):389–399, September 1986. CODEN AUJSA3. ISSN 0004-9581.

Schmid:1986:MOP

[858] Wolfgang Schmid. The multiple outlier problem in time series analysis. Australian Journal of Statistics, 28(3):400–413, September 1986. CO-DEN AUJSA3. ISSN 0004-9581.

Kourouklis:1986:SCP

[859] Stavros Kourouklis. On a simultaneous characterization of the Poisson and Bernoulli distributions. Australian Journal of Statistics, 28(3):414– 417, September 1986. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1986:BRc

[860] Anonymous. Book reviews. Australian Journal of Statistics, 28(3):418– 423, September 1986. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1986:BNb

[861] Anonymous. Book notices. Australian Journal of Statistics, 28(3):423, September 1986. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1986:C

[862] Anonymous. Correction. Australian Journal of Statistics, 28(3):424, September 1986. CODEN AUJSA3. ISSN 0004-9581.

Hettmansperger:1987:WTR

[863] Thomas P. Hettmansperger. Why not try a robust regression? Australian Journal of Statistics, 29(1):1–18, April 1987. CODEN AUJSA3. ISSN 0004-9581.

Hall:1987:SSM

[864] Peter Hall and D. M. Titterington. On smoothing sparse multinomial data. Australian Journal of Statistics, 29(1):19–37, April 1987. CODEN AUJSA3. ISSN 0004-9581.

Rao:1987:CPD

[865] B. L. S. Prakasa Rao and M. Sreehari. On a characterization of Poisson distribution through inequalities of Chernoff-type. Australian Journal of Statistics, 29(1):38–41, April 1987. CODEN AUJSA3. ISSN 0004-9581.

Peiris:1987:NPD

[866] M. Shelton Peiris. A note on the predictors of differenced sequences. Australian Journal of Statistics, 29(1):42–48, April 1987. CODEN AUJSA3. ISSN 0004-9581.

SenGupta:1987:TEC

[867] Ashis SenGupta. On tests for equicorrelation coefficient of a standard symmetric multivariate normal distribution. Australian Journal of Statistics, 29(1):49–59, April 1987. CODEN AUJSA3. ISSN 0004-9581.

Rogers:1987:DER

[868] Alan J. Rogers. On the derivation of efficiency results in linear estimation. Australian Journal of Statistics, 29(1):60–65, April 1987. CODEN AUJSA3. ISSN 0004-9581.

Sheather:1987:NME

[869] Simon J. Sheather. A new method of estimating the asymptotic standard error of the Hodges–Lehmann estimator based on generalized least squares. Australian Journal of Statistics, 29(1):66–83, April 1987. CO-DEN AUJSA3. ISSN 0004-9581.

Russell:1987:COB

[870] K. G. Russell and J. A. Eccleston. The construction of optimal balanced incomplete block designs when adjacent observations are correlated. Australian Journal of Statistics, 29(1):84–90, April 1987. CODEN AUJSA3. ISSN 0004-9581.

Oja:1987:PTM

[871] Hannu Oja. On permutation tests in multiple regression and analysis of covariance problems. Australian Journal of Statistics, 29(1):91–100, April 1987. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1987:FVC

[872] H. O. Lancaster. Finiteness of the variances in characterizations of the normal distribution. Australian Journal of Statistics, 29(1):101–106, April 1987. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1987:BRa

[873] Anonymous. Book reviews. Australian Journal of Statistics, 29(1):107– 112, April 1987. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1987:BNa

[874] Anonymous. Book notices. Australian Journal of Statistics, 29(1):112, April 1987. CODEN AUJSA3. ISSN 0004-9581.

Laurence:1987:STV

[875] Alexa F. Laurence and Byron J. T. Morgan. Selection of the transformation variable in the Laplace transform method of estimation. Australian Journal of Statistics, 29(2):113–127, June 1987. CODEN AUJSA3. ISSN 0004-9581.

Okafor:1987:SST

[876] Fabian C. Okafor and Raghunath Arnab. Some strategies of two-stage sampling for estimating population ratios over two occasions. Australian Journal of Statistics, 29(2):128–142, June 1987. CODEN AUJSA3. ISSN 0004-9581.

Becker:1987:NIP

[877] Niels Becker and Paul Yip. Nonparametric inference for a partially observed compartmental process. Australian Journal of Statistics, 29(2): 143–150, June 1987. CODEN AUJSA3. ISSN 0004-9581.

Hirotsu:1987:OAO

[878] C. Hirotsu and Agnes M. Herzberg. Optimal allocation of observations for inference on k ordered normal population means. Australian Journal of Statistics, 29(2):151–165, June 1987. CODEN AUJSA3. ISSN 0004-9581.

Csorgo:1987:OET

[879] Miklós Csörgö, Lnos Horváth, and Pál Révész. On the optimality of estimating the tail index and a naive estimator. Australian Journal of Statistics, 29(2):166–178, June 1987. CODEN AUJSA3. ISSN 0004-9581.

Kokic:1987:TUR

[880] P. N. Kokic. On tests of uniformity for randomly distributed arcs on a circle. Australian Journal of Statistics, 29(2):179–187, June 1987. CO-DEN AUJSA3. ISSN 0004-9581.

Deshpande:1987:GMS

[881] M. N. Deshpande and S. G. Prabhu-Ajgaonkar. A generalisation of Midzuno’s sampling procedure. Australian Journal of Statistics, 29(2): 188–192, June 1987. CODEN AUJSA3. ISSN 0004-9581.

Kijima:1987:SRU

[882] Masaaki Kijima. Some results for uniformizable semi-Markov processes. Australian Journal of Statistics, 29(2):193–207, June 1987. CODEN AU-JSA3. ISSN 0004-9581.

Adegboye:1987:CRW

[883] O. S. Adegboye. A classification rule whose probability of misclassification is independent of the variance. Australian Journal of Statistics, 29 (2):208–213, June 1987. CODEN AUJSA3. ISSN 0004-9581.

Wei:1987:OSR

[884] L. J. Wei and M. W. Knuiman. A one-sided rank test for multivariate censored data. Australian Journal of Statistics, 29(2):214–219, June 1987. CODEN AUJSA3. ISSN 0004-9581.

Robinson:1987:LEI

[885] G. K. Robinson. Letter to the Editor: Instructions to referees — do we need some? Australian Journal of Statistics, 29(2):220–221, June 1987. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1987:BRb

[886] Anonymous. Book reviews. Australian Journal of Statistics, 29(2):222– 230, June 1987. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1987:BNb

[887] Anonymous. Book notices. Australian Journal of Statistics, 29(2):231, June 1987. CODEN AUJSA3. ISSN 0004-9581.

Lee:1987:DDA

[888] Andy H. Lee. Diagnostic displays for assessing leverage and influence in generalized linear models. Australian Journal of Statistics, 29(3):233– 243, September 1987. CODEN AUJSA3. ISSN 0004-9581.

Mukerjee:1987:RTE

[889] Rahul Mukerjee, T. J. Rao, and K. Vijayan. Regression type estimators using multiple auxiliary information. Australian Journal of Statistics, 29 (3):244–254, September 1987. CODEN AUJSA3. ISSN 0004-9581.

Jackway:1987:RIT

[890] Paul T. Jackway and Rosalie A. Boyce. Response inducing techniques for Mail surveys. Australian Journal of Statistics, 29(3):255–263, September 1987. CODEN AUJSA3. ISSN 0004-9581.

Copas:1987:RSM

[891] J. B. Copas and M. C. Jones. Regression shrinkage methods and autoregressive time series prediction. Australian Journal of Statistics, 29(3): 264–277, September 1987. CODEN AUJSA3. ISSN 0004-9581.

Morgenthaler:1987:CIS

[892] Stephan Morgenthaler. Confidence intervals for scale. Australian Journal of Statistics, 29(3):278–292, September 1987. CODEN AUJSA3. ISSN 0004-9581.

Russell:1987:COI

[893] K. G. Russell and J. A. Eccleston. The construction of optimal imcomplete block designs when observations within a block are correlated. Australian Journal of Statistics, 29(3):293–302, September 1987. CO-DEN AUJSA3. ISSN 0004-9581.

Collins:1987:PTP

[894] Mark F. Collins. A permutation test for planar regression. Australian Journal of Statistics, 29(3):303–308, September 1987. CODEN AUJSA3. ISSN 0004-9581.

Williams:1987:NCD

[895] E. R. Williams. A note on change-over designs. Australian Journal of Statistics, 29(3):309–316, September 1987. CODEN AUJSA3. ISSN 0004-9581.

Knuiman:1987:IPR

[896] M. W. Knuiman and T. P. Speed. Incorporating previous results into the analysis of general balanced designed experiments. Australian Journal of Statistics, 29(3):317–333, September 1987. CODEN AUJSA3. ISSN 0004-9581.

Poskitt:1987:BAM

[897] D. S. Poskitt. Bayes ARMA model determination: Some empirical evidence. Australian Journal of Statistics, 29(3):334–347, September 1987. CODEN AUJSA3. ISSN 0004-9581.

Bofinger:1987:SPC

[898] Eve Bofinger. Step down procedures for comparison with a control. Australian Journal of Statistics, 29(3):348–364, September 1987. CODEN AUJSA3. ISSN 0004-9581.

Kageyama:1987:LER

[899] Sanpei Kageyama. Letter to the Editor: Recommendation of binary designs. Australian Journal of Statistics, 29(3):365–366, September 1987. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1987:BRBa

[900] Anonymous. Books reviews: Books reviews. Australian Journal of Statistics, 29(3):367–370, September 1987. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1987:BNc

[901] Anonymous. Book notices. Australian Journal of Statistics, 29(3):370– 371, September 1987. CODEN AUJSA3. ISSN 0004-9581.

Duffy:1988:RRV

[902] John C. Duffy and Jennifer J. Waterton. Randomised response vs. direct questioning: Estimating the prevalence of alcohol related problems in a field survey. Australian Journal of Statistics, 30(1):1–14, March 1988. CODEN AUJSA3. ISSN 0004-9581.

Speed:1988:RSC

[903] T. P. Speed. The role of statisticians in CSIRO: Past, present and future. Australian Journal of Statistics, 30(1):15–34, March 1988. CODEN AUJSA3. ISSN 0004-9581.

Lloyd:1988:SIA

[904] C. J. Lloyd. Some issues arising from the analysis of 2×2 contingency tables. Australian Journal of Statistics, 30(1):35–46, March 1988. CODEN AUJSA3. ISSN 0004-9581.

Ghosh:1988:SSE

[905] D. K. Ghosh and P. K. Karmakar. Some series of efficiency balanced designs. Australian Journal of Statistics, 30(1):47–51, March 1988. CO-DEN AUJSA3. ISSN 0004-9581.

Smith:1988:APL

[906] Peter James Smith. Asymptotic properties of linear regression estimators under a fixed censorship model. Australian Journal of Statistics, 30(1): 52–66, March 1988. CODEN AUJSA3. ISSN 0004-9581.

Arsham:1988:NRE

[907] H. Arsham. Nonparametric reliability estimation based on a few ordered observations. Australian Journal of Statistics, 30(1):67–77, March 1988. CODEN AUJSA3. ISSN 0004-9581.

Eccleston:1988:AON

[908] J. A. Eccleston and J. A. John. Adjusted orthogonal nested row-column designs. Australian Journal of Statistics, 30(1):78–84, March 1988. CO-DEN AUJSA3. ISSN 0004-9581.

McGilchrist:1988:PPV

[909] C. A. McGilchrist and B. R. Cullis. Properties of the patterned vech operator. Australian Journal of Statistics, 30(1):85–92, March 1988. CODEN AUJSA3. ISSN 0004-9581.

Phatarfod:1988:SSW

[910] R. M. Phatarfod and Aidan Sudbury. A simple sequential Wilcoxon test. Australian Journal of Statistics, 30(1):93–106, March 1988. CODEN AUJSA3. ISSN 0004-9581.

Arnab:1988:VEM

[911] R. Arnab. Variance estimation in multi-stage sampling. Australian Journal of Statistics, 30(1):107–110, March 1988. CODEN AUJSA3. ISSN 0004-9581.

Kourouklis:1988:AOL

[912] Stavros Kourouklis. Asymptotic optimality of likelihood ratio tests for exponential distributions under Type II censoring. Australian Journal of Statistics, 30(1):111–114, March 1988. CODEN AUJSA3. ISSN 0004-9581.

Anderson:1988:LER

[913] Oliver D. Anderson. Letter to the Editor: Referees are also human. Australian Journal of Statistics, 30(1):115–116, March 1988. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1988:BRa

[914] Anonymous. Book reviews. Australian Journal of Statistics, 30(1):117– 121, March 1988. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1988:BNa

[915] Anonymous. Book notices. Australian Journal of Statistics, 30(1):121– 123, March 1988. CODEN AUJSA3. ISSN 0004-9581.

Anderson:1988:SMO

[916] Dorothy A. Anderson. Some models for overdispersed binomial data. Australian Journal of Statistics, 30(2):125–148, June 1988. CODEN AU-JSA3. ISSN 0004-9581.

Chambers:1988:MLE

[917] R. L. Chambers, L. Woyzbun, and R. Pillig. Maximum likelihood estimation of Gross flows. Australian Journal of Statistics, 30(2):149–162, June 1988. CODEN AUJSA3. ISSN 0004-9581.

Lloyd:1988:AEB

[918] Christopher J. Lloyd. Applications and extensions of Basu’s results on maximal ancillarity. Australian Journal of Statistics, 30(2):163–173, June 1988. CODEN AUJSA3. ISSN 0004-9581.

Meister:1988:RRS

[919] H. Meister and O. Moeschlin. The reduction of risks for set-valued estimators. Australian Journal of Statistics, 30(2):174–182, June 1988. CODEN AUJSA3. ISSN 0004-9581.

Klebaner:1988:URW

[920] Fima C. Klebaner. The unemployment rate when the labour market is modelled by a Markov chain. Australian Journal of Statistics, 30(2): 183–188, June 1988. CODEN AUJSA3. ISSN 0004-9581.

Anh:1988:HHE

[921] V. V. Anh. On the Hildreth-houck estimator for random coefficient regression models. Australian Journal of Statistics, 30(2):189–195, June 1988. CODEN AUJSA3. ISSN 0004-9581.

Sadooghi-Alvandi:1988:NEP

[922] S. M. Sadooghi-Alvandi. A note on the efficiency of proportional stratification. Australian Journal of Statistics, 30(2):196–199, June 1988. CO-DEN AUJSA3. ISSN 0004-9581.

Giaccotto:1988:JTH

[923] Carmelo Giaccotto and Subhash C. Sharma. Jackknife tests for heteroscedasticity in the general linear model. Australian Journal of Statistics, 30(2):200–216, June 1988. CODEN AUJSA3. ISSN 0004-9581.

Nakano:1988:DEH

[924] Junji Nakano and Shigemi Tagami. Delay estimation by a Hilbert transform method. Australian Journal of Statistics, 30(2):217–227, June 1988. CODEN AUJSA3. ISSN 0004-9581.

Lianto:1988:CDV

[925] S. Lianto and C. A. McGilchrist. Cholesky decomposition of a variance matrix in repeated measures analysis. Australian Journal of Statistics, 30(2):228–234, June 1988. CODEN AUJSA3. ISSN 0004-9581.

Smith:1988:CSE

[926] Murray D. Smith. Convergent series expressions for inverse moments of quadratic forms in normal variables. Australian Journal of Statistics, 30 (2):235–246, June 1988. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1988:BRb

[927] Anonymous. Book reviews. Australian Journal of Statistics, 30(2):247– 259, June 1988. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1988:BNb

[928] Anonymous. Book notices. Australian Journal of Statistics, 30(2):259– 261, June 1988. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1988:C

[929] Anonymous. Correction. Australian Journal of Statistics, 30(2):262, June 1988. CODEN AUJSA3. ISSN 0004-9581.

Danaher:1988:PEA

[930] P. J. Danaher. Parameter estimation and applications for a generalisation of the beta-binomial distribution. Australian Journal of Statistics, 30(3): 263–275, September 1988. CODEN AUJSA3. ISSN 0004-9581.

Brown:1988:KTC

[931] B. M. Brown. Kendall’s Tau and contingency tables. Australian Journal of Statistics, 30(3):276–291, September 1988. CODEN AUJSA3. ISSN 0004-9581.

Kshirsagar:1988:NML

[932] Anant M. Kshirsagar. A note on multivariate linear models with nonadditivity. Australian Journal of Statistics, 30(3):292–298, September 1988. CODEN AUJSA3. ISSN 0004-9581.

Yip:1988:IAM

[933] Paul Yip. Inference about the mean of a Poisson distribution in the presence of a nuisance parameter. Australian Journal of Statistics, 30 (3):299–306, September 1988. CODEN AUJSA3. ISSN 0004-9581.

Gupta:1988:NDL

[934] A. K. Gupta and D. K. Nagar. Nonnull distribution of likelihood ratio criterion for testing multisample sphericity in the complex case. Australian Journal of Statistics, 30(3):307–318, September 1988. CODEN AUJSA3. ISSN 0004-9581.

MacGillivray:1988:RBS

[935] H. L. MacGillivray and K. P. Balanda. The relationships between skewness and kurtosis. Australian Journal of Statistics, 30(3):319–337, September 1988. CODEN AUJSA3. ISSN 0004-9581.

Bolfarine:1988:SPP

[936] Heleno Bolfarine and Josemar Rodrigues. On the simple projection predictor in finite populations. Australian Journal of Statistics, 30(3):338– 341, September 1988. CODEN AUJSA3. ISSN 0004-9581.

Yandell:1988:BMD

[937] Brian S. Yandell and Lajos Horváth. Bootstrapped multi-dimensional product limit process. Australian Journal of Statistics, 30(3):342–358, September 1988. CODEN AUJSA3. ISSN 0004-9581.

Gupta:1988:GDT

[938] A. K. Gupta and J. Tang. On a general distribution theory for a class of likelihood ratio criteria. Australian Journal of Statistics, 30(3):359–366, September 1988. CODEN AUJSA3. ISSN 0004-9581.

Okafor:1988:RSE

[939] Fabian C. Okafor. Ratio-to-size estimators of mean per subunit in twostage sampling over two occasions. Australian Journal of Statistics, 30 (3):367–378, September 1988. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1988:BRc

[940] Anonymous. Book reviews. Australian Journal of Statistics, 30(3):379– 387, September 1988. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1988:BNc

[941] Anonymous. Book notices. Australian Journal of Statistics, 30(3):388– 389, September 1988. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1988:PJG

[942] Anonymous. Publications of J. Gani. Australian Journal of Statistics, 30A(1):1–13, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Moran:1988:CES

[943] P. A. P. Moran. Combinations of events: Some notes on permanents. Australian Journal of Statistics, 30A(1):17–20, May 1988. CODEN AU-JSA3. ISSN 0004-9581.

Raade:1988:CEG

[944] Lennart Råde. Combinations of events: On games with D- and T -events. Australian Journal of Statistics, 30A(1):21–26, May 1988. CODEN AU-JSA3. ISSN 0004-9581.

Seneta:1988:CED

[945] E. Seneta. Combinations of events: Degree, iteration and permutation in improving Bonferroni-type bounds. Australian Journal of Statistics, 30A(1):27–38, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Bailey:1988:ESP

[946] Norman T. J. Bailey. Epidemics: Statistical problems in the modelling and prediction of HIV/AIDS. Australian Journal of Statistics, 30A(1): 41–55, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Dietz:1988:EFE

[947] K. Dietz. Epidemics: The first epidemic model: a historical note on P. D. En’ko. Australian Journal of Statistics, 30A(1):56–65, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Blake:1988:FMM

[948] J. R. Blake. Fitting and modelling: Mathematical theory of cavitation bubble dynamics. Australian Journal of Statistics, 30A(1):69–89, May 1988. CODEN AUJSA3. ISSN 0004-9581.

deHoog:1988:FMC

[949] F. R. de Hoog and R. S. Anderssen. Fitting and modelling: Convergence of kernel functions for cubic smoothing splines on non-equispaced grids. Australian Journal of Statistics, 30A(1):90–99, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Ewens:1988:FMP

[950] W. J. Ewens. Fitting and modelling: Problems of statistical modelling in human genetics. Australian Journal of Statistics, 30A(1):100–106, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Gates:1988:FMP

[951] David Gates and Mark Westcott. Fitting and modelling: Point processes with a clustering transition. Australian Journal of Statistics, 30A(1): 107–122, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Vere-Jones:1988:FMV

[952] D. Vere-Jones. Fitting and modelling: On the variance properties of stress release models. Australian Journal of Statistics, 30A(1):123–135, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Bock:1988:QFC

[953] Mary Ellen Bock and Herbert Solomon. Quadratic forms and chisquared: Distributions of quadratic forms. Australian Journal of Statistics, 30A(1):139–149, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Buckley:1988:QFC

[954] M. J. Buckley and G. K. Eagleson. Quadratic forms and chi-squared: an approximation to the distribution of quadratic forms in normal random variables. Australian Journal of Statistics, 30A(1):150–159, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Lewis:1988:QFC

[955] Toby Lewis. Quadratic forms and chi-squared: a simple improvedaccuracy normal approximation for χ^{2}. Australian Journal of Statistics, 30A(1):160–167, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Becker:1988:SIR

[956] Niels Becker. Statistical inference: The role of conditional inference. Australian Journal of Statistics, 30A(1):171–178, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Hall:1988:SIB

[957] Peter Hall and Michael Martin. Statistical inference: On the bootstrap and two-sample problems. Australian Journal of Statistics, 30A(1):179– 192, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Heathcote:1988:SIP

[958] C. R. Heathcote. Statistical inference: The problem of misclassification of mortality statistics. Australian Journal of Statistics, 30A(1):193–204, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Johnson:1988:SIE

[959] Norman L. Johnson and Samuel Kotz. Statistical inference: Effects of errors in inspection on a binary method for isolating nonconforming items. Australian Journal of Statistics, 30A(1):205–214, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Morton:1988:SIA

[960] Richard Morton. Statistical inference: Analysis of generalized linear models with nested strata of variation. Australian Journal of Statistics, 30A(1):215–224, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Wood:1988:SIN

[961] J. T. Wood, E. R. Williams, and T. P. Speed. Statistical inrefence: Nonorthogonal block structure in two-phase designs. Australian Journal of Statistics, 30A(1):225–237, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Hida:1988:SFW

[962] Takewki Hida. Stochastic functionals: White noise analysis and stochastic functional differential equations. Australian Journal of Statistics, 30A (1):241–246, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Ito:1988:SFD

[963] Yoshifusa Ito. Stochastic functionals: Differential operators arising from translation of Poisson functionals. Australian Journal of Statistics, 30A (1):247–258, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Resnick:1988:SDA

[964] Sidney I. Resnick. Structure of distributions: Association and multivariate extreme value distributions. Australian Journal of Statistics, 30A(1): 261–271, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Shanbhag:1988:SDD

[965] D. N. Shanbhag. Structure of distributions: The Davidson–Kendall problem and related results on the structure of the Wishart distribution. Australian Journal of Statistics, 30A(1):272–280, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Hannan:1988:TSR

[966] E. J. Hannan and C. H. Hesse. Time series: Rates of convergence for the quantile function of a linear process. Australian Journal of Statistics, 30A(1):283–295, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Nicholls:1988:TSB

[967] D. F. Nicholls and A. L. Pope. Time series: Bias in the estimation of multivariate autoregressions. Australian Journal of Statistics, 30A(1): 296–309, May 1988. CODEN AUJSA3. ISSN 0004-9581.

Heyde:1988:PBH

[968] C. C. Heyde and E. Seneta. Preface: The bicentennial history issue of the Australian Journal of Statistics. Australian Journal of Statistics, 30B (1):1, August 1988. CODEN AUJSA3. ISSN 0004-9581.

Seneta:1988:SEA

[969] E. Seneta. Silhouettes in early Australian statistics. Australian Journal of Statistics, 30B(1):2–22, August 1988. CODEN AUJSA3. ISSN 0004-9581.

Heyde:1988:OSL

[970] C. C. Heyde. Official statistics in the Late Colonial Period leading on to the work of the first Commonwealth statistician, G. H. Knibbs. Australian Journal of Statistics, 30B(1):23–43, August 1988. CODEN AU-JSA3. ISSN 0004-9581.

Kerr:1988:ISD

[971] John D. Kerr. Introduction of statistical design and analysis by the Queensland Bureau of Sugar Experiment Stations. Australian Journal of Statistics, 30B(1):44–53, August 1988. CODEN AUJSA3. ISSN 0004-9581.

Field:1988:BC

[972] J. B. F. Field, F. E. Speed, T. P. Speed, and J. M. Williams. Biometrics in the CSIR: 1930–1940. Australian Journal of Statistics, 30B(1):54–76, August 1988. CODEN AUJSA3. ISSN 0004-9581.

Duncan:1988:ABM

[973] D. B. Duncan. Australian biometry and multiple comparisons. Australian Journal of Statistics, 30B(1):77–98, August 1988. CODEN AUJSA3. ISSN 0004-9581.

Lancaster:1988:SSN

[974] H. O. Lancaster. Statistical Society of New South Wales. Australian Journal of Statistics, 30B(1):99–109, August 1988. CODEN AUJSA3. ISSN 0004-9581.

Williams:1988:SED

[975] E. J. Williams. A survey of experimental design in Australia. Australian Journal of Statistics, 30B(1):110–130, August 1988. CODEN AUJSA3. ISSN 0004-9581.

Dunstan:1989:EDF

[976] R. Dunstan and R. L. Chambers. Estimating distribution functions from survey data with limited benchmark information. Australian Journal of Statistics, 31(1):1–11, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Cunningham:1989:ENG

[977] R. B. Cunningham and C. R. Heathcote. Estimating a non-Gaussian regression model with multicollinearity. Australian Journal of Statistics, 31(1):12–17, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Solomon:1989:CVM

[978] P. J. Solomon. On components of variance and modelling exceedances over a threshold. Australian Journal of Statistics, 31(1):18–24, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Wise:1989:EPC

[979] Keith Wise and Paul Kabaila. The estimation of plutonium concentrations in soil. Australian Journal of Statistics, 31(1):25–41, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Becker:1989:AVI

[980] Niels Becker and Paul Yip. Analysis of variations in an infection rate. Australian Journal of Statistics, 31(1):42–52, March 1989. CODEN AU-JSA3. ISSN 0004-9581.

Rogers:1989:MDI

[981] Alan J. Rogers. Minimum distance and induced statistics for testing inequalities. Australian Journal of Statistics, 31(1):53–61, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Kiivefu:1989:CML

[982] H. T. Kiivefu and N. A. Campbell. Covariance models for lattice data. Australian Journal of Statistics, 31(1):62–77, March 1989. CODEN AU-JSA3. ISSN 0004-9581.

Edgeman:1989:IGC

[983] Rick L. Edgeman. Inverse Gaussian control charts. Australian Journal of Statistics, 31(1):78–84, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Ipinyomi:1989:PED

[984] R. A. Ipinyomi. Partially equineighboured designs. Australian Journal of Statistics, 31(1):85–94, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Govindarajulu:1989:UAE

[985] Z. Govindarajulu and H. C. Howard. Uniform asymptotic expansions applied to sequential t- and t^{2}-tests. Australian Journal of Statistics, 31 (1):95–104, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Chen:1989:RTS

[986] Shande Chen and Govind S. Mudholkar. A remark on testing significance of an observed correlation matrix. Australian Journal of Statistics, 31 (1):105–110, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Nguyen:1989:CAC

[987] Nam-Ky Nguyen and Aloke Dey. Computer aided-construction of Doptimal 2^{m} fractional designs of resolution V. Australian Journal of Statistics, 31(1):111–117, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Jarrett:1989:RBE

[988] Richard Jarrett. A review of bounds for the efficiency factor of block designs. Australian Journal of Statistics, 31(1):118–129, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Hagwood:1989:RTM

[989] Charles Hagwood. A renewal theorem in multidimensional time. Australian Journal of Statistics, 31(1):130–137, March 1989. CODEN AU-JSA3. ISSN 0004-9581.

Williams:1989:CIA

[990] Evan J. Williams and Paul Yip. Conditional inference about an equicorrelation coefficient. Australian Journal of Statistics, 31(1):138–142, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Rossiter:1989:STD

[991] M. H. Rossiter. The sojourn time distribution of an alternating renewal process. Australian Journal of Statistics, 31(1):143–152, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Huggins:1989:STS

[992] R. M. Huggins. The sign test for stochastic processes. Australian Journal of Statistics, 31(1):153–165, March 1989. CODEN AUJSA3. ISSN 0004-

Diggle:1989:SSA

[993] Peter J. Diggle and Michael F. Hutchinson. On spline smoothing with autocorrelated errors. Australian Journal of Statistics, 31(1):166–182, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Diggle:1989:SBA

[994] Peter J. Diggle and John B. Donnelly. A selected bibliography on the analysis of repeated measurements and related areas. Australian Journal of Statistics, 31(1):183–193, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Morton:1989:EQL

[995] R. Morton. On the efficiency of the quasi-likelihood estimators for exponential families with extra variation. Australian Journal of Statistics, 31 (1):194–199, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Robinson:1989:LTS

[996] J. Robinson. Limit theorems for standardized partial sums of weakly exchangeable arrays. Australian Journal of Statistics, 31(1):200–214, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Welsh:1989:CSE

[997] A. H. Welsh. Concomitant scale estimation in regression problems with increasing dimension. Australian Journal of Statistics, 31(1):215–227, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1989:BRa

[998] Anonymous. Book reviews. Australian Journal of Statistics, 31(1):228– 247, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1989:BNa

[999] Anonymous. Book notices. Australian Journal of Statistics, 31(1):247– 250, March 1989. CODEN AUJSA3. ISSN 0004-9581.

Hannah:1989:ECP

[1000] Murray Hannah, Niels Becker, Jacob Rich, and Charles Castle. Estimation of the conditional probability of tuberculosis given the response of a Mantoux test. Australian Journal of Statistics, 31(2):253–267, June 1989. CODEN AUJSA3. ISSN 0004-9581.

Singh:1989:MES

[1001] Randhir Singh. Method of estimation for sampling from incomplete frames. Australian Journal of Statistics, 31(2):269–276, June 1989. CO-DEN AUJSA3. ISSN 0004-9581.

Katulska:1989:ETW

[1002] Krystyna Katulska. On the estimation of total weight in singular spring balance weighing designs under the covariance matrix of errors s^{2}G. Australian Journal of Statistics, 31(2):277–286, June 1989. CODEN AU-JSA3. ISSN 0004-9581.

Tomizawa:1989:DCS

[1003] Sadao Tomizawa. Decompositions for conditional symmetry model into palindromic symmetry and modified marginal homogeneity models. Australian Journal of Statistics, 31(2):287–296, June 1989. CODEN AU-JSA3. ISSN 0004-9581.

Mengersen:1989:SBP

[1004] Kerrie Mengersen and Eve Bofinger. Selecting the t best populations — scale parameter case. Australian Journal of Statistics, 31(2):297–307, June 1989. CODEN AUJSA3. ISSN 0004-9581. See corrigendum [1047].

Kijima:1989:TPS

[1005] Masaaki Kijima. On transition probabilities of skip-free Markov chains. Australian Journal of Statistics, 31(2):309–314, June 1989. CODEN AU-JSA3. ISSN 0004-9581.

Karunamuni:1989:EBS

[1006] Rohan Karunamuni. On empirical Bayes stopping times. Australian Journal of Statistics, 31(2):315–326, June 1989. CODEN AUJSA3. ISSN 0004-9581.

Rohatgi:1989:UEP

[1007] Vijay K. Rohatgi. Unbiased estimation of parametric functions in sampling from two one-truncation parameter families. Australian Journal of Statistics, 31(2):327–332, June 1989. CODEN AUJSA3. ISSN 0004-9581.

Zergaw:1989:SMC

[1008] Getachew Zergaw. A sequential method of constructing optimal block designs. Australian Journal of Statistics, 31(2):333–342, June 1989. CO-DEN AUJSA3. ISSN 0004-9581.

Kokic:1989:BEB

[1009] P. N. Kokic. A Berry–Esseen bound for the ratio estimator. Australian Journal of Statistics, 31(2):343–348, June 1989. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1989:BRb

[1010] Anonymous. Book reviews. Australian Journal of Statistics, 31(2):349– 360, June 1989. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1989:BNb

[1011] Anonymous. Book notices. Australian Journal of Statistics, 31(2):360– 362, June 1989. CODEN AUJSA3. ISSN 0004-9581.

Adams:1989:DTP

[1012] Geoff Adams and Ray Watson. A discrete time parametric model for the analysis of failure time data. Australian Journal of Statistics, 31(3): 365–384, September 1989. CODEN AUJSA3. ISSN 0004-9581.

Hoque:1989:ECO

[1013] Asraul Hoque. The efficiency of the Cochrane–Orcutt procedure. Australian Journal of Statistics, 31(3):385–392, September 1989. CODEN AUJSA3. ISSN 0004-9581.

Groenewald:1989:BTS

[1014] P. C. N. Groenewald and D. J. de Waal. Bayesian tests for some precise hypotheses on multi-normal means. Australian Journal of Statistics, 31 (3):393–408, September 1989. CODEN AUJSA3. ISSN 0004-9581.

Adegboye:1989:TAR

[1015] O. Sola Adegboye and A. K. Gupta. On testing against restricted alternatives for the variances of Gaussian models. Australian Journal of Statistics, 31(3):409–415, September 1989. CODEN AUJSA3. ISSN 0004-9581.

Schucany:1989:SSV

[1016] William R. Schucany and Daniel M. Bankson. Small sample variance estimators for U-statistics. Australian Journal of Statistics, 31(3):417– 426, September 1989. CODEN AUJSA3. ISSN 0004-9581.

Pogany:1989:AST

[1017] Tibor Pogány. An approach to the sampling theorem for continuous time processes. Australian Journal of Statistics, 31(3):427–432, September 1989. CODEN AUJSA3. ISSN 0004-9581.

Ganeshanandam:1989:SVA

[1018] S. Ganeshanandam and W. J. Krzanowski. On selecting variables and assessing their performance in linear discriminant analysis. Australian Journal of Statistics, 31(3):433–447, September 1989. CODEN AUJSA3. ISSN 0004-9581.

Becker:1989:ODF

[1019] Niels Becker, Barry McDonald, and Celestine Khoo. Optimal designs for fitting a proportional hazards regression model to data subject to censoring. Australian Journal of Statistics, 31(3):449–468, September 1989. CODEN AUJSA3. ISSN 0004-9581.

Liu:1989:QSE

[1020] Jingyu Liu and H. A. David. Quantiles of sums and expected values of ordered sums. Australian Journal of Statistics, 31(3):469–474, September 1989. CODEN AUJSA3. ISSN 0004-9581.

Maesono:1989:LBN

[1021] Yoshihiko Maesono. Lower bounds for the normal approximation of a sum of independent random variables. Australian Journal of Statistics, 31(3):475–485, September 1989. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1989:BRc

[1022] Anonymous. Book reviews. Australian Journal of Statistics, 31(3):487– 496, September 1989. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1989:BNc

[1023] Anonymous. Book notices. Australian Journal of Statistics, 31(3):496, September 1989. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1989:II

[1024] Anonymous. Introduction: Introduction. Australian Journal of Statistics, 31A(1):1–7, August 1989. CODEN AUJSA3. ISSN 0004-9581.

Smith:1989:DAM

[1025] D. J. Smith and A. J. Callaghan. Definitional aspects of measurement of unemployment — a comparative study of the Australian Longitudinal Survey and other sources of youth unemployment data. Australian Journal of Statistics, 31A(1):9–24, August 1989. CODEN AUJSA3. ISSN 0004-9581.

Trewin:1989:AID

[1026] Dennis Trewin, Garth Bode, Peter Boal, and Debbie Newton. An automatic interaction detection analysis on unemployed youth. Australian Journal of Statistics, 31A(1):25–45, August 1989. CODEN AUJSA3. ISSN 0004-9581.

Miller:1989:SIE

[1027] Paul Miller and Paul Volker. Socioeconomic influences on educational attainment: Evidence and implications for the tertiary education finance debate. Australian Journal of Statistics, 31A(1):47–70, August 1989. CODEN AUJSA3. ISSN 0004-9581.

Bradbury:1989:FEL

[1028] Bruce Bradbury, Ian McRae, and Lyn Woyzbun. Families and early labour market experience: an analysis of siblings. Australian Journal of Statistics, 31A(1):71–97, August 1989. CODEN AUJSA3. ISSN 0004-9581.

Prior:1989:IFB

[1029] Heather Prior and John J. Beggs. Influence of family background on the educational and labour-force outcomes of year 12 school-leavers. Australian Journal of Statistics, 31A(1):99–124, August 1989. CODEN AU-JSA3. ISSN 0004-9581.

Larum:1989:WDA

[1030] John Larum and John J. Beggs. What drives Australian teenage labour force participation? Australian Journal of Statistics, 31A(1):125–142, August 1989. CODEN AUJSA3. ISSN 0004-9581.

Sawyer:1989:NPE

[1031] K. R. Sawyer. Non-parametric estimation and testing of econometric models of duration: an application to the Australian Longitudinal Survey. Australian Journal of Statistics, 31A(1):143–164, August 1989. CO-DEN AUJSA3. ISSN 0004-9581.

Dunsmuir:1989:MTB

[1032] William Dunsmuir, Richard Tweedie, Lloyd Flack, and Kerrie Mengersen. Modelling of transitions between employment states for young Australians. Australian Journal of Statistics, 31A(1):165–196, August 1989. CODEN AUJSA3. ISSN 0004-9581.

James:1989:UDM

[1033] Ian R. James. Unemployment duration: Modelling and estimation with particular reference to the Australian Longitudinal Survey. Australian Journal of Statistics, 31A(1):197–212, August 1989. CODEN AUJSA3. ISSN 0004-9581.

Eyland:1989:TFT

[1034] E. Ann Eyland and Lester W. Johnson. The transition from full-time study to work. Australian Journal of Statistics, 31A(1):213–225, August 1989. CODEN AUJSA3. ISSN 0004-9581.

Kijima:1990:UTP

[1035] Masaaki Kijima. On the unimodality of transition probabilities in Markov chains. Australian Journal of Statistics, 32(1):1–10, March 1990. CODEN AUJSA3. ISSN 0004-9581.

Weissfeld:1990:IDN

[1036] L. A. Weissfeld and H. Schneider. Influence diagnostics for the normal linear model with censored data. Australian Journal of Statistics, 32(1): 11–20, March 1990. CODEN AUJSA3. ISSN 0004-9581.

King:1990:PST

[1037] Maxwell L. King. The power of Student’s t test: Can a non-similar test do better? Australian Journal of Statistics, 32(1):21–27, March 1990. CODEN AUJSA3. ISSN 0004-9581.

Carlin:1990:AAB

[1038] J. B. Carlin. An algorithmic approach to Bayesian linear model calculations. Australian Journal of Statistics, 32(1):29–43, March 1990. CODEN AUJSA3. ISSN 0004-9581.

Kochar:1990:DFT

[1039] Subhash C. Kochar and R. P. Gupta. Distribution-free tests based on sub-sample extrema for testing against positive dependence. Australian Journal of Statistics, 32(1):45–51, March 1990. CODEN AUJSA3. ISSN 0004-9581.

Shao:1990:TTE

[1040] Jun Shao and Shein-Chung Chow. Test for treatment effect based on binary data with random sample sizes. Australian Journal of Statistics, 32(1):53–70, March 1990. CODEN AUJSA3. ISSN 0004-9581.

Pervaiz:1990:MCC

[1041] Muhammad Khalid Pervaiz and C. J. Skinner. A Monte Carlo comparison of elliptical-theory and other tests for equality of covariance matrices. Australian Journal of Statistics, 32(1):71–86, March 1990. CODEN AU-JSA3. ISSN 0004-9581.

Bhatti:1990:BOT

[1042] Muhammad I. Bhatti and Maxwell L. King. A beta-optimal test of the equicorrelation coefficient. Australian Journal of Statistics, 32(1):87–97, March 1990. CODEN AUJSA3. ISSN 0004-9581.

Dollinger:1990:CED

[1043] Michael B. Dollinger and Robert G. Staudte. The construction of equileverage designs for multiple linear regression. Australian Journal of Statistics, 32(1):99–118, March 1990. CODEN AUJSA3. ISSN 0004-9581.

Sharpe:1990:BRT

[1044] Ken Sharpe. Book review: The Teaching of Practical Statistics, by C. W. Anderson and R. M. Loynes. Australian Journal of Statistics, 32(1): 119–121, March 1990. CODEN AUJSA3. ISSN 0004-9581.

Saunders:1990:BRA

[1045] Ian W. Saunders. Book review: Analysis of Infectious Disease Data. by Niels G. Becker. Australian Journal of Statistics, 32(1):121–123, March 1990. CODEN AUJSA3. ISSN 0004-9581.

Cohn:1990:BRA

[1046] H. Cohn. Book review: Advanced Probability Theory. by Janos Galambos. Australian Journal of Statistics, 32(1):124–125, March 1990. CODEN AUJSA3. ISSN 0004-9581.

Mengersen:1990:CSB

[1047] Kerrie Mengersen and Eve Bofinger. Corrigendum: Selecting the t best populations — scale parameter case. Australian Journal of Statistics, 32 (1):127, March 1990. CODEN AUJSA3. ISSN 0004-9581. See [1004].

Mudholkar:1990:EAC

[1048] Govind S. Mudholkar and Philip A. Smethurst. Empirically accelerated convergence of the spacings statistics to normality. Australian Journal of Statistics, 32(2):129–138, June 1990. CODEN AUJSA3. ISSN 0004-9581.

Hall:1990:RCN

[1049] Peter Hall and A. H. Welsh. The rate of convergence in the normal approximation for the Wilcoxon R-estimator. Australian Journal of Statistics, 32(2):139–150, June 1990. CODEN AUJSA3. ISSN 0004-9581.

Anderson:1990:TST

[1050] Oliver D. Anderson. Teaching statistics: a talk about time series. Australian Journal of Statistics, 32(2):151–161, June 1990. CODEN AU-JSA3. ISSN 0004-9581.

Danaher:1990:MCM

[1051] Peter J. Danaher. A Markov chain model for magazine exposure. Australian Journal of Statistics, 32(2):163–176, June 1990. CODEN AU-JSA3. ISSN 0004-9581.

Fisher:1990:BHT

[1052] Nicholas I. Fisher and Peter Hall. On bootstrap hypothesis testing. Australian Journal of Statistics, 32(2):177–190, June 1990. CODEN AU-JSA3. ISSN 0004-9581.

Pollard:1990:MDA

[1053] Graham Pollard. A method for determining the asymptotic efficiency of some sequential probability ratio tests. Australian Journal of Statistics, 32(2):191–204, June 1990. CODEN AUJSA3. ISSN 0004-9581.

Haslett:1990:CFI

[1054] Stephen Haslett. Covariances for fixed interval smoothed Kalman filter parameter estimates. Australian Journal of Statistics, 32(2):205–216, June 1990. CODEN AUJSA3. ISSN 0004-9581.

He:1990:ERN

[1055] Xuming He. Estimating the root of a nonparametric regression function in a robust fashion. Australian Journal of Statistics, 32(2):217–225, June 1990. CODEN AUJSA3. ISSN 0004-9581.

Verbyla:1990:CDR

[1056] A. P. Verbyla. A conditional derivation of residual maximum likelihood. Australian Journal of Statistics, 32(2):227–230, June 1990. CODEN AU-JSA3. ISSN 0004-9581.

Tjur:1990:NUB

[1057] Tue Tjur. A new upper bound for the efficiency factor of a block design. Australian Journal of Statistics, 32(2):231–237, June 1990. CODEN AU-JSA3. ISSN 0004-9581.

Khattree:1990:ERS

[1058] Ravindra Khattree and Rameshwar D. Gupta. Estimation of realized signal to noise ratio for a pair of multivariate signals. Australian Journal of Statistics, 32(2):239–246, June 1990. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1990:BRa

[1059] Anonymous. Book reviews. Australian Journal of Statistics, 32(2):247– 253, June 1990. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1990:BN

[1060] Anonymous. Book notices. Australian Journal of Statistics, 32(2):253, June 1990. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1990:BRB

[1061] Anonymous. Books received. Australian Journal of Statistics, 32(2): 253–254, June 1990. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1990:PMA

[1062] Anonymous. Pitman Medal awarded to C. C. Heyde. Australian Journal of Statistics, 32(3):259–260, September 1990. CODEN AUJSA3. ISSN 0004-9581.

Irwig:1990:CME

[1063] Les M. Irwig, Hennie T. Groeneveld, and Judy M. Simpson. Correcting for measurement error in an exposure-response relationship based on dichotomising a continuous dependent variable. Australian Journal of Statistics, 32(3):261–269, September 1990. CODEN AUJSA3. ISSN 0004-9581.

Gerontidis:1990:VCM

[1064] Ioannis I. Gerontidis. On the variance-covariance matrix in Markovian manpower systems in continuous time. Australian Journal of Statistics, 32(3):271–280, September 1990. CODEN AUJSA3. ISSN 0004-9581.

Akahira:1990:LIA

[1065] Masafumi Akahira and Kei Takeuchi. Loss of information associated with the order statistics and related estimators in the double exponential distribution case. Australian Journal of Statistics, 32(3):281–291, September 1990. CODEN AUJSA3. ISSN 0004-9581.

Michels:1990:TSR

[1066] Paul Michels and Götz Trenkler. Testing the stability of regression coefficients using generalized recursive residuals. Australian Journal of Statistics, 32(3):293–312, September 1990. CODEN AUJSA3. ISSN 0004-9581.

Konijn:1990:NPI

[1067] H. S. Konijn. Note on point and interval estimation and prediction for normal equicorrelated data. Australian Journal of Statistics, 32(3):313– 316, September 1990. CODEN AUJSA3. ISSN 0004-9581.

Scott:1990:IME

[1068] David J. Scott and Wang Dong Qian. An iterative method for estimation of parameters in a clustered sampling model. Australian Journal of Statistics, 32(3):317–325, September 1990. CODEN AUJSA3. ISSN 0004-9581.

Robinson:1990:RV

[1069] G. K. Robinson. A role for variograms. Australian Journal of Statistics, 32(3):327–335, September 1990. CODEN AUJSA3. ISSN 0004-9581.

Gupta:1990:IBD

[1070] Sudhir Gupta and Rahul Mukerjee. On incomplete block designs for symmetric parallel line assays. Australian Journal of Statistics, 32(3): 337–344, September 1990. CODEN AUJSA3. ISSN 0004-9581.

Street:1990:TSO

[1071] Deborah J. Street, J. A. Eccleston, and William H. Wilson. Tables of small optimal repeated measurements designs. Australian Journal of Statistics, 32(3):345–359, September 1990. CODEN AUJSA3. ISSN 0004-9581.

Jones:1990:VKD

[1072] M. C. Jones. Variable kernel density estimates and variable kernel density estimates. Australian Journal of Statistics, 32(3):361–371, September 1990. CODEN AUJSA3. ISSN 0004-9581.

Mehra:1990:CNT

[1073] K. L. Mehra, N. G. N. Prasad, and K. S. Madhava Rao. A class of nonparametric tests for the one-sample location problem. Australian Journal of Statistics, 32(3):373–392, September 1990. CODEN AUJSA3. ISSN 0004-9581.

Sadooghi-Alvandi:1990:EPP

[1074] S. M. Sadooghi-Alvandi. Estimation of the parameter of a Poisson distribution using a LINEX loss function. Australian Journal of Statistics, 32(3):393–398, September 1990. CODEN AUJSA3. ISSN 0004-9581.

Nguyen:1990:CAC

[1075] Nam-Ky Nguyen and Aloke Dey. Computer-aided construction of small (M, S)-optimal incomplete block designs. Australian Journal of Statistics, 32(3):399–410, September 1990. CODEN AUJSA3. ISSN 0004-9581.

Brewer:1990:AUI

[1076] K. R. W. Brewer and S. M. Tam. Is the assumption of uniform intra-class correlation ever needed? Australian Journal of Statistics, 32(3):411–423, September 1990. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1990:BRc

[1077] Anonymous. Book reviews. Australian Journal of Statistics, 32(3):425– 442, September 1990. CODEN AUJSA3. ISSN 0004-9581.

Tweedie:1991:PMA

[1078] R. L. Tweedie. Pitman Medal awarded to E. J. Hannan. Australian Journal of Statistics, 33(1):1–4, March 1991. CODEN AUJSA3. ISSN 0004-9581.

Mcgilchrist:1991:SMM

[1079] C. A. Mcgilchrist and L. J. Hills. A semi-Markov model for ear infection. Australian Journal of Statistics, 33(1):5–16, March 1991. CODEN AUJSA3. ISSN 0004-9581.

Adekola:1991:NAP

[1080] O. A. Adekola. A note on asymptotic posterior normality for stochastic processes. Australian Journal of Statistics, 33(1):17–21, March 1991. CODEN AUJSA3. ISSN 0004-9581.

Hurvich:1991:VES

[1081] Clifford M. Hurvich, Jeffrey S. Simonoff, and Scott L. Zeger. Variance estimation for sample autocovariances: Direct and resampling approaches. Australian Journal of Statistics, 33(1):23–42, March 1991. CODEN AU-JSA3. ISSN 0004-9581.

Nanayakkara:1991:CTU

[1082] Nuwan Nanayakkara and Noel Cressie. Combining two unbiased estimators of a common mean of two normal populations. Australian Journal of Statistics, 33(1):43–56, March 1991. CODEN AUJSA3. ISSN 0004-9581.

Smyth:1991:EDM

[1083] Gordon K. Smyth. Exponential dispersion models and the Gauss–Newton algorithm. Australian Journal of Statistics, 33(1):57–64, March 1991. CODEN AUJSA3. ISSN 0004-9581.

Ghosh:1991:ARE

[1084] Subir Ghosh and Yi-Jing Duh. Adjusting residuals in estimation of dispersion effects. Australian Journal of Statistics, 33(1):65–74, March 1991. CODEN AUJSA3. ISSN 0004-9581.

Wei:1991:SSO

[1085] Bo-Cheng Wei. Some second order asymptotics in nonlinear regression. Australian Journal of Statistics, 33(1):75–84, March 1991. CODEN AU-JSA3. ISSN 0004-9581.

Anh:1991:SRE

[1086] V. V. Anh and K. E. Lunney. Spectral representation and ergodicity of asymptotically stationary processes. Australian Journal of Statistics, 33 (1):85–94, March 1991. CODEN AUJSA3. ISSN 0004-9581.

Prabhakaran:1991:IFM

[1087] V. T. Prabhakaran and V. K. Mahajan. On an identity for finding moments of sample moments of bivariate normal random variables. Australian Journal of Statistics, 33(1):95–102, March 1991. CODEN AU-JSA3. ISSN 0004-9581.

Anonymous:1991:BRa

[1088] Anonymous. Book reviews. Australian Journal of Statistics, 33(1):103– 117, March 1991. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1991:C

[1089] Anonymous. Corrigendum. Australian Journal of Statistics, 33(1):119,

March 1991. CODEN AUJSA3. ISSN 0004-9581.

Jarrett:1991:PMA

[1090] R. G. Jarrett. Pitman Medal awarded to P. G. Hall. Australian Journal of Statistics, 33(2):122–124, June 1991. CODEN AUJSA3. ISSN 0004-9581.

Becker:1991:ULL

[1091] Niels G. Becker and Mads Melbye. Use of a log-linear model to compute the empirical survival curve from interval-censored data, with application to data on tests for HIV positivity. Australian Journal of Statistics, 33 (2):125–133, June 1991. CODEN AUJSA3. ISSN 0004-9581.

Adekola:1991:APN

[1092] O. A. Adekola. Asymptotic posterior normality for stochastic processes: a non-ergodic example. Australian Journal of Statistics, 33(2):135–144, June 1991. CODEN AUJSA3. ISSN 0004-9581.

Singh:1991:JSR

[1093] Radhey S. Singh. James–Stein rule estimators in linear regression models with multivariate-t distributed error. Australian Journal of Statistics, 33 (2):145–158, June 1991. CODEN AUJSA3. ISSN 0004-9581.

Diamond:1991:UCF

[1094] Neil T. Diamond. The use of a class of foldover designs as search designs. Australian Journal of Statistics, 33(2):159–166, June 1991. CODEN AU-JSA3. ISSN 0004-9581.

Deshmukh:1991:BS

[1095] Shailaja R. Deshmukh. Bernoulli sampling. Australian Journal of Statistics, 33(2):167–176, June 1991. CODEN AUJSA3. ISSN 0004-9581.

Sengupta:1991:SSA

[1096] Sukalyan Sengupta. Some simple approximations for the doubly noncentral z distribution. Australian Journal of Statistics, 33(2):177–181, June 1991. CODEN AUJSA3. ISSN 0004-9581.

Bofinger:1991:SDB

[1097] Eve Bofinger. Selecting “demonstrably best” or “demonstrably worst” exponential populations. Australian Journal of Statistics, 33(2):183–190, June 1991. CODEN AUJSA3. ISSN 0004-9581.

Marschner:1991:HAS

[1098] Ian C. Marschner. Heterogeneity among susceptibles in epidemic models. Australian Journal of Statistics, 33(2):191–207, June 1991. CODEN AUJSA3. ISSN 0004-9581.

Huggins:1991:RAB

[1099] Richard M. Huggins and Ian C. Marschner. Robust analysis of the bifurcating autoregressive model in cell lineage studies. Australian Journal of Statistics, 33(2):209–220, June 1991. CODEN AUJSA3. ISSN 0004-9581.

Marschner:1991:REE

[1100] Ian C. Marschner. Robust estimation for epidemic models. Australian Journal of Statistics, 33(2):221–240, June 1991. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1991:BRb

[1101] Anonymous. Book reviews. Australian Journal of Statistics, 33(2):241– 248, June 1991. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1991:BNa

[1102] Anonymous. Book notices. Australian Journal of Statistics, 33(2):248, June 1991. CODEN AUJSA3. ISSN 0004-9581.

Lo:1991:AAM

[1103] Sing Kai Lo. On the analysis and application of measures of linkage disequilibrium. Australian Journal of Statistics, 33(3):249–259, September 1991. CODEN AUJSA3. ISSN 0004-9581.

Clark:1991:RSB

[1104] R. M. Clark and A. G. Harcourt. Randomisation and the 1975 Senate ballot draw. Australian Journal of Statistics, 33(3):261–278, September 1991. CODEN AUJSA3. ISSN 0004-9581.

Catchpole:1991:EBV

[1105] W. R. Catchpole and E. A. Catchpole. Estimating biomass in a vegetation Mosaic using double sampling with regression. Australian Journal of Statistics, 33(3):279–289, September 1991. CODEN AUJSA3. ISSN 0004-9581.

Kourouklis:1991:BEL

[1106] Stavros Kourouklis. Bahadur efficiency of likelihood ratio and related tests in nonregular models. Australian Journal of Statistics, 33(3):291– 298, September 1991. CODEN AUJSA3. ISSN 0004-9581.

Martin:1991:EBD

[1107] R. J. Martin and J. A. Eccleston. Efficient block designs for correlated observations. Australian Journal of Statistics, 33(3):299–311, September 1991. CODEN AUJSA3. ISSN 0004-9581.

Davison:1991:ECT

[1108] A. C. Davison, P. W. F. Smith, and J. Whittaker. An exact conditional test for covariance selection models. Australian Journal of Statistics, 33 (3):313–318, September 1991. CODEN AUJSA3. ISSN 0004-9581.

Silvapulle:1991:RRR

[1109] Mervyn J. Silvapulle. Robust ridge regression based on an M-estimator. Australian Journal of Statistics, 33(3):319–333, September 1991. CO-DEN AUJSA3. ISSN 0004-9581.

Hayter:1991:DFM

[1110] Anthony J. Hayter and Glenn Stone. Distribution free multiple comparisons for monotonically ordered treatment effects. Australian Journal of Statistics, 33(3):335–346, September 1991. CODEN AUJSA3. ISSN 0004-9581.

Gill:1991:SPS

[1111] A. N. Gill and G. P. Mehta. Selection procedures for scale parameters using two-sample U-statistics. Australian Journal of Statistics, 33(3): 347–362, September 1991. CODEN AUJSA3. ISSN 0004-9581.

Brown:1991:GPD

[1112] B. M. Brown and M. J. Pella. The grouping problem in distributionfree planar regression. Australian Journal of Statistics, 33(3):363–372, September 1991. CODEN AUJSA3. ISSN 0004-9581.

Anderson:1991:AGW

[1113] O. D. Anderson. All the Gaussian white noise serial covariance moments to order four. Australian Journal of Statistics, 33(3):373–396, September 1991. CODEN AUJSA3. ISSN 0004-9581.

Bai:1991:NAM

[1114] Jun Bai, Anthony J. Jakeman, and Michael McAleer. A new approach to maximum likelihood estimation of the three-parameter gamma and Weibull distributions. Australian Journal of Statistics, 33(3):397–410, September 1991. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1991:BRc

[1115] Anonymous. Book reviews. Australian Journal of Statistics, 33(3):411– 426, September 1991. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1991:BNb

[1116] Anonymous. Book notices. Australian Journal of Statistics, 33(3):426, September 1991. CODEN AUJSA3. ISSN 0004-9581.

Pithuncharurnlap:1992:NNA

[1117] M. Pithuncharurnlap, K. E. Basford, and B. R. Cullis. Nearest neighbour analysis of unequally replicated trials. Australian Journal of Statistics, 34(1):1–9, March 1992. CODEN AUJSA3. ISSN 0004-9581.

Lewgoy:1992:ESS

[1118] Flavio Lewgoy and Alan E. Stark. Experimental search for sexual isolation in flies. Australian Journal of Statistics, 34(1):11–17, March 1992. CODEN AUJSA3. ISSN 0004-9581.

Ruppert:1992:CKD

[1119] D. Ruppert and M. P. Wand. Correcting for kurtosis in density estimation. Australian Journal of Statistics, 34(1):19–29, March 1992. CODEN AUJSA3. ISSN 0004-9581.

Puri:1992:EDF

[1120] Prem S. Puri and Harshinder Singh. Estimation of a distribution function dominating stochastically a known distribution function. Australian Journal of Statistics, 34(1):31–38, March 1992. CODEN AUJSA3. ISSN 0004-9581.

Kumar:1992:ECM

[1121] Somesh Kumar and Divakar Sharma. Estimating the common mean of a bivariate normal population. Australian Journal of Statistics, 34(1): 39–46, March 1992. CODEN AUJSA3. ISSN 0004-9581.

Patel:1992:UEP

[1122] S. R. Patel and M. B. Bhatt. Unbiased estimation of parametre functions in sampling from one- and two-truncation parameter families. Australian Journal of Statistics, 34(1):47–54, March 1992. CODEN AUJSA3. ISSN 0004-9581.

Goria:1992:FRT

[1123] M. N. Goria. On the fourth root transformation of chi-square. Australian Journal of Statistics, 34(1):55–64, March 1992. CODEN AUJSA3. ISSN 0004-9581.

Bofinger:1992:CST

[1124] Eve Bofinger. Comparisons and selection of two-parameter exponential populations. Australian Journal of Statistics, 34(1):65–75, March 1992. CODEN AUJSA3. ISSN 0004-9581.

Saunders:1992:EDC

[1125] Ian W. Saunders and John A. Eccleston. Experimental design for continuous processes. Australian Journal of Statistics, 34(1):77–89, March 1992. CODEN AUJSA3. ISSN 0004-9581.

Gleeson:1992:DFE

[1126] Alan C. Gleeson and John A. Eccleston. On the design of field experiments under a one-dimensional neighbour model. Australian Journal of Statistics, 34(1):91–97, March 1992. CODEN AUJSA3. ISSN 0004-9581.

Kulkarni:1992:REH

[1127] P. M. Kulkarni. Robust estimation and hypothesis tests for first-order threshold autoregressive models. Australian Journal of Statistics, 34(1): 99–104, March 1992. CODEN AUJSA3. ISSN 0004-9581.

Kim:1992:EBE

[1128] Young-Won Kim and I. V. Basawa. Empirical Bayes estimation for firstorder autoregressive processes. Australian Journal of Statistics, 34(1): 105–114, March 1992. CODEN AUJSA3. ISSN 0004-9581.

Adke:1992:RCP

[1129] S. R. Adke and N. Balakrishna. Renewal counting process induced by a discrete Markov chain. Australian Journal of Statistics, 34(1):115–121, March 1992. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1992:BRa

[1130] Anonymous. Book reviews. Australian Journal of Statistics, 34(1):123– 125, March 1992. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1992:BN

[1131] Anonymous. Book notices. Australian Journal of Statistics, 34(1):126– 127, March 1992. CODEN AUJSA3. ISSN 0004-9581.

Becker:1992:SCA

[1132] Niels G. Becker. Statistical challenges of AIDS: (the 1990 Knibbs Lecture, Canberra, 27 November 1990). Australian Journal of Statistics, 34(2): 129–144, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Cullis:1992:NRM

[1133] B. R. Cullis and A. P. Verbyla. Nonlinear regression modelling and time dependent covariates in repeated measures experiments. Australian Journal of Statistics, 34(2):145–160, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Weissfeld:1992:MGC

[1134] L. A. Weissfeld and Anant M. Kshirsagar. A modified growth curve model and its application to clinical studies. Australian Journal of Statistics, 34(2):161–168, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Kuk:1992:SMM

[1135] Anthony Y. C. Kuk. A semiparametric mixture model for the analysis of competing risks data. Australian Journal of Statistics, 34(2):169–180, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Marschner:1992:RMD

[1136] Lan C. Marschner. Robust methods for dependent cell pedigree data. Australian Journal of Statistics, 34(2):181–198, June 1992. CODEN AU-JSA3. ISSN 0004-9581.

Heyde:1992:QLM

[1137] C. C. Heyde and Y.-X. Lin. On quasi-likelihood methods and estimation for bunching processes and heteroscedastic regression models. Australian Journal of Statistics, 34(2):199–206, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Khattree:1992:EGT

[1138] Ravindra Khattree. Estimation of guarantee time and mean life after warranty for two-parameter exponential failure model. Australian Journal of Statistics, 34(2):207–215, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Tobing:1992:RRM

[1139] H. Tobing and C. A. McGilchrist. Recursive residuals for multivariate regression models. Australian Journal of Statistics, 34(2):217–232, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Jones:1992:FFM

[1140] P. N. Jones and G. J. McLachlan. Fitting finite mixture models in a regression context. Australian Journal of Statistics, 34(2):233–240, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Lloyd:1992:EC

[1141] Christopher Lloyd. Effective conditioning. Australian Journal of Statistics, 34(2):241–260, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Saunders:1992:EFS

[1142] L. R. Saunders. An exact formula for the symmetrical incomplete beta function where the parameter is an integer or half-integer. Australian Journal of Statistics, 34(2):261–264, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Yip:1992:GFT

[1143] Paul Yip. Goodness of fit tests for heterogeneity in a capture-recapture experiment. Australian Journal of Statistics, 34(2):265–271, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Pollard:1992:OTS

[1144] Graham H. Pollard. The optimal test for selecting the greater of two binomial probabilities. Australian Journal of Statistics, 34(2):273–284, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Weerakkody:1992:THA

[1145] Govinda J. Weerakkody. Testing hypotheses about the common mean of two normal populations. Australian Journal of Statistics, 34(2):285–293, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Pezvaiz:1992:MAC

[1146] Muhammad Khalid Pezvaiz and C. J. Skinner. Multivariate analysis of clustered survey data: Testing equality of covariance matrices. Australian Journal of Statistics, 34(2):295–306, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Pakes:1992:LBC

[1147] Anthony G. Pakes and Ravindra Khattree. Length-biasing, characterizations of laws and the moment problem. Australian Journal of Statistics, 34(2):307–322, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Pakes:1992:CTM

[1148] Anthony G. Pakes. On characterizations through mixed sums. Australian Journal of Statistics, 34(2):323–340, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1992:BRb

[1149] Anonymous. Book reviews. Australian Journal of Statistics, 34(2):341– 344, June 1992. CODEN AUJSA3. ISSN 0004-9581.

Brown:1992:CEJ

[1150] Bruce M. Brown and Evan J. Williams. Conversations with Edwin J. G. Pitman. Australian Journal of Statistics, 34(3):345–360, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Kiiveri:1992:ARS

[1151] H. T. Kiiveri and N. A. Campbell. Allocation of remotely sensed data using Markov models for image data and pixel labels. Australian Journal of Statistics, 34(3):361–374, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Sparks:1992:QCM

[1152] Ross S. Sparks. Quality control with multivariate data. Australian Journal of Statistics, 34(3):375–390, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Chatterjee:1992:NAD

[1153] Sangit Chatterjee and A. Narayanan. A new approach to discrimination and classification using a Hausdorff type distance. Australian Journal of Statistics, 34(3):391–406, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Gupta:1992:PNC

[1154] Rameshwar D. Gupta and Harshinder Singh. Pitman nearness comparisons of estimates of two ordered normal means. Australian Journal of Statistics, 34(3):407–414, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Gill:1992:BCD

[1155] P. S. Gill. Balanced change-over designs for autocorrelated observations. Australian Journal of Statistics, 34(3):415–420, September 1992. CO-DEN AUJSA3. ISSN 0004-9581.

Wu:1992:AMP

[1156] Eden K. H. Wu and Denis H. Y. Leung. Asymptotic minimax properties of L-estimators of scale. Australian Journal of Statistics, 34(3):421–432, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Evans:1992:BCV

[1157] Merran A. Evans. Burr critical value approximations for tests of autocorrelation and heteroscedasticity. Australian Journal of Statistics, 34 (3):433–442, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Lee:1992:DCC

[1158] Yoong-Sin Lee and Moy Chee Lee. On the derivation and computation of the Cornish–Fisher expansion. Australian Journal of Statistics, 34(3): 443–450, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Amey:1992:TSU

[1159] Alphonse K. A. Amey and A. K. Gupta. Testing sphericity under a mixture model. Australian Journal of Statistics, 34(3):451–460, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Brahim:1992:IDG

[1160] Joseph G. Brahim and Sanford Weisberg. Incomplete data in generalized linear models with continuous covariates. Australian Journal of Statistics, 34(3):461–470, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Carroll:1992:SCR

[1161] R. J. Carroll and Peter Hall. Semiparametric comparison of regression curves via normal likelihoods. Australian Journal of Statistics, 34(3): 471–487, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Balasubramanian:1992:OMO

[1162] K. Balasubramanian, N. Balakrishnan, and H. J. Malik. Operator methods in order statistics. Australian Journal of Statistics, 34(3):489–496, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Kiiveri:1992:FSC

[1163] H. T. Kiiveri. Fitting spatial correlation models: Approximating a likelihood approximation. Australian Journal of Statistics, 34(3):497–512, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Huang:1992:ESA

[1164] D. Huang and V. V. Anh. Estimation of spatial ARMA models. Australian Journal of Statistics, 34(3):513–530, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Pollett:1992:NEP

[1165] P. K. Pollett and D. Vere Jones. A note on evanescent processes. Australian Journal of Statistics, 34(3):531–536, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Mousa:1992:BRL

[1166] M. A. M. Ali Mousa. The bootstrap risk of linear empirical Bayes estimates. Australian Journal of Statistics, 34(3):537–544, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1992:BRc

[1167] Anonymous. Book reviews. Australian Journal of Statistics, 34(3):545– 558, September 1992. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1993:PMA

[1168] Anonymous. Pitman Medal awarded to A. T. James. Australian Journal of Statistics, 35(1):1–4, March 1993. CODEN AUJSA3. ISSN 0004-9581.

Maller:1993:FAR

[1169] R. A. Maller. Factorial analysis of recidivist data. Australian Journal of Statistics, 35(1):5–18, March 1993. CODEN AUJSA3. ISSN 0004-9581.

Belanger:1993:CPE

[1170] J. Bélanger and D. Gagnon. Classification and proportion estimation. Australian Journal of Statistics, 35(1):19–28, March 1993. CODEN AU-JSA3. ISSN 0004-9581.

Dorfman:1993:CDB

[1171] Alan H. Dorfman. A comparison of design-based and model-based estimators of the finite population distribution function. Australian Journal of Statistics, 35(1):29–41, March 1993. CODEN AUJSA3. ISSN 0004-9581.

Huggins:1993:RAV

[1172] R. M. Huggins. On the robust analysis of variance components models for pedigree data. Australian Journal of Statistics, 35(1):43–57, March 1993. CODEN AUJSA3. ISSN 0004-9581.

Luong:1993:MQD

[1173] Andrew Luong and José Garrido. Minimum quadratic distance estimation for a parametric family of discrete distributions defined recursively. Australian Journal of Statistics, 35(1):59–67, March 1993. CODEN AU-JSA3. ISSN 0004-9581.

Wood:1993:ECP

[1174] Andrew T. A. Wood. Estimation of the concentration parameters of the Fisher matrix distribution on 50(3) and the Bingham distribution on S_{q}, q ≥ 2. Australian Journal of Statistics, 35(1):69–79, March 1993. CODEN AUJSA3. ISSN 0004-9581.

Huggins:1993:AIC

[1175] R. M. Huggins. Asymptotic inference for a class of chain binomial models. Australian Journal of Statistics, 35(1):81–87, March 1993. CODEN AUJSA3. ISSN 0004-9581.

Jones:1993:DWH

[1176] M. C. Jones. Do not weight for heteroscedasticity in nonparametric regression. Australian Journal of Statistics, 35(1):89–92, March 1993. CODEN AUJSA3. ISSN 0004-9581.

Wang:1993:SMB

[1177] Suojin Wang. Saddlepoint methods for bootstrap confidence bands in nonparametric regression. Australian Journal of Statistics, 35(1):93–101, March 1993. CODEN AUJSA3. ISSN 0004-9581.

Segal:1993:PEL

[1178] Mark R. Segal, Bernard Rosner, Albert G. S. Liou, and Jan P. Schouten. Parameter estimation for low-order auto-regressive models with missing values. Australian Journal of Statistics, 35(1):103–118, March 1993. CO-DEN AUJSA3. ISSN 0004-9581.

Anonymous:1993:BRa

[1179] Anonymous. Book reviews. Australian Journal of Statistics, 35(1):119– 127, March 1993. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1993:ARS

[1180] Anonymous. Acknowledgement of referees’ services. Australian Journal of Statistics, 35(1):128, March 1993. CODEN AUJSA3. ISSN 0004-9581.

Edwards:1993:MGF

[1181] A. W. F. Edwards. Mendel, Galton, Fisher. Australian Journal of Statistics, 35(2):129–140, June 1993. CODEN AUJSA3. ISSN 0004-9581.

Bednarski:1993:TLE

[1182] T. Bednarski and B. R. Clarke. Trimmed likelihood estimation of location and scale of the normal distribution. Australian Journal of Statistics, 35 (2):141–153, June 1993. CODEN AUJSA3. ISSN 0004-9581.

Koenker:1993:AFW

[1183] Roger Koenker, José A. F. Machado, Christopher L. Skeels, and A. H. Welsh. Amemiya’s form of the weighted least squares estimator. Australian Journal of Statistics, 35(2):155–174, June 1993. CODEN AU-JSA3. ISSN 0004-9581.

Froda:1993:NEC

[1184] S. Froda and A. Luong. Nonparametric estimation for competing risks. Australian Journal of Statistics, 35(2):175–183, June 1993. CODEN AU-JSA3. ISSN 0004-9581.

Adegboye:1993:OCR

[1185] O. Sola Adegboye. The optimal classification rule for exponential populations. Australian Journal of Statistics, 35(2):185–194, June 1993. CO-DEN AUJSA3. ISSN 0004-9581.

Bolfarine:1993:PFP

[1186] Heleno Bolfarine and Mönica C. Sandoval. Prediction of the finite population distribution function under Gaussian superpopulation models. Australian Journal of Statistics, 35(2):195–204, June 1993. CODEN AU-JSA3. ISSN 0004-9581.

Kabaila:1993:SPP

[1187] Paul Kabaila. Some properties of profile bootstrap confidence intervals. Australian Journal of Statistics, 35(2):205–214, June 1993. CODEN AU-JSA3. ISSN 0004-9581.

Davison:1993:SBM

[1188] A. C. Davison and Peter Hall. On Studentizing and blocking methods for implementing the bootstrap with dependent data. Australian Journal of Statistics, 35(2):215–224, June 1993. CODEN AUJSA3. ISSN 0004-9581.

Liang:1993:CGD

[1189] TaChen Liang and N. Balakrishnan. A characterization of geometric distributions through conditional independence. Australian Journal of Statistics, 35(2):225–228, June 1993. CODEN AUJSA3. ISSN 0004-9581.

Williams:1993:UBL

[1190] E. R. Williams and J. A. John. Upper bounds for latinized designs. Australian Journal of Statistics, 35(2):229–236, June 1993. CODEN AU-JSA3. ISSN 0004-9581.

John:1993:CRR

[1191] J. A. John and D. Whitaker. Construction of resolvable row-column designs using simulated annealing. Australian Journal of Statistics, 35 (2):237–245, June 1993. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1993:OLJ

[1192] Anonymous. Obituaries: S. Lipton, E. J. G. Pitman. Australian Journal of Statistics, 35(2):246, June 1993. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1993:BRb

[1193] Anonymous. Book reviews. Australian Journal of Statistics, 35(2):247– 256, June 1993. CODEN AUJSA3. ISSN 0004-9581.

Best:1993:EAR

[1194] D. J. Best. Extended analysis for ranked data. Australian Journal of Statistics, 35(3):257–262, September 1993. CODEN AUJSA3. ISSN 0004-9581.

Pithuncharurnlap:1993:NAA

[1195] M. Pithuncharurnlap, K. E. Basford, and W. T. Federer. Neighbour analysis with adjustment for interplot competition. Australian Journal of Statistics, 35(3):263–270, September 1993. CODEN AUJSA3. ISSN 0004-9581.

Smith:1993:ERQ

[1196] Murray D. Smith. Expectations of ratios of quadratic forms in normal variables: Evaluating some top-order invariant polynomials. Australian Journal of Statistics, 35(3):271–282, September 1993. CODEN AUJSA3. ISSN 0004-9581.

Horgan:1993:NEA

[1197] G. W. Horgan and J. Haslett. Nonparametric estimation of angular distributions from Cartesian observations with error. Australian Journal of Statistics, 35(3):283–292, September 1993. CODEN AUJSA3. ISSN 0004-9581.

Mukerjee:1993:ISD

[1198] Rahul Mukerjee and Sujit K. Basu. Inverse sampling for domain estimation in a stratified population. Australian Journal of Statistics, 35(3): 293–302, September 1993. CODEN AUJSA3. ISSN 0004-9581.

Akritas:1993:REU

[1199] Michael G. Akritas, Indrani Basak, and Myung Hwi Lee. Robust estimation under Type II censoring. Australian Journal of Statistics, 35(3): 303–318, September 1993. CODEN AUJSA3. ISSN 0004-9581.

Jones:1993:KDE

[1200] M. C. Jones. Kernel density estimation when the bandwidth is large. Australian Journal of Statistics, 35(3):319–326, September 1993. CO-DEN AUJSA3. ISSN 0004-9581.

Hayakawa:1993:IBI

[1201] Yu Hayakawa. Interrelationships between l_{P} -isotropic densities and l_{P} - isotropic survival functions, and De Finetti representations of Schurconcave survival functions. Australian Journal of Statistics, 35(3):327– 332, September 1993. CODEN AUJSA3. ISSN 0004-9581.

Campbell:1993:CVA

[1202] N. A. Campbell and H. T. Kiiveri. Canonical variate analysis with spatially-correlated data. Australian Journal of Statistics, 35(3):333–344, September 1993. CODEN AUJSA3. ISSN 0004-9581.

Albert:1993:RRG

[1203] Jeffrey M. Albert and Anant M. Kshirsagar. The reduced-rank growth curve model for discriminant analysis of longitudinal data. Australian Journal of Statistics, 35(3):345–357, September 1993. CODEN AUJSA3. ISSN 0004-9581.

Williams:1993:TSG

[1204] E. R. Williams and J. T. Wood. Testing the significance of genotype– environment interaction. Australian Journal of Statistics, 35(3):359–362, September 1993. CODEN AUJSA3. ISSN 0004-9581.

Nguyen:1993:ACO

[1205] Nam-Ky Nguyen and E. R. Williams. An algorithm for constructing optimal resolvable row-column designs. Australian Journal of Statistics, 35(3):363–370, September 1993. CODEN AUJSA3. ISSN 0004-9581.

Venables:1993:RSS

[1206] W. N. Venables and J. A. Eccleston. Randomized search strategies for finding optimal or near optimal block and row-column designs. Australian Journal of Statistics, 35(3):371–382, September 1993. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1993:BRc

[1207] Anonymous. Book reviews. Australian Journal of Statistics, 35(3):383– 389, September 1993. CODEN AUJSA3. ISSN 0004-9581.

Anderson:1993:C

[1208] O. D. Anderson. Corrigendum. Australian Journal of Statistics, 35(3): 390, September 1993. CODEN AUJSA3. ISSN 0004-9581.

Gani:1994:OEJ

[1209] Joe Gani. Obituary: Edward James Hannan. Australian Journal of Statistics, 36(1):1–8, March 1994. CODEN AUJSA3. ISSN 0004-9581.

Kerkvliet:1994:ELM

[1210] Joe Kerkvliet. Estimating a logit model with randomized data: the case of cocaine use. Australian Journal of Statistics, 36(1):9–20, March 1994. CODEN AUJSA3. ISSN 0004-9581.

Mathers:1994:CSB

[1211] C. D. Mathers, R. S. Harris, and P. A. L. Lancaster. A Cusum scheme based on the exponential distribution for surveillance of rare congenital malformations. Australian Journal of Statistics, 36(1):21–30, March 1994. CODEN AUJSA3. ISSN 0004-9581.

Richardson:1994:APR

[1212] A. M. Richardson and A. H. Welsh. Asymptotic properties of restricted maximum likelihood (REML) estimates for hierarchical mixed linear models. Australian Journal of Statistics, 36(1):31–43, March 1994. CO-DEN AUJSA3. ISSN 0004-9581.

Clarke:1994:EEA

[1213] Brenton R. Clarke. Empirical evidence for adaptive confidence intervals and identification of outliers using methods of trimming. Australian Journal of Statistics, 36(1):45–58, March 1994. CODEN AUJSA3. ISSN 0004-9581.

Bofinger:1994:SWE

[1214] Eve Bofinger. Should we estimate the probability of correct selection? Australian Journal of Statistics, 36(1):59–66, March 1994. CODEN AU-JSA3. ISSN 0004-9581.

vanderLinde:1994:CVS

[1215] Angelika van der Linde. On cross-validation for smoothing splines in the case of dependent observations. Australian Journal of Statistics, 36(1): 67–73, March 1994. CODEN AUJSA3. ISSN 0004-9581.

Moyeed:1994:RCS

[1216] R. A. Moyeed and P. J. Diggle. Rates of convergence in semi-parametric modelling of longitudinal data. Australian Journal of Statistics, 36(1): 75–93, March 1994. CODEN AUJSA3. ISSN 0004-9581.

Young:1994:ISM

[1217] Dean M. Young, Patrick L. Odell, and John W. Seaman, Jr. On the independence of the sample mean and translation-invariant statistics for matrix normal distributions. Australian Journal of Statistics, 36(1):95– 100, March 1994. CODEN AUJSA3. ISSN 0004-9581.

Govindarajulu:1994:AEA

[1218] Z. Govindarajulu and H. C. Howard. Asymptotic expansions applied to F -test criteria. Australian Journal of Statistics, 36(1):101–113, March 1994. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1994:C

[1219] Anonymous. Corrigenda. Australian Journal of Statistics, 36(1):114, March 1994. CODEN AUJSA3. ISSN 0004-9581.

Mukerjee:1994:ORO

[1220] Rahul Mukerjee and S. Sengupta. A-optimal run orders with a linear trend. Australian Journal of Statistics, 36(1):115–122, March 1994. CO-DEN AUJSA3. ISSN 0004-9581.

Anonymous:1994:BRa

[1221] Anonymous. Book reviews. Australian Journal of Statistics, 36(1):123– 128, March 1994. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1994:PMAa

[1222] Anonymous. Pitman Medal awarded to E. J. Williams. Australian Journal of Statistics, 36(1):129–132, March 1994. CODEN AUJSA3. ISSN 0004-9581.

Laslett:1994:ESI

[1223] G. M. Laslett, C. J. Lloyd, and G. K. Robinson. Encounters with statistical inference — an interview with evan Williams. Australian Journal of Statistics, 36(2):133–152, June 1994. CODEN AUJSA3. ISSN 0004-9581.

Hopper:1994:MNM

[1224] John L. Hopper and John D. Mathews. A multivariate normal model for pedigree and longitudinal data and the software ‘Fisher’. Australian Journal of Statistics, 36(2):153–176, June 1994. CODEN AUJSA3. ISSN 0004-9581.

Campbell:1994:VSA

[1225] N. A. Campbell and S. L. Furby. Variable selection along canonical vectors. Australian Journal of Statistics, 36(2):177–183, June 1994. CODEN AUJSA3. ISSN 0004-9581.

Yeo:1994:SSS

[1226] A. J. Yeo and G. F. Yeo. Selecting satisfactory secretaries. Australian Journal of Statistics, 36(2):185–198, June 1994. CODEN AUJSA3. ISSN 0004-9581.

Gordon:1994:SST

[1227] Ian Gordon. Sample size for two independent proportions: a review. Australian Journal of Statistics, 36(2):199–209, June 1994. CODEN AU-JSA3. ISSN 0004-9581.

Gulati:1994:NQE

[1228] Sneh Gulati and W. J. Padgett. Nonparametric quantile estimation from record-breaking data. Australian Journal of Statistics, 36(2):211–223, June 1994. CODEN AUJSA3. ISSN 0004-9581.

Coakley:1994:MRT

[1229] Clint W. Coakley and Thomas P. Hettmansperger. The maximum resistance of tests. Australian Journal of Statistics, 36(2):225–233, June 1994. CODEN AUJSA3. ISSN 0004-9581.

Saunders:1994:ACS

[1230] Ian W. Saunders. An algorithm for choosing the sampling times for an experiment on a continuous process. Australian Journal of Statistics, 36 (2):235–244, June 1994. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1994:BRb

[1231] Anonymous. Book reviews. Australian Journal of Statistics, 36(2):245– 251, June 1994. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1994:FPa

[1232] Anonymous. Forthcoming papers. Australian Journal of Statistics, 36 (2):252, June 1994. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1994:PMAb

[1233] Anonymous. Pitman Medal awarded to J. M. Gani. Australian Journal of Statistics, 36(3):253–255, September 1994. CODEN AUJSA3. ISSN 0004-9581.

Tallis:1994:SMT

[1234] G. M. Tallis. A stochastic model for team golf competitions with applications to handicapping. Australian Journal of Statistics, 36(3):257–269, September 1994. CODEN AUJSA3. ISSN 0004-9581.

Huggins:1994:ARE

[1235] Richard M. Huggins and Danuta Z. Loesch. Applications of robust estimation for variance components models:the detection of major gene effects in finger ridge counts in normal and fragile X families. Australian Journal of Statistics, 36(3):271–286, September 1994. CODEN AUJSA3. ISSN 0004-9581.

Bansal:1994:REI

[1236] Naveen K. Bansal and M. Bhandary. Robust M-estimation of the intraclass correlation coefficient. Australian Journal of Statistics, 36(3): 287–301, September 1994. CODEN AUJSA3. ISSN 0004-9581.

Lin:1994:RBQ

[1237] Yan-Xia Lin. The relationship between quasi-score and locally Esufficient estimating functions. Australian Journal of Statistics, 36(3): 303–311, September 1994. CODEN AUJSA3. ISSN 0004-9581.

Wan:1994:SPI

[1238] Alan T. K. Wan. The sampling performance of inequality restricted and pre-test estimators in a mis-specified linear model. Australian Journal of Statistics, 36(3):313–325, September 1994. CODEN AUJSA3. ISSN 0004-9581.

Wei:1994:CRE

[1239] Bo-Cheng Wei. On confidence regions of embedded models in regular parametric families (a geometric approach). Australian Journal of Statistics, 36(3):327–338, September 1994. CODEN AUJSA3. ISSN 0004-9581.

Quine:1994:PAD

[1240] M. P. Quine. Probability approximations for divisible discrete distributions. Australian Journal of Statistics, 36(3):339–349, September 1994. CODEN AUJSA3. ISSN 0004-9581.

Wilms:1994:CSI

[1241] Roel J. G. Wilms and Jan G. F. Thiemann. Characterizations of shiftinvariant distributions based on summation modulo one. Australian Journal of Statistics, 36(3):351–354, September 1994. CODEN AUJSA3. ISSN 0004-9581.

Jing:1994:BSC

[1242] Bing-Yi Jing. Bahadur slopes for comparisons of tests based on resampling. Australian Journal of Statistics, 36(3):355–362, September 1994. CODEN AUJSA3. ISSN 0004-9581.

Piepho:1994:TIN

[1243] Hans-Peter Piepho. On tests for interaction in a nonreplicated two-way layout. Australian Journal of Statistics, 36(3):363–369, September 1994. CODEN AUJSA3. ISSN 0004-9581.

Eccleston:1994:ACO

[1244] John A. Eccleston and Deborah J. Street. An algorithm for the construction of optimal or near-optimal change-over designs. Australian Journal of Statistics, 36(3):371–378, September 1994. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1994:BRc

[1245] Anonymous. Book reviews. Australian Journal of Statistics, 36(3):379– 385, September 1994. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1994:FPb

[1246] Anonymous. Forthcoming papers. Australian Journal of Statistics, 36 (3):386, September 1994. CODEN AUJSA3. ISSN 0004-9581.

Yu:1995:SMD

[1247] Philip L. H. Yu and Richard Cowan. A statistical model for duplicate tournaments in games of luck and skill. Australian Journal of Statistics, 37(1):1–17, March 1995. CODEN AUJSA3. ISSN 0004-9581.

Mengersen:1995:IMC

[1248] K. L. Mengersen, R. L. Tweedie, and B. Biggerstaff. The impact of method choice on meta-analysis. Australian Journal of Statistics, 37(1): 19–44, March 1995. CODEN AUJSA3. ISSN 0004-9581.

Lenth:1995:RTB

[1249] R. V. Lenth, M. Markatou, and J. Tsimikas. Robust tests based on the sample characteristic function. Australian Journal of Statistics, 37(1): 45–60, March 1995. CODEN AUJSA3. ISSN 0004-9581.

Zhou:1995:SAT

[1250] S. Zhou, R. A. Maller, and T. P. Speed. On a shared allele test of random mating. Australian Journal of Statistics, 37(1):61–72, March 1995. CODEN AUJSA3. ISSN 0004-9581.

Wan:1995:OCV

[1251] Alan T. K. Wan. The optimal critical value of a pre-test for an inequality restriction in a mis-specified regression model. Australian Journal of Statistics, 37(1):73–81, March 1995. CODEN AUJSA3. ISSN 0004-9581.

Pusz:1995:RCM

[1252] J. Pusz. A regression characterisation of the Meixner hypergeometric distribution. Australian Journal of Statistics, 37(1):83–87, March 1995. CODEN AUJSA3. ISSN 0004-9581.

Fosam:1995:CSI

[1253] E. B. Fosam and T. Sapatinas. Characterisations of some income distributions based on multiplicative damage models. Australian Journal of Statistics, 37(1):89–93, March 1995. CODEN AUJSA3. ISSN 0004-9581.

Bello:1995:MVI

[1254] A. L. Bello. Missing values, imputation and error rate estimators in discriminant analysis. Australian Journal of Statistics, 37(1):95–104, March 1995. CODEN AUJSA3. ISSN 0004-9581.

Hannan:1995:NAP

[1255] E. J. Hannan and L. Kavalieris. A note on ARMA parameterisation*. Australian Journal of Statistics, 37(1):105–110, March 1995. CODEN AUJSA3. ISSN 0004-9581.

Uddin:1995:RBN

[1256] Nizam Uddin and Dexter C. Whittinghill. The robustness of binary and non-binary balanced nested row-column designs under the unavailability of blocks: a comparison. Australian Journal of Statistics, 37(1):111–124, March 1995. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1995:BRa

[1257] Anonymous. Book reviews. Australian Journal of Statistics, 37(1):125– 128, March 1995. CODEN AUJSA3. ISSN 0004-9581.

Kokic:1995:ETF

[1258] Philip N. Kokic, Lynelle Moon, Jane Gooday, and Ray L. Chambers. Estimating temporal farm income distributions using spatial smoothing techniques. Australian Journal of Statistics, 37(2):129–143, June 1995. CODEN AUJSA3. ISSN 0004-9581.

Barnett:1995:LOS

[1259] Neil S. Barnett, Ian S. Gomm, and Len Armour. Location of the optimal sampling point for the quality assessment of continuous streams. Australian Journal of Statistics, 37(2):145–152, June 1995. CODEN AU-JSA3. ISSN 0004-9581.

Maritz:1995:PPT

[1260] J. S. Maritz. A permutation paired test allowing for missing values. Australian Journal of Statistics, 37(2):153–159, June 1995. CODEN AU-JSA3. ISSN 0004-9581.

Boularan:1995:LPP

[1261] Joel Boularan, Louis Ferre, and Philippe Vieu. Location of particular points in non-parametric regression analysis. Australian Journal of Statistics, 37(2):161–168, June 1995. CODEN AUJSA3. ISSN 0004-9581.

Vijayan:1995:NNU

[1262] K. Vijayan, P. Mukhopadhyay, and S. Bhattacharyya. On non-negative unbiased estimation of quadratic forms in finite population sampling. Australian Journal of Statistics, 37(2):169–178, June 1995. CODEN AU-JSA3. ISSN 0004-9581.

Lewis:1995:EAB

[1263] Toby Lewis and Nicholas I. Fisher. Estimating the angle between the mean directions of two spherical distributions. Australian Journal of Statistics, 37(2):179–191, June 1995. CODEN AUJSA3. ISSN 0004-9581.

Moyeed:1995:SSD

[1264] R. A. Moyeed. Spline smoother as a dynamic linear model. Australian Journal of Statistics, 37(2):193–204, June 1995. CODEN AUJSA3. ISSN 0004-9581.

VandeVen:1995:LLM

[1265] R. Van de Ven and N. C. Weber. Log-linear models for mean and dispersion in mixed Poisson regression models. Australian Journal of Statistics, 37(2):205–216, June 1995. CODEN AUJSA3. ISSN 0004-9581.

Yanez:1995:CQL

[1266] N. David Yanez III and Jeffrey R. Wilson. Comparison of quasi-likelihood models for overdispersion. Australian Journal of Statistics, 37(2):217– 231, June 1995. CODEN AUJSA3. ISSN 0004-9581.

Rayner:1995:STB

[1267] J. C. W. Rayner and D. J. Best. Smooth tests for the bivariate Poisson distribution. Australian Journal of Statistics, 37(2):233–245, June 1995. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1995:BRb

[1268] Anonymous. Book reviews. Australian Journal of Statistics, 37(2):247– 252, June 1995. CODEN AUJSA3. ISSN 0004-9581.

Daley:1995:TMS

[1269] D. J. Daley. Two-moment scaling formulae for aggregating examination marks. Australian Journal of Statistics, 37(3):253–272, September 1995. CODEN AUJSA3. ISSN 0004-9581.

Biggins:1995:LAS

[1270] J. D. Biggins and K. K. Yue. A likelihood approach to scaling examination marks. Australian Journal of Statistics, 37(3):273–282, September 1995. CODEN AUJSA3. ISSN 0004-9581.

Sparks:1995:QCM

[1271] Ross Sparks and Roy Ballantyne. Quality control methods in largescale assessment procedures using ‘double-marking’ or ‘partial doublemarking’. Australian Journal of Statistics, 37(3):283–298, September 1995. CODEN AUJSA3. ISSN 0004-9581.

James:1995:ADC

[1272] Ian R. James, Stephen J. Hogan, and Simon A. Mallal. Analysis of doubly censored data and the HIV incubation period distribution in Western Australia. Australian Journal of Statistics, 37(3):299–312, September 1995. CODEN AUJSA3. ISSN 0004-9581.

Chen:1995:OST

[1273] Ling Chen. A one-sided trimmed mean estimator for the mean of positively skewed distributions. Australian Journal of Statistics, 37(3):313– 322, September 1995. CODEN AUJSA3. ISSN 0004-9581.

Sommer:1995:RVS

[1274] Suzanne Sommer and Robert G. Staudte. Robust variable selection in regression in the presence of outliers and leverage points. Australian Journal of Statistics, 37(3):323–336, September 1995. CODEN AUJSA3. ISSN 0004-9581.

Jing:1995:SRP

[1275] Bing-Yi Jing. Some resampling procedures under symmetry. Australian Journal of Statistics, 37(3):337–344, September 1995. CODEN AUJSA3. ISSN 0004-9581.

Diamond:1995:SPF

[1276] Neil T. Diamond. Some properties of a foldover design. Australian Journal of Statistics, 37(3):345–352, September 1995. CODEN AUJSA3. ISSN 0004-9581.

Saunders:1995:ADF

[1277] I. W. Saunders, J. A. Eccleston, and R. J. Martin. An algorithm for the design of 2 factorial experiments on continuous processes. Australian Journal of Statistics, 37(3):353–365, September 1995. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1995:BRc

[1278] Anonymous. Book reviews. Australian Journal of Statistics, 37(3):367– 374, September 1995. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1995:LE

[1279] Anonymous. Letter to the Editor. Australian Journal of Statistics, 37 (3):375, September 1995. CODEN AUJSA3. ISSN 0004-9581.

McLennan:1996:PAB

[1280] W. McLennan. The product of the Australian Bureau of Statistics: the 1995 Knibbs Lecture, Canberra, 28 November 1995. Australian Journal of Statistics, 38(1):1–14, April 1996. CODEN AUJSA3. ISSN 0004-9581.

Steel:1996:QIT

[1281] David Steel, Joe Vella, and Peter Harrington. Quality issues in telephone surveys: Coverage, non-response and quota sampling. Australian Journal of Statistics, 38(1):15–34, April 1996. CODEN AUJSA3. ISSN 0004-9581.

Chaudhuri:1996:NCV

[1282] Arijit Chaudhuri, Tapabrata Maiti, and Debesh Roy. A note on competing variance estimators in randomised response surveys. Australian Journal of Statistics, 38(1):35–42, April 1996. CODEN AUJSA3. ISSN 0004-9581.

JinhChang:1996:TSU

[1283] Horng JinhChang and Der HsinLiang. A two-stage unrelated randomised response procedure. Australian Journal of Statistics, 38(1):43–51, April 1996. CODEN AUJSA3. ISSN 0004-9581.

McGilchrist:1996:SAT

[1284] C. A. McGilchrist and K. K. W. Yau. Survival analysis with time dependent frailty using a longitudinal model. Australian Journal of Statistics, 38(1):53–60, April 1996. CODEN AUJSA3. ISSN 0004-9581.

Henze:1996:PRC

[1285] Norbert Henze and Bernhard Klar. Properly rescaled components of smooth tests of fit are diagnostic. Australian Journal of Statistics, 38(1): 61–74, April 1996. CODEN AUJSA3. ISSN 0004-9581.

Paek:1996:CMC

[1286] Pil S. Paek and Anant M. Kshirsagar. Chance of misclassification in the case of several normal populations. Australian Journal of Statistics, 38 (1):75–82, April 1996. CODEN AUJSA3. ISSN 0004-9581.

Kshirsagar:1996:NMR

[1287] Anant M. Kshirsagar and Joanna Chung in Cheng. A new measure of rotatability of response surface designs. Australian Journal of Statistics, 38(1):83–89, April 1996. CODEN AUJSA3. ISSN 0004-9581.

Monod:1996:VRF

[1288] H. Monod, J-M. Azaïs, and R. A. Bailey. Valid randomisation for the first-difference analysis. Australian Journal of Statistics, 38(1):91–106, April 1996. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1996:BRa

[1289] Anonymous. Book reviews. Australian Journal of Statistics, 38(1):107– 111, April 1996. CODEN AUJSA3. ISSN 0004-9581.

Hoque:1996:MFA

[1290] Asraul Hoque. Multiperiod forecasting analysis of a dynamic model for a small sample. Australian Journal of Statistics, 38(2):113–129, June 1996. CODEN AUJSA3. ISSN 0004-9581.

Ng:1996:TSD

[1291] T. M. Ng and W. K. Li. Tests for seasonal differencing with an unknown break-point. Australian Journal of Statistics, 38(2):131–153, June 1996. CODEN AUJSA3. ISSN 0004-9581.

Schick:1996:CER

[1292] Anton Schick. \sqrt{n}-consistent estimation in a random coefficient autoregressive model. Australian Journal of Statistics, 38(2):155–160, June 1996. CODEN AUJSA3. ISSN 0004-9581.

Pan:1996:DFC

[1293] Guohua Pan. Distribution-free confidence procedure for umbrella orderings. Australian Journal of Statistics, 38(2):161–172, June 1996. CODEN AUJSA3. ISSN 0004-9581.

Wang:1996:DDA

[1294] J. Z. Wang, P. Cooke, and S. Li. Determination of domains of attraction based on a sequence of maxima. Australian Journal of Statistics, 38(2): 173–181, June 1996. CODEN AUJSA3. ISSN 0004-9581.

Taylor:1996:CVM

[1295] Jeremy M. G. Taylor, William G. Cumberland, and Xiangyi Meng. Components of variance models with transformations. Australian Journal of Statistics, 38(2):183–191, June 1996. CODEN AUJSA3. ISSN 0004-9581.

Young:1996:IDP

[1296] Dean M. Young, Leah M. Lehman, and Laurie M. Meaux. Independence distribution-preserving nonnegative-definite covariance structures for the sample variance. Australian Journal of Statistics, 38(2):193–200, June 1996. CODEN AUJSA3. ISSN 0004-9581.

Singh:1996:RAU

[1297] Sarjinder Singh, A. H. Joarder, and Maxwell L. King. Regression analysis using scrambled responses. Australian Journal of Statistics, 38(2):201– 211, June 1996. CODEN AUJSA3. ISSN 0004-9581.

Lefebvre:1996:HLC

[1298] Mario Lefebvre and Richard Labib. Hitting lines and circles with diffusion processes. Australian Journal of Statistics, 38(2):213–222, June 1996. CODEN AUJSA3. ISSN 0004-9581.

Preece:1996:MFB

[1299] D. A. Preece. Multi-factor balanced block designs with complete adjusted orthogonality for all pairs of treatment factors. Australian Journal of Statistics, 38(2):223–230, June 1996. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1996:BRb

[1300] Anonymous. Book reviews. Australian Journal of Statistics, 38(2):231– 234, June 1996. CODEN AUJSA3. ISSN 0004-9581.

Tang:1996:DCMa

[1301] Pak F. Tang and Neil S. Barnett. Dispersion control for multivariate processes. Australian Journal of Statistics, 38(3):235–251, September 1996. CODEN AUJSA3. ISSN 0004-9581. See also [1302].

Tang:1996:DCMb

[1302] Pak F. Tang and Neil S. Barnett. Dispersion control for multivariate processes — some comparisons. Australian Journal of Statistics, 38(3): 253–273, September 1996. CODEN AUJSA3. ISSN 0004-9581. See also

Manchester:1996:EIR

[1303] Lise Manchester. Empirical influence for robust smoothing. Australian Journal of Statistics, 38(3):275–290, September 1996. CODEN AUJSA3. ISSN 0004-9581.

Hasofer:1996:EQF

[1304] A. M. Hasofer and J. Z. Wang. Estimation of quantiles for a Fréchet-type distribution. Australian Journal of Statistics, 38(3):291–298, September 1996. CODEN AUJSA3. ISSN 0004-9581.

Sapatinas:1996:GRR

[1305] Theofanis Sapatinas. On the generalized Rao–Rubin condition and some variants. Australian Journal of Statistics, 38(3):299–306, September 1996. CODEN AUJSA3. ISSN 0004-9581.

Paula:1996:IDP

[1306] Gilberto A. Paula. Influence diagnostics in proper dispersion models. Australian Journal of Statistics, 38(3):307–316, September 1996. CO-DEN AUJSA3. ISSN 0004-9581.

Chaiyapong:1996:IIR

[1307] Y. Chaiyapong and C. J. Lloyd. Improved inference from recapture experiments with behavioural response through modelling and auxiliary experimentation. Australian Journal of Statistics, 38(3):317–331, September 1996. CODEN AUJSA3. ISSN 0004-9581.

Longford:1996:MBV

[1308] N. T. Longford. Model-based variance estimation in surveys with stratified clustered design. Australian Journal of Statistics, 38(3):333–352, September 1996. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1996:BRc

[1309] Anonymous. Book reviews. Australian Journal of Statistics, 38(3):353– 358, September 1996. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1997:PMA

[1310] Anonymous. Pitman Medal awarded to W. J. Ewens. Australian Journal of Statistics, 39(1):1–3, March 1997. CODEN AUJSA3. ISSN 0004-9581.

Saei:1997:RTM

[1311] Ayoub Saei and C. A. McGilchrist. Random threshold models applied to inflated zero class data. Australian Journal of Statistics, 39(1):5–16, March 1997. CODEN AUJSA3. ISSN 0004-9581.

Wang:1997:EGS

[1312] You-Gan Wang and Ming Yin. Estimation for the general sample selection models. Australian Journal of Statistics, 39(1):17–24, March 1997. CODEN AUJSA3. ISSN 0004-9581.

Jing:1997:TSN

[1313] Bing-Yi Jing and John Robinson. Two-sample nonparametric tilting method. Australian Journal of Statistics, 39(1):25–34, March 1997. CO-DEN AUJSA3. ISSN 0004-9581.

Krzanowski:1997:AER

[1314] W. J. Krzanowski and D. J. Hand. Assessing error rate estimators: the leave-one-out method reconsidered. Australian Journal of Statistics, 39 (1):35–46, March 1997. CODEN AUJSA3. ISSN 0004-9581.

Chen:1997:ELB

[1315] Song Xi Chen. Empirical likelihood-based kernel density estimation. Australian Journal of Statistics, 39(1):47–56, March 1997. CODEN AUJSA3. ISSN 0004-9581.

Jones:1997:FSE

[1316] M. C. Jones and R. J. Karunamuni. Fourier series estimation for length biased data. Australian Journal of Statistics, 39(1):57–68, March 1997. CODEN AUJSA3. ISSN 0004-9581.

Wesolowski:1997:CDL

[1317] Jacek Wesolowski and Mohammad Ahsanullah. On characterizing distributions via linearity of regression for order statistics. Australian Journal of Statistics, 39(1):69–78, March 1997. CODEN AUJSA3. ISSN 0004-9581.

Liu:1997:MTD

[1318] W. Liu. On a multiple three-decision procedure for comparing several treatments with a control. Australian Journal of Statistics, 39(1):79–92, March 1997. CODEN AUJSA3. ISSN 0004-9581.

Jensen:1997:CCP

[1319] D. R. Jensen. Conditioning and concentration of principal components. Australian Journal of Statistics, 39(1):93–104, March 1997. CODEN AUJSA3. ISSN 0004-9581.

Moffatt:1997:GLC

[1320] Peter G. Moffatt. Global log-concavity of the likelihood in models with grouped discrete data. Australian Journal of Statistics, 39(1):105–112, March 1997. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1997:BRa

[1321] Anonymous. Book reviews. Australian Journal of Statistics, 39(1):113– 118, March 1997. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1997:ARSa

[1322] Anonymous. Acknowledgement of referees’ services. Australian Journal of Statistics, 39(1):119–120, March 1997. CODEN AUJSA3. ISSN 0004-9581.

Vere-Jones:1997:BVG

[1323] David Vere-Jones. Boris Vladimirovich Gnedenko, 1912–1995 a personal tribute. Australian Journal of Statistics, 39(2):121–128, June 1997. CO-DEN AUJSA3. ISSN 0004-9581.

Wei:1997:SSO

[1324] Bo-Cheng Wei and Hong-Tu Zhu. Some second order asymptotics in exponential family non-linear regression models (a geometric approach). Australian Journal of Statistics, 39(2):129–148, June 1997. CODEN AU-JSA3. ISSN 0004-9581.

Morton:1997:EEN

[1325] Richard Morton, Mike Patefield, and Phil Bowtell. Estimating equations and non-linear functional relationships. Australian Journal of Statistics, 39(2):149–167, June 1997. CODEN AUJSA3. ISSN 0004-9581.

Liu:1997:STW

[1326] Wei Liu. Stepwise tests when the test statistics are independent. Australian Journal of Statistics, 39(2):169–177, June 1997. CODEN AU-JSA3. ISSN 0004-9581.

Pakes:1997:NRN

[1327] Anthony G. Pakes and Fred W. Steutel. On the number of records near the maximum. Australian Journal of Statistics, 39(2):179–192, June 1997. CODEN AUJSA3. ISSN 0004-9581.

Kabaila:1997:TUC

[1328] Paul Kabaila and Chris J. Lloyd. Tight upper confidence limits from discrete data. Australian Journal of Statistics, 39(2):193–204, June 1997. CODEN AUJSA3. ISSN 0004-9581.

Hirotsu:1997:TWC

[1329] Chihiro Hirotsu. Two-way change-point model and its application. Australian Journal of Statistics, 39(2):205–218, June 1997. CODEN AU-JSA3. ISSN 0004-9581.

Sato:1997:SRM

[1330] Michikazu Sato. Some remarks on the mean, median, mode and skewness. Australian Journal of Statistics, 39(2):219–224, June 1997. CODEN AUJSA3. ISSN 0004-9581.

Wang:1997:EBC

[1331] You-Gan Wang. Error bounds for calculation of the Gittins indices. Australian Journal of Statistics, 39(2):225–233, June 1997. CODEN AU-JSA3. ISSN 0004-9581.

Zou:1997:LET

[1332] Guohua Zou. Letters to the Editor: Two-stage randomized response procedures as single stage procedures. Australian Journal of Statistics, 39(2):235–236, June 1997. CODEN AUJSA3. ISSN 0004-9581.

Draper:1997:LEK

[1333] Norman R. Draper and Friedrich Pukelsheim. Letters to the Editor: Kshirsagar & Cheng’s rotatability measure. Australian Journal of Statistics, 39(2):236–238, June 1997. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1997:BRb

[1334] Anonymous. Book reviews. Australian Journal of Statistics, 39(2):239– 244, June 1997. CODEN AUJSA3. ISSN 0004-9581.

Sheather:1997:EAN

[1335] Simon J. Sheather. Editorial announcement: New journal structure. Australian Journal of Statistics, 39(2):245–246, June 1997. CODEN AU-JSA3. ISSN 0004-9581.

Lin:1997:OAC

[1336] G. D. Lin and J. S. Huang. Orginal article: The cube of a logistic distribution is indeterminate. Australian Journal of Statistics, 39(3): 247–252, September 1997. CODEN AUJSA3. ISSN 0004-9581.

Kakarala:1997:OAL

[1337] Ramakrishna Kakarala and Geoffrey S. Watson. Orginal article: Lower bounds for the divergence of directional and axial estimators. Australian Journal of Statistics, 39(3):253–260, September 1997. CODEN AUJSA3. ISSN 0004-9581.

Kauermann:1997:OAN

[1338] Goran Kauermann. Orginal article: a note on multivariate logistic models for contingency tables. Australian Journal of Statistics, 39(3):261–276, September 1997. CODEN AUJSA3. ISSN 0004-9581.

Chaturvedi:1997:OAB

[1339] Anoop Chaturvedi, Hikaru Hasegawa, and Sheel Asthana. Orginal article: Bayesian analysis of the linear regression model with non-normal disturbances. Australian Journal of Statistics, 39(3):277–293, September 1997. CODEN AUJSA3. ISSN 0004-9581.

Ravishanker:1997:OAB

[1340] Nalini Ravishanker and Bonnie K. Ray. Orginal article: Bayesian analysis of vector ARFIMA processes. Australian Journal of Statistics, 39(3): 295–311, September 1997. CODEN AUJSA3. ISSN 0004-9581. See corrigendum [2289].

Hyndman:1997:OAN

[1341] R. J. Hyndman and M. P. Wand. Orginal article: Nonparametric autocovariance function estimation. Australian Journal of Statistics, 39(3): 313–324, September 1997. CODEN AUJSA3. ISSN 0004-9581.

Gimenez:1997:OAC

[1342] Patricia Gimenez and Heleno Bolfarine. Orginal article: Corrected score functions in classical error-in-variables and incidental parameter models. Australian Journal of Statistics, 39(3):325–344, September 1997. CO-DEN AUJSA3. ISSN 0004-9581.

Tang:1997:OAC

[1343] M. K. Tang and Wing K. Fung. Orginal article: Case-deletion diagnostics for test statistics in multivariate regression. Australian Journal of Statistics, 39(3):345–353, September 1997. CODEN AUJSA3. ISSN 0004-9581.

Best:1997:OAG

[1344] D. J. Best and J. C. W. Rayner. Orginal article: Goodness of fit for the binomial distribution. Australian Journal of Statistics, 39(3):355–364, September 1997. CODEN AUJSA3. ISSN 0004-9581.

Robinson:1997:LE

[1345] Geoff Robinson. Letter to the Editor. Australian Journal of Statistics, 39(3):365–366, September 1997. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1997:BRc

[1346] Anonymous. Book reviews. Australian Journal of Statistics, 39(3):367– 370, September 1997. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1997:ARSb

[1347] Anonymous. Acknowledgment of referees’ services. Australian Journal of Statistics, 39(3):371, September 1997. CODEN AUJSA3. ISSN 0004-9581.

Anonymous:1997:FP

[1348] Anonymous. Forthcoming papers. Australian Journal of Statistics, 39 (3):372, September 1997. CODEN AUJSA3. ISSN 0004-9581.

Chan:1998:AMC

[1349] Jennifer S. K. Chan, Anthony Y. C. Kuk, James Bell, and Charles Mc Gilchrist. The analysis of methadone clinic data using marginal and conditional logistic models with mixture or random effects. Australian & New Zealand Journal of Statistics, 40(1):1–10, March 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chetwynd:1998:ERS

[1350] A. G. Chetwynd and P. J. Diggle. On estimating the reduced second moment measure of a stationary spatial point process. Australian & New Zealand Journal of Statistics, 40(1):11–15, March 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Sato:1998:LDR

[1351] Toshiyuki Sato, Yoshihide Kakizawa, and Masanobu Taniguchi. Large deviation results for statistics of short- and long-memory Gaussian processes. Australian & New Zealand Journal of Statistics, 40(1):17–29, March 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hay:1998:AVP

[1352] J. L. Hay and A. N. Pettitt. The added variable plot for a time series of counts. Australian & New Zealand Journal of Statistics, 40(1):31– 42, March 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Leung:1998:BRU

[1353] Denis Heng-Yan Leung and You-Gan Wang. Bias reduction using stochastic approximation. Australian & New Zealand Journal of Statistics, 40(1):43–52, March 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Field:1998:RCI

[1354] Christopher A. Field and A. H. Welsh. Robust confidence intervals for regression parameters. Australian & New Zealand Journal of Statistics, 40(1):53–64, March 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kim:1998:DAR

[1355] Soon-Kwi Kim and Richard Huggins. Diagnostics for autocorrelated regression models. Australian & New Zealand Journal of Statistics, 40(1): 65–71, March 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gao:1998:SBD

[1356] Hongsheng Gao and Peter Smith. Some bounds on the distribution of certain quadratic forms in normal random variables. Australian & New Zealand Journal of Statistics, 40(1):73–81, March 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Barakat:1998:WLS

[1357] H. M. Barakat. Weak limit of the sample extremal quotient. Australian & New Zealand Journal of Statistics, 40(1):83–93, March 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gharib:1998:SER

[1358] M. Gharib. Some estimates for the rate of convergence in the multidimensional central limit theorem for homogeneous Markov chains. Australian & New Zealand Journal of Statistics, 40(1):95–102, March 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Nomakuchi:1998:GCM

[1359] Kentaro Nomakuchi. A generalization of the Childs–Moran result for orthoscheme probabilities. Australian & New Zealand Journal of Statistics, 40(1):103–109, March 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1998:FPa

[1360] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 40(1):110, March 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

John:1998:LD

[1361] J. A. John and E. R. Williams. t-latinized designs. Australian & New Zealand Journal of Statistics, 40(1):111–118, March 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1998:BRa

[1362] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 40(1):119–128, March 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Huggins:1998:CMF

[1363] R. M. Huggins, D. Z. Loesch, and N. H. Hoang. A comparison of methods of fitting models to twin data. Australian & New Zealand Journal of Statistics, 40(2):129–140, June 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chatterjee:1998:HCE

[1364] Srikanta Chatterjee and Claudio Michelini. Household consumption equivalence scales: Some estimates from New Zealand household expenditure and income survey data. Australian & New Zealand Journal of Statistics, 40(2):141–150, June 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bui:1998:RIB

[1365] Quang M. Bui and Richard M. Huggins. Robust inference for the bivariate bifurcating autoregressive model. Australian & New Zealand Journal of Statistics, 40(2):151–163, June 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hu:1998:EAP

[1366] Xiaomi Hu. An exact algorithm for projection onto a polyhedral cone. Australian & New Zealand Journal of Statistics, 40(2):165–170, June 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hasegawa:1998:RCI

[1367] Hikaru Hasegawa. Risk comparison of inequality constrained estimators in the heteroscedastic linear model. Australian & New Zealand Journal of Statistics, 40(2):171–180, June 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ohtani:1998:SPI

[1368] Kazuhiro Ohtani and Alan T. K. Wan. On the sampling performance of an improved Stein inequality restricted estimator. Australian & New Zealand Journal of Statistics, 40(2):181–187, June 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kabaila:1998:CSW

[1369] Paul Kabaila. The choice of statistic on which to base tight upper confidence limits. Australian & New Zealand Journal of Statistics, 40(2): 189–196, June 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Konijn:1998:CIM

[1370] H. S. Konijn. Certain inverse multivariate hypergeometric and multinomial distributions. Australian & New Zealand Journal of Statistics, 40 (2):197–204, June 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cuculescu:1998:MSU

[1371] Ioan Cuculescu and Radu Theodorescu. Multiplicative strong unimodality. Australian & New Zealand Journal of Statistics, 40(2):205–214, June 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Yeo:1998:ICS

[1372] Geoffrey F. Yeo. Interview costs in the secretary problem. Australian & New Zealand Journal of Statistics, 40(2):215–219, June 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pham-Gia:1998:DST

[1373] T. Pham-Gia. Distribution of the stopping time in Bayesian sequential sampling. Australian & New Zealand Journal of Statistics, 40(2):221– 227, June 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Datta:1998:NAP

[1374] Somnath Datta, William P. McCormick, and George Mathew. Nonlinear autoregression with positive innovations. Australian & New Zealand Journal of Statistics, 40(2):229–239, June 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bose:1998:DRM

[1375] Mausumi Bose. Designs with repeated measurements on any number of units over varying periods. Australian & New Zealand Journal of Statistics, 40(2):241–248, June 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1998:BRb

[1376] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 40(2):249–255, June 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1998:FPb

[1377] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 40(2):256, June 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Fisher:1998:TOT

[1378] Nicholas I. Fisher and Peter Hall. Tributes and obituary: Tribute: Geoffrey Stuart Watson (3 December 1921–3 January 1998). Australian & New Zealand Journal of Statistics, 40(3):257–268, September 1998. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wells:1998:AOT

[1379] J. Wells. Applications: Oversampling through households or other clusters: Comparisons of methods for weighting the oversampled elements. Australian & New Zealand Journal of Statistics, 40(3):269–278, September 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Strachan:1998:TML

[1380] Rodney Strachan, Maxwell King, and Sarjinder Singh. Theory and methods: Likelihood-based estimation of the regression model with scrambled responses. Australian & New Zealand Journal of Statistics, 40(3):279– 290, September 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Singh:1998:TME

[1381] Sarjinder Singh, Stephen Horn, and Sadeq Chowdhury. Theory and methods: Estimation of stigmatized characteristics of a hidden gang in a finite population. Australian & New Zealand Journal of Statistics, 40 (3):291–298, September 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wang:1998:TMG

[1382] You-Gan Wang. Theory and methods: Growth curves with explanatory variables and estimation of the effect of tagging. Australian & New Zealand Journal of Statistics, 40(3):299–304, September 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

FitzGerald:1998:TME

[1383] Patrick E. B. FitzGerald and Matthew W. Knuiman. Theory and methods: Estimation in regressive logistic regression analyses of familial data with missing outcomes. Australian & New Zealand Journal of Statistics, 40(3):305–316, September 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kozek:1998:TMM

[1384] Andrzej S. Kozek. Theory and methods: on minimum distance estimation using Kolmogorov–Lévy type metrics. Australian & New Zealand Journal of Statistics, 40(3):317–334, September 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Givaruangsawat:1998:TMH

[1385] Sumalee Givaruangsawat, Govinda J. Weerakkody, and Patrick D. Gerard. Theory and methods: Hypothesis testing in two-stage cluster sampling. Australian & New Zealand Journal of Statistics, 40(3):335–344, September 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Zhao:1998:TMI

[1386] Yuejen Zhao and Andy H. Lee. Theory and methods: Influence diagnostics for simultaneous equations models. Australian & New Zealand Journal of Statistics, 40(3):345–358, September 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chan:1998:TMO

[1387] L. Y. Chan, J. H. Meng, Y. C. Jiang, and Y. N. Guan. Theory and methods: D-optimal axial designs for quadratic and cubic additive mixture models. Australian & New Zealand Journal of Statistics, 40(3):359–372, September 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ahmed:1998:TMN

[1388] M. S. Ahmed. Theory and methods: a note on regression-type estimators using multiple auxiliary information. Australian & New Zealand Journal of Statistics, 40(3):373–376, September 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See rejoinder [1472].

Anonymous:1998:BRc

[1389] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 40(3):377–383, September 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1998:FPc

[1390] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 40(3):384, September 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1998:ANP

[1391] Anonymous. Announcement notice: Pitman Medal awarded to E. Seneta. Australian & New Zealand Journal of Statistics, 40(4):385–387, December 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Catchpole:1998:ACP

[1392] E. A. Catchpole and B. R. Anderson. Applications: Comparing the performances of NSW and Victorian students in a first-year mathematics course (with discussion). Australian & New Zealand Journal of Statistics, 40(4):389–406, December 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Feltz:1998:TMG

[1393] Carol J. Feltz. Theory & methods: Generalizations of the δ-corrected Kolmogorov–Smirnov goodness-of-fit test. Australian & New Zealand Journal of Statistics, 40(4):407–413, December 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ozturk:1998:TMR

[1394] Ömer Öztürk. Theory & methods: a robust and almost fully efficient M-estimator. Australian & New Zealand Journal of Statistics, 40(4):415– 424, December 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Raghavarao:1998:TME

[1395] Damaraju Raghavarao and James B. Wiley. Theory & methods: Estimating main effects with Pareto optimal subsets. Australian & New Zealand Journal of Statistics, 40(4):425–432, December 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pigeot:1998:TMJ

[1396] I. Pigeot and H. Strugholtz. Theory & methods: Jackknifing the Mantel– Haenszel estimator in sparse contingency tables. Australian & New Zealand Journal of Statistics, 40(4):433–441, December 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Futschik:1998:TMD

[1397] Andreas Futschik and Georg Ch. Pflug. Theory & methods: Distance testing for selecting the best population. Australian & New Zealand Journal of Statistics, 40(4):443–464, December 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Beh:1998:TMP

[1398] Eric J. Beh and Pamela J. Davy. Theory & methods: Partitioning Pearson’s chi-squared statistic for a completely ordered three-way contingency table. Australian & New Zealand Journal of Statistics, 40(4):465–477, December 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ma:1998:TMS

[1399] Chunsheng Ma and John Robinson. Theory & methods: Saddlepoint approximation for sample and bootstrap quantiles. Australian & New Zealand Journal of Statistics, 40(4):479–486, December 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Draper:1998:TMR

[1400] N. R. Draper and J. A. John. Theory & methods: Response surface designs where levels of some factors are difficult to change. Australian & New Zealand Journal of Statistics, 40(4):487–495, December 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1998:BRd

[1401] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 40(4):497–505, December 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1998:FPd

[1402] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 40(4):506, December 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1998:IIA

[1403] Anonymous. Index: Index to articles and book reviews in volume 40. Australian & New Zealand Journal of Statistics, 40(4):507–512, December 1998. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Sheather:1999:ERD

[1404] Simon Sheather. Editorial: Retirement of Daryl Daley as Technical Editor. Australian & New Zealand Journal of Statistics, 41(1):1, March 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Meyer:1999:AOE

[1405] Denny Meyer and Tim Napier-Munn. Applications: Optimal experiments for time-dependent mineral processes. Australian & New Zealand Journal of Statistics, 41(1):3–17, March 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

McBride:1999:AET

[1406] Graham B. McBride. Applications: Equivalence tests can enhance environmental science and management. Australian & New Zealand Journal of Statistics, 41(1):19–29, March 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1999:ARS

[1407] Anonymous. Acknowledgment of referees’ services. Australian & New Zealand Journal of Statistics, 41(1):30, March 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Vignaux:1999:TMS

[1408] V. A. Vignaux and J. L. Scott. Theory & methods: Simplifying regression models using dimensional analysis. Australian & New Zealand Journal of Statistics, 41(1):31–41, March 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1999:FPa

[1409] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 41(1):42, March 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Balemi:1999:TMS

[1410] Andrew Balemi and Alan Lee. Theory & methods: Some properties of the Liang–Zeger method applied to clustered binary regression. Australian & New Zealand Journal of Statistics, 41(1):43–58, March 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Guo:1999:TMN

[1411] Jiin-Huarng Guo. Theory & methods: a nonparametric test for the parallelism of two first-order autoregressive processes. Australian & New Zealand Journal of Statistics, 41(1):59–65, March 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Buning:1999:TMA

[1412] Herbert Büning and Wolfgang Kössler. Theory & methods: The asymptotic power of Jonckheere-type tests for ordered alternatives. Australian & New Zealand Journal of Statistics, 41(1):67–77, March 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Stefanov:1999:TME

[1413] V. T. Stefanov and A. G. Pakes. Theory & methods: Explicit distributional results in pattern formation II. Australian & New Zealand Journal of Statistics, 41(1):79–90, March 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See corrigenda [1428].

Dollinger:1999:TMGa

[1414] M. B. Dollinger, E. Kulinskaya, and R. G. Staudte. Theory & methods: Guarded weights of evidence and acceptability profiles based on signs: I. Permutation arguments (with discussion). Australian & New Zealand Journal of Statistics, 41(1):91–110, March 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See also Part II [1447].

John:1999:TMR

[1415] J. A. John, K. G. Russell, E. R. Williams, and D. Whitaker. Theory & methods: Resolvable designs with unequal block sizes. Australian & New Zealand Journal of Statistics, 41(1):111–116, March 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1999:BRa

[1416] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 41(1):117–126, March 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Nicholls:1999:OLP

[1417] Des Nicholls. Oration lecture: Presidential address: Statistics into the 21st century (with discussion). Australian & New Zealand Journal of Statistics, 41(2):127–139, June 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lee:1999:AMR

[1418] Alan Lee. Applications: Modelling rugby league data via bivariate negative binomial regression. Australian & New Zealand Journal of Statistics, 41(2):141–152, June 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hunt:1999:TMM

[1419] Lynette Hunt and Murray Jorgensen. Theory & methods: Mixture model clustering using the multimix program. Australian & New Zealand Journal of Statistics, 41(2):154–171, June 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gupta:1999:TMG

[1420] Rameshwar D. Gupta and Debasis Kundu. Theory & methods: Generalized exponential distributions. Australian & New Zealand Journal of Statistics, 41(2):173–188, June 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cordeiro:1999:TMS

[1421] Gauss M. Cordeiro and Klaus L. P. Vasconcellos. Theory & methods: Second-order biases of the maximum likelihood estimates in von Mises regression models. Australian & New Zealand Journal of Statistics, 41(2): 189–198, June 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Singh:1999:TMC

[1422] Sarjinder Singh, Stephen Horn, Sadeq Chowdhury, and Frank Yu. Theory & methods: Calibration of the estimators of variance. Australian & New Zealand Journal of Statistics, 41(2):199–212, June 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Adamidis:1999:TME

[1423] Konstantinos Adamidis. Theory & methods: an EM algorithm for estimating negative binomial parameters. Australian & New Zealand Journal of Statistics, 41(2):213–221, June 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hasofer:1999:TME

[1424] A. M. Hasofer and S. Li. Theory & methods: Estimation for Type II domain of attraction based on the W statistic. Australian & New Zealand Journal of Statistics, 41(2):223–232, June 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Beh:1999:TMP

[1425] Eric J. Beh and Pamela J. Davy. Theory & methods: Partitioning Pearson’s chi-squared statistic for a partially ordered three-way contingency table. Australian & New Zealand Journal of Statistics, 41(2):233–246, June 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1999:BRb

[1426] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 41(2):247–253, June 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1999:FPb

[1427] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 41(2):253, June 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Stefanov:1999:CED

[1428] V. T. Stefanov and A. G. Pakes. Corrigenda: Explicit distributional results in pattern formation II. Australian & New Zealand Journal of Statistics, 41(2):254, June 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1413].

Mackinnon:1999:ADM

[1429] Murray J. Mackinnon and Ned Glick. Applications: Data mining and knowledge discovery in databases — an overview. Australian & New Zealand Journal of Statistics, 41(3):255–275, September 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Williams:1999:AEB

[1430] E. R. Williams, J. A. John, and D. Whitaker. Applications: Example of block designs for plant and tree breeding trials. Australian & New Zealand Journal of Statistics, 41(3):277–284, September 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Taplin:1999:TMA

[1431] Ross H. Taplin. Theory & methods: an analysis of two-dimensional agricultural field trials in the presence of outliers and fertility jumps. Australian & New Zealand Journal of Statistics, 41(3):285–297, September 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lloyd:1999:TMC

[1432] Christopher J. Lloyd and Yongyuth Chaiyapong. Theory & methods: Choosing sampling effort for recapture experiments with behavioural response. Australian & New Zealand Journal of Statistics, 41(3):299–304, September 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lee:1999:TMP

[1433] Andy H. Lee and John S. Yick. Theory & methods: a perturbation approach to outlier detection in two-way contingency tables. Australian & New Zealand Journal of Statistics, 41(3):305–315, September 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Smith:1999:TMF

[1434] Murray D. Smith and Peter G. Moffatt. Theory & methods: Fisher’s information on the correlation coefficient in bivariate logistic models. Australian & New Zealand Journal of Statistics, 41(3):315–330, September 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kabaila:1999:TME

[1435] Paul Kabaila. Theory & methods: an efficient simulation method for the computation of a class of conditional expectations. Australian & New Zealand Journal of Statistics, 41(3):331–336, September 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Klar:1999:TMM

[1436] Bernhard Klar. Theory & methods: Model selection for a family of discrete failure time distributions. Australian & New Zealand Journal of Statistics, 41(3):337–352, September 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

King:1999:TMS

[1437] Robert A. R. King and H. L. MacGillivray. Theory & methods: a starship estimation method for the generalized λ distributions. Australian & New Zealand Journal of Statistics, 41(3):353–374, September 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1999:BRc

[1438] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 41(3):375–382, September 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1999:FPc

[1439] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 41(3):382, September 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1999:IV

[1440] Anonymous. Indexes for volumes 40 and 41. Australian & New Zealand Journal of Statistics, 41(4), December 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Millar:1999:AER

[1441] Russell B. Millar and Trevor J. Willis. Applications: Estimating the relative density of snapper in and around a marine reserve using a log-linear mixed-effects model. Australian & New Zealand Journal of Statistics, 41 (4):383–394, December 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Brown:1999:TMC

[1442] J. A. Brown. Theory & methods: a comparison of two adaptive sampling designs. Australian & New Zealand Journal of Statistics, 41(4):395–403, December 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Sahoo:1999:TMC

[1443] L. N. Sahoo and Pournamasi Panda. Theory & methods: a class of estimators using auxiliary information in two-stage sampling. Australian & New Zealand Journal of Statistics, 41(4):405–410, December 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Manly:1999:TMA

[1444] Bryan F. J. Manly and R. I. C. Chris Francis. Theory & methods: Analysis of variance by randomization when variances are unequal. Australian & New Zealand Journal of Statistics, 41(4):411–429, December 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Tzavelas:1999:TMC

[1445] George Tzavelas. Theory & methods: a characterization of the normal family of distributions based on the bias of the maximum likelihood estimator. Australian & New Zealand Journal of Statistics, 41(4):431–438, December 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Liu:1999:TMS

[1446] Zhihui Liu. Theory & methods: Set compound estimation under entropy loss. Australian & New Zealand Journal of Statistics, 41(4):439–450, December 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Dollinger:1999:TMGb

[1447] M. B. Dollinger, E. Kulinskaya, and R. G. Staudte. Theory & methods: Guarded weights of evidence and acceptability profiles based on signs: II. Asymptotic results. Australian & New Zealand Journal of Statistics, 41(4):451–461, December 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See also Part I [1414].

Anh:1999:TML

[1448] Vo Anh, Rodney Wolff, Jiti Gao, and Quang Tieng. Theory & methods: Local linear kernel regression with long-range dependent errors. Australian & New Zealand Journal of Statistics, 41(4):463–479, December 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bofinger:1999:TMH

[1449] Eve Bofinger. Theory & methods: Homogeneity of two straight lines: Equivalence approach using fitted regressions. Australian & New Zealand Journal of Statistics, 41(4):481–491, December 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1999:BRd

[1450] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 41(4):493–501, December 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:1999:FPd

[1451] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 41(4):502, December 1999. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Durbin:2000:FLS

[1452] J. Durbin. The Foreman Lecture: The state space approach to time series analysis and its potential for official statistics (with discussion). Australian & New Zealand Journal of Statistics, 42(1):1–23, March 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Sparks:2000:ACC

[1453] Ross S. Sparks. Applications: Cusum charts for Ar1 data: are they worth the effort? Australian & New Zealand Journal of Statistics, 42(1): 25–42, March 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hunt:2000:TMW

[1454] Lesley Hunt. Theory & methods: Why do we do it? Statisticians and the practice of statistics. Australian & New Zealand Journal of Statistics, 42 (1):43–58, March 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Koudou:2000:TMC

[1455] A. E. Koudou and D. Pommeret. Theory & methods: a construction of Lancaster probabilities with margins in the multidimensional Meixner class. Australian & New Zealand Journal of Statistics, 42(1):59–66, March 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kabaila:2000:TMP

[1456] Paul Kabaila and Chris J. Lloyd. Theory & methods: Profile upper confidence limits from discrete data. Australian & New Zealand Journal of Statistics, 42(1):67–79, March 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

ONeill:2000:TMW

[1457] Michael E. O’Neill and Ky Mathews. Theory & methods: a weighted least squares approach to Levene’s test of homogeneity of variance. Australian & New Zealand Journal of Statistics, 42(1):81–100, March 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Blest:2000:TMR

[1458] David C. Blest. Theory & methods: Rank correlation — an alternative measure. Australian & New Zealand Journal of Statistics, 42(1):101– 111, March 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Clarke:2000:TMN

[1459] Brenton R. Clarke, David K. Gamble, and Tadeusz Bednarski. Theory & methods: a note on robustness of the β-trimmed mean. Australian & New Zealand Journal of Statistics, 42(1):113–117, March 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See corrigendum [1990].

Anonymous:2000:BRa

[1460] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 42(1):119–126, March 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2000:FPa

[1461] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 42(1):126, March 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2000:FPb

[1462] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 42(2), June 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2000:ARS

[1463] Anonymous. Acknowledgment of referees’ services. Australian & New Zealand Journal of Statistics, 42(2), June 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Michelini:2000:ANZ

[1464] Claudio Michelini. Applications: New Zealand consumption equivalence scales from quasi-unit record data. Australian & New Zealand Journal of Statistics, 42(2):127–144, June 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hyndman:2000:AGA

[1465] Rob J. Hyndman and Gary K. Grunwald. Applications: Generalized additive modelling of mixed distribution Markov models with application to Melbourne’s rainfall. Australian & New Zealand Journal of Statistics, 42(2):145–158, June 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bishop:2000:AGE

[1466] Janet Bishop, David Die, and You-Gan Wang. Applications: a generalized estimating equations approach for analysis of the impact of new technology on a trawl fishery. Australian & New Zealand Journal of Statistics, 42(2):159–177, June 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Baringhaus:2000:TMW

[1467] Ludwig Baringhaus, Nora Gürtler, and Norbert Henze. Theory & methods: Weighted integral test statistics and components of smooth tests of fit. Australian & New Zealand Journal of Statistics, 42(2):179–192, June 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lloyd:2000:TMF

[1468] Chris J. Lloyd. Theory & methods: Fitting ROC curves using non-linear binomial regression. Australian & New Zealand Journal of Statistics, 42 (2):193–204, June 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ruppert:2000:TMS

[1469] David Ruppert and Raymond J. Carroll. Theory & methods: Spatiallyadaptive penalties for spline fitting. Australian & New Zealand Journal of Statistics, 42(2):205–223, June 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Constantine:2000:TMP

[1470] A. G. Constantine, J. B. F. Field, and N. I. Robinson. Theory & methods: Probabilities of failure in mixed acceptance sampling schemes. Australian & New Zealand Journal of Statistics, 42(2):225–233, June 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bose:2000:TMC

[1471] Mausumi Bose and Bhramar Mukherjee. Theory & methods: Cross-over designs in the presence of higher order carry-overs. Australian & New Zealand Journal of Statistics, 42(2):235–244, June 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Mukerjee:2000:LRA

[1472] Rahul Mukerjee, T. J. Rao, and K. Vijayan. Letters: Rejoinder to Ahmed, M. S. (1998). “A note on regression-type estimators using multiple auxiliary information”. Australian & New Zealand Journal of Statistics, 42(2):245, June 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1388].

Anonymous:2000:BRb

[1473] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 42(2):247–254, June 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2000:FPc

[1474] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 42(3):0, September 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bennett:2000:ELS

[1475] Don J. Bennett and David Steel. An evaluation of a large-scale CATI household survey using random digit dialling. Australian & New Zealand Journal of Statistics, 42(3):255–270, September 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Parisi:2000:SRP

[1476] Francis Parisi and Robert Lund. Seasonality and return periods of landfalling Atlantic Basin hurricanes. Australian & New Zealand Journal of Statistics, 42(3):271–282, September 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Baddeley:2000:PMP

[1477] Adrian Baddeley and Rolf Turner. Practical maximum pseudolikelihood for spatial point patterns: (with discussion). Australian & New Zealand Journal of Statistics, 42(3):283–322, September 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See addendum [1599].

Chen:2000:PEB

[1478] Di Chen, Jye-Chyi Lu, Chin-Shang Li, and Jinho Park. Parameter estimation for bivariate shock models with singular distribution for censored data with concomitant order statistics. Australian & New Zealand Journal of Statistics, 42(3):323–336, September 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Deshmukh:2000:MS

[1479] S. R. Deshmukh. Markov sampling. Australian & New Zealand Journal of Statistics, 42(3):337–345, September 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Artes:2000:ACL

[1480] Rinaldo Artes, Gilberto A. Paula, and Ronald Ranvaud. Analysis of circular longitudinal data based on generalized estimating equations. Australian & New Zealand Journal of Statistics, 42(3):347–358, September 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hu:2000:TMI

[1481] Xiaomi Hu. Tests for monotonic intensity in a Poisson process. Australian & New Zealand Journal of Statistics, 42(3):359–365, September 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Yip:2000:PLE

[1482] Paul S. F. Yip and Qizhi Chen. A partial likelihood estimator of vaccine efficacy. Australian & New Zealand Journal of Statistics, 42(3):367–374, September 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2000:BRc

[1483] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 42(3):375–382, September 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chotikapanich:2000:APD

[1484] Duangkamon Chotikapanich and William E. Griffiths. Applications: Posterior distributions for the Gini coefficient using grouped data. Australian & New Zealand Journal of Statistics, 42(4):383–392, December 2000. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lee:2000:TME

[1485] S. J. Lee. Theory & methods: On estimators of scale and dispersion in neighbourhoods of parametric families. Australian & New Zealand Journal of Statistics, 42(4):393–405, December 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Montanari:2000:TMC

[1486] Giorgio E. Montanari. Theory & methods: Conditioning on auxiliary variable means in finite population inference. Australian & New Zealand Journal of Statistics, 42(4):407–421, December 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Huggins:2000:TMI

[1487] R. M. Huggins and I. V. Basawa. Theory & methods: Inference for the extended bifurcating autoregressive model for cell lineage studies. Australian & New Zealand Journal of Statistics, 42(4):423–432, December 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Rodrigues:2000:TME

[1488] Josemar Rodrigues, José Galvão Leite, and Luis A. Milan. Theory & methods: an empirical Bayes inference for the von Mises distribution. Australian & New Zealand Journal of Statistics, 42(4):433–440, December 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Altman:2000:TMK

[1489] Naomi Altman. Theory & methods: Krige, smooth, both or neither?: (with discussion). Australian & New Zealand Journal of Statistics, 42 (4):441–461, December 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Fraccaro:2000:TMR

[1490] Richard Fraccaro, Rob J. Hyndman, and Alan Veevers. Theory & methods: Residual diagnostic plots for checking for model mis-specification in time series regression. Australian & New Zealand Journal of Statistics, 42(4):463–477, December 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Grunwald:2000:TMN

[1491] Gary K. Grunwald, Rob J. Hyndman, Leanna Tedesco, and Richard L. Tweedie. Theory & methods: Non-Gaussian conditional linear Ar(1) models. Australian & New Zealand Journal of Statistics, 42(4):479–495, December 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2000:LE

[1492] Anonymous. Letter to the Editor. Australian & New Zealand Journal of Statistics, 42(4):497–498, December 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2000:BRd

[1493] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 42(4):499–506, December 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2000:IIV

[1494] Anonymous. Index: Index to volume 42, 2000. Australian & New Zealand Journal of Statistics, 42(4):507–509, December 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2000:FPd

[1495] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 42(4):510, December 2000. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2001:ANP

[1496] Anonymous. Announcement notice: Pitman Medal for 2000 awarded to Graham Wilkinson. Australian & New Zealand Journal of Statistics, 43 (1):1–3, March 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2001:ARS

[1497] Anonymous. Acknowledgment of referees’ services. Australian & New Zealand Journal of Statistics, 43(1):4, March 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Sheather:2001:EA

[1498] Simon Sheather. Editorial announcement. Australian & New Zealand Journal of Statistics, 43(1):5–6, March 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Manda:2001:CMA

[1499] Samuel O. M. Manda. A comparison of methods for analysing a nested frailty model to child survival in Malawi. Australian & New Zealand Journal of Statistics, 43(1):7–16, March 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lin:2001:NCP

[1500] Gwo Dong Lin. A note on the characterization of positive Linnik laws. Australian & New Zealand Journal of Statistics, 43(1):17–20, March 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Dunkl:2001:CGD

[1501] Charles F. Dunkl and Donald E. Ramirez. Computation of the generalized F distribution. Australian & New Zealand Journal of Statistics, 43(1):21–31, March 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Singh:2001:MEU

[1502] Sarjinder Singh, Anwar H. Joarder, and D. S. Tracy. Median estimation using double sampling. Australian & New Zealand Journal of Statistics, 43(1):33–46, March 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Huang:2001:CTM

[1503] Fuchun Huang and Yosihiko Ogata. Comparison of two methods for calculating the partition functions of various spatial statistical models. Australian & New Zealand Journal of Statistics, 43(1):47–65, March 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ozturk:2001:GAC

[1504] Ömer Öztürk. A generalization of Ahmad’s class of Mann–Whitney– Wilcoxon statistics. Australian & New Zealand Journal of Statistics, 43 (1):67–74, March 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anderson:2001:PTL

[1505] Marti J. Anderson and John Robinson. Permutation tests for linear models. Australian & New Zealand Journal of Statistics, 43(1):75–88, March 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kulasekera:2001:TER

[1506] K. B. Kulasekera and J. Wang. A test of equality of regression curves using Gâteaux scores. Australian & New Zealand Journal of Statistics, 43(1):89–99, March 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pakes:2001:CUD

[1507] Anthony G. Pakes, Wen-Liang Hung, and Jong-Wuu Wu. Criteria for the unique determination of probability distributions by moments. Australian & New Zealand Journal of Statistics, 43(1):101–111, March 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2001:FPa

[1508] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 43(1):112, March 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chai:2001:TPF

[1509] F. S. Chai and C. T. Liao. Three parallel flats designs for two-level factorial experiments. Australian & New Zealand Journal of Statistics, 43(1):113–119, March 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2001:BRa

[1510] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 43(1):121–128, March 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Smith:2001:AAC

[1511] Alison Smith, Brian Cullis, and Arthur Gilmour. Applications: The analysis of crop variety evaluation data in Australia. Australian & New Zealand Journal of Statistics, 43(2):129–145, June 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cook:2001:TMS

[1512] R. Dennis Cook and Xiangrong Yin. Theory & methods: Special invited paper: Dimension reduction and visualization in discriminant analysis (with discussion). Australian & New Zealand Journal of Statistics, 43 (2):147–199, June 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ng:2001:TMJ

[1513] M. P. Ng and E. R. Williams. Theory & methods: Joint-regression analysis for incomplete two-way tables. Australian & New Zealand Journal of Statistics, 43(2):201–206, June 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Feltz:2001:TMP

[1514] Carol J. Feltz and Gerald A. Goldin. Theory & methods: Partitionbased goodness-of-fit tests on the line and the circle. Australian & New Zealand Journal of Statistics, 43(2):207–220, June 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Berman:2001:TMA

[1515] Simeon M. Berman. Theory & methods: Approximate identifiability, moments and censored data. Australian & New Zealand Journal of Statistics, 43(2):221–230, June 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Mosler:2001:TMT

[1516] Karl Mosler and Wilfried Seidel. Theory & methods: Testing for homogeneity in an exponential mixture model. Australian & New Zealand Journal of Statistics, 43(2):231–247, June 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2001:BRb

[1517] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 43(2):249–255, June 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2001:FPb

[1518] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 43(2):256, June 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lloyd:2001:ANS

[1519] Chris Lloyd. Announcement notice: Simon Sheather completes term as Managing Editor and Theory & Methods Editor. Australian & New Zealand Journal of Statistics, 43(3):257–258, September 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

DeSilva:2001:AEM

[1520] Basil M. De Silva, Greg R. Pond, and Tim B. Swartz. Applications: Estimation of the magnitude of victory in one-day cricket RMIT University, Mayo Clinic Rochester and Simon Fraser University. Australian & New Zealand Journal of Statistics, 43(3):259–268, September 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Zhang:2001:AEP

[1521] Xichuan (Mark) Zhang, Craig H. McLaren, and Caleb C. S. Leung. Applications: an Easter proximity effect: Modelling and adjustment. Australian & New Zealand Journal of Statistics, 43(3):269–280, September 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Salehi:2001:TMN

[1522] Mohammad M. Salehi and George A. F. Seber. Theory & methods: a new proof of Murthy’s estimator which applies to sequential sampling. Australian & New Zealand Journal of Statistics, 43(3):281–286, September 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

OGorman:2001:TMA

[1523] Thomas W. O’Gorman. Theory & methods: Adaptive estimation using weighted least squares. Australian & New Zealand Journal of Statistics, 43(3):287–297, September 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kabaila:2001:TMO

[1524] Paul Kabaila. Theory & methods: On the order of data reductions by ancillarity and by sufficiency. Australian & New Zealand Journal of Statistics, 43(3):299–301, September 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kabaila:2001:TME

[1525] Paul Kabaila and John Byrne. Theory & methods: Exact short confidence intervals from discrete data. Australian & New Zealand Journal of Statistics, 43(3):303–309, September 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hossain:2001:TMS

[1526] S. S. Hossain and H. A. Muttlak. Theory & methods: Selected ranked set sampling. Australian & New Zealand Journal of Statistics, 43(3):311– 325, September 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Beh:2001:TMP

[1527] Eric J. Beh. Theory & methods: Partitioning Pearson’s chi-squared statistic for singly ordered two-way contingency tables. Australian & New Zealand Journal of Statistics, 43(3):327–333, September 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Tomizawa:2001:TMM

[1528] Sadao Tomizawa, Nobuko Miyamoto, and Yusuke Hatanaka. Theory & methods: Measure of asymmetry for square contingency tables having ordered categories. Australian & New Zealand Journal of Statistics, 43 (3):335–349, September 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chowdhury:2001:TMA

[1529] Sadeq R. Chowdhury and C. A. McGilchrist. Theory & methods: Analysis of contingency tables with clustered observations. Australian & New Zealand Journal of Statistics, 43(3):351–358, September 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Shinjikashvili:2001:TMR

[1530] Eka Shinjikashvili. Theory & methods: ϵ-repetitions of the maximum residuals in an Ar(1) model. Australian & New Zealand Journal of Statistics, 43(3):359–365, September 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Nguyen:2001:TMC

[1531] Nam-Ky Nguyen. Theory & methods: Cutting experimental designs into blocks. Australian & New Zealand Journal of Statistics, 43(3):367–374, September 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2001:BRc

[1532] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 43(3):375–381, September 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2001:FPc

[1533] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 43(3):382, September 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2001:AEA

[1534] Anonymous. Announcement: Editorial announcement: Inaugural Editor’s award. Australian & New Zealand Journal of Statistics, 43(4):383, December 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Doherty:2001:ASH

[1535] Mike Doherty. Applications: The surrogate Henderson filters in X-11. Australian & New Zealand Journal of Statistics, 43(4):385–392, December 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Vere-Jones:2001:AHS

[1536] David Vere-Jones, Murray J. MacKinnon, and Brian Silverstone. Applications: Harold Silverstone: a perspective. Australian & New Zealand Journal of Statistics, 43(4):393–398, December 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Terpstra:2001:TMW

[1537] Jeffrey T. Terpstra, Joseph W. McKean, and Joshua D. Naranjo. Theory & methods: Weighted Wilcoxon estimates for autoregression. Australian & New Zealand Journal of Statistics, 43(4):399–419, December 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bofinger:2001:TMS

[1538] Eve Bofinger and Wei Liu. Theory & methods: Selecting closest to control. Australian & New Zealand Journal of Statistics, 43(4):421–430, December 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Dobbie:2001:TMM

[1539] Melissa J. Dobbie and A. H. Welsh. Theory & methods: Modelling correlated zero-inflated count data. Australian & New Zealand Journal of Statistics, 43(4):431–444, December 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chen:2001:TMC

[1540] Gemai Chen and Min Chen. Theory & methods: On a class of nonlinear Ar(P) models with nonlinear arch errors. Australian & New Zealand Journal of Statistics, 43(4):445–454, December 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Jensen:2001:TMR

[1541] D. R. Jensen. Theory & methods: Recovering information from incremental responses. Australian & New Zealand Journal of Statistics, 43 (4):455–459, December 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

SenGupta:2001:TMS

[1542] Ashis SenGupta and Arnab Kumar Laha. Theory & methods: The slippage problem for the circular normal distribution. Australian & New Zealand Journal of Statistics, 43(4):461–471, December 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Marriott:2001:TMN

[1543] John M. Marriott and Nancy M. Spencer. Theory & methods: a note on Bayesian prediction from the regression model with informative priors. Australian & New Zealand Journal of Statistics, 43(4):473–480, December 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chen:2001:TMT

[1544] Lin-An Chen, Kuo-Yuan Liang, and Chwen-Chi Liu. Theory & methods: Two-stage Welsh’s trimmed mean for the simultaneous equations model. Australian & New Zealand Journal of Statistics, 43(4):481–492, December 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gupta:2001:LE

[1545] Rameshwar D. Gupta and Debasis Kundu. Letter to the Editor. Australian & New Zealand Journal of Statistics, 43(4):493, December 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2001:BRd

[1546] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 43(4):495–499, December 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2001:FPd

[1547] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 43(4):500, December 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2001:IIV

[1548] Anonymous. Index: Index to volume 43, 2001. Australian & New Zealand Journal of Statistics, 43(4):501–503, December 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2001:IKI

[1549] Anonymous. Index: Keyword index for volumes 42 and 43. Australian & New Zealand Journal of Statistics, 43(4):506–510, December 2001. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lloyd:2002:EA

[1550] Chris J. Lloyd. Editorial announcements. Australian & New Zealand Journal of Statistics, 44(1):1–2, March 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2002:ARS

[1551] Anonymous. Acknowledgment of referees’ services. Australian & New Zealand Journal of Statistics, 44(1):2, March 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Tee:2002:OLA

[1552] Garry J. Tee. Oration lecture: Applications: Nineteenth and early twentieth century statistics: Some New Zealand connections. Australian & New Zealand Journal of Statistics, 44(1):3–12, March 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Welsh:2002:TMI

[1553] A. H. Welsh. Theory & methods: Incomplete detection in enumeration surveys: Whither distance sampling? Australian & New Zealand Journal of Statistics, 44(1):13–22, March 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lee:2002:TMT

[1554] Thomas C. M. Lee. Theory & methods: Tree-based wavelet regression for correlated data using the minimum description length principle. Australian & New Zealand Journal of Statistics, 44(1):23–39, March 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hwang:2002:TMC

[1555] Wen-Han Hwang, Anne Chao, and Paul S. F. Yip. Theory & methods: Continuous-time capture-recapture models with time variation and behavioural response. Australian & New Zealand Journal of Statistics, 44 (1):41–54, March 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Huggins:2002:TMP

[1556] Richard M. Huggins. Theory & methods: a parametric empirical Bayes approach to the analysis of capture-recapture experiments. Australian & New Zealand Journal of Statistics, 44(1):55–62, March 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Salehi:2002:TMU

[1557] Mohammad M. Salehi and George A. F. Seber. Theory & methods: Unbiased estimators for restricted adaptive cluster sampling. Australian & New Zealand Journal of Statistics, 44(1):63–74, March 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lloyd:2002:TMS

[1558] Chris J. Lloyd. Theory & methods: Semi-parametric estimation of ROC curves based on binomial regression modelling. Australian & New Zealand Journal of Statistics, 44(1):75–86, March 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Qiu:2002:TMF

[1559] Peihua Qiu. Theory & methods: Fitting a semi-parametric model based on two sources of information. Australian & New Zealand Journal of Statistics, 44(1):87–97, March 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bian:2002:TMB

[1560] Guorui Bian. Theory & methods: Bayesian estimates in a one-way ANOVA random effects model. Australian & New Zealand Journal of Statistics, 44(1):99–108, March 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Liu:2002:TMA

[1561] Wei Liu, Nan Wang, and Suojin Wang. Theory & methods: Approximations of distributions for some standardized partial sums in sequential analysis. Australian & New Zealand Journal of Statistics, 44(1):109–119, March 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2002:BRa

[1562] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 44(1):121–127, March 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2002:FPa

[1563] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 44(1):128, March 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Shao:2002:AMT

[1564] Quanxi Shao and N. A. Campbell. Applications: Modelling trends in groundwater levels by segmented regression with constraints. Australian & New Zealand Journal of Statistics, 44(2):129–141, June 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Brown:2002:TMI

[1565] B. M. Brown. Theory & methods: On the importance of being smooth. Australian & New Zealand Journal of Statistics, 44(2):143–154, June 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gerlach:2002:TMB

[1566] Richard Gerlach, Ron Bird, and Anthony Hall. Theory & methods: Bayesian variable selection in logistic regression: predicting company earnings direction. Australian & New Zealand Journal of Statistics, 44 (2):155–168, June 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bremner:2002:TMM

[1567] Alexandra P. Bremner and Ross H. Taplin. Theory & methods: Modified classification and regression tree splitting criteria for data with interactions. Australian & New Zealand Journal of Statistics, 44(2):169–176, June 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Yau:2002:TMZ

[1568] Kelvin K. W. Yau, Andy H. Lee, and Angus S. K. Ng. Theory & methods: a zero-augmented gamma mixed model for longitudinal data with many zeros. Australian & New Zealand Journal of Statistics, 44(2):177– 183, June 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Tang:2002:TMD

[1569] Nian-Sheng Tang, Bo-Cheng Wei, and Xue-Ren Wang. Theory & methods: Detection of influential observations in nonlinear reproductive dispersion models. Australian & New Zealand Journal of Statistics, 44(2): 185–192, June 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lu:2002:TMS

[1570] Jye-Chyi Lu, Di Chen, and Nianci Gan. Theory & methods: Semiparametric modelling and likelihood estimation with estimating equations. Australian & New Zealand Journal of Statistics, 44(2):193–212, June 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Willink:2002:TME

[1571] R. Willink. Theory & methods: Effect on probabilities and quantiles of adding a quantity with small variance. Australian & New Zealand Journal of Statistics, 44(2):213–220, June 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Al-Saleh:2002:TME

[1572] Mohammad Fraiwan Al-Saleh and Gang Zheng. Theory & methods: Estimation of bivariate characteristics using ranked set sampling. Australian & New Zealand Journal of Statistics, 44(2):221–232, June 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Koutrouvelis:2002:TME

[1573] Ioannis A. Koutrouvelis and Simos Meintanis. Theory & methods: Estimating the parameters of Poisson-exponential models. Australian & New Zealand Journal of Statistics, 44(2):233–245, June 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2002:FPb

[1574] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 44(2):246, June 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2002:BRb

[1575] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 44(2):247–255, June 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Westbrooke:2002:AIM

[1576] Ian Westbrooke and Lisa Jones. Applications: Imputation of Maori descent for electoral calculations in New Zealand. Australian & New Zealand Journal of Statistics, 44(3):257–265, September 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Westbrooke:2002:AME

[1577] Ian Westbrooke and Michael Ryan. Applications: The mathematics of electoral district allocation in New Zealand. Australian & New Zealand Journal of Statistics, 44(3):267–276, September 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Claeskens:2002:TMD

[1578] Gerda Claeskens and Peter Hall. Theory & methods: Data sharpening for hazard rate estimation. Australian & New Zealand Journal of Statistics, 44(3):277–283, September 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Neuhaus:2002:TMB

[1579] John M. Neuhaus. Theory & methods: Bias due to ignoring the sample design in case-control studies. Australian & New Zealand Journal of Statistics, 44(3):285–293, September 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Taplin:2002:TMP

[1580] Ross H. Taplin. Theory & methods: Pseudo–modelling of outliers for estimation of regression and time series models. Australian & New Zealand Journal of Statistics, 44(3):295–310, September 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Jung:2002:TMD

[1581] Kang-Mo Jung. Theory & methods: a diagnostic method for principal components tests. Australian & New Zealand Journal of Statistics, 44 (3):311–318, September 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Knight:2002:TME

[1582] John L. Knight, Stephen E. Satchell, and Jun Yu. Theory & methods: Estimation of the stochastic volatility model by the empirical characteristic function method. Australian & New Zealand Journal of Statistics, 44(3):319–335, September 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kumagai:2002:TME

[1583] Etsuo Kumagai, Nobuo Inagaki, and Kiyoshi Inoue. Theory & methods: On Efron’s parameterization. Australian & New Zealand Journal of Statistics, 44(3):337–343, September 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

DeSouza:2002:TMD

[1584] Francisco A. M. De Souza and Gilberto A. Paula. Theory & methods: Deviance residuals for an angular response. Australian & New Zealand Journal of Statistics, 44(3):345–356, September 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ghosh:2002:TMC

[1585] Subir Ghosh and Colleen Burns. Theory & methods: Comparison of four new general classes of search designs. Australian & New Zealand Journal of Statistics, 44(3):357–366, September 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chi:2002:TMO

[1586] Yunchan Chi. Theory & methods: Ordered tests for right-censored survival data. Australian & New Zealand Journal of Statistics, 44(3):367– 380, September 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2002:BRc

[1587] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 44(3):381–384, September 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2002:FPc

[1588] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 44(3):384, September 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Seneta:2002:MEP

[1589] Eugene Seneta. In memoriam: Emeritus Professor Henry Oliver Lancaster, AO FAA, 1 February 1913–2 December 2001. Australian & New Zealand Journal of Statistics, 44(4):385–400, December 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Li:2002:ABH

[1590] Ning Li, Guoqi Qian, and Richard Huggins. Analysis of betweenhousehold heterogeneity in disease transmission from data on outbreak sizes. Australian & New Zealand Journal of Statistics, 44(4):401–411, December 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Clarke:2002:ANI

[1591] Paul S. Clarke and P. F. Tate. An application of non-ignorable nonresponse models for Gross flows estimation in the British labour force survey. Australian & New Zealand Journal of Statistics, 44(4):413–425, December 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Brown:2002:KWM

[1592] B. M. Brown and T. P. Hettmansperger. Kruskal–Wallis, multiple comparisons and Efron dice. Australian & New Zealand Journal of Statistics, 44(4):427–438, December 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Zitikis:2002:ADG

[1593] Ričardas Zitikis and Joseph L. Gastwirth. The asymptotic distribution of the S–Gini index. Australian & New Zealand Journal of Statistics, 44(4):439–446, December 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Littlejohn:2002:PAC

[1594] Roger P. Littlejohn and James R. Webster. Periodogram analysis of censored time series from a designed experiment. Australian & New Zealand Journal of Statistics, 44(4):447–455, December 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

John:2002:D

[1595] J. A. John, K. Ruggiero, and E. R. Williams. α_{n}-designs. Australian & New Zealand Journal of Statistics, 44(4):457–465, December 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ng:2002:DTF

[1596] Meei Pyng Ng and Ray Watson. Dealing with ties in failure time data. Australian & New Zealand Journal of Statistics, 44(4):467–478, December 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Munk:2002:MRE

[1597] Axel Munk and Frits Ruymgaart. Minimax rates for estimating the variance and its derivatives in non-parametric regression. Australian & New Zealand Journal of Statistics, 44(4):479–488, December 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Vu:2002:ESM

[1598] Hien T. V. Vu and Matthew W. Knuiman. Estimation in semiparametric marginal shared gamma frailty models. Australian & New Zealand Journal of Statistics, 44(4):489–501, December 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Baddeley:2002:ABR

[1599] Adrian Baddeley and Rolf Turner. Addendum to A. Baddeley & R. Turner (2000). Australian & New Zealand Journal of Statistics, 44(4): 503, December 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1477].

Anonymous:2002:FPd

[1600] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 44(4):504, December 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2002:BRd

[1601] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 44(4):505–506, December 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2002:IV

[1602] Anonymous. Indexes, volume 44. Australian & New Zealand Journal of Statistics, 44(4):507–509, December 2002. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2003:PMA

[1603] Anonymous. Pitman Medal for 2002 awarded to Terry Speed. Australian & New Zealand Journal of Statistics, 45(1):1–4, March 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lloyd:2003:IEA

[1604] Chris Lloyd. Inaugural Editor’s award. Australian & New Zealand Journal of Statistics, 45(1):5, March 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Tallis:2003:ANS

[1605] G. M. Tallis, P. Leppard, and T. J. O’Neill. Applications: a note on the sensitivities of self-reporting and screen detection of primary breast tumours. Australian & New Zealand Journal of Statistics, 45(1):7–18, March 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Martin:2003:ADA

[1606] R. J. Martin, L. M. Platts, A. B. Seddon, and E. C. Stillman. Applications: The design and analysis of a mixture experiment on glass durability. Australian & New Zealand Journal of Statistics, 45(1):19–27, March 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Sain:2003:TMN

[1607] Stephan R. Sain. Theory & methods: a new characterization and estimation of the zero-bias bandwidth. Australian & New Zealand Journal of Statistics, 45(1):29–42, March 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

OMalley:2003:TMB

[1608] A. James O’Malley and John J. Deely. Theory & methods: Bayesian measures of the minimum detectable concentration of an immunoassay. Australian & New Zealand Journal of Statistics, 45(1):43–65, March 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Nyangoma:2003:TMB

[1609] Stephen O. Nyangoma. Theory & methods: Bias in multinomial models. Australian & New Zealand Journal of Statistics, 45(1):67–82, March 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Danielak:2003:TME

[1610] Katarzyna Danielak and Tomasz Rychlik. Theory & methods: Exact bounds for the bias of trimmed means. Australian & New Zealand Journal of Statistics, 45(1):83–96, March 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chan:2003:TMC

[1611] Barbara S. P. Chan and J. A. Eccleston. Theory & methods: On the construction of nearest–neighbour balanced row–column designs. Australian & New Zealand Journal of Statistics, 45(1):97–106, March 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Khan:2003:TMO

[1612] M. G. M. Khan, E. A. Khan, and M. J. Ahsan. Theory & methods: an optimal multivariate stratified sampling design using dynamic programming. Australian & New Zealand Journal of Statistics, 45(1):107–113, March 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Menendez:2003:TMT

[1613] M. L. Menéndez, J. A. Pardo, and L. Pardo. Theory & methods: Tests for bivariate symmetry against ordered alternatives in square contingency tables. Australian & New Zealand Journal of Statistics, 45(1): 115–123, March 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2003:BRa

[1614] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 45(1):125–126, March 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2003:ARS

[1615] Anonymous. Acknowledgment of referees’ services. Australian & New Zealand Journal of Statistics, 45(1):127, March 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2003:FPa

[1616] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 45(1):128, March 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Griffiths:2003:ASC

[1617] David Griffiths, Chandra Gulati, and Jim Ollis. Applications: Statistical control for road pavements. Australian & New Zealand Journal of Statistics, 45(2):129–140, June 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Taplin:2003:ATS

[1618] Ross H. Taplin. Applications: Teaching statistical consulting before statistical methodology. Australian & New Zealand Journal of Statistics, 45 (2):141–152, June 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Larocque:2003:TMA

[1619] Denis Larocque, Serge Tardif, and Constance van Eeden. Theory & methods: an affine-invariant generalization of the Wilcoxon signed-rank test for the bivariate location problem. Australian & New Zealand Journal of Statistics, 45(2):153–165, June 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lloyd:2003:TMO

[1620] Chris J. Lloyd and Paul Kabaila. Theory & methods: On the optimality and limitations of Buehler bounds. Australian & New Zealand Journal of Statistics, 45(2):167–174, June 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Blest:2003:TMN

[1621] David C. Blest. Theory & methods: a new measure of kurtosis adjusted for skewness. Australian & New Zealand Journal of Statistics, 45(2): 175–179, June 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Smith:2003:TMU

[1622] Narelle F. Smith and Deborah J. Street. Theory & methods: The use of balanced incomplete block designs in designing randomized response surveys. Australian & New Zealand Journal of Statistics, 45(2):181– 194, June 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Oktaba:2003:TMG

[1623] Wiktor Oktaba. Theory & methods: The general multivariate Gauss– Markov model of the incomplete block design. Australian & New Zealand Journal of Statistics, 45(2):195–205, June 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Yip:2003:TMS

[1624] Paul S. F. Yip, Xiangzhong Fang, Yong Zhou, and Yan Wang. Theory & methods: Sequential procedure for fixed accuracy estimation of the population size in recapture sampling. Australian & New Zealand Journal of Statistics, 45(2):207–216, June 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Huang:2003:TMP

[1625] Fuchun Huang. Theory & methods: Prediction error property of the lasso estimator and its generalization. Australian & New Zealand Journal of Statistics, 45(2):217–228, June 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See corrected proof [1695].

Aoki:2003:TMM

[1626] Satoshi Aoki and Akimichi Takemura. Theory & methods: Minimal basis for a connected Markov chain over 3 × 3 × K contingency tables with fixed two-dimensional marginals. Australian & New Zealand Journal of Statistics, 45(2):229–249, June 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2003:FPb

[1627] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 45(2):250, June 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2003:LE

[1628] Anonymous. Letter to the Editor. Australian & New Zealand Journal of Statistics, 45(2):251–252, June 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2003:BRb

[1629] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 45(2):253–256, June 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lin:2003:AMD

[1630] Tsung I. Lin and Jack C. Lee. Applications: On modelling data from degradation sample paths over time. Australian & New Zealand Journal of Statistics, 45(3):257–270, September 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hazelton:2003:TMV

[1631] Martin L. Hazelton. Theory & methods: Variable kernel density estimation. Australian & New Zealand Journal of Statistics, 45(3):271–284, September 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cohen:2003:TME

[1632] Arthur Cohen, David Madigan, and Harold B. Sackrowitz. Theory & methods: Effective directed tests for models with ordered categorical data. Australian & New Zealand Journal of Statistics, 45(3):285–300, September 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anderson:2003:TMG

[1633] Marti J. Anderson and John Robinson. Theory & methods: Generalized discriminant analysis based on distances. Australian & New Zealand Journal of Statistics, 45(3):301–318, September 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Berger:2003:TMT

[1634] Yves G. Berger, Mohammed E. H. Tirari, and Yves Tillé. Theory & methods: Towards optimal regression estimation in sample surveys. Australian & New Zealand Journal of Statistics, 45(3):319–329, September 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Baek:2003:TMC

[1635] Jong-Il Baek, Tae-Sung Kim, and Han-Ying Liang. Theory & methods: On the convergence of moving average processes under dependent conditions. Australian & New Zealand Journal of Statistics, 45(3):331– 342, September 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Shiau:2003:TMM

[1636] Jyh-Jen Horng Shiau and Lin-An Chen. Theory & methods: a multivariate parallelogram and its application to multivariate trimmed means.

Australian & New Zealand Journal of Statistics, 45(3):343–352, September 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

DeOliveira:2003:TMN

[1637] Victor De Oliveira. Theory & methods: a note on the correlation structure of transformed Gaussian random fields. Australian & New Zealand Journal of Statistics, 45(3):353–366, September 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bandyopadhyay:2003:TMN

[1638] Uttam Bandyopadhyay and Atanu Biswas. Theory & methods: Nonparametric group sequential designs in randomized clinical trials. Australian & New Zealand Journal of Statistics, 45(3):367–376, September 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Whitaker:2003:LE

[1639] David Whitaker. Letter to the Editor. Australian & New Zealand Journal of Statistics, 45(3):377–378, September 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2003:BRc

[1640] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 45(3):379–381, September 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2003:FPc

[1641] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 45(3):382, September 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Sparks:2003:AMI

[1642] Ross Sparks. Applications: Monitoring for increases in process variance. Australian & New Zealand Journal of Statistics, 45(4):383–394, December 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Stephens:2003:TMA

[1643] Matthew Stephens and Peter Donnelly. Theory and methods: Ancestral inference in population genetics models with selection (with discussion). Australian & New Zealand Journal of Statistics, 45(4):395–430, December 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Henderson:2003:TMM

[1644] Brent Henderson and Richard Jarrett. Theory and methods: Models with errors due to misreported measurements. Australian & New Zealand Journal of Statistics, 45(4):431–444, December 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Thompson:2003:TMS

[1645] Robin Thompson, Brian Cullis, Alison Smith, and Arthur Gilmour. Theory and methods: a sparse implementation of the average information algorithm for factor analytic and reduced rank variance models. Australian & New Zealand Journal of Statistics, 45(4):445–459, December 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Laib:2003:TMN

[1646] Naâmane Laïb. Theory and methods: Non-parametric testing of conditional variance functions in time series. Australian & New Zealand Journal of Statistics, 45(4):461–475, December 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Li:2003:TML

[1647] Chin-Shang Li. Theory and methods: a lack-of-fit test for heteroscedastic regression models via cosine-series smoothers. Australian & New Zealand Journal of Statistics, 45(4):477–489, December 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hayakawa:2003:TMB

[1648] Yu Hayakawa, Philip D. O’Neill, Darren Upton, and Paul S. F. Yip. Theory and methods: Bayesian inference for a stochastic epidemic model with uncertain numbers of susceptibles of several types. Australian & New Zealand Journal of Statistics, 45(4):491–502, December 2003. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Khan:2003:LE

[1649] M. G. M. Khan. Letter to the Editor. Australian & New Zealand Journal of Statistics, 45(4):503–504, December 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2003:FPd

[1650] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 45(4):505–506, December 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2003:IV

[1651] Anonymous. Indexes for volume 45. Australian & New Zealand Journal of Statistics, 45(4):507–510, December 2003. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lloyd:2004:IEA

[1652] Chris Lloyd. Inaugural Editor’s award. Australian & New Zealand Journal of Statistics, 46(1):1–2, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pollett:2004:F

[1653] Phil Pollett and Peter Taylor. Foreword. Australian & New Zealand Journal of Statistics, 46(1):3, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gani:2004:OPS

[1654] Joe Gani. On the occasion of the presentation of the SSAI Festschrift to Daryl John Daley. Australian & New Zealand Journal of Statistics, 46 (1):5–11, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gani:2004:VBD

[1655] J. Gani. A versatile birth–death model applicable to four distinct problems. Australian & New Zealand Journal of Statistics, 46(1):13–21, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Whittle:2004:DSN

[1656] Peter Whittle. The distribution of species numbers in a controlled environment. Australian & New Zealand Journal of Statistics, 46(1):23–27, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pakes:2004:CDR

[1657] Anthony G. Pakes. Convergence and divergence of random series. Australian & New Zealand Journal of Statistics, 46(1):29–40, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bohm:2004:KES

[1658] Stephan Böhm, Lothar Heinrich, and Volker Schmidt. Kernel estimation of the spectral density of stationary random closed sets. Australian & New Zealand Journal of Statistics, 46(1):41–51, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Heyde:2004:ACC

[1659] C. C. Heyde. Asymptotics and criticality for a correlated Bernoulli process. Australian & New Zealand Journal of Statistics, 46(1):53–57, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Laslett:2004:EAC

[1660] Geoff M. Laslett, J. Paige Eveson, and Tom Polacheck. Estimating the age at capture in capture-recapture studies of fish growth. Australian & New Zealand Journal of Statistics, 46(1):59–66, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Stoyan:2004:SAM

[1661] Dietrich Stoyan, Helga Stoyan, and Gunter Döge. Statistical analyses and modelling of the mixing process of Euro coins in Germany and Europe. Australian & New Zealand Journal of Statistics, 46(1):67–77, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kingman:2004:PPR

[1662] J. F. C. Kingman. Powers and products of regenerative phenomena. Australian & New Zealand Journal of Statistics, 46(1):79–86, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cornell:2004:UEP

[1663] Stephen J. Cornell and Valerie S. Isham. Ultimate extinction of the promiscuous bisexual Galton–Watson metapopulation. Australian & New Zealand Journal of Statistics, 46(1):87–98, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hall:2004:PRB

[1664] Peter Hall and Nader Tajvidi. Prediction regions for bivariate extreme events. Australian & New Zealand Journal of Statistics, 46(1):99–112, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pollett:2004:EUP

[1665] Phil Pollett and Hanjun Zhang. Existence and uniqueness of Q-processes with a given finite μ-invariant measure. Australian & New Zealand Journal of Statistics, 46(1):113–120, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Miyoshi:2004:RTC

[1666] Naoto Miyoshi and Tomasz Rolski. Ross-type conjectures on monotonicity of queues. Australian & New Zealand Journal of Statistics, 46(1): 121–131, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Vere-Jones:2004:RMP

[1667] David Vere-Jones and Frederic Paik Schoenberg. Rescaling marked point processes. Australian & New Zealand Journal of Statistics, 46(1):133– 143, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Baltrunas:2004:SDL

[1668] Aleksandras Baltrūnas and Claudia Klüppelberg. Subexponential distributions — large deviations with applications to insurance and queueing models. Australian & New Zealand Journal of Statistics, 46(1):145–154, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Vesilo:2004:LRD

[1669] R. A. Vesilo. Long-range dependence of Markov renewal processes. Australian & New Zealand Journal of Statistics, 46(1):155–171, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Leadbetter:2004:RRF

[1670] M. R. Leadbetter and G. V. Spaniolo. Reflections on Rice’s formulae for level crossings — history, extensions and use. Australian & New Zealand Journal of Statistics, 46(1):173–180, March 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ball:2004:ASP

[1671] Frank G. Ball and Robin K. Milne. Applications of simple point process methods to superpositions of aggregated stationary processes. Australian & New Zealand Journal of Statistics, 46(1):181–196, March 2004. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Brockhoff:2004:AAE

[1672] P. B. Brockhoff, D. J. Best, and J. C. W. Rayner. Applications: an application of extended analysis for ranked data with ties. Australian & New Zealand Journal of Statistics, 46(2):197–204, June 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Jorgensen:2004:AUM

[1673] Murray Jorgensen. Applications: Using multinomial mixture models to cluster Internet traffic. Australian & New Zealand Journal of Statistics, 46(2):205–218, June 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Insley:2004:TMI

[1674] Robin Insley, Lucia Mok, and Tim Swartz. Theory & methods: Issues related to sports gambling. Australian & New Zealand Journal of Statistics, 46(2):219–232, June 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cressie:2004:TME

[1675] Noel Cressie, Sylvia Richardson, and Isabelle Jaussent. Theory & methods: Ecological bias: Use of maximum-entropy approximations. Australian & New Zealand Journal of Statistics, 46(2):233–255, June 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cavanaugh:2004:TMC

[1676] Joseph E. Cavanaugh. Theory & methods: Criteria for linear model selection based on Kullback’s symmetric divergence. Australian & New Zealand Journal of Statistics, 46(2):257–274, June 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anh:2004:TMC

[1677] V. V. Anh, V. P. Knopova, and N. N. Leonenko. Theory & methods: Continuous-time stochastic processes with cyclical long-range dependence. Australian & New Zealand Journal of Statistics, 46(2):275– 296, June 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wesolowski:2004:TMS

[1678] Jacek Wesołowski and Mohammad Ahsanullah. Theory & methods: Switching order statistics through random power contractions. Australian & New Zealand Journal of Statistics, 46(2):297–303, June 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Peng:2004:TME

[1679] Liang Peng and Yongcheng Qi. Theory & methods: Estimating the firstand second-order parameters of a heavy-tailed distribution. Australian & New Zealand Journal of Statistics, 46(2):305–312, June 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Iksanov:2004:TMN

[1680] Aleksander M. Iksanov and Che Soong Kim. Theory & methods: New explicit examples of fixed points of Poisson shot noise transforms. Australian & New Zealand Journal of Statistics, 46(2):313–321, June 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2004:FPa

[1681] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 46(2):322–323, June 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Welham:2004:APL

[1682] Sue Welham, Brian Cullis, Beverley Gogel, Arthur Gilmour, and Robin Thompson. Applications: Prediction in linear mixed models. Australian & New Zealand Journal of Statistics, 46(3):325–347, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

ONeill:2004:AFC

[1683] Michael E. O’Neill, Peter C. Thomson, Brent C. Jacobs, Phil Brain, Ruth C. Butler, Heather Turner, and Bernadetha Mitakda. Applications: Fitting and comparing seed germination models with a focus on the inverse normal distribution. Australian & New Zealand Journal of Statistics, 46(3):349–366, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bremner:2004:TMPa

[1684] Alexandra P. Bremner and Ross H. Taplin. Theory & methods: Performance of localized regression tree splitting criteria on data with discontinuities. Australian & New Zealand Journal of Statistics, 46(3):367–381, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kennedy:2004:TMT

[1685] Nyongesa L. Kennedy. Theory & methods: Testing for the presence of disease by pooling samples. Australian & New Zealand Journal of Statistics, 46(3):383–390, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hepworth:2004:TMM

[1686] Graham Hepworth. Theory & methods: Mid-P confidence intervals based on the likelihood ratio for proportions estimated by group testing.

Australian & New Zealand Journal of Statistics, 46(3):391–405, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Zhang:2004:TMA

[1687] Biao Zhang. Theory & methods: Assessing goodness-of-fit of categorical regression models based on case-control data. Australian & New Zealand Journal of Statistics, 46(3):407–423, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Huh:2004:TMD

[1688] J. Huh and B. U. Park. Theory & methods: Detection of a change point with local polynomial fits for the random design case. Australian & New Zealand Journal of Statistics, 46(3):425–441, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ozturk:2004:TMM

[1689] Omer Ozturk, Douglas A. Wolfe, and Roxana Alexandridis. Theory & methods: Multi-sample inference for simple-tree alternatives with ranked-set samples. Australian & New Zealand Journal of Statistics, 46(3):443–455, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Huang:2004:TME

[1690] J. S. Huang and G. D. Lin. Theory & methods: Estimating the population median using only a central subset of observations. Australian & New Zealand Journal of Statistics, 46(3):457–462, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kabaila:2004:TMB

[1691] Paul Kabaila and Chris J. Lloyd. Theory & methods: Buehler confidence limits and nesting. Australian & New Zealand Journal of Statistics, 46 (3):463–469, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ahmad:2004:TMG

[1692] Ibrahim A. Ahmad and Ibrahim A. Alwasel. Theory & methods: Goodness-of-fit testing of a weak memoryless property of life distributions. Australian & New Zealand Journal of Statistics, 46(3):471–481, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

M:2004:TMG

[1693] Mohammad Salehi M. and George A. F. Seber. Theory & methods: a general inverse sampling scheme and its application to adaptive cluster sampling. Australian & New Zealand Journal of Statistics, 46(3):483– 494, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Willink:2004:TMA

[1694] R. Willink. Theory & methods: Approximating the difference of two tvariables for all degrees of freedom using truncated variables. Australian & New Zealand Journal of Statistics, 46(3):495–504, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Rosset:2004:TMC

[1695] Saharon Rosset and Ji Zhu. Theory & methods: Corrected proof of the result of ‘A prediction error property of the Lasso estimator and its generalization’ by Huang (2003). Australian & New Zealand Journal of Statistics, 46(3):505–510, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1625].

Martin:2004:BRa

[1696] Michael A. Martin. Book review. Australian & New Zealand Journal of Statistics, 46(3):511–512, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Martin:2004:BRb

[1697] Gael M. Martin. Book review. Australian & New Zealand Journal of Statistics, 46(3):512–514, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2004:FPb

[1698] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 46(3):515, September 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Mayo:2004:KLW

[1699] Oliver Mayo. Knibbs Lecture for 2002: To what extent has Fisher’s research program been fulfilled in Australia? Australian & New Zealand Journal of Statistics, 46(4):517–529, December 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lloyd:2004:AEF

[1700] Chris J. Lloyd and Donald Frommer. Applications: Estimating the false negative fraction for a multiple screening test for bowel cancer when negatives are not verified. Australian & New Zealand Journal of Statistics, 46(4):531–542, December 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Millar:2004:ASM

[1701] Russell B. Millar. Applications: Simulated maximum likelihood applied to non-Gaussian and nonlinear mixed effects and state–space models. Australian & New Zealand Journal of Statistics, 46(4):543–554, December 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Groenewald:2004:TMB

[1702] P. C. N. Groenewald and A. C. Schoeman. Theory & methods: Bayesian detection and analysis of changing transition matrices of stationary Markov chains. Australian & New Zealand Journal of Statistics, 46(4): 555–567, December 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Salganik:2004:TMC

[1703] M. P. Salganik, M. P. Wand, and N. Lange. Theory & methods: Comparison of feature significance quantile approximations. Australian & New Zealand Journal of Statistics, 46(4):569–581, December 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bremner:2004:TMPb

[1704] Alexandra P. Bremner and Ross H. Taplin. Theory & methods: Performance of a localized tree splitting criterion in tree averaging. Australian & New Zealand Journal of Statistics, 46(4):583–599, December 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Huggins:2004:TMR

[1705] Richard M. Huggins and Danuta Z. Loesch. Theory & methods: Robust non-parametric estimation of genotype–phenotype relationships in human pedigrees. Australian & New Zealand Journal of Statistics, 46 (4):601–613, December 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Legendre:2004:TMA

[1706] Pierre Legendre and François-Joseph Lapointe. Theory & methods: Assessing congruence among distance matrices: Single-malt Scotch whiskies revisited. Australian & New Zealand Journal of Statistics, 46(4):615–629, December 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ng:2004:TMW

[1707] H. K. T. Ng and N. Balakrishnan. Theory & methods: Wilcoxon-type rank-sum precedence tests. Australian & New Zealand Journal of Statistics, 46(4):631–648, December 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Clarke:2004:TMS

[1708] Brenton R. Clarke and Christopher J. Milne. Theory & methods: Small sample bias correction for Huber’s proposal-2 scale M-estimator. Australian & New Zealand Journal of Statistics, 46(4):649–656, December 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Baba:2004:TMP

[1709] Kunihiro Baba, Ritei Shibata, and Masaaki Sibuya. Theory & methods: Partial correlation and conditional correlation as measures of conditional independence. Australian & New Zealand Journal of Statistics, 46(4): 657–664, December 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Danielak:2004:TMS

[1710] Katarzyna Danielak and Mohammad Z. Raqab. Theory & methods: Sharp bounds for expectations of k th record increments. Australian & New Zealand Journal of Statistics, 46(4):665–673, December 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Biswas:2004:TMR

[1711] Atanu Biswas and Anup Dewanji. Theory & methods: a randomized longitudinal play-the-winner design for repeated binary data. Australian & New Zealand Journal of Statistics, 46(4):675–684, December 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Naskar:2004:TMI

[1712] Malay Naskar and Kalyan Das. Theory & methods: Inference in Dirichlet process mixed generalized linear models by using Monte Carlo EM. Australian & New Zealand Journal of Statistics, 46(4):685–701, December 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2004:FPc

[1713] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 46(4):702, December 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2004:I

[1714] Anonymous. Indexes. Australian & New Zealand Journal of Statistics, 46(4):703–708, December 2004. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2005:PMA

[1715] Anonymous. Pitman Medal for 2004 awarded to Adrian Baddeley. Australian & New Zealand Journal of Statistics, 47(1):1–2, March 2005. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cripps:2005:AST

[1716] Edward Cripps, David Nott, William T. M. Dunsmuir, and Christopher Wikle. Applications: Space–time modelling of Sydney Harbour winds. Australian & New Zealand Journal of Statistics, 47(1):3–17, March 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Thomas:2005:TMU

[1717] Len Thomas, Stephen T. Buckland, Ken B. Newman, and John Harwood. Theory & methods: a unified framework for modelling wildlife population dynamics. Australian & New Zealand Journal of Statistics, 47(1):19– 34, March 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Besbeas:2005:TMP

[1718] P. Besbeas, S. N. Freeman, and B. J. T. Morgan. Theory & methods: The potential of integrated population modelling. Australian & New Zealand Journal of Statistics, 47(1):35–48, March 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lebreton:2005:TMD

[1719] Jean-Dominique Lebreton. Theory & methods: Dynamical and statistical models for exploited populations. Australian & New Zealand Journal of Statistics, 47(1):49–63, March 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Mackenzie:2005:TMW

[1720] Darryl I. Mackenzie. Theory & methods: Was it there? Dealing with imperfect detection for species presence/absence data. Australian & New Zealand Journal of Statistics, 47(1):65–74, March 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Caswell:2005:TMS

[1721] Hal Caswell. Theory & methods: Sensitivity analysis of the stochastic growth rate: Three extensions. Australian & New Zealand Journal of Statistics, 47(1):75–85, March 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hyndman:2005:TML

[1722] Rob J. Hyndman, Maxwell L. King, Ivet Pitrun, and Baki Billah. Theory & methods: Local linear forecasts using cubic smoothing splines. Australian & New Zealand Journal of Statistics, 47(1):87–99, March 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Flynn:2005:TMP

[1723] David B. Flynn, Simone D. Grose, Gael M. Martin, and Vance L. Martin. Theory & methods: Pricing Australian S&P200 options: a Bayesian approach based on generalized distributional forms. Australian & New Zealand Journal of Statistics, 47(1):101–117, March 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wright:2005:TMC

[1724] C. L. Wright, J. A. John, D. Whitaker, and E. R. Williams. Theory & methods: Construction of resolvable row-column designs based on contractions. Australian & New Zealand Journal of Statistics, 47(1): 119–124, March 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2005:BRa

[1725] Anonymous. Book reviews. Australian & New Zealand Journal of Statistics, 47(1):125–127, March 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2005:FPa

[1726] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 47(1):127–128, March 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ridall:2005:ABH

[1727] P. G. Ridall and A. N. Pettitt. Applications: Bayesian hidden Markov models for longitudinal counts. Australian & New Zealand Journal of Statistics, 47(2):129–145, June 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Yin:2005:TMN

[1728] Xiangrong Yin. Theory & methods: Non-parametric estimation of direction in single-index models with categorical predictors. Australian & New Zealand Journal of Statistics, 47(2):147–161, June 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Stamey:2005:TMM

[1729] James D. Stamey and Dean M. Young. Theory & methods: Maximum likelihood estimation for a Poisson rate parameter with misclassified counts. Australian & New Zealand Journal of Statistics, 47(2): 163–172, June 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chen:2005:TMAa

[1730] Di Chen, Jye-Chyi Lu, and Shu Chuan Lin. Theory & methods: Asymptotic distributions of semiparametric maximum likelihood estimators with estimating equations for group-censored data. Australian & New Zealand Journal of Statistics, 47(2):173–192, June 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ganguli:2005:TMA

[1731] Bhaswati Ganguli, John Staudenmayer, and M. P. Wand. Theory & methods: Additive models with predictors subject to measurement error. Australian & New Zealand Journal of Statistics, 47(2):193–202, June 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Baklizi:2005:TMC

[1732] Ayman Baklizi. Theory & methods: a continuously adaptive rank test for shift in location. Australian & New Zealand Journal of Statistics, 47(2): 203–209, June 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Rayner:2005:TMM

[1733] J. C. W. Rayner and O. Thas. Theory & methods: More informative testing for bivariate symmetry. Australian & New Zealand Journal of Statistics, 47(2):211–217, June 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Headrick:2005:TMO

[1734] Todd C. Headrick and Bruno D. Zumbo. Theory & methods: On optimizing multi-level designs: Power under budget constraints. Australian & New Zealand Journal of Statistics, 47(2):219–229, June 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Arnab:2005:TMN

[1735] Raghunath Arnab and Sarjinder Singh. Theory & methods: a note on variance estimation for the generalized regression predictor. Australian & New Zealand Journal of Statistics, 47(2):231–234, June 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Froda:2005:TME

[1736] Sorana Froda and Gino Colavita. Theory & methods: Estimating predator-prey systems via ordinary differential equations with closed orbits. Australian & New Zealand Journal of Statistics, 47(2):235–254, June 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2005:BRB

[1737] Anonymous. Book review. Australian & New Zealand Journal of Statistics, 47(2):255, June 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2005:FPb

[1738] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 47(2):256, June 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Jasso:2005:MNL

[1739] G. Jasso and S. Kotz. In memoriam: Norman Lloyd Johnson: 9 January 1917–18 November 2004. Australian & New Zealand Journal of Statistics, 47(3):257–258, September 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Russell:2005:TML

[1740] K. G. Russell and David Griffiths. Theory & methods: a lotto systems problem. Australian & New Zealand Journal of Statistics, 47(3):259– 267, September 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Poskitt:2005:TME

[1741] D. S. Poskitt and Jing Zhang. Theory & methods: Estimating components in finite mixtures and hidden Markov models. Australian & New Zealand Journal of Statistics, 47(3):269–286, September 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wang:2005:TMS

[1742] Yan Wang. Theory & methods: a semiparametric regression model with missing covariates in continuous-time capture-recapture studies. Australian & New Zealand Journal of Statistics, 47(3):287–297, September 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Yip:2005:TMB

[1743] Paul S. F. Yip, Liqun Xi, Richard Arnold, and Yu Hayakawa. Theory & methods: a beta-binomial model for estimating the size of a heterogeneous population. Australian & New Zealand Journal of Statistics, 47 (3):299–308, September 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Nur:2005:TMP

[1744] Darfiana Nur, Kerrie L. Mengersen, and Rodney C. Wolff. Theory & methods: Phase randomization: a convergence diagnostic test for MCMC. Australian & New Zealand Journal of Statistics, 47(3):309– 323, September 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Walker:2005:TMN

[1745] Stephen Walker. Theory & methods: a note on Bayes factors. Australian & New Zealand Journal of Statistics, 47(3):325–328, September 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lim:2005:TMT

[1746] Jae-Hak Lim, Dae-Kyung Kim, and Dong Ho Park. Theory & methods: Tests for detecting more nbu-ness at a specific age. Australian & New Zealand Journal of Statistics, 47(3):329–337, September 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Saei:2005:TMM

[1747] Ayoub Saei and Ray Chambers. Theory & methods: Modelling truncated and clustered count data. Australian & New Zealand Journal of Statistics, 47(3):339–349, September 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Calle:2005:TMS

[1748] M. Luz Calle and Guadalupe Gómez. Theory & methods: a semiparametric hierarchical method for a regression model with an interval-censored covariate. Australian & New Zealand Journal of Statistics, 47(3):351– 364, September 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Berger:2005:TMV

[1749] Yves G. Berger. Theory & methods: Variance estimation with highly stratified sampling designs with unequal probabilities. Australian & New Zealand Journal of Statistics, 47(3):365–373, September 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Farrell:2005:TMM

[1750] Patrick J. Farrell and Sarjinder Singh. Theory & methods: Modelassisted higher-order calibration of estimators of variance. Australian & New Zealand Journal of Statistics, 47(3):375–383, September 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2005:FPc

[1751] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 47(3):384, September 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Brewer:2005:HGI

[1752] Ken Brewer. Historical & general interest: Anomalies, probings, insights: Ken Foreman’s role in the sampling inference controversy of the Late 20th Century. Australian & New Zealand Journal of Statistics, 47(4): 385–399, December 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Graham:2005:AHB

[1753] Patrick Graham. Applications: Hierarchical Bayesian modelling of social variation in the age dependence of disability prevalence. Australian & New Zealand Journal of Statistics, 47(4):401–424, December 2005. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Heathcote:2005:TME

[1754] Christopher R. Heathcote and Borek D. Puza. Theory & methods: Estimating processes of counts from cross-sectional aggregate data, with an application to multistate life tables and health expectancies. Australian & New Zealand Journal of Statistics, 47(4):425–439, December 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wood:2005:TME

[1755] G. R. Wood, C. D. Lai, and C. G. Qiao. Theory & methods: Estimation of a proportion using several independent samples of binomial mixtures. Australian & New Zealand Journal of Statistics, 47(4):441–448, December 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Taplin:2005:TMR

[1756] Ross H. Taplin. Theory & methods: a robust Bayes factor for linear models. Australian & New Zealand Journal of Statistics, 47(4):449–462, December 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Muller:2005:TMT

[1757] Samuel Müller. Theory & methods: Two-stage support estimation. Australian & New Zealand Journal of Statistics, 47(4):463–472, December 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chen:2005:TMAb

[1758] Cathy W. S. Chen, Mike K. P. So, and Richard H. Gerlach. Theory & methods: Assessing and testing for threshold nonlinearity in stock returns. Australian & New Zealand Journal of Statistics, 47(4):473–488, December 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Mokkadem:2005:TML

[1759] Abdelkader Mokkadem, Mariane Pelletier, and Julien Worms. Theory & methods: Large and moderate deviations principles for kernel estimation of a multivariate density and its partial derivatives. Australian & New Zealand Journal of Statistics, 47(4):489–502, December 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Vonta:2005:TME

[1760] Filia Vonta. Theory & methods: Efficient estimation in regression frailty or transformation models based on an algorithm. Australian & New Zealand Journal of Statistics, 47(4):503–514, December 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

PintodaCosta:2005:TMW

[1761] Joaquim Pinto da Costa and Carlos Soares. Theory & methods: a weighted rank measure of correlation. Australian & New Zealand Journal of Statistics, 47(4):515–529, December 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See letters [1821, 1822].

Fry:2005:TMD

[1762] Tim R. L. Fry and Mark N. Harris. Theory & methods: The Dogit ordered generalized extreme value model. Australian & New Zealand Journal of Statistics, 47(4):531–542, December 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bairamov:2005:TMC

[1763] I. Bairamov, M. Ahsanullah, and Anthony G. Pakes. Theory & methods: a characterization of continuous distributions via regression on pairs of record values. Australian & New Zealand Journal of Statistics, 47(4): 543–547, December 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kabaila:2005:TMC

[1764] Paul Kabaila. Theory & methods: On the coverage probability of confidence intervals in regression after variable selection. Australian & New Zealand Journal of Statistics, 47(4):549–562, December 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Byrne:2005:TMN

[1765] John Byrne. Theory & methods: a new short exact geometric confidence interval. Australian & New Zealand Journal of Statistics, 47(4):563– 569, December 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2005:FPd

[1766] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 47(4):570, December 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2005:I

[1767] Anonymous. Index. Australian & New Zealand Journal of Statistics, 47 (4):571–576, December 2005. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hu:2006:HZD

[1768] Chin-Yuan Hu, Aleksander M. Iksanov, Gwo Dong Lin, and Oleg K. Zakusylo. The Hurwitz zeta distribution. Australian & New Zealand Journal of Statistics, 48(1):1–6, March 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cuevas:2006:PEG

[1769] Antonio Cuevas, Wenceslao González-Manteiga, and Alberto Rodríguez-Casal. Plug-in estimation of general level sets. Australian & New Zealand Journal of Statistics, 48(1):7–19, March 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Liang:2006:WSN

[1770] Han-Ying Liang and Jong-Il Baek. Weighted sums of negatively associated random variables. Australian & New Zealand Journal of Statistics, 48(1):21–31, March 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lund:2006:PPT

[1771] Robert Lund, Qin Shao, and Ishwar Basawa. Parsimonious periodic time series modeling. Australian & New Zealand Journal of Statistics, 48(1): 33–47, March 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Woods:2006:ABD

[1772] David Woods and Susan Lewis. All-bias designs for polynomial spline regression models. Australian & New Zealand Journal of Statistics, 48(1): 49–58, March 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Peng:2006:NCM

[1773] Liang Peng and Yongcheng Qi. A new calibration method of constructing empirical likelihood-based confidence intervals for the tail index. Australian & New Zealand Journal of Statistics, 48(1):59–66, March 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hankin:2006:NFN

[1774] Robin K. S. Hankin and Alan Lee. A new family of non-negative distributions. Australian & New Zealand Journal of Statistics, 48(1):67–78, March 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Zhang:2006:MLE

[1775] Haimeng Zhang and M. Bhaskara Rao. Maximum likelihood estimation in linear models with equi-correlated random errors. Australian & New Zealand Journal of Statistics, 48(1):79–93, March 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bhandari:2006:RIP

[1776] Subir K. Bhandari, Ayanendranath Basu, and Sahadeb Sarkar. Robust inference in parametric models using the family of generalized negative exponential disparities. Australian & New Zealand Journal of Statistics, 48(1):95–114, March 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2006:FPa

[1777] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 48(1):115, March 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chao:2006:ALB

[1778] Anne Chao, Tsung-Jen Shen, and Wen-Han Hwang. Application of Laplace’s boundary-mode approximations to estimate species and shared species richness. Australian & New Zealand Journal of Statistics, 48(2): 117–128, June 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Neumeyer:2006:BTE

[1779] Natalie Neumeyer, Holger Dette, and Eva-Renate Nagel. Bootstrap tests for the error distribution in linear and nonparametric regression models. Australian & New Zealand Journal of Statistics, 48(2):129–156, June 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lee:2006:ERT

[1780] Shen-Ming Lee, Cathy W. S. Chen, Richard H. Gerlach, and Li-Hui Huang. Estimation in Ricker’s two-release method: a Bayesian approach. Australian & New Zealand Journal of Statistics, 48(2):157–169, June 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Corr:2006:CPB

[1781] Patrick Corr and Rebecca Kippen. The case for parity and birth-order statistics. Australian & New Zealand Journal of Statistics, 48(2):171– 200, June 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Stamey:2006:BHM

[1782] James D. Stamey, Dean M. Young, and Doyle Boese. A Bayesian hierarchical model for Poisson rate and reporting-probability inference using double sampling. Australian & New Zealand Journal of Statistics, 48(2): 201–212, June 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Berzel:2006:ENV

[1783] Andreas Berzel, Gillian Z. Heller, and Walter Zucchini. Estimating the number of visits to the doctor. Australian & New Zealand Journal of Statistics, 48(2):213–224, June 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wang:2006:BAA

[1784] Jenting Wang and John S. J. Hsu. Bayesian analysis of the additive mixed model for randomized block designs. Australian & New Zealand Journal of Statistics, 48(2):225–236, June 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cripps:2006:BSS

[1785] Edward Cripps, Robert Kohn, and David Nott. Bayesian subset selection and model averaging using a centred and dispersed prior for the error variance. Australian & New Zealand Journal of Statistics, 48(2):237– 252, June 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Takagi:2006:SOA

[1786] Yoshiji Takagi. Second order asymptotic comparison of estimators under universal domination criterion. Australian & New Zealand Journal of Statistics, 48(2):253–267, June 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2006:FPb

[1787] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 48(2):269, June 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Zadlo:2006:PTV

[1788] Tomasz Ża̧dło. On prediction of total value in incompletely specified domains. Australian & New Zealand Journal of Statistics, 48(3):269– 283, September 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Prendergast:2006:DIO

[1789] Luke A. Prendergast. Detecting influential observations in sliced inverse regression analysis. Australian & New Zealand Journal of Statistics, 48 (3):285–304, September 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Mukherjee:2006:NSD

[1790] Bhramar Mukherjee. A note on sampling designs for random processes with no quadratic mean derivative. Australian & New Zealand Journal of Statistics, 48(3):305–319, September 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lu:2006:ERB

[1791] Zeng-Hua Lu and Bruce M. Brown. An exclusive regressors binary mixture model with an application to labour supply. Australian & New Zealand Journal of Statistics, 48(3):321–333, September 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Griffiths:2006:GSS

[1792] William E. Griffiths and Ma. Rebecca Valenzuela. Gibbs samplers for a set of seemingly unrelated regressions. Australian & New Zealand Journal of Statistics, 48(3):335–351, September 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Clarke:2006:ATL

[1793] Brenton R. Clarke and Daniel D. Schubert. An adaptive trimmed likelihood algorithm for identification of multivariate outliers. Australian & New Zealand Journal of Statistics, 48(3):353–371, September 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Domijan:2006:SMM

[1794] Katarina Domijan, Murray Jorgensen, and Jeff Reid. Semi-mechanistic modelling in nonlinear regression: a case study. Australian & New Zealand Journal of Statistics, 48(3):373–392, September 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See discussion [1795].

Haslett:2006:IDP

[1795] Stephen Haslett, R. B. Millar, K. B. Newman, Martyn Plummer, and A. H. Welsh. Invited discussion of the paper by Domijan, Jorgensen & Reid. Australian & New Zealand Journal of Statistics, 48(3):393–413, September 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1794].

Anonymous:2006:FPc

[1796] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 48(3):415, September 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Byth:2006:AAR

[1797] K. Byth, D. R. Cox, and P. Forder. Applications: Assessing the relationship between symptoms of allergic rhinoconjunctivitis and pollen counts. Australian & New Zealand Journal of Statistics, 48(4):417–428, December 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Saha:2006:TMG

[1798] Amitava Saha. Theory & methods: a generalized two-stage randomized response procedure in complex sample surveys. Australian & New Zealand Journal of Statistics, 48(4):429–443, December 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wood:2006:TMC

[1799] Simon N. Wood. Theory & methods: On confidence intervals for generalized additive models based on penalized regression splines. Australian & New Zealand Journal of Statistics, 48(4):445–464, December 2006. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Taylor:2006:TMA

[1800] Julian D. Taylor and Arūnas P. Verbyla. Theory & methods: Asymptotic likelihood approximations using a partial Laplace approximation. Australian & New Zealand Journal of Statistics, 48(4):465–476, December 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Elmore:2006:TMF

[1801] R. T. Elmore, T. P. Hettmansperger, and F. Xuan. Theory & methods: a fully nonparametric test for one-way layouts. Australian & New Zealand Journal of Statistics, 48(4):477–490, December 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Choi:2006:TML

[1802] H. Choi, Y. K. Lee, B. U. Park, and K. S. Yu. Theory & methods: Local polynomial quasi-likelihood regression on random fields. Australian & New Zealand Journal of Statistics, 48(4):491–506, December 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2006:FPd

[1803] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 48(4):507, December 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2006:IV

[1804] Anonymous. Indexes for volume 48. Australian & New Zealand Journal of Statistics, 48(4):i–iii, December 2006. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Welham:2007:CMM

[1805] Sue J. Welham, Brian R. Cullis, Michael G. Kenward, and Robin Thompson. A comparison of mixed model splines for curve fitting. Australian & New Zealand Journal of Statistics, 49(1):1–23, March 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gurrin:2007:CIB

[1806] Lyle C. Gurrin and Paul R. Burton. A comparison of the imprecise beta class, the randomized play-the-winner rule and the triangular test for clinical trials with binary responses. Australian & New Zealand Journal of Statistics, 49(1):25–35, March 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gonzalez:2007:MEC

[1807] Miguel González, Manuel Mota, and Alfonso Ramos. Moment estimation in the class of bisexual branching processes with population-size dependent mating. Australian & New Zealand Journal of Statistics, 49(1): 37–50, March 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Raqab:2007:BBW

[1808] Mohammad Z. Raqab. Bounds on the bias of Winsorized means. Australian & New Zealand Journal of Statistics, 49(1):51–60, March 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Williams:2007:CCD

[1809] E. R. Williams and J. A. John. Construction of crossover designs with correlated errors. Australian & New Zealand Journal of Statistics, 49(1): 61–68, March 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Zhang:2007:LME

[1810] Jin Zhang. Likelihood moment estimation for the generalized Pareto distribution. Australian & New Zealand Journal of Statistics, 49(1):69– 77, March 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pewsey:2007:WFC

[1811] Arthur Pewsey, Toby Lewis, and M. C. Jones. The wrapped t family of circular distributions. Australian & New Zealand Journal of Statistics, 49(1):79–91, March 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Dehay:2007:LED

[1812] Dominique Dehay and Jian-Feng Yao. On likelihood estimation for discretely observed Markov jump processes. Australian & New Zealand Journal of Statistics, 49(1):93–107, March 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

RuizEspejo:2007:BRA

[1813] Mariano Ruiz Espejo. Book reviews: Applied Optimal Designs. Australian & New Zealand Journal of Statistics, 49(1):109–110, March 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Dobbie:2007:BR

[1814] Dr Melissa Dobbie. Book reviews. Australian & New Zealand Journal of Statistics, 49(1):111–113, March 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2007:FPa

[1815] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 49(1):114, March 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pakes:2007:DCT

[1816] Anthony G. Pakes and Jorge Navarro. Distributional characterizations through scaling relations. Australian & New Zealand Journal of Statistics, 49(2):115–135, June 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Grillenzoni:2007:PRR

[1817] Carlo Grillenzoni. Pattern recognition via robust smoothing with application to laser data. Australian & New Zealand Journal of Statistics, 49 (2):137–153, June 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Leng:2007:AFT

[1818] Chenlei Leng and Shuangge Ma. Accelerated failure time models with nonlinear covariates effects. Australian & New Zealand Journal of Statistics, 49(2):155–172, June 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wager:2007:MSP

[1819] Carrie Wager, Florin Vaida, and Göran Kauermann. Model selection for penalized spline smoothing using Akaike information criteria. Australian & New Zealand Journal of Statistics, 49(2):173–190, June 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hwang:2007:ASR

[1820] Wen-Han Hwang and Richard Huggins. Application of semiparametric regression models in the analysis of capture-recapture experiments. Australian & New Zealand Journal of Statistics, 49(2):191–202, June 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Genest:2007:LER

[1821] Christian Genest and Jean-François Plante. Letter to the Editor: Re: Pinto da Costa, J. & Soares, C. (2005). A weighted rank measure of correlation. Australian & New Zealand Journal of Statistics, 49(2):203– 204, June 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1761, 1822].

PintodaCosta:2007:LER

[1822] Joaquim Pinto da Costa and Carlos Soares. Letter to the Editor: Rejoinder to letter to the Editor from C. Genest and J-F. Plante concerning ‘Pinto da Costa, J. & Soares, C. (2005). A weighted rank measure of correlation.’. Australian & New Zealand Journal of Statistics, 49(2): 205–207, June 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1761, 1821].

Anonymous:2007:FPb

[1823] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 49(2):208, June 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lijoi:2007:CRN

[1824] Antonio Lijoi, Igor Prünster, and Stephen G. Walker. On convergence rates for nonparametric posterior distributions. Australian & New Zealand Journal of Statistics, 49(3):209–219, September 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Queen:2007:EDA

[1825] Catriona M. Queen, Ben J. Wright, and Casper J. Albers. Eliciting a directed acyclic graph for a multivariate time series of vehicle counts in a traffic network. Australian & New Zealand Journal of Statistics, 49 (3):221–239, September 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pardoe:2007:GDV

[1826] Iain Pardoe and R. Dennis Cook. A graphical diagnostic for variance functions. Australian & New Zealand Journal of Statistics, 49(3):241– 250, September 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bebbington:2007:BTC

[1827] Mark Bebbington, Chin-Diew Lai, and Ričardas Zitikis. Bathtub-type curves in reliability and beyond. Australian & New Zealand Journal of Statistics, 49(3):251–265, September 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ferraty:2007:NRF

[1828] Frédéric Ferraty, André Mas, and Philippe Vieu. Nonparametric regression on functional data: Inference and practical aspects. Australian & New Zealand Journal of Statistics, 49(3):267–286, September 2007. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Branscum:2007:BBR

[1829] Adam J. Branscum, Wesley O. Johnson, and Mark C. Thurmond. Bayesian beta regression: Applications to household expenditure data and genetic distance between foot-and-mouth disease viruses. Australian & New Zealand Journal of Statistics, 49(3):287–301, September 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Liu:2007:NME

[1830] Danping Liu, Paul S. F. Yip, and Richard M. Huggins. Nonparametric method for estimating the size of an open population using marginal data from repeated multiple lists. Australian & New Zealand Journal of Statistics, 49(3):303–320, September 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kim:2007:SEE

[1831] Gunky Kim, Mervyn J. Silvapulle, and Paramsothy Silvapulle. Semiparametric estimation of the error distribution in multivariate regression using copulas. Australian & New Zealand Journal of Statistics, 49(3): 321–336, September 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2007:FPc

[1832] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 49(3):337, September 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Baklizi:2007:TSU

[1833] Ayman Baklizi. Testing symmetry using a trimmed longest run statistic. Australian & New Zealand Journal of Statistics, 49(4):339–347, December 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Quang:2007:SLL

[1834] Nguyen Van Quang and Le Van Thanh. On the strong law of large numbers under rearrangements for sequences of blockwise orthogonal random elements in Banach spaces. Australian & New Zealand Journal of Statistics, 49(4):349–357, December 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Claeskens:2007:PFM

[1835] Gerda Claeskens, Christophe Croux, and Johan Van Kerckhoven. Prediction-focused model selection for autoregressive models. Australian & New Zealand Journal of Statistics, 49(4):359–379, December 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Dey:2007:PBN

[1836] Partha Sarathi Dey and Ayanendranath Basu. On the power breakdown of the negative exponential disparity tests. Australian & New Zealand Journal of Statistics, 49(4):381–384, December 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ghazali:2007:TWS

[1837] Syed Shakir Ali Ghazali, Chris J. Skinner, and M. H. Tahir. Threeway stratification sampling design when one of the stratifying variables is time. Australian & New Zealand Journal of Statistics, 49(4):385–395, December 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wheldon:2007:ITE

[1838] M. C. Wheldon, M. J. Anderson, and B. W. Johnson. Identifying treatment effects in multi-channel measurements in electroencephalographic studies: Multivariate permutation tests and multiple comparisons. Australian & New Zealand Journal of Statistics, 49(4):397–413, December 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ausin:2007:BER

[1839] M. Concepcion Ausin and Hedibert F. Lopes. Bayesian estimation of ruin probabilities with a heterogeneous and heavy-tailed insurance claim-size distribution. Australian & New Zealand Journal of Statistics, 49(4):415– 434, December 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ahmed:2007:SPA

[1840] S. Ejaz Ahmed, Kjell A. Doksum, S. Hossain, and Jinhong You. Shrinkage, pretest and absolute penalty estimators in partially linear models. Australian & New Zealand Journal of Statistics, 49(4):435–454, December 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Robinson:2007:BRG

[1841] Andrew Robinson. Book review: Generalized Additive Models: an Introduction in R, by Simon N. Wood. Australian & New Zealand Journal of Statistics, 49(4):455–456, December 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2007:FPd

[1842] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 49(4):457, December 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2007:I

[1843] Anonymous. Indexes. Australian & New Zealand Journal of Statistics, 49(4):459–461, December 2007. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Mcelroy:2008:FTE

[1844] Tucker Mcelroy and Dimitris N. Politis. A fine-tuned estimator of a general convergence rate. Australian & New Zealand Journal of Statistics, 50 (1):1–11, March 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

McGree:2008:PBO

[1845] J. M. McGree and J. A. Eccleston. Probability-based optimal design. Australian & New Zealand Journal of Statistics, 50(1):13–28, March 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chen:2008:HTD

[1846] Cathy W. S. Chen, F. C. Liu, and Mike K. P. So. Heavy-tailed-distributed threshold stochastic volatility models in financial time series. Australian & New Zealand Journal of Statistics, 50(1):29–51, March 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Crane:2008:CEF

[1847] Glenis J. Crane and John van der Hoek. Conditional expectation formulae for copulas. Australian & New Zealand Journal of Statistics, 50(1): 53–67, March 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Giri:2008:CPC

[1848] Khageswor Giri and Paul Kabaila. The coverage probability of confidence intervals in 2^{r} factorial experiments after preliminary hypothesis testing. Australian & New Zealand Journal of Statistics, 50(1):69–79, March 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Vos:2008:PBT

[1849] Paul W. Vos and Suzanne Hudson. Problems with binomial two-sided tests and the associated confidence intervals. Australian & New Zealand Journal of Statistics, 50(1):81–89, March 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2008:PM

[1850] Anonymous. Pitman Medal for 2005 and 2006. Australian & New Zealand Journal of Statistics, 50(1):91–94, March 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2008:FPa

[1851] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 50(1):95, March 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Aneiros-Perez:2008:SPA

[1852] Germán Aneiros-Pérez. Semi-parametric analysis of covariance under dependence conditions within each group. Australian & New Zealand Journal of Statistics, 50(2):97–123, June 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Mao:2008:LBP

[1853] Chang Xuan Mao. Lower bounds to the population size when capture probabilities vary over individuals. Australian & New Zealand Journal of Statistics, 50(2):125–134, June 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Choy:2008:SMD

[1854] S. T. Boris Choy and Jennifer S. K. Chan. Scale mixtures distributions in statistical modelling. Australian & New Zealand Journal of Statistics, 50(2):135–146, June 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gani:2008:ACB

[1855] J. Gani and R. J. Swift. Approximating carrier-borne epidemics. Australian & New Zealand Journal of Statistics, 50(2):147–159, June 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

OMalley:2008:RML

[1856] A. James O’Malley, Murray H. Smith, and William A. Sadler. A restricted maximum likelihood procedure for estimating the variance function of an immunoassay. Australian & New Zealand Journal of Statistics, 50(2):161–177, June 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wand:2008:SRO

[1857] M. P. Wand and J. T. Ormerod. On semiparametric regression with O’Sullivan penalized splines. Australian & New Zealand Journal of Statistics, 50(2):179–198, June 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See corrigendum [1923].

Willink:2008:CFE

[1858] R. Willink. A closed-form expression for the Pearson Type IV distribution function. Australian & New Zealand Journal of Statistics, 50(2): 199–205, June 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2008:FPb

[1859] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 50(2):206, June 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bairamov:2008:JBP

[1860] Ismihan Bairamov and Serkan Eryilmaz. Joint behaviour of precedences and exceedances in random threshold models. Australian & New Zealand Journal of Statistics, 50(3):209–216, September 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Stepanova:2008:AEB

[1861] Natalia Stepanova and Shu Wang. Asymptotic efficiency of the Blest-type tests for independence. Australian & New Zealand Journal of Statistics, 50(3):217–233, September 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Rayner:2008:GER

[1862] J. C. W. Rayner, O. Thas, and B. De Boeck. A generalized Emerson recurrence relation. Australian & New Zealand Journal of Statistics, 50 (3):235–240, September 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wang:2008:CPD

[1863] Lihong Wang. Change-point detection with rank statistics in longmemory time-series models. Australian & New Zealand Journal of Statistics, 50(3):241–256, September 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Berthelsen:2008:NPB

[1864] Kasper K. Berthelsen and Jesper Møller. Non-parametric Bayesian inference for inhomogeneous Markov point processes. Australian & New Zealand Journal of Statistics, 50(3):257–272, September 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wang:2008:URT

[1865] Gaowen Wang and Wei-Lin Mao. Unit root testing in the presence of heavy-tailed GARCH errors. Australian & New Zealand Journal of Statistics, 50(3):273–292, September 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2008:FPc

[1866] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 50(3):293, September 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Butler:2008:AOC

[1867] David G. Butler, John A. Eccleston, and Brian R. Cullis. On an approximate optimality criterion for the design of field experiments under spatial dependence. Australian & New Zealand Journal of Statistics, 50 (4):295–307, December 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kabaila:2008:CIU

[1868] Paul Kabaila and Jarrod Tuck. Confidence intervals utilizing prior information in the Behrens–Fisher problem. Australian & New Zealand Journal of Statistics, 50(4):309–328, December 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lloyd:2008:EVD

[1869] Chris J. Lloyd. Exact P -values for discrete models obtained by estimation and maximization. Australian & New Zealand Journal of Statistics, 50 (4):329–345, December 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Liang:2008:ELB

[1870] Hua Liang and Yongsong Qin. Empirical likelihood-based inferences for partially linear models with missing covariates. Australian & New Zealand Journal of Statistics, 50(4):347–359, December 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gay:2008:OSI

[1871] R. Gay and D. S. Poskitt. Optimal semiparametric inference for the tail index based on ratios of the largest extremes. Australian & New Zealand Journal of Statistics, 50(4):361–380, December 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Li:2008:PBM

[1872] Deyuan Li. On the probability of being maximal. Australian & New Zealand Journal of Statistics, 50(4):381–394, December 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Singh:2008:RAE

[1873] Housila P. Singh and Sunil Kumar. A regression approach to the estimation of the finite population mean in the presence of non-response. Australian & New Zealand Journal of Statistics, 50(4):395–408, December 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2008:FPd

[1874] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 50(4):409, December 2008. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2009:PM

[1875] Anonymous. Pitman Medal for 2008. Australian & New Zealand Journal of Statistics, 51(1):1–7, March 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wand:2009:IPS

[1876] M. P. Wand. Invited paper: Semiparametric regression and graphical models. Australian & New Zealand Journal of Statistics, 51(1):9–41, March 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Foster:2009:CPE

[1877] Scott D. Foster, Arūnas P. Verbyla, and Wayne S. Pitchford. Contributed papers: Estimation, prediction and inference for the lasso random effects model. Australian & New Zealand Journal of Statistics, 51(1):43–61, March 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Salmond:2009:CPW

[1878] Rob Salmond. Contributed papers: Weekday-only polling and partisan support levels: Evidence from New Zealand. Australian & New Zealand Journal of Statistics, 51(1):63–76, March 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Stephenson:2009:CPH

[1879] Alec G. Stephenson. Contributed papers: High-dimensional parametric modelling of multivariate extreme events. Australian & New Zealand Journal of Statistics, 51(1):77–88, March 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bieniek:2009:CPN

[1880] Mariusz Bieniek. Contributed papers: a note on characterizations of distributions by regressions of non-adjacent generalized order statistics. Australian & New Zealand Journal of Statistics, 51(1):89–99, March 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wang:2009:CPM

[1881] Yan Wang and Joanne Thandrayen. Contributed papers: Multiple-record systems estimation using latent class models. Australian & New Zealand Journal of Statistics, 51(1):101–111, March 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2009:FPa

[1882] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 51(1):113–114, March 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Young:2009:IPR

[1883] G. Alastair Young. Invited paper: Routes to higher-order accuracy in parametric inference. Australian & New Zealand Journal of Statistics, 51(2):115–126, June 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hidiroglou:2009:CPV

[1884] M. A. Hidiroglou, J. N. K. Rao, and David Haziza. Contributed papers: Variance estimation in two-phase sampling. Australian & New Zealand Journal of Statistics, 51(2):127–141, June 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hamza:2009:CPE

[1885] Kais Hamza. Contributed papers: an exact test for hazard similarity. Australian & New Zealand Journal of Statistics, 51(2):143–159, June 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gao:2009:CPR

[1886] Jiti Gao, Degui Li, and Zhengyan Lin. Contributed papers: Robust estimation in parametric time series models under long- and short-range-dependent structures. Australian & New Zealand Journal of Statistics, 51(2):161–181, June 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Muller:2009:CPP

[1887] Samuel Müller and Céline Vial. Contributed papers: Partially linear model selection by the bootstrap. Australian & New Zealand Journal of Statistics, 51(2):183–200, June 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Penzer:2009:CPA

[1888] Jeremy Penzer, Mingjin Wang, and Qiwei Yao. Contributed papers: Approximating volatilities by asymmetric power GARCH functions. Australian & New Zealand Journal of Statistics, 51(2):201–225, June 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

McGrory:2009:CPV

[1889] C. A. McGrory and D. M. Titterington. Contributed papers: Variational Bayesian analysis for hidden Markov models. Australian & New Zealand Journal of Statistics, 51(2):227–244, June 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2009:FPb

[1890] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 51(2):245–246, June 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Dubnicka:2009:KDE

[1891] Suzanne R. Dubnicka. Kernel density estimation with missing data and auxiliary variables. Australian & New Zealand Journal of Statistics, 51 (3):247–270, September 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kabaila:2009:UBM

[1892] Paul Kabaila and Khageswor Giri. Upper bounds on the minimum coverage probability of confidence intervals in regression after model selection. Australian & New Zealand Journal of Statistics, 51(3):271–287, September 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Duc:2009:ABA

[1893] Hiep Duc, Bin Jalaludin, and Geoff Morgan. Associations between air pollution and hospital visits for cardiovascular diseases in the elderly in Sydney using Bayesian statistical methods. Australian & New Zealand Journal of Statistics, 51(3):289–303, September 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Xu:2009:NCB

[1894] Xingzhong Xu, Xiaobo Ding, and Shuran Zhao. A new confidence band for continuous cumulative distribution functions. Australian & New Zealand Journal of Statistics, 51(3):305–318, September 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Nguyen:2009:CAR

[1895] Danh V. Nguyen and Damla Şentürk. Covariate-adjusted regression for longitudinal data incorporating correlation between repeated measurements. Australian & New Zealand Journal of Statistics, 51(3):319–333, September 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Beh:2009:ENI

[1896] Eric J. Beh and Thomas B. Farver. An evaluation of non-iterative methods for estimating the linear-by-linear parameter of ordinal log-linear models. Australian & New Zealand Journal of Statistics, 51(3):335–352, September 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See corrigendum [1924].

VanDenHout:2009:ANC

[1897] Ardo Van Den Hout and Irene Klugkist. Accounting for non-compliance in the analysis of randomized response data. Australian & New Zealand Journal of Statistics, 51(3):353–372, September 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2009:FPc

[1898] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 51(3):373, September 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pakes:2009:NON

[1899] Anthony G. Pakes. Numbers of observations near order statistics. Australian & New Zealand Journal of Statistics, 51(4):375–395, December 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Weatherall:2009:GSA

[1900] M. Weatherall, R. M. Pickering, and Scott Harris. Graphical sensitivity analysis with different methods of imputation for a trial with probable non-ignorable missing data. Australian & New Zealand Journal of Statistics, 51(4):397–413, December 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Akram:2009:ESN

[1901] Muhammad Akram, Rob J. Hyndman, and J. Keith Ord. Exponential smoothing and non-negative data. Australian & New Zealand Journal of Statistics, 51(4):415–432, December 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Smyth:2009:LAD

[1902] Gordon K. Smyth and Arūnas P. Verbyla. Leverage adjustments for dispersion modelling in generalized nonlinear models. Australian & New Zealand Journal of Statistics, 51(4):433–448, December 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gordon:2009:EPW

[1903] Ian R. Gordon and Peter Hall. Estimating a parameter when it is known that the parameter exceeds a given value. Australian & New Zealand Journal of Statistics, 51(4):449–460, December 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wouters:2009:DDS

[1904] Heidi Wouters, Olivier Thas, and Jean-Pierre Ottoy. Data-driven smooth tests and a diagnostic tool for lack-of-fit for circular data. Australian & New Zealand Journal of Statistics, 51(4):461–480, December 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Heathcote:2009:UAD

[1905] Christopher R. Heathcote, Borek D. Puza, and Steven P. Roberts. The use of aggregate data to estimate Gompertz-type old-age mortality in heterogeneous populations. Australian & New Zealand Journal of Statistics, 51(4):481–497, December 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Haslett:2009:CDG

[1906] S. J. Haslett and K. Govindaraju. Cloning data: Generating datasets with exactly the same multiple linear regression fit. Australian & New Zealand Journal of Statistics, 51(4):499–503, December 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2009:FPd

[1907] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 51(4):505, December 2009. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lopez-Diaz:2010:SOR

[1908] Miguel López-Díaz. A stochastic ordering for random variables with applications. Australian & New Zealand Journal of Statistics, 52(1):1– 16, March 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Yip:2010:ENP

[1909] Paul S. F. Yip, Ray Watson, K. S. Chan, Eric H. Y. Lau, Feng Chen, Ying Xu, Liqun Xi, Derek Y. T. Cheung, Brian Y. T. Ip, and Danping Liu. Estimation of the number of people in a demonstration. Australian & New Zealand Journal of Statistics, 52(1):17–26, March 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Beh:2010:NSC

[1910] Eric J. Beh and Luigi D’Ambra. Non-symmetrical correspondence analysis with concatenation and linear constraints. Australian & New Zealand Journal of Statistics, 52(1):27–44, March 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Martin:2010:DDJ

[1911] Michael A. Martin, Steven Roberts, and Letian Zheng. Delete-2 and delete-3 jackknife procedures for unmasking in regression. Australian & New Zealand Journal of Statistics, 52(1):45–60, March 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chen:2010:SJT

[1912] Pu Chen and Chih-Ying Hsiao. Subsampling the Johansen test with stable innovations. Australian & New Zealand Journal of Statistics, 52 (1):61–73, March 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bandyopadhyay:2010:NEC

[1913] Dipankar Bandyopadhyay and Amalia Jácome Pumar. Nonparametric estimation of conditional cumulative hazards for missing population marks. Australian & New Zealand Journal of Statistics, 52(1):75–91, March 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Tao:2010:SCI

[1914] Jian Tao, Man-Lai Tang, and Ning-Zhong Shi. A stepwise confidence interval procedure based on an asymmetric loss function with applications to toxicological evaluation. Australian & New Zealand Journal of Statistics, 52(1):93–107, March 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Shen:2010:CPM

[1915] Yan Shen, Min Xie, and Loon Ching Tang. On the change point of the mean residual life of series and parallel systems. Australian & New Zealand Journal of Statistics, 52(1):109–121, March 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2010:FPa

[1916] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 52(1):123, March 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Welham:2010:APC

[1917] Sue J. Welham, Beverley J. Gogel, Alison B. Smith, Robin Thompson, and Brian R. Cullis. Applications paper: a comparison of analysis methods for late-stage variety evaluation trials. Australian & New Zealand Journal of Statistics, 52(2):125–149, June 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Paroli:2010:TMP

[1918] R. Paroli and L. Spezia. Theory and methods papers: Reversible jump Markov chain Monte Carlo methods and segmentation algorithms in hidden Markov models. Australian & New Zealand Journal of Statistics, 52 (2):151–166, June 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Tzavidis:2010:TMP

[1919] Nikos Tzavidis, Stefano Marchetti, and Ray Chambers. Theory and methods papers: Robust estimation of small-area means and quantiles. Australian & New Zealand Journal of Statistics, 52(2):167–186, June 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Mattheou:2010:TMP

[1920] K. Mattheou and A. Karagrigoriou. Theory and methods papers: a new family of divergence measures for tests of fit. Australian & New Zealand Journal of Statistics, 52(2):187–200, June 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kulasekera:2010:TMP

[1921] K. B. Kulasekera and Wei Lin. Theory and methods papers: Error variance estimation for the single-index model. Australian & New Zealand Journal of Statistics, 52(2):201–219, June 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Maheu:2010:TMP

[1922] John M. Maheu, Jonathan J. Reeves, and Xuan Xie. Theory and methods papers: Forecasting volatility in the presence of model instability. Australian & New Zealand Journal of Statistics, 52(2):221–237, June 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wand:2010:CSR

[1923] M. P. Wand and J. T. Ormerod. Corrigendum: On semiparametric regression with O’Sullivan penalised splines. Australian & New Zealand Journal of Statistics, 52(2):239, June 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1857].

Beh:2010:CEN

[1924] Eric J. Beh and Thomas B. Farver. Corrigendum: An evaluation of noniterative methods for estimating the linear-by-linear parameter of ordinal log-linear models. Australian & New Zealand Journal of Statistics, 52(2): 241–242, June 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1896].

Anonymous:2010:FPb

[1925] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 52(2):243, June 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Sparks:2010:APT

[1926] Ross Sparks. Applications papers: Two-pass Cusum to identify agecluster outbreaks. Australian & New Zealand Journal of Statistics, 52 (3):245–260, September 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Voinov:2010:APP

[1927] Vassilly Voinov, Evgeniy Voinov, and Roza Rakhimova. Applications papers: Poisson versus binomial: Appointment of judges to the U.S. Supreme Court. Australian & New Zealand Journal of Statistics, 52(3): 261–274, September 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ghosh:2010:TMPa

[1928] Pulak Ghosh and Timothy Hanson. Theory and methods papers: a semiparametric Bayesian approach to multivariate longitudinal data. Australian & New Zealand Journal of Statistics, 52(3):275–288, September 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ghosh:2010:TMPb

[1929] Pulak Ghosh, Paramjit Gill, Saman Muthukumarana, and Tim Swartz. Theory and methods papers: a semiparametric Bayesian approach to network modelling using Dirichlet process prior distributions. Australian & New Zealand Journal of Statistics, 52(3):289–302, September 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Haziza:2010:TMP

[1930] David Haziza, Guillaume Chauvet, and Jean-Claude Deville. Theory and methods papers: Sampling and estimation in the presence of cut-off sampling. Australian & New Zealand Journal of Statistics, 52(3):303– 319, September 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gani:2010:TMP

[1931] J. Gani. Theory and methods papers: Modelling epidemic diseases. Australian & New Zealand Journal of Statistics, 52(3):321–329, September 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See responses [1932, 1933].

Daley:2010:TMP

[1932] D. J. Daley. Theory and methods papers: Comment on the paper by Gani. Australian & New Zealand Journal of Statistics, 52(3):330–334, September 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1934].

Welsh:2010:TMP

[1933] A. H. Welsh. Theory and methods papers: Comment on the paper by Gani. Australian & New Zealand Journal of Statistics, 52(3):335–336, September 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1931].

Anonymous:2010:TMP

[1934] Anonymous. Theory and methods papers: Gani’s response to Daley and Welsh. Australian & New Zealand Journal of Statistics, 52(3):337, September 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1931, 1932, 1933].

Anonymous:2010:FPc

[1935] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 52(3):339, September 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Haslett:2010:APS

[1936] S. J. Haslett and G. Jones. Application papers: Small-area estimation of poverty: the aid industry standard and its alternatives. Australian & New Zealand Journal of Statistics, 52(4):341–362, December 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Garthwaite:2010:APS

[1937] Paul H. Garthwaite and Emmanuel Mubwandarikwa. Application papers: Selection of weights for weighted model averaging. Australian & New Zealand Journal of Statistics, 52(4):363–382, December 2010. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Beutner:2010:APN

[1938] E. Beutner and E. Cramer. Application papers: Nonparametric metaanalysis for minimal-repair systems. Australian & New Zealand Journal of Statistics, 52(4):383–401, December 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hall:2010:TMP

[1939] Peter Hall and Hugh Miller. Theory and methods papers: Sequential, bottom-up variable selection for high-dimensional classification. Australian & New Zealand Journal of Statistics, 52(4):403–421, December 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bai:2010:TMP

[1940] Zhidong Bai, Jiaqi Chen, and Jianfeng Yao. Theory and methods papers: On estimation of the population spectral distribution from a highdimensional sample covariance matrix. Australian & New Zealand Journal of Statistics, 52(4):423–437, December 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bedrick:2010:TMP

[1941] Edward J. Bedrick and Winston K. Crandall. Theory and methods papers: Model selection criteria for loglinear models. Australian & New Zealand Journal of Statistics, 52(4):439–449, December 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ogata:2010:TMP

[1942] Hiroaki Ogata and Masanobu Taniguchi. Theory and methods papers: an empirical likelihood approach for non-Gaussian vector stationary processes and its application to minimum contrast estimation. Australian & New Zealand Journal of Statistics, 52(4):451–468, December 2010. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Venables:2010:BRI

[1943] Bill Venables. Book review: Introduction to Data Technologies. by Paul Murrell. Australian & New Zealand Journal of Statistics, 52(4):469– 470, December 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2010:FPd

[1944] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 52(4):471, December 2010. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Steele:2011:IPDa

[1945] Fiona Steele. Invited paper with discussion: Multilevel discrete-time event history models with applications to the analysis of recurrent employment transitions. Australian & New Zealand Journal of Statistics, 53 (1):1–20, March 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See discussion [1946, 1947, 1948].

Haynes:2011:IPD

[1946] Michele Haynes. Invited paper with discussion: Discussion on the paper by Steele. Australian & New Zealand Journal of Statistics, 53(1):20– 23, March 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1945, 1948].

Brown:2011:IPD

[1947] James Brown. Invited paper with discussion: Discussion on the paper by Steele. Australian & New Zealand Journal of Statistics, 53(1):23– 24, March 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1945, 1948].

Steele:2011:IPDb

[1948] Fiona Steele. Invited paper with discussion: Steele’s response to Haynes and Brown. Australian & New Zealand Journal of Statistics, 53(1):24– 26, March 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1947, 1946, 1945].

Lloyd:2011:TMP

[1949] Chris J. Lloyd and Max V. Moldovan. Theory and methods papers: Constructing more powerful exact tests of equivalence from binary matched pairs. Australian & New Zealand Journal of Statistics, 53(1):27–42, March 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

McCabe:2011:TMP

[1950] Brendan P. M. McCabe, Gael M. Martin, and Keith Freeland. Theory and methods papers: a quasi-locally most powerful test for correlation in the conditional variance of positive data. Australian & New Zealand Journal of Statistics, 53(1):43–62, March 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Zhang:2011:TMP

[1951] Yanyu Zhang and Jialiang Li. Theory and methods papers: Combining multiple markers for multi-category classification: an ROC surface approach. Australian & New Zealand Journal of Statistics, 53(1):63–78, March 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Rolls:2011:TMP

[1952] David A. Rolls and Owen D. Jones. Theory and methods papers: Testing for continuous local martingales using the crossing tree. Australian & New Zealand Journal of Statistics, 53(1):79–107, March 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pledger:2011:APM

[1953] Megan J. Pledger and R. Hugh Morton. Applications paper: Modelling the 2004 Super 12 Rugby Union Competition. Australian & New Zealand Journal of Statistics, 53(1):109–121, March 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kenward:2011:BRA

[1954] Mike Kenward. Book review: Antedependence Models for Longitudinal Data, by Dale L. Zimmerman and Vicente A. Núñez–Antón. Australian & New Zealand Journal of Statistics, 53(1):123–124, March 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Liu:2011:BRA

[1955] Ivy Liu. Book review: Analysis of Ordinal Categorical Data, 2nd edn by Alan Agresti. Australian & New Zealand Journal of Statistics, 53(1): 124–126, March 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Liu:2011:BRT

[1956] Wei Liu. Book review: Testing Statistical Hypotheses of Equivalence and Noninferiority, 2nd edn, by Stefan Wellek. Australian & New Zealand Journal of Statistics, 53(1):126–127, March 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wood:2011:BRE

[1957] J. T. Wood. Book review: Environmental and Ecological Statistics with R, by Song S. Qian. Australian & New Zealand Journal of Statistics, 53 (1):127–128, March 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2011:FPa

[1958] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 53(1):129, March 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Peng:2011:TMP

[1959] Liang Peng. Theory and methods papers: Empirical likelihood methods for the Gini index. Australian & New Zealand Journal of Statistics, 53(2): 131–139, June 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Torabi:2011:TMP

[1960] Mahmoud Torabi. Theory and methods papers: Small area estimation using survey weights with functional measurement error in the covariate. Australian & New Zealand Journal of Statistics, 53(2):141–155, June 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Genest:2011:TMP

[1961] Christian Genest, Johanna Nešlehová, and Noomen Ben Ghorbal. Theory and methods papers: Estimators based on Kendall’s Tau in multivariate copula models. Australian & New Zealand Journal of Statistics, 53(2):157–177, June 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chua:2011:TMP

[1962] Chew Lian Chua, G. C. Lim, and Penelope Smith. Theory and methods papers: a Bayesian simulation approach to inference on a multi-state latent factor intensity model. Australian & New Zealand Journal of Statistics, 53(2):179–195, June 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Rublik:2011:TMP

[1963] František Rublík and Ján Somorčík. Theory and methods papers: Affineequivariant spatial median and its use in the multivariate multi-sample location problem. Australian & New Zealand Journal of Statistics, 53(2): 197–215, June 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Rohan:2011:TMP

[1964] Neelabh Rohan and T. V. Ramanathan. Theory and methods papers: Order selection in ARMA models using the focused information criterion. Australian & New Zealand Journal of Statistics, 53(2):217–231, June 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hewson:2011:APA

[1965] Paul J. Hewson. Applications papers: Analysis of head injuries using a Bayesian vector generalized additive model. Australian & New Zealand Journal of Statistics, 53(2):233–246, June 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Tonda:2011:APN

[1966] Tetsuji Tonda, Kenichi Satoh, Teruyuki Nakayama, Kota Katanoda, Tomotaka Sobue, and Megu Ohtaki. Applications papers: a nonparametric mixed-effects model for cancer mortality. Australian & New Zealand Journal of Statistics, 53(2):247–256, June 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Young:2011:LER

[1967] Jim Young. Letter to Editor: Re: Duc, H., Jalaludin, B. & Morgan, G. (2009). Associations between air pollution and hospital visits for cardiovascular diseases in the elderly in Sydney using Bayesian statistical methods. Australian & New Zealand Journal of Statistics, 53(2):257– 259, June 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1968].

Duc:2011:LER

[1968] Hiep Duc, Bin Jalaludin, and Geoff Morgan. Letter to Editor: Reply to comment on ‘Associations between air pollution and hospital visits for cardiovascular diseases in the elderly in Sydney using Bayesian statistical methods’. Australian & New Zealand Journal of Statistics, 53(2):259– 261, June 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1967].

Welsh:2011:BRS

[1969] A. H. Welsh. Book review: Statistical Inference: an Integrated Bayesian/ Likelihood Approach, by Murray Aitkin. Australian & New Zealand Journal of Statistics, 53(2):263–264, June 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Scott:2011:BRI

[1970] David Scott. Book review: Introduction to Scientific Programming and Simulation Using R, by O. Jones, R. Maillardet and A. Robinson. Australian & New Zealand Journal of Statistics, 53(2):264–265, June 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Alliot:2011:BRH

[1971] Pierre Alliot. Book review: Hidden Markov Models for Time Series: an Introduction Using R, by W. Zucchini and I. L. Macdonald. Australian & New Zealand Journal of Statistics, 53(2):265–266, June 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gebski:2011:BRD

[1972] Val Gebski. Book review: Design and Analysis of Clinical Trials with Time-to-Event Endpoints edited, by Karl E. Peace. Australian & New Zealand Journal of Statistics, 53(2):266–267, June 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2011:FPb

[1973] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 53(2):269, June 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Mnatsakanov:2011:TMP

[1974] Robert M. Mnatsakanov, Shengqiao Li, and E. James Harner. Theory and methods papers: Estimation of multivariate Shannon entropy using moments. Australian & New Zealand Journal of Statistics, 53(3):271– 288, September 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Baksalary:2011:TMP

[1975] Oskar Maria Baksalary and Götz Trenkler. Theory and methods papers: Between Olse and Blue. Australian & New Zealand Journal of Statistics, 53(3):289–303, September 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Neville:2011:TMP

[1976] Sarah E. Neville, M. J. Palmer, and M. P. Wand. Theory and methods papers: Generalized extreme value additive model analysis via mean field variational Bayes. Australian & New Zealand Journal of Statistics, 53 (3):305–330, September 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chacon:2011:TMP

[1977] José E. Chacón and Tarn Duong. Theory and methods papers: Unconstrained pilot selectors for smoothed cross-validation. Australian & New Zealand Journal of Statistics, 53(3):331–351, September 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Fisher:2011:APG

[1978] N. I. Fisher and A. J. Lee. Applications papers: Getting the ‘correct’ answer from survey responses: a simple application of the EM algorithm. Australian & New Zealand Journal of Statistics, 53(3):353–364, September 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chia:2011:APE

[1979] Joanne L. C. Chia and Adrian Baddeley. Applications papers: Estimating a ratio of means from bivariate counts with applications in stereology. Australian & New Zealand Journal of Statistics, 53(3):365–387, September 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Beh:2011:BRB

[1980] Eric J. Beh. Book review: Biplots in Practice, by Michael Greenacre. Australian & New Zealand Journal of Statistics, 53(3):389–391, September 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hudson:2011:BRA

[1981] Malcolm Hudson. Book review: Analysing Seasonal Health Data, by Adrian G. Barnett and Annette J. Dobson. Australian & New Zealand Journal of Statistics, 53(3):391–393, September 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Penev:2011:BRL

[1982] Spiridon Penev. Book review: Linear Causal Modeling with Structural Equations, by Stanley Mulaik. Australian & New Zealand Journal of Statistics, 53(3):393–395, September 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Walsh:2011:BRS

[1983] Daniel Walsh. Book review: Statistical Image Processing and Multidimensional Modeling, by Paul Fieguth. Australian & New Zealand Journal of Statistics, 53(3):395–396, September 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2011:FPc

[1984] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 53(3):397, September 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2011:CA

[1985] Anonymous. Campbell Award 2010. Australian & New Zealand Journal of Statistics, 53(4):399–401, December 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hwang:2011:TMP

[1986] Wen-Han Hwang and Richard Huggins. Theory and methods papers: a semiparametric model for a functional behavioural response to capture in capture–recapture experiments. Australian & New Zealand Journal of Statistics, 53(4):403–421, December 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Griffin:2011:TMP

[1987] Jim E. Griffin and Philip J. Brown. Theory and methods papers: Bayesian hyper-lassos with non-convex penalization. Australian & New Zealand Journal of Statistics, 53(4):423–442, December 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gubbiotti:2011:TMP

[1988] Stefania Gubbiotti and Fulvio De Santis. Theory and methods papers: a Bayesian method for the choice of the sample size in equivalence trials. Australian & New Zealand Journal of Statistics, 53(4):443–460, December 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Clarke:2011:APO

[1989] G. Peter Y. Clarke and Katia T. Stefanova. Applications papers: Optimal design for early-generation plant-breeding trials with unreplicated or partially replicated test lines. Australian & New Zealand Journal of Statistics, 53(4):461–480, December 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Clarke:2011:CNR

[1990] B. R. Clarke, D. K. Gamble, and T. Bednarski. Corrigendum: On a Note on Robustness of the β-Trimmed Mean. Australian & New Zealand Journal of Statistics, 53(4):481, December 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1459].

Staudte:2011:BRR

[1991] Robert Staudte and Luke Prendergast. Book review: Robust Methods in Biostatistics, by Stephane Heritier, Eva Cantoni, Samuel Copt and Maria-Pia Victoria-Feser. Australian & New Zealand Journal of Statistics, 53(4):483–486, December 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Williams:2011:BRC

[1992] Emlyn Williams. Book review: Computer-Assisted Statistics Textbooks — Experiments in Agriculture and Biology, by W. Douglas Stirling. Australian & New Zealand Journal of Statistics, 53(4):486–487, December 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Marschner:2011:BRM

[1993] Ian C. Marschner. Book review: Monte Carlo Simulation for the Pharmaceutical Industry: Concepts, Algorithms and Case Studies by Mark Chang. Australian & New Zealand Journal of Statistics, 53(4):487–489, December 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Sisson:2011:BRH

[1994] Scott A. Sisson. Book review: Handbook of Spatial Statistics, by A. E. Gelfand, P. Diggle, P. Guttorp and M. Fuentes. Australian & New Zealand Journal of Statistics, 53(4):489–491, December 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Fletcher:2011:BRR

[1995] David Fletcher. Book review: R Through Excel: a Spreadsheet Interface for Statistics, Data Analysis, and Graphics, by Richard M. Heiberger and Erich Neuwirth. Australian & New Zealand Journal of Statistics, 53(4):491–492, December 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wood:2011:BRF

[1996] G. Wood. Book review: The Fundamentals of Modern Statistical Genetics. by Nan M. Laird and Christoph Lange. Australian & New Zealand Journal of Statistics, 53(4):492–493, December 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Beh:2011:BRU

[1997] Eric J. Beh. Book review: Understanding Biplots, by John Gower, Sugnet Gardner-Lubbe and Niël Le Roux. Australian & New Zealand Journal of Statistics, 53(4):493–494, December 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2011:FPd

[1998] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 53(4):495, December 2011. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gurrin:2012:APC

[1999] Lyle C. Gurrin and Lidija Turkovic. Applications papers: Combining individual participant data and summary statistics from both continuously valued and binary variables to estimate regression parameters. Australian & New Zealand Journal of Statistics, 54(1):1–21, March 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hurlbert:2012:APL

[2000] Stuart H. Hurlbert and Celia M. Lombardi. Applications papers: Lopsided reasoning on lopsided tests and multiple comparisons. Australian & New Zealand Journal of Statistics, 54(1):23–42, March 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hosseini:2012:APB

[2001] Fatemeh Hosseini and Mohsen Mohammadzadeh. Applications papers: Bayesian prediction for spatial generalised linear mixed models with closed skew normal latent variables. Australian & New Zealand Journal of Statistics, 54(1):43–62, March 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Huang:2012:TMP

[2002] Xin Huang, Gengsheng Qin, Yan Yuan, and Xiao hua Zhou. Theory and methods papers: Confidence intervals for the difference between two partial AUCs. Australian & New Zealand Journal of Statistics, 54(1): 63–79, March 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Diaconis:2012:TMP

[2003] Persi Diaconis and Robert Griffiths. Theory and methods papers: Exchangeable pairs of Bernoulli random variables, Krawtchouk polynomials, and Ehrenfest urns. Australian & New Zealand Journal of Statistics, 54(1):81–101, March 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Salvati:2012:TMP

[2004] Nicola Salvati, Hukum Chandra, and Ray Chambers. Theory and methods papers: Model-based direct estimation of small-area distributions. Australian & New Zealand Journal of Statistics, 54(1):103–123, March 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lloyd:2012:BRL

[2005] Chris Lloyd. Book review: Logistic Regression Models, by J. M. Hilbe. Australian & New Zealand Journal of Statistics, 54(1):125–126, March 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Brown:2012:BRP

[2006] Bruce Brown. Book review: Permutation Tests for Complex Data: Theory, Applications and Software, by F. Pesarin and L. Salmaso. Australian & New Zealand Journal of Statistics, 54(1):126–127, March 2012. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Brittain:2012:BRD

[2007] Erica Brittain. Book review: Design and Analysis of Non-inferiority Trials, by Mark D. Rothmann, Brian L. Wiens and Ivan S. F. Chan. Australian & New Zealand Journal of Statistics, 54(1):127–129, March 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2012:FPa

[2008] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 54(1):131, March 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2012:CA

[2009] Anonymous. Campbell Award 2011. Australian & New Zealand Journal of Statistics, 54(2):133–136, June 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Beh:2012:TMPa

[2010] Eric J. Beh and Rosaria Lombardo. Theory and methods papers: a genealogy of correspondence analysis. Australian & New Zealand Journal of Statistics, 54(2):137–168, June 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kasza:2012:TMP

[2011] Jessica Kasza, Gary Glonek, and Patty Solomon. Theory and methods papers: Estimating Bayesian networks for high-dimensional data with complex mean structure and random effects. Australian & New Zealand Journal of Statistics, 54(2):169–187, June 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Albatineh:2012:TMP

[2012] Ahmed N. Albatineh, Hafiz M. R. Khan, and Magdalena Niewiadomska-Bugaj. Theory and methods papers: On the equivalence of some indices of similarity: Implication for binary presence/absence data. Australian & New Zealand Journal of Statistics, 54(2):189–198, June 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Dembinska:2012:TMP

[2013] Anna Dembińska. Theory and methods papers: Limit theorems for proportions of observations falling into random regions determined by order statistics. Australian & New Zealand Journal of Statistics, 54(2): 199–210, June 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Zhang:2012:TMP

[2014] Chongqi Zhang, Wengkee Wong, and Heng Peng. Theory and methods papers: Dual-objective optimal mixture designs. Australian & New Zealand Journal of Statistics, 54(2):211–222, June 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chipperfield:2012:TMP

[2015] James Chipperfield, Julia Chessman, and Russell Lim. Theory and methods papers: Combining household surveys using mass imputation to estimate population totals. Australian & New Zealand Journal of Statistics, 54(2):223–238, June 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kim:2012:TMP

[2016] Ji Young Kim. Theory and methods papers: Parametric fractional imputation for non-ignorable categorical missing data with follow-up. Australian & New Zealand Journal of Statistics, 54(2):239–250, June 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Galbraith:2012:BRA

[2017] Sally Galbraith. Book review: Applied Missing Data Analysis, by Craig K. Enders. Australian & New Zealand Journal of Statistics, 54(2):251, June 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cumming:2012:BRA

[2018] Geoff Cumming. Book review: Applied Meta-Analysis for Social Science Research, by Noel A. Card. Australian & New Zealand Journal of Statistics, 54(2):252–253, June 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Martin:2012:BRI

[2019] Michael A. Martin. Book review: An Introduction to Bootstrap Methods with Applications to R, by M. R. Chernick and R. A. LaBudde. Australian & New Zealand Journal of Statistics, 54(2):253–255, June 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lu:2012:BRS

[2020] Zudi Lu. Book review: Short-Memory Linear Processes and Econometric Applications, by Kairat T. Mynbaev. Australian & New Zealand Journal of Statistics, 54(2):255–257, June 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2012:FPb

[2021] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 54(2):259, June 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Langrock:2012:APF

[2022] Roland Langrock. Applications papers: Flexible latent-state modelling of Old Faithful’s eruption inter-arrival times in 2009. Australian & New Zealand Journal of Statistics, 54(3):261–279, September 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Decrouez:2012:APC

[2023] Geoffrey Decrouez and Andrew P. Robinson. Applications papers: Confidence intervals for the weighted sum of two independent binomial proportions. Australian & New Zealand Journal of Statistics, 54(3):281–299, September 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cao:2012:APG

[2024] Ricardo Cao, José A. Vilar, and Juan M. Vilar. Applications papers: Generalised variance function estimation for binary variables in largescale sample surveys. Australian & New Zealand Journal of Statistics, 54(3):301–324, September 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Grillenzoni:2012:APE

[2025] Carlo Grillenzoni. Applications papers: Evaluation of recursive detection methods for turning points in financial time series. Australian & New Zealand Journal of Statistics, 54(3):325–342, September 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kiermeier:2012:APE

[2026] Andreas Kiermeier, Arūnas P. Verbyla, and Richard G. Jarrett. Applications papers: Estimating a single shelf-life for multiple batches. Australian & New Zealand Journal of Statistics, 54(3):343–358, September 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hossain:2012:APS

[2027] Shakhawat Hossain and Ejaz Ahmed. Applications papers: Shrinkage and penalty estimators of a Poisson regression model. Australian & New Zealand Journal of Statistics, 54(3):359–373, September 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kokic:2012:APC

[2028] Philip Kokic, Kenton Lawson, Lisa Elliston, Chun Liang, and Walter Shafron. Applications papers: a comparison of two methods of linking spatial and economic data from farm surveys. Australian & New Zealand Journal of Statistics, 54(3):375–390, September 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

James:2012:BRD

[2029] Ian James. Book review: Dynamic Mixed Models for Familial Longitudinal Data. Australian & New Zealand Journal of Statistics, 54(3):391–392, September 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Sen:2012:BRB

[2030] Śaunak Sen. Book review: Bayesian modeling in bioinformatics. Australian & New Zealand Journal of Statistics, 54(3):392–393, September 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Jamsen:2012:BRA

[2031] Kris M. Jamsen and Lyle C. Gurrin. Book review: Applied Bayesian Hierarchical Methods. Australian & New Zealand Journal of Statistics, 54(3):393, September 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Martin:2012:BRC

[2032] Andrew Martin. Book review: Controversial Statistical Issues in Clinical Trials. Australian & New Zealand Journal of Statistics, 54(3):393–394, September 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kabaila:2012:BRI

[2033] Paul Kabaila. Book review: Introduction to the Theory of Statistical Inference. Australian & New Zealand Journal of Statistics, 54(3):394– 395, September 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2012:FPc

[2034] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 54(3):397, September 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2012:PMP

[2035] Anonymous. 2012 Pitman Medal: Professor Alan Welsh, FAA. Australian & New Zealand Journal of Statistics, 54(4):399–400, December 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Radice:2012:APB

[2036] Rosalba Radice. Applications paper: a Bayesian approach to modelling reticulation events with application to the ribosomal protein gene rps11 of flowering plants. Australian & New Zealand Journal of Statistics, 54 (4):401–426, December 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Taylor:2012:TMP

[2037] Julian D. Taylor, Arūnas P. Verbyla, Colin Cavanagh, and Marcus Newberry. Theory and methods papers: Variable selection in linear mixed models using an extended class of penalties. Australian & New Zealand Journal of Statistics, 54(4):427–449, December 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bergman:2012:TMP

[2038] Jakob Bergman and Björn Holmquist. Theory and methods papers: a measure of dependence between two compositions. Australian & New Zealand Journal of Statistics, 54(4):451–461, December 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Li:2012:TMP

[2039] Zhouping Li, Xiping Wang, and Wang Zhou. Theory and methods papers: Empirical likelihood for compound Poisson processes. Australian & New Zealand Journal of Statistics, 54(4):463–474, December 2012. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Beh:2012:TMPb

[2040] Eric J. Beh and Thomas B. Farver. Theory and methods papers: The Box–Cox transformation and non-iterative estimation methods for ordinal log-linear models. Australian & New Zealand Journal of Statistics, 54(4):475–484, December 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Plante:2012:TMP

[2041] Jean-François Plante and Gitte Bjørg Windfeldt. Theory and methods papers: Using predictive risk for process control. Australian & New Zealand Journal of Statistics, 54(4):485–504, December 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Welsh:2012:BRE

[2042] A. H. Welsh. Book review: The Epistemology of Statistical Science. Australian & New Zealand Journal of Statistics, 54(4):505–507, December 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wand:2012:BRS

[2043] M. P. Wand. Book review: Smoothing Splines: Methods and Applications. Australian & New Zealand Journal of Statistics, 54(4):507–508, December 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Roberts:2012:BRS

[2044] Steven Roberts. Book review: Statistical Thinking in Epidemiology. Australian & New Zealand Journal of Statistics, 54(4):508–509, December 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Broemeling:2012:BRS

[2045] Lyle Broemeling. Book review: Statistical Evaluation of Diagnostic Performance: Topics in ROC Analysis. Australian & New Zealand Journal of Statistics, 54(4):509–510, December 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2012:FPd

[2046] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 54(4):511, December 2012. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kiesse:2013:APW

[2047] Tristan Senga Kiessé and Dominique Mizère. Application paper: Weighted Poisson and semiparametric kernel models applied to parasite growth. Australian & New Zealand Journal of Statistics, 55(1):1–13, March 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Rayner:2013:TMPa

[2048] John C. W. Rayner, Olivier Thas, Peter Pipelers, and Eric J. Beh. Theory and methods papers: Calculating bivariate orthonormal polynomials by recurrence. Australian & New Zealand Journal of Statistics, 55(1): 15–24, March 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Paige:2013:TMP

[2049] Robert L. Paige and A. Alexandre Trindade. Theory and methods papers: Fast and accurate inference for the smoothing parameter in semiparametric models. Australian & New Zealand Journal of Statistics, 55 (1):25–41, March 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gao:2013:TMP

[2050] Shesheng Gao, Yongmin Zhong, and Chengfan Gu. Theory and methods papers: Random weighting estimation of confidence intervals for quantiles. Australian & New Zealand Journal of Statistics, 55(1):43–53, March 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kuhnert:2013:BRR

[2051] Petra Kuhnert. Book review: Recursive Partitioning and Applications. Australian & New Zealand Journal of Statistics, 55(1):55–56, March 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Heritier:2013:BRD

[2052] Stephane Heritier. Book review: Design and Analysis of Quality of Life Studies in Clinical Trials. Australian & New Zealand Journal of Statistics, 55(1):56–57, March 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Maller:2013:BRC

[2053] Ross Maller. Book review: Competing Risks and Multistate Models with R. Australian & New Zealand Journal of Statistics, 55(1):57–60, March 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Robinson:2013:BRM

[2054] Andrew Robinson. Book review: Maximum Likelihood Estimation for Sample Surveys. Australian & New Zealand Journal of Statistics, 55(1): 60–61, March 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wang:2013:BRG

[2055] You-Gan Wang. Book review: Generalized Estimating Equations. Australian & New Zealand Journal of Statistics, 55(1):61, March 2013. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2013:FPa

[2056] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 55(1):63, March 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2013:CA

[2057] Anonymous. Campbell Award 2012. Australian & New Zealand Journal of Statistics, 55(2):65–67, June 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anacleto:2013:APF

[2058] Osvaldo Anacleto, Catriona Queen, and Casper J. Albers. Application paper: Forecasting multivariate road traffic flows using Bayesian dynamic graphical models, splines and other traffic variables. Australian & New Zealand Journal of Statistics, 55(2):69–86, June 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Khan:2013:TMP

[2059] M. Atikur Rahman Khan and D. S. Poskitt. Theory and methods papers: a note on window length selection in singular spectrum analysis. Australian & New Zealand Journal of Statistics, 55(2):87–108, June 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Qin:2013:TMP

[2060] Yongsong Qin, Qingzhu Lei, and Lijun Luo. Theory and methods papers: Empirical likelihood for nonparametric models under linear process errors. Australian & New Zealand Journal of Statistics, 55(2):109–128, June 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Mercatanti:2013:TMP

[2061] Andrea Mercatanti. Theory and methods papers: a likelihood-based analysis for relaxing the exclusion restriction in randomized experiments with noncompliance. Australian & New Zealand Journal of Statistics, 55 (2):129–153, June 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cai:2013:TMP

[2062] Yuzhi Cai. Theory and methods papers: a quantile survival model for censored data. Australian & New Zealand Journal of Statistics, 55(2): 155–172, June 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ma:2013:TMP

[2063] Xiangmei Ma and Liming Xiang. Theory and methods papers: Testing for the presence of a cure fraction in clustered interval-censored survival data. Australian & New Zealand Journal of Statistics, 55(2):173–190, June 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pakes:2013:BRA

[2064] Anthony G. Pakes. Book review: Abraham De Moivre: Setting the Stage for Classical Probability and its Applications. Australian & New Zealand Journal of Statistics, 55(2):191–197, June 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Heller:2013:BRG

[2065] Gillian Z. Heller. Book review: Generalized Linear Mixed Models: Modern Concepts, Methods and Applications. Australian & New Zealand Journal of Statistics, 55(2):197–198, June 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Jones:2013:BRJ

[2066] Geoff Jones. Book review: Joint Models for Longitudinal and Time-to-Event Data: with Applications in R. Australian & New Zealand Journal of Statistics, 55(2):198–199, June 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

White:2013:BRC

[2067] Gentry White. Book review: Criminal Justice Forecasts of Risk: a Machine Learning Approach. Australian & New Zealand Journal of Statistics, 55(2):199–201, June 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2013:FPb

[2068] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 55(2):203, June 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Holland:2013:IPR

[2069] Barbara R. Holland. Invited paper: The rise of statistical phylogenetics. Australian & New Zealand Journal of Statistics, 55(3):205–220, September 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Abeysekera:2013:APC

[2070] Waruni Abeysekera, Paul Kabaila, and Oguzhan Yilmaz. Application papers: The coverage probability of confidence intervals in one-way analysis of covariance after two F tests. Australian & New Zealand Journal of Statistics, 55(3):221–234, September 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Karavarsamis:2013:APC

[2071] Natalie Karavarsamis, Andrew P. Robinson, Graham Hepworth, Andrew J. Hamilton, and Geoffrey W. Heard. Application papers: Comparison of four bootstrap-based interval estimators of species occupancy and detection probabilities. Australian & New Zealand Journal of Statistics, 55(3):235–252, September 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Williams:2013:TMP

[2072] E. R. Williams and H. P. Piepho. Theory and methods papers: a comparison of spatial designs for field variety trials. Australian & New Zealand Journal of Statistics, 55(3):253–258, September 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kabaila:2013:TMP

[2073] Paul Kabaila and Khageswor Giri. Theory and methods papers: Further properties of frequentist confidence intervals in regression that utilize uncertain prior information. Australian & New Zealand Journal of Statistics, 55(3):259–270, September 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Jung:2013:TMP

[2074] Sangah Jung and Sangun Park. Theory and methods papers: Simplified maximum likelihood inference based on the likelihood decomposition for missing data. Australian & New Zealand Journal of Statistics, 55(3):271– 283, September 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wood:2013:TMP

[2075] Graham R. Wood and David J. Saville. Theory and methods papers: a geometric examination of linear model assumptions. Australian & New Zealand Journal of Statistics, 55(3):285–303, September 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Rayner:2013:TMPb

[2076] J. C. W. Rayner and D. J. Best. Theory and methods papers: Extended ANOVA and rank transform procedures. Australian & New Zealand Journal of Statistics, 55(3):305–319, September 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Marra:2013:TMP

[2077] Giampiero Marra, Georgios Papageorgiou, and Rosalba Radice. Theory and methods papers: Estimation of a semiparametric recursive bivariate probit model with nonparametric mixing. Australian & New Zealand Journal of Statistics, 55(3):321–342, September 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Baade:2013:BRH

[2078] Peter Baade. Book review: Handbook of Statistics in Clinical Oncology. Australian & New Zealand Journal of Statistics, 55(3):343, September 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Browne:2013:BRB

[2079] William Browne. Book review: Bayesian item response modeling: Theory and applications. Australian & New Zealand Journal of Statistics, 55(3): 344, September 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bebbington:2013:BRE

[2080] Mark Bebbington. Book review: Event History Analysis with R. Australian & New Zealand Journal of Statistics, 55(3):344–346, September 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Scarrott:2013:BRE

[2081] Carl Scarrott. Book review: Extreme Value Methods with Applications to Finance. Australian & New Zealand Journal of Statistics, 55(3):346– 348, September 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2013:FPc

[2082] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 55(3):349, September 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hsieh:2013:APB

[2083] Jeff Ching-Fu Hsieh, Susanna M. Cramb, James M. McGree, Peter D. Baade, Nathan A. M. Dunn, and Kerrie L. Mengersen. Application papers: Bayesian spatial analysis for the evaluation of breast cancer detection methods. Australian & New Zealand Journal of Statistics, 55 (4):351–367, December 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Khan:2013:APT

[2084] Shahedul A. Khan, Grace S. Chiu, and Joel A. Dubin. Application papers: Therapeutic hypothermia: Quantification of the transition of core body temperature using the flexible mixture bent-cable model for longitudinal data. Australian & New Zealand Journal of Statistics, 55 (4):369–385, December 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kim:2013:TMP

[2085] Tae Yoon Kim, Jeongcheol Ha, Sun Young Hwang, Cheolyong Park, and Zhi-Ming Luo. Theory and methods papers: Central limit theorems for reduced U-statistics under dependence and their usefulness. Australian & New Zealand Journal of Statistics, 55(4):387–399, December 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Abebe:2013:TMP

[2086] Asheber Abebe and Joseph W. McKean. Theory and methods papers: Weighted Wilcoxon estimators in nonlinear regression. Australian & New Zealand Journal of Statistics, 55(4):401–420, December 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Liu:2013:TMP

[2087] Wei Liu, Sanyu Zhou, and Frank Bretz. Theory and methods papers: Exact simultaneous confidence bands for quadratic and cubic polynomial regression with applications in dose response study. Australian & New Zealand Journal of Statistics, 55(4):421–434, December 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Qian:2013:TMP

[2088] Guoqi Qian, Xiaoping Shi, and Yuehua Wu. Theory and methods papers: a statistical test of change-point in mean that almost surely has zero error probabilities. Australian & New Zealand Journal of Statistics, 55(4):435– 454, December 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Siray:2013:TMP

[2089] Gülesen Üstündağ Şiray and Selma Toker. Theory and methods papers: Restricted two parameter ridge estimator. Australian & New Zealand Journal of Statistics, 55(4):455–469, December 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Welsh:2013:TMP

[2090] A. H. Welsh, Dan Hedlin, and Markus G. Šova. Theory and methods papers: The stabilisation of model parameter estimates from repeated surveys with rare observations. Australian & New Zealand Journal of Statistics, 55(4):471–491, December 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Stone:2013:BRS

[2091] Glenn Stone. Book review: Statistics for High-Dimensional Data: Methods, Theory and Applications. Australian & New Zealand Journal of Statistics, 55(4):493, December 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bolstad:2013:BRB

[2092] William M. Bolstad. Book review: The BUGS Book: a Practical Introduction to Bayesian Analysis. Australian & New Zealand Journal of Statistics, 55(4):493–495, December 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Jorgensen:2013:BRS

[2093] Murray Jorgensen. Book review: A Statistical Guide for the Ethically Perplexed. Australian & New Zealand Journal of Statistics, 55(4):495– 497, December 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Clarke:2013:BRM

[2094] Brenton R. Clarke. Book review: Methodology in Robust and Nonparametric Statistics. Australian & New Zealand Journal of Statistics, 55 (4):497–499, December 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2013:FPd

[2095] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 55(4):501, December 2013. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lumley:2014:IPT

[2096] Thomas Lumley and Alastair Scott. Invited paper: Tests for regression models fitted to survey data. Australian & New Zealand Journal of Statistics, 56(1):1–14, March 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hussein:2014:APN

[2097] Abdulkadir A. Hussein, Sévérien Nkurunziza, and Katrina Tomanelli. Applications paper: Nonparametric shrinkage estimation for Aalen’s additive hazards model. Australian & New Zealand Journal of Statistics, 56(1):15–26, March 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Samart:2014:APL

[2098] Klairung Samart and Ray Chambers. Applications paper: Linear regression with nested errors using probability-linked data. Australian & New Zealand Journal of Statistics, 56(1):27–46, March 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Withers:2014:TMP

[2099] Christopher S. Withers and Saralees Nadarajah. Theory and methods papers: Confidence intervals for linear combinations of Poisson means. Australian & New Zealand Journal of Statistics, 56(1):47–58, March 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lee:2014:TMPa

[2100] Keunbaik Lee and Jae Keun Yoo. Theory and methods papers: Canonical correlation analysis through linear modeling. Australian & New Zealand Journal of Statistics, 56(1):59–72, March 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

You:2014:TMP

[2101] Chong You, John T. Ormerod, and Samuel Müller. Theory and methods papers: On variational Bayes estimation and variational information criteria for linear regression models. Australian & New Zealand Journal of Statistics, 56(1):73–87, March 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Williams:2014:TMP

[2102] E. R. Williams, J. A. John, and D. Whitaker. Theory and methods papers: Construction of more flexible and efficient P –rep designs. Australian & New Zealand Journal of Statistics, 56(1):89–96, March 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Brown:2014:BRC

[2103] Bruce Brown. Book review: Confidence Intervals for Proportions and Related Measures of Effect Size. Australian & New Zealand Journal of Statistics, 56(1):97–98, March 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2014:FPa

[2104] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 56(1):99, March 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2014:NAC

[2105] Anonymous. NZSA Awards 2013: Campbell Award 2013. Australian & New Zealand Journal of Statistics, 56(2):101–102, June 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2014:NAN

[2106] Anonymous. NZSA Awards 2013: NZSA Research Awards 2013. Australian & New Zealand Journal of Statistics, 56(2):103, June 2014. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Delaigle:2014:IPN

[2107] Aurore Delaigle. Invited paper: Nonparametric kernel methods with errors-in-variables: Constructing estimators, computing them, and avoiding common mistakes. Australian & New Zealand Journal of Statistics, 56(2):105–124, June 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gambacorta:2014:APD

[2108] R. Gambacorta, M. Iannario, and R. Valliant. Applications paper: Design-based inference in a mixture model for ordinal variables for a two stage stratified design. Australian & New Zealand Journal of Statistics, 56(2):125–143, June 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Scealy:2014:TMP

[2109] J. L. Scealy and A. H. Welsh. Theory and methods papers: Colours and cocktails: Compositional data analysis 2013 Lancaster Lecture. Australian & New Zealand Journal of Statistics, 56(2):145–169, June 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

vandenBoogaart:2014:TMP

[2110] Karl Gerald van den Boogaart, Juan José Egozcue, and Vera Pawlowsky-Glahn. Theory and methods papers: Bayes Hilbert spaces. Australian & New Zealand Journal of Statistics, 56(2):171–194, June 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Patterson:2014:BRD

[2111] Scott D. Patterson. Book review: Design and Analysis of Bridging Studies. Australian & New Zealand Journal of Statistics, 56(2):195–196, June 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Jones:2014:BRL

[2112] Geoff Jones. Book review: Latent Markov Models for Longitudinal Data. Australian & New Zealand Journal of Statistics, 56(2):196–197, June 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2014:FPb

[2113] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 56(2):199, June 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Iwata:2014:APD

[2114] Takaki Iwata. Applications papers: Decomposition of seasonality and long-term trend in seismological data: a Bayesian modelling of earthquake detection capability. Australian & New Zealand Journal of Statistics, 56(3):201–215, September 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Li:2014:APS

[2115] Haifen Li, Jiajia Zhang, and Yincai Tang. Applications papers: Smooth semi-nonparametric analysis for mixture cure models and its application to breast cancer. Australian & New Zealand Journal of Statistics, 56 (3):217–235, September 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Fan:2014:APR

[2116] Chunpeng Fan and Donghui Zhang. Applications papers: Robust small sample inference for generalised estimating equations: an application of the ANOVA-type test. Australian & New Zealand Journal of Statistics, 56(3):237–255, September 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Melville:2014:APM

[2117] Gavin J. Melville and Alan H. Welsh. Applications papers: Model-based prediction in ecological surveys including those with incomplete detection. Australian & New Zealand Journal of Statistics, 56(3):257–281, September 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lee:2014:TMPb

[2118] Taewook Lee, Cheolwoo Park, and Young Joo Yoon. Theory and methods papers: Bridge estimation for linear regression models with mixing properties. Australian & New Zealand Journal of Statistics, 56(3):283– 302, September 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Robinson:2014:BRU

[2119] Andrew Robinson. Book review: Understanding Advanced Statistical Methods. Australian & New Zealand Journal of Statistics, 56(3):303, September 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Murrell:2014:BRR

[2120] Paul Murrell. Book review: The R Student Companion. Australian & New Zealand Journal of Statistics, 56(3):304, September 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Milne:2014:BRS

[2121] Barry Milne. Book review: Statistical Disclosure Control. Australian & New Zealand Journal of Statistics, 56(3):305, September 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2014:FPc

[2122] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 56(3):307, September 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Prangle:2014:IPD

[2123] D. Prangle, M. G. B. Blum, G. Popovic, and S. A. Sisson. Invited papers: Diagnostic tools for approximate Bayesian computation using the coverage property. Australian & New Zealand Journal of Statistics, 56(4):309–329, December 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Neuhaus:2014:IPC

[2124] John M. Neuhaus and Charles E. McCulloch. Invited papers: Covariate decomposition methods for longitudinal missing-at-random data and predictors associated with subject-specific effects. Australian & New Zealand Journal of Statistics, 56(4):331–345, December 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Moradi:2014:APA

[2125] Mohammad Moradi, Jennifer A. Brown, and Weilong Guo. Applications paper: Adaptive cluster sampling in two-stage sampling. Australian & New Zealand Journal of Statistics, 56(4):347–357, December 2014. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lloyd:2014:TMP

[2126] Chris J. Lloyd. Theory and methods papers: On the exact size of tests of treatment effects in multi-arm clinical trials. Australian & New Zealand Journal of Statistics, 56(4):359–369, December 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kabaila:2014:TMP

[2127] Paul Kabaila and Dilshani Tissera. Theory and methods papers: Confidence intervals in regression that utilize uncertain prior information about a vector parameter. Australian & New Zealand Journal of Statistics, 56(4):371–383, December 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bairamov:2014:TMP

[2128] Ismihan Bairamov and Mahdi Tavangar. Theory and methods papers: More on progressively Type-II censored order statistics for multivariate observations. Australian & New Zealand Journal of Statistics, 56(4):385– 394, December 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Talebi:2014:TMP

[2129] Hooshang Talebi and Elham Jalali. Theory and methods papers: a series of single array 2^{m} factorial search designs for even m. Australian & New Zealand Journal of Statistics, 56(4):395–405, December 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Donald:2014:TMP

[2130] Margaret R. Donald and Kerrie L. Mengersen. Theory and methods papers: Methods for constructing uncertainty intervals for queries of Bayesian nets. Australian & New Zealand Journal of Statistics, 56(4): 407–427, December 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gillam:2014:BRM

[2131] Marianne H. Gillam. Book review: Multivariate Survival Analysis and Competing Risks. Australian & New Zealand Journal of Statistics, 56 (4):429–430, December 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Jorgensen:2014:BRA

[2132] Murray Jorgensen. Book review: Algorithmic Probability and Friends. Bayesian Prediction and Artificial Intelligence. Australian & New Zealand Journal of Statistics, 56(4):430–431, December 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bergman:2014:BRP

[2133] Jakob Bergman. Book review: Practical Multivariate Analysis, Fifth Edition. Australian & New Zealand Journal of Statistics, 56(4):431– 432, December 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Salvado:2014:BRS

[2134] Olivier Salvado. Book review: Statistical and Computational Methods in brain Image Analysis. Australian & New Zealand Journal of Statistics, 56(4):432–433, December 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Marschner:2014:BRU

[2135] Ian C. Marschner. Book review: Understanding the New Statistics: Effect Sizes, Confidence Intervals, and Meta-Analysis. Australian & New Zealand Journal of Statistics, 56(4):433–435, December 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2014:FPd

[2136] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 56(4):437, December 2014. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hunter:2015:HPS

[2137] Jeffrey J. Hunter. History paper: Shayle R. Searle: Pioneer in linear modelling. Australian & New Zealand Journal of Statistics, 57(1):1– 14, March 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Porter:2015:APS

[2138] Aaron T. Porter, Christopher K. Wikle, and Scott H. Holan. Applications papers: Small area estimation via multivariate Fay–Herriot models with latent spatial dependence. Australian & New Zealand Journal of Statistics, 57(1):15–29, March 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gaiselmann:2015:APS

[2139] Gerd Gaiselmann, Rafal Kulik, and Volker Schmidt. Applications papers: Statistical inference for curved fibrous objects in 3D — based on multiple short observations of multivariate autoregressive processes. Australian & New Zealand Journal of Statistics, 57(1):31–54, March 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Davis:2015:APS

[2140] Jack Davis, Harsha Perera, and Tim B. Swartz. Applications papers: a simulator for Twenty20 Cricket. Australian & New Zealand Journal of Statistics, 57(1):55–71, March 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cao:2015:TMP

[2141] Lei Cao, Jian Tao, Ning-Zhong Shi, and Wei Liu. Theory and methods papers: a stepwise confidence interval procedure under unknown variances based on an asymmetric loss function for toxicological evaluation. Australian & New Zealand Journal of Statistics, 57(1):73–98, March 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kabaila:2015:TMP

[2142] Paul Kabaila and Gayan Dharmarathne. Theory and methods papers: a comparison of Bayesian and frequentist interval estimators in regression that utilize uncertain prior information. Australian & New Zealand Journal of Statistics, 57(1):99–118, March 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Menictas:2015:TMP

[2143] Marianne Menictas and Matt P. Wand. Theory and methods papers: Variational inference for heteroscedastic semiparametric regression. Australian & New Zealand Journal of Statistics, 57(1):119–138, March 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cai:2015:TMP

[2144] Yuzhi Cai and Tao Jiang. Theory and methods papers: Estimation of non-crossing quantile regression curves. Australian & New Zealand Journal of Statistics, 57(1):139–162, March 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ma:2015:BRI

[2145] Jun Ma. Book review: Interval-Censored Time-to-Event Data: Methods and Applications. Australian & New Zealand Journal of Statistics, 57(1):

163–164, March 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Beh:2015:BRA

[2146] Eric J. Beh. Book review: Applied Categorical and Count Data Analysis. Australian & New Zealand Journal of Statistics, 57(1):164–165, March 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2015:FPa

[2147] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 57(1):167, March 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lago:2015:API

[2148] Luise Patricia Lago and Robert Graham Clark. Applications papers: Imputation of household survey data using linear mixed models. Australian & New Zealand Journal of Statistics, 57(2):169–187, June 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Salehi:2015:API

[2149] Mohammad Salehi, Mohammad Moradi, Jassim A. Al Khayat, Jennifer Brown, and Adil Eltayeb Mohamed Yousif. Applications papers: Inverse adaptive cluster sampling with unequal selection probabilities: Case studies on crab holes and arsenic pollution. Australian & New Zealand Journal of Statistics, 57(2):189–201, June 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Williams:2015:APO

[2150] Emlyn Williams and Hans-Peter Piepho. Applications papers: Optimality and contrasts in block designs with unequal treatment replication. Australian & New Zealand Journal of Statistics, 57(2):203–209, June 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Rayner:2015:TMP

[2151] J. C. W. Rayner, D. J. Best, and O. Thas. Theory and methods papers: Extended analysis of at least partially ordered multi-factor ANOVA. Australian & New Zealand Journal of Statistics, 57(2):211–224, June 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Godolphin:2015:TMP

[2152] J. D. Godolphin and E. J. Godolphin. Theory and methods papers: The use of treatment concurrences to assess robustness of binary block designs against the loss of whole blocks. Australian & New Zealand Journal of Statistics, 57(2):225–239, June 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hall:2015:TMP

[2153] Daniel B. Hall and Jing Shen. Theory and methods papers: Marginal projected multivariate linear models for clustered angular data. Australian & New Zealand Journal of Statistics, 57(2):241–257, June 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wang:2015:TMP

[2154] Xinghui Wang and Shuhe Hu. Theory and methods papers: On the strong consistency of M-estimates in linear models for negatively superadditive dependent errors. Australian & New Zealand Journal of Statistics, 57(2):259–274, June 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Namazi-Rad:2015:TMP

[2155] Mohammad-Reza Namazi-Rad and David Steel. Theory and methods papers: What level of statistical model should we use in small area estimation? Australian & New Zealand Journal of Statistics, 57(2):275–298, June 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2015:FPb

[2156] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 57(2):299, June 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Huang:2015:APH

[2157] Shih-Feng Huang and Chan-Yi Tsai. Applications papers: Hedging barrier options in GARCH models with transaction costs. Australian & New Zealand Journal of Statistics, 57(3):301–324, September 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kang:2015:APC

[2158] Su Yun Kang, James McGree, Peter Baade, and Kerrie Mengersen. Applications papers: a case study for modelling cancer incidence using Bayesian spatio-temporal models. Australian & New Zealand Journal of Statistics, 57(3):325–345, September 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pace:2015:TMP

[2159] Luigi Pace, Alessandra Salvan, and Nicola Sartori. Theory and methods papers: Inference from pseudolikelihoods with plug-in estimates. Australian & New Zealand Journal of Statistics, 57(3):347–361, September 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Sarnacchiaro:2015:TMP

[2160] Pasquale Sarnacchiaro, Luigi D’Ambra, and Ida Camminatiello. Theory and methods papers: Measures of association and visualization of log odds ratio structure for a two way contingency table. Australian & New Zealand Journal of Statistics, 57(3):363–376, September 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ullah:2015:TMP

[2161] Insha Ullah and Beatrix Jones. Theory and methods papers: Regularised Manova for high-dimensional data. Australian & New Zealand Journal of Statistics, 57(3):377–389, September 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Merola:2015:TMP

[2162] Giovanni Maria Merola. Theory and methods papers: Least squares sparse principal component analysis: a backward elimination approach to attain large loadings. Australian & New Zealand Journal of Statistics, 57(3):391–429, September 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Graham:2015:BRB

[2163] Patrick Graham. Book review: Bayesian methods in health economics. Australian & New Zealand Journal of Statistics, 57(3):431–433, September 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Quinn:2015:BRS

[2164] Barry Quinn. Book review: Stationary Stochastic Processes Theory and Applications. Australian & New Zealand Journal of Statistics, 57(3):433– 434, September 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2015:FPc

[2165] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 57(3):435, September 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Marschner:2015:IPR

[2166] Ian C. Marschner. Invited paper: Relative risk regression for binary outcomes: Methods and recommendations. Australian & New Zealand Journal of Statistics, 57(4):437–462, December 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wong:2015:ASA

[2167] T. S. T. Wong. Applications: Statistical analysis of heat waves in the state of Victoria in Australia. Australian & New Zealand Journal of Statistics, 57(4):463–480, December 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Thas:2015:TMG

[2168] O. Thas, J. C. W. Rayner, and J. De Neve. Theory and methods: Generalised smooth tests of goodness of fit utilising L-moments. Australian & New Zealand Journal of Statistics, 57(4):481–499, December 2015. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Dres:2015:TME

[2169] Konstantine A. Dres, Graham Hepworth, and Ray Watson. Theory and methods: Exact confidence intervals for proportions estimated by group testing with different group sizes. Australian & New Zealand Journal of Statistics, 57(4):501–516, December 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bludowsky:2015:TMO

[2170] Andrea Bludowsky, Joachim Kunert, and John Stufken. Theory and methods: Optimal designs for the carryover model with random interactions between subjects and treatments. Australian & New Zealand Journal of Statistics, 57(4):517–533, December 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Karaman:2015:TMG

[2171] Filiz Karaman and Erhan Cene. Theory and methods: Generalisation of Mood’s test to factorial designs to detect heterogeneity of variances. Australian & New Zealand Journal of Statistics, 57(4):535–544, December 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Li:2015:TME

[2172] Rui Li, Xiaoli Li, and Xian Zhou. Theory and methods: Efficient inference for longitudinal data varying-coefficient regression models. Australian & New Zealand Journal of Statistics, 57(4):545–570, December 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Vignes:2015:BRI

[2173] Matthieu Vignes. Book review: Inference Principles for Biostatisticians. Australian & New Zealand Journal of Statistics, 57(4):571–572, December 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Stoklosa:2015:BRA

[2174] Jakub Stoklosa. Book review: Analysis of Capture–Recapture Data. Australian & New Zealand Journal of Statistics, 57(4):572–574, December 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Prvan:2015:BRE

[2175] Tania Prvan. Book review: Essentials of Multivariate Data Analysis. Australian & New Zealand Journal of Statistics, 57(4):574–575, December 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2015:FPd

[2176] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 57(4):577, December 2015. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Huggins:2016:ANE

[2177] Richard M. Huggins, Paul S. F. Yip, and Jakub Stoklosa. Applications: Nonparametric estimation of the number of drug users in Hong Kong using repeated multiple lists. Australian & New Zealand Journal of Statistics, 58(1):1–13, March 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Sengupta:2016:API

[2178] Aritra Sengupta, Noel Cressie, Brian H. Kahn, and Richard Frey. Applications: Predictive inference for big, spatial, non-Gaussian data: MODIS cloud data and its change-of-support. Australian & New Zealand Journal of Statistics, 58(1):15–45, March 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Han:2016:TMC

[2179] Zifei Han and Victor De Oliveira. Theory and methods: On the correlation structure of Gaussian copula models for geostatistical count data. Australian & New Zealand Journal of Statistics, 58(1):47–69, March 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Khan:2016:TMS

[2180] Atikur R. Khan and D. S. Poskitt. Theory and methods: Signal identification in singular spectrum analysis. Australian & New Zealand Journal of Statistics, 58(1):71–98, March 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Garthwaite:2016:TME

[2181] Paul H. Garthwaite and Inge Koch. Theory and methods: Evaluating the contributions of individual variables to a quadratic form. Australian & New Zealand Journal of Statistics, 58(1):99–119, March 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Turkmen:2016:TMG

[2182] A. S. Turkmen and O. Ozturk. Theory and methods: Generalised rank regression estimator with standard error adjusted lasso. Australian & New Zealand Journal of Statistics, 58(1):121–135, March 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Meyer:2016:BRD

[2183] Renate Meyer. Book review: Dependence Modelling with Copulas. Australian & New Zealand Journal of Statistics, 58(1):137–138, March 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wang:2016:BRU

[2184] Kevin Wang. Book review: Using R for Introductory Statistics. Australian & New Zealand Journal of Statistics, 58(1):138–139, March 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Miyano:2016:BRC

[2185] Hisao Miyano. Book review: Correspondence Analysis: Theory, Practice and New Strategies. Australian & New Zealand Journal of Statistics, 58 (1):139–141, March 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Liu:2016:BRA

[2186] Ivy Liu. Book review: Analysis of Categorical Data with R. Australian & New Zealand Journal of Statistics, 58(1):141–142, March 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2016:FPa

[2187] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 58(1):143, March 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2016:ACA

[2188] Anonymous. Awards: Campbell Award. Australian & New Zealand Journal of Statistics, 58(2):145–148, June 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2016:APM

[2189] Anonymous. Awards: Pitman Medal. Australian & New Zealand Journal of Statistics, 58(2):149–152, June 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Preedalikit:2016:AJM

[2190] Kemmawadee Preedalikit, Ivy Liu, Yuichi Hirose, Nokuthaba Sibanda, and Daniel Fernández. Applications: Joint modeling of survival and longitudinal ordered data using a semiparametric approach. Australian & New Zealand Journal of Statistics, 58(2):153–172, June 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bryant:2016:ACF

[2191] Fred B. Bryant and Karl G. Jöreskog. Applications: Confirmatory factor analysis of ordinal data using full-information adaptive quadrature. Australian & New Zealand Journal of Statistics, 58(2):173–196, June 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Zhu:2016:ADT

[2192] Hong Zhu, Yingchao Xie, and Maochao Xu. Applications: Discrete truncated power-law distributions. Australian & New Zealand Journal of Statistics, 58(2):197–209, June 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wang:2016:TMB

[2193] You-Gan Wang, Yudong Zhao, and Liya Fu. Theory and methods: The Buckley–James estimator and induced smoothing. Australian & New Zealand Journal of Statistics, 58(2):211–225, June 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Zhu:2016:TMB

[2194] Weixuan Zhu, J. Miguel Marin, and Fabrizio Leisen. Theory and methods: a bootstrap likelihood approach to Bayesian computation. Australian & New Zealand Journal of Statistics, 58(2):227–244, June 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Rohart:2016:TMM

[2195] Florian Rohart. Theory and methods: Multiple hypothesis testing for variable selection. Australian & New Zealand Journal of Statistics, 58(2): 245–267, June 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ramirez-Aldana:2016:TML

[2196] Ricardo Ramírez-Aldana and Guillermina Eslava-Gómez. Theory and methods: Label- and level-invariant graphical log-linear models. Australian & New Zealand Journal of Statistics, 58(2):269–291, June 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kabaila:2016:TML

[2197] Paul Kabaila and Yining Kong. Theory and methods: Lower bounds on integrated risk, subject to inequality constraints. Australian & New Zealand Journal of Statistics, 58(2):293–315, June 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2016:FPb

[2198] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 58(2):317, June 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Sabo:2016:APC

[2199] R. T. Sabo. Applications: Providing consulting experiences through role playing in a graduate statistics course. Australian & New Zealand Journal of Statistics, 58(3):319–333, September 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pang:2016:TMR

[2200] Zhen Pang, Bingqing Lin, and Jiming Jiang. Theory and methods: Regularisation parameter selection via bootstrapping. Australian & New Zealand Journal of Statistics, 58(3):335–356, September 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pham:2016:TME

[2201] Hoa Pham and Alan Branford. Theory and methods: Exploring parameter relations for multi-stage models in stage-wise constant and time dependent hazard rates. Australian & New Zealand Journal of Statistics, 58(3):357–376, September 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Li:2016:TMM

[2202] Haocheng Li and Grace Y. Yi. Theory and methods: Missing data mechanisms for analysing longitudinal data with incomplete observations in both responses and covariates. Australian & New Zealand Journal of Statistics, 58(3):377–396, September 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Sutradhar:2016:TMI

[2203] Brajendra C. Sutradhar, K. V. Vineetha Warriyar, and Nan Zheng. Theory and methods: Inference in semi-parametric dynamic models for repeated count data. Australian & New Zealand Journal of Statistics, 58 (3):397–434, September 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2016:FPc

[2204] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 58(3):435, September 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Fernandez:2016:AMS

[2205] D. Fernández and R. Arnold. Applications: Model selection for mixturebased clustering for ordinal data. Australian & New Zealand Journal of Statistics, 58(4):437–472, December 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ji:2016:ASM

[2206] Tieming Ji and Jie Chen. Applications: Statistical models for DNA copy number variation detection using read-depth data from next generation sequencing experiments. Australian & New Zealand Journal of Statistics, 58(4):473–491, December 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Thomson:2016:TME

[2207] T. Thomson, S. Hossain, and M. Ghahramani. Theory and methods: Efficient estimation for time series following generalized linear models.

Australian & New Zealand Journal of Statistics, 58(4):493–513, December 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Prendergast:2016:TMR

[2208] L. A. Prendergast and A. L. Garnham. Theory and methods: Response and predictor folding to counter symmetric dependency in dimension reduction. Australian & New Zealand Journal of Statistics, 58(4):515– 532, December 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ross:2016:BR

[2209] Joshua V. Ross. Book review. Australian & New Zealand Journal of Statistics, 58(4):533–534, December 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2016:FPd

[2210] Anonymous. Forthcoming papers. Australian & New Zealand Journal of Statistics, 58(4):535, December 2016. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chipperfield:2017:AAC

[2211] James Chipperfield, James J. Brown, and Nicole Watson. Applications: The Australian Census Longitudinal Dataset: using record linkage to create a longitudinal sample from a series of cross-sections. Australian & New Zealand Journal of Statistics, 59(1):1–16, March 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Elsawah:2017:APE

[2212] A. M. Elsawah. Applications: a powerful and efficient algorithm for breaking the links between aliased effects in asymmetric designs. Australian & New Zealand Journal of Statistics, 59(1):17–41, March 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Zhang:2017:TMB

[2213] Jin Zhang. Theory and methods: Bayesian (mean) most powerful tests. Australian & New Zealand Journal of Statistics, 59(1):43–56, March 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anderson:2017:TMS

[2214] Marti J. Anderson, Daniel C. I. Walsh, K. Robert Clarke, Ray N. Gorley, and Edlin Guerra-Castro. Theory and methods: Some solutions to the multivariate Behrens–Fisher problem for dissimilarity-based analyses. Australian & New Zealand Journal of Statistics, 59(1):57–79, March 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Yang:2017:TMO

[2215] Po Yang and Chang-Yun Lin. Theory and methods: Optimal split-plot orthogonal arrays. Australian & New Zealand Journal of Statistics, 59(1): 81–94, March 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Killiches:2017:TME

[2216] Matthias Killiches, Daniel Kraus, and Claudia Czado. Theory and methods: Examination and visualisation of the simplifying assumption for vine copulas in three dimensions. Australian & New Zealand Journal of Statistics, 59(1):95–117, March 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Rumcheva:2017:TMI

[2217] Pavlina Rumcheva and Brett Presnell. Theory and methods: an improved test of equality of mean directions for the Langevin–von Mises– Fisher distribution. Australian & New Zealand Journal of Statistics, 59 (1):119–135, March 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Borges:2017:TMG

[2218] Patrick Borges, Marcelo Bourguignon, and Fabio Fajardo Molinares. Theory and methods: a generalised NGINAR(1) process with inflatedparameter geometric counting series. Australian & New Zealand Journal of Statistics, 59(1):137–150, March 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Stoklosa:2017:BR

[2219] Jakub Stoklosa. Book review. Australian & New Zealand Journal of Statistics, 59(1):151–152, March 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2017:FPF

[2220] Anonymous. Forthcoming papers: Forthcoming papers. Australian & New Zealand Journal of Statistics, 59(1):153, March 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bersimis:2017:AUL

[2221] S. Bersimis and P. Economou. Applications: The use of length-biased distributions in statistical monitoring. Australian & New Zealand Journal of Statistics, 59(2):155–167, June 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

DAmbra:2017:TMA

[2222] A. D’Ambra, A. Crisci, and L. D’Ambra. Theory and methods: Association models in the weighted log ratio analysis for rates. Australian & New Zealand Journal of Statistics, 59(2):169–185, June 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lumley:2017:TMP

[2223] Thomas Lumley. Theory and methods: Pseudo-R^{2} statistics under complex sampling. Australian & New Zealand Journal of Statistics, 59(2): 187–194, June 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Huang:2017:TMG

[2224] A. Huang. Theory and methods: On generalised estimating equations for vector regression. Australian & New Zealand Journal of Statistics, 59(2): 195–213, June 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wang:2017:TME

[2225] Hongxia Wang, Yuehua Wu, and Elton Chan. Theory and methods: Efficient estimation of nonparametric spatial models with general correlation structures. Australian & New Zealand Journal of Statistics, 59(2): 215–233, June 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Vignes:2017:BR

[2226] Matthieu Vignes. Book review. Australian & New Zealand Journal of Statistics, 59(2):235–236, June 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Beath:2017:BR

[2227] Ken J. Beath. Book review. Australian & New Zealand Journal of Statistics, 59(2):236–238, June 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kateri:2017:TMG

[2228] Maria Kateri, Anna Gottard, and Claudia Tarantola. Theory and methods: Generalised quasi-symmetry models for ordinal contingency tables. Australian & New Zealand Journal of Statistics, 59(3):239–253, September 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Mandal:2017:TMO

[2229] S. Mandal, B. Torsney, and M. Chowdhury. Theory and methods: Optimal designs for minimising covariances among parameter estimators in a linear model. Australian & New Zealand Journal of Statistics, 59(3):255– 273, September 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lloyd:2017:TMS

[2230] Chris J. Lloyd. Theory and methods: The size accuracy of combination tests. Australian & New Zealand Journal of Statistics, 59(3):275–287, September 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

He:2017:TMJ

[2231] Yong He, Xinsheng Zhang, Jiadong Ji, and Bin Liu. Theory and methods: Joint estimation of multiple high-dimensional Gaussian copula graphical models. Australian & New Zealand Journal of Statistics, 59(3):289–310, September 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Muggeo:2017:TMI

[2232] Vito M. R. Muggeo. Theory and methods: Interval estimation for the breakpoint in segmented regression: a smoothed score-based approach. Australian & New Zealand Journal of Statistics, 59(3):311–322, September 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Visscher:2017:BR

[2233] Peter M. Visscher. Book review. Australian & New Zealand Journal of Statistics, 59(3):323–324, September 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Schuckers:2017:BR

[2234] Michael E. Schuckers. Book reviews. Australian & New Zealand Journal of Statistics, 59(3):324–325, September 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Brien:2017:IRM

[2235] C. J. Brien. Invited review: Multiphase experiments in practice: A look back. Australian & New Zealand Journal of Statistics, 59(4):327– 352, December 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Manderson:2017:SCM

[2236] A. A. Manderson, E. Cripps, K. Murray, and B. A. Turlach. Statistical computing: Monotone polynomials using BUGS and Stan. Australian & New Zealand Journal of Statistics, 59(4):353–370, December 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Moltchanova:2017:ACA

[2237] Elena Moltchanova, Daniel Gerhard, Fathimath Mohamed, Sally Gaw, and Chris N. Glover. Applications: On correlation analysis of many-to-many observations: an alternative to Pearson’s correlation coefficient and its application to an ecotoxicological study. Australian & New Zealand Journal of Statistics, 59(4):371–387, December 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Coretto:2017:ANE

[2238] Pietro Coretto and Francesco Giordano. Applications: Nonparametric estimation of the dynamic range of music signals. Australian & New Zealand Journal of Statistics, 59(4):389–412, December 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bowden:2017:AUM

[2239] Ross S. Bowden and Brenton R. Clarke. Applications: Using multivariate time series methods to estimate location and climate change effects on temperature readings employed in electricity demand simulation. Australian & New Zealand Journal of Statistics, 59(4):413–431, December 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Diffey:2017:TMN

[2240] S. M. Diffey, A. B. Smith, A. H. Welsh, and B. R. Cullis. Theory and methods: a new REML (parameter expanded) EM algorithm for linear mixed models. Australian & New Zealand Journal of Statistics, 59(4): 433–448, December 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ullah:2017:TMI

[2241] Insha Ullah, Matthew D. M. Pawley, Adam N. H. Smith, and Beatrix Jones. Theory and methods: Improving the detection of unusual observations in high-dimensional settings. Australian & New Zealand Journal of Statistics, 59(4):449–462, December 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Suesse:2017:TMA

[2242] Thomas Suesse, John C. W. Rayner, and Olivier Thas. Theory and methods: Assessing the fit of finite mixture distributions. Australian & New Zealand Journal of Statistics, 59(4):463–483, December 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Rayner:2017:TMU

[2243] J. C. W. Rayner and D. J. Best. Theory and methods: Unconditional analogues of Cochran–Mantel–Haenszel tests. Australian & New Zealand Journal of Statistics, 59(4):485–494, December 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hazelton:2017:TMP

[2244] Martin L. Hazelton and Timothy P. Bilton. Theory and methods: Polytope samplers for network tomography. Australian & New Zealand Journal of Statistics, 59(4):495–511, December 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Clarke:2017:TMI

[2245] Brenton R. Clarke, Alexandra Höller, Christine H. Müller, and Karuru Wamahiu. Theory and methods: Investigation of the performance of trimmed estimators of life time distributions with censoring. Australian & New Zealand Journal of Statistics, 59(4):513–525, December 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lim:2017:TMC

[2246] Yaeji Lim and Hee-Seok Oh. Theory and methods: Confidence intervals for nonparametric quantile regression: an emphasis on smoothing splines approach. Australian & New Zealand Journal of Statistics, 59(4):527– 543, December 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Vignes:2017:BRB

[2247] Matthieu Vignes, Mélodie Sammarro, Anthime Schrefheere, Alex White, and Stephen Marsland. Book review: Bayesian Networks: With Examples in R. Australian & New Zealand Journal of Statistics, 59(4):545–546, December 2017. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2018:IIa

[2248] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 60(1):i–iv, March 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Martin:2018:P

[2249] Michael A. Martin. Preface. Australian & New Zealand Journal of Statistics, 60(1):1–3, March 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Aoshima:2018:SHD

[2250] Makoto Aoshima, Dan Shen, Haipeng Shen, Kazuyoshi Yata, Yi-Hui Zhou, and J. S. Marron. A survey of high dimension low sample size asymptotics. Australian & New Zealand Journal of Statistics, 60(1):4– 19, March 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cheng:2018:GFR

[2251] Ming-Yen Cheng, Sanying Feng, Gaorong Li, and Heng Lian. Greedy forward regression for variable screening. Australian & New Zealand Journal of Statistics, 60(1):20–42, March 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hua:2018:AKB

[2252] Hairui Hua, Dimitrios Bagkavos, and Prakash N. Patil. An L_{1} analysis of a kernel-based hazard rate estimator. Australian & New Zealand Journal of Statistics, 60(1):43–64, March 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Johnstone:2018:TSW

[2253] Iain M. Johnstone. Tail sums of Wishart and Gaussian eigenvalues beyond the bulk edge. Australian & New Zealand Journal of Statistics, 60 (1):65–74, March 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kim:2018:EPG

[2254] Andy S. I. Kim and Matt P. Wand. On expectation propagation for generalised, linear and mixed models. Australian & New Zealand Journal of Statistics, 60(1):75–102, March 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kuffner:2018:CHB

[2255] Todd A. Kuffner, Stephen M. S. Lee, and G. A. Young. Consistency of a hybrid block bootstrap for distribution and variance estimation for sample quantiles of weakly dependent sequences. Australian & New Zealand Journal of Statistics, 60(1):103–114, March 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lee:2018:KEM

[2256] Seonmi Lee, Woncheol Jang, and Byeong U. Park. Kernel excess mass test for multimodality. Australian & New Zealand Journal of Statistics, 60(1):115–131, March 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Samonenko:2018:NPT

[2257] I. Samonenko and J. Robinson. A new permutation test statistic for K-sample multivariate designs. Australian & New Zealand Journal of Statistics, 60(1):132–139, March 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Tajvidi:2018:GAG

[2258] N. Tajvidi and B. A. Turlach. A general approach to generate random variates for multivariate copulae. Australian & New Zealand Journal of Statistics, 60(1):140–155, March 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2018:IIb

[2259] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 60(2):i–iv, June 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Boys:2018:BIP

[2260] Richard J. Boys, Holly F. Ainsworth, and Colin S. Gillespie. Bayesian inference for a partially observed birth-death process using data on proportions. Australian & New Zealand Journal of Statistics, 60(2):157–173, June 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Manderson:2018:DBF

[2261] A. A. Manderson, K. Murray, and B. A. Turlach. Dynamic Bayesian forecasting of AFL match results using the Skellam distribution. Australian & New Zealand Journal of Statistics, 60(2):174–187, June 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Yang:2018:ORP

[2262] Po Yang and Jingjing Hu. Optimal robust parameter designs with qualitative and quantitative factors. Australian & New Zealand Journal of Statistics, 60(2):188–198, June 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Godolphin:2018:NRP

[2263] Janet D. Godolphin. A note on the robustness of PBIBD(2)s against breakdown in the event of observation loss. Australian & New Zealand Journal of Statistics, 60(2):199–208, June 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Schumacher:2018:IDC

[2264] Fernanda L. Schumacher, Victor H. Lachos, Filidor E. Vilca-Labra, and Luis M. Castro. Influence diagnostics for censored regression models with autoregressive errors. Australian & New Zealand Journal of Statistics, 60(2):209–229, June 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hattab:2018:LFT

[2265] Mohammad W. Hattab and Ronald Christensen. Lack of fit tests based on sums of ordered residuals for linear models. Australian & New Zealand Journal of Statistics, 60(2):230–257, June 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Liu:2018:FAU

[2266] Yu Liu and Yong Wang. A fast algorithm for univariate log-concave density estimation. Australian & New Zealand Journal of Statistics, 60 (2):258–275, June 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Curran:2018:BRF

[2267] J. M. Curran. Book review: Foundations of Statistical Algorithms With References to R Packages, By C. Weihs, O. Mersmann and U. Ligges. Boca Raton, Florida: CRC Press. 2015. 500 pages. UK£44.99 (hardback). ISBN 978-1-4398-7885-9. Australian & New Zealand Journal of Statistics, 60(2):276–277, June 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2018:IIc

[2268] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 60(3):i–iv, September 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pham:2018:GAM

[2269] Tung H. Pham and Matt P. Wand. Generalised additive mixed models analysis via gammSlice. Australian & New Zealand Journal of Statistics, 60(3):279–300, September 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Orfanogiannaki:2018:MPH

[2270] K. Orfanogiannaki and D. Karlis. Multivariate Poisson hidden Markov models with a case study of modelling seismicity. Australian & New Zealand Journal of Statistics, 60(3):301–322, September 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Shokoohi:2018:SPS

[2271] Farhad Shokoohi and Mahmoud Torabi. Semi-parametric small-area estimation by combining time-series and cross-sectional data methods. Australian & New Zealand Journal of Statistics, 60(3):323–342, September 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Zheng:2018:GQL

[2272] Nan Zheng and Brajendra C. Sutradhar. Generalised quasi-likelihood inference in a semi-parametric binary dynamic mixed logit model. Australian & New Zealand Journal of Statistics, 60(3):343–373, September 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gong:2018:NKE

[2273] Xiaodong Gong and Jiti Gao. Nonparametric kernel estimation of the impact of tax policy on the demand for private health insurance in Australia. Australian & New Zealand Journal of Statistics, 60(3):374–393, September 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Li:2018:HPL

[2274] Haocheng Li and Yukun Zhang. Hybrid pairwise-likelihood estimation methods for incomplete longitudinal binary data. Australian & New Zealand Journal of Statistics, 60(3):394–404, September 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chambers:2018:SAC

[2275] Ray Chambers. Book review: Statistics in Action: A Canadian Outlook, Edited by J. F. Lawless. Boca Raton, Florida: CRC Press 2014 384 pages. $185.00 (hardback). ISBN 978-1-4822-3623-1. Australian & New Zealand Journal of Statistics, 60(3):405, September 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bulger:2018:MDA

[2276] David Bulger. Book review: Music Data Analysis: Foundations and Applications, Edited By C. Weihs D. Jannach I. Vatolkin and G. Rudolph. Boca Raton, Florida: CRC Press. 2017. 675 pages. AUD$94.99 (hardback). ISBN 978-1-4987-1956-8. Australian & New Zealand Journal of Statistics, 60(3):405–406, September 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2018:IId

[2277] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 60(4):i–iv, December 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wang:2018:SMA

[2278] Jingli Wang and Jialiang Li. Semiparametric model averaging prediction: a Bayesian approach. Australian & New Zealand Journal of Statistics, 60(4):407–422, December 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kadhem:2018:UHM

[2279] Safaa. K. Kadhem, Paul Hewson, and Irene Kaimi. Using hidden Markov models to model spatial dependence in a network. Australian & New Zealand Journal of Statistics, 60(4):423–446, December 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Zhang:2018:DAT

[2280] Mingjuan Zhang, Cheng Zhou, Yong He, and Bin Liu. Data-adaptive test for high-dimensional multivariate analysis of variance problem. Australian & New Zealand Journal of Statistics, 60(4):447–470, December 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Karunanayaka:2018:ECO

[2281] Ruwan C. Karunanayaka and Boxin Tang. On the existence and constructions of orthogonal designs. Australian & New Zealand Journal of Statistics, 60(4):471–480, December 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ghosh:2018:VCE

[2282] Subir Ghosh, Li Guo, and Luyao Peng. Variance component estimators OPE, NOPE and AOPE in linear mixed effects models. Australian & New Zealand Journal of Statistics, 60(4):481–505, December 2018. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2019:IIa

[2283] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 61(1):i–iv, March 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

OGorman:2019:CNC

[2284] Thomas W. O’Gorman. Constructing narrower confidence intervals by inverting adaptive tests. Australian & New Zealand Journal of Statistics, 61(1):1–17, March 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kabaila:2019:CIC

[2285] Paul Kabaila and Christeen Wijethunga. Confidence intervals centred on bootstrap smoothed estimators. Australian & New Zealand Journal of Statistics, 61(1):19–38, March 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Verbyla:2019:NMS

[2286] Arunas Petras Verbyla. A note on model selection using information criteria for general linear models estimated using REML. Australian & New Zealand Journal of Statistics, 61(1):39–50, March 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hepworth:2019:BCE

[2287] Graham Hepworth. Bias correction of estimated proportions using inverse binomial group testing. Australian & New Zealand Journal of Statistics, 61(1):51–60, March 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hui:2019:TRE

[2288] F. K. C. Hui, Samuel Müller, and A. H. Welsh. Testing random effects in linear mixed models: another look at the F -test (with discussion). Australian & New Zealand Journal of Statistics, 61(1):61–84, March 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Doppelt:2019:PST

[2289] Ross Doppelt and Keith O’Hara. Posterior sampling in two classes of multivariate fractionally integrated models: corrigendum to Ravishanker, N. and B. K. Ray (1997) Australian Journal of Statistics 39 (3), 295– 311. Australian & New Zealand Journal of Statistics, 61(1):85–87, March 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic). See [1340].

Anonymous:2019:IIb

[2290] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 61(2):i–iv, June 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Peyrard:2019:EAI

[2291] N. Peyrard, M.-J. Cros, S. de Givry, A. Franc, S. Robin, R. Sabbadin, T. Schiex, and M. Vignes. Exact or approximate inference in graphical models: why the choice is dictated by the treewidth, and how variable elimination can be exploited. Australian & New Zealand Journal of Statistics, 61(2):89–133, June 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Tian:2019:LRA

[2292] Guo-Liang Tian, Yin Liu, and Man-Lai Tang. Logistic regression analysis of non-randomized response data collected by the parallel model in sensitive surveys. Australian & New Zealand Journal of Statistics, 61(2): 134–151, June 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

McElroy:2019:MEE

[2293] Tucker McElroy and Richard Penny. Maximum entropy extreme-value seasonal adjustment. Australian & New Zealand Journal of Statistics, 61 (2):152–174, June 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

McGrory:2019:CRS

[2294] Clare A. McGrory, Daniel C. Ahfock, and Ricardo T. Lemos. Climate regime shift detection with a trans-dimensional, sequential Monte Carlo, variational Bayes method. Australian & New Zealand Journal of Statistics, 61(2):175–188, June 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gelling:2019:RPR

[2295] Nicholas Gelling, Matthew R. Schofield, and Richard J. Barker. R package rjmcmc: reversible jump MCMC using post-processing. Australian & New Zealand Journal of Statistics, 61(2):189–212, June 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Beesley:2019:SIM

[2296] Lauren J. Beesley, Jeremy M. G. Taylor, and Roderick J. A. Little. Sequential imputation for models with latent variables assuming latent ignorability. Australian & New Zealand Journal of Statistics, 61(2):213– 233, June 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Brien:2019:MEL

[2297] C. J. Brien. Multiphase experiments with at least one later laboratory phase. II. Nonorthogonal designs. Australian & New Zealand Journal of Statistics, 61(2):234–268, June 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Downes:2019:BR

[2298] Stephanie M. Downes. Book reviews. Australian & New Zealand Journal of Statistics, 61(2):269–270, June 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Parry:2019:BRI

[2299] Matthew Parry. Book review: Introduction to Bayesian Statistics, Third Edition. Australian & New Zealand Journal of Statistics, 61(2):271, June 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2019:IIc

[2300] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 61(3):i–iv, September 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Trevenen:2019:UHM

[2301] Michelle L. Trevenen, Berwin A. Turlach, Peter R. Eastwood, Leon M. Straker, and Kevin Murray. Using hidden Markov models with raw, triaxial wrist accelerometry data to determine sleep stages. Australian & New Zealand Journal of Statistics, 61(3):273–298, September 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Dong:2019:SES

[2302] Chaohua Dong, Jiti Gao, and Bin Peng. Series estimation for singleindex models under constraints. Australian & New Zealand Journal of Statistics, 61(3):299–335, September 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Papageorgiou:2019:BDR

[2303] Georgios Papageorgiou. Bayesian density regression for discrete outcomes. Australian & New Zealand Journal of Statistics, 61(3):336–359, September 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ramirez-Hassan:2019:FEN

[2304] Andrés Ramírez-Hassan and Manuel Correa-Giraldo. Focused estimation for noisy and small data sets: a Bayesian minimum expected loss estimator approach. Australian & New Zealand Journal of Statistics, 61 (3):360–379, September 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wickramarachchi:2019:RFS

[2305] D. C. Wickramarachchi, B. L. Robertson, M. Reale, C. J. Price, and J. A. Brown. A reflected feature space for CART. Australian & New Zealand Journal of Statistics, 61(3):380–391, September 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Welsh:2019:BR

[2306] A. H. Welsh. Book review. Australian & New Zealand Journal of Statistics, 61(3):392–395, September 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2019:IId

[2307] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 61(4):i–iv, December 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Millar:2019:RFR

[2308] Russell Millar. Rachel Fewster: Recipient of NZSA Campbell Award 2018. Australian & New Zealand Journal of Statistics, 61(4):397–400, December 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Koch:2019:KNB

[2309] Inge Koch, Kanta Naito, and Hiroaki Tanaka. Kernel naive Bayes discrimination for high-dimensional pattern recognition. Australian & New Zealand Journal of Statistics, 61(4):401–428, December 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kabaila:2019:ACI

[2310] Paul Kabaila and Rheanna Mainzer. An alternative to confidence intervals constructed after a Hausman pretest in panel data. Australian & New Zealand Journal of Statistics, 61(4):429–444, December 2019. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wang:2019:FAE

[2311] Kevin Yx Wang, Garth Tarr, Jean Yh Yang, and Samuel Mueller. Fast and approximate exhaustive variable selection for generalised linear models with APES. Australian & New Zealand Journal of Statistics, 61(4): 445–465, December 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Williams:2019:EVB

[2312] E. R. Williams and H. P. Piepho. Error variance bias in neighbour balance and evenness of distribution designs. Australian & New Zealand Journal of Statistics, 61(4):466–473, December 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Maindonald:2019:BR

[2313] John Maindonald. Book review. Australian & New Zealand Journal of Statistics, 61(4):474–476, December 2019. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2020:IIa

[2314] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 62(1):i–iv, March 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2020:PM

[2315] Anonymous. Pitman Medal 2016 and 2018. Australian & New Zealand Journal of Statistics, 62(1):1–4, March 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Booth:2020:GRE

[2316] James G. Booth and Alan H. Welsh. Generalised regression estimation via the bootstrap. Australian & New Zealand Journal of Statistics, 62 (1):5–24, March 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bakouch:2020:NSB

[2317] Hassan S. Bakouch, Y. Sunecher, N. Mamode Khan, and V. Jowaheer. A non-stationary bivariate INAR(1) process with a simple crossdependence: Estimation with some properties. Australian & New Zealand Journal of Statistics, 62(1):25–48, March 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Clark:2020:MLE

[2318] Robert Graham Clark. Maximum likelihood estimation for outcomedependent samples. Australian & New Zealand Journal of Statistics, 62 (1):49–70, March 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gunawan:2020:BWI

[2319] David Gunawan, Anastasios Panagiotelis, William Griffiths, and Duangkamon Chotikapanich. Bayesian weighted inference from surveys. Australian & New Zealand Journal of Statistics, 62(1):71–94, March 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Karavarsamis:2020:STT

[2320] N. Karavarsamis, G. Guillera-Arroita, R. M. Huggins, and B. J. T. Morgan. The score test for the two-sample occupancy model. Australian & New Zealand Journal of Statistics, 62(1):95–115, March 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Yee:2020:VPN

[2321] Thomas W. Yee. The VGAM package for negative binomial regression. Australian & New Zealand Journal of Statistics, 62(1):116–131, March 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2020:IIb

[2322] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 62(2):i–iv, June 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Walter:2020:DCT

[2323] Scott R. Walter, Bruce M. Brown, and William T. M. Dunsmuir. Detecting changes in task length due to task-switching in the presence of repeated length-biased sampling. Australian & New Zealand Journal of Statistics, 62(2):133–152, June 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Elliott:2020:MTT

[2324] Tom Elliott and Thomas Lumley. Modelling the travel time of transit vehicles in real-time through a GTFS-based road network using GPS vehicle locations. Australian & New Zealand Journal of Statistics, 62(2): 153–167, June 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Forbes:2020:SMT

[2325] Jeremy Forbes, Dianne Cook, and Rob J. Hyndman. Spatial modelling of the two-party preferred vote in Australian federal elections: 2001–2016. Australian & New Zealand Journal of Statistics, 62(2):168–185, June 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Vernic:2020:CBM

[2326] Raluca Vernic. On a class of bivariate mixed Sarmanov distributions. Australian & New Zealand Journal of Statistics, 62(2):186–211, June 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Neuhaus:2020:REL

[2327] John M. Neuhaus and Charles E. McCulloch. Robust estimation for longitudinal data under outcome-dependent visit processes. Australian & New Zealand Journal of Statistics, 62(2):212–231, June 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Shi:2020:MNL

[2328] Jianhong Shi, Jie Zhang, Xiaorui Wang, and Weixing Song. Modal nonlinear regression in the presence of Laplace measurement error. Australian & New Zealand Journal of Statistics, 62(2):232–248, June 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Tsagris:2020:FMC

[2329] Michail Tsagris and Connie Stewart. A folded model for compositional data analysis. Australian & New Zealand Journal of Statistics, 62(2): 249–277, June 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Wang:2020:VSF

[2330] Xinyang Wang. Variable selection for first-order Poisson integer-valued autoregressive model with covariables. Australian & New Zealand Journal of Statistics, 62(2):278–295, June 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2020:IIc

[2331] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 62(3):i–iv, September 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Lu:2020:SVE

[2332] Renjie Lu, Philip L. H. Yu, and Xiaohang Wang. Sparse vector error correction models with application to cointegration-based trading. Australian & New Zealand Journal of Statistics, 62(3):297–321, September 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gamage:2020:ISM

[2333] Ramadha D. Piyadi Gamage and Wei Ning. Inference for short-memory time series models based on modified empirical likelihood. Australian & New Zealand Journal of Statistics, 62(3):322–339, September 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Kurosawa:2020:GFM

[2334] Takeshi Kurosawa, Francis K. C. Hui, A. H. Welsh, Kousuke Shinmura, and Nobuoki Eshima. On goodness-of-fit measures for Poisson regression models. Australian & New Zealand Journal of Statistics, 62(3):340–366, September 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Tsai:2020:ATS

[2335] Shin-Fu Tsai. Approximate two-sided tolerance intervals for normal mixture distributions. Australian & New Zealand Journal of Statistics, 62 (3):367–382, September 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Reddy:2020:SRP

[2336] K. G. Reddy and M. G. M. Khan. stratifyR: an R package for optimal stratification and sample allocation for univariate populations. Australian & New Zealand Journal of Statistics, 62(3):383–405, September 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2020:IId

[2337] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 62(4):i–ii, December 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Alomair:2020:PPT

[2338] Mohammed A. Alomair, Stelios D. Georgiou, and Manohar Aggarwal. Projection properties of three-level screening designs. Australian & New Zealand Journal of Statistics, 62(4):407–425, December 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Bilton:2020:RTP

[2339] Penelope Bilton, Geoff Jones, Siva Ganesh, and Stephen Haslett. Regression trees for poverty mapping. Australian & New Zealand Journal of Statistics, 62(4):426–443, December 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hinoveanu:2020:LBP

[2340] Laurentiu Catalin Hinoveanu, Fabrizio Leisen, and Cristiano Villa. A loss-based prior for Gaussian graphical models. Australian & New Zealand Journal of Statistics, 62(4):444–466, December 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Li:2020:EEV

[2341] Zhijian Li and Wei Lin. Efficient error variance estimation in nonparametric regression. Australian & New Zealand Journal of Statistics, 62(4):467–484, December 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pandhare:2020:FIC

[2342] S. C. Pandhare and T. V. Ramanathan. The focussed information criterion for generalised linear regression models for time series. Australian & New Zealand Journal of Statistics, 62(4):485–507, December 2020. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Sengupta:2020:EPM

[2343] Debjit Sengupta, Tathagata Banerjee, and Surupa Roy. Estimation of Poisson mean with under-reported counts: a double sampling approach. Australian & New Zealand Journal of Statistics, 62(4):508–535, December 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Assmann:2020:BEM

[2344] Christian Aßmann and Marcel Preising. Bayesian estimation and model comparison for linear dynamic panel models with missing values. Australian & New Zealand Journal of Statistics, 62(4):536–557, December 2020. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2021:IIa

[2345] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 63(1):i–iv, March 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hazelton:2021:FAB

[2346] Martin L. Hazelton and R. Turner. A Festschrift for Adrian Baddeley. Australian & New Zealand Journal of Statistics, 63(1):1–5, March 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Jensen:2021:SIM

[2347] Eva B. Vedel Jensen. Stereological inference on mean particle shape from vertical sections. Australian & New Zealand Journal of Statistics, 63(1): 6–18, March 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Stoyan:2021:DRM

[2348] Dietrich Stoyan, Viktor Benes, and Filip Seitl. Dependent radius marks of Laguerre tessellations: a case study. Australian & New Zealand Journal of Statistics, 63(1):19–32, March 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Christoffersen:2021:MCP

[2349] Andreas Dyreborg Christoffersen, Jesper Møller, and Heidi Søgaard Christensen. Modelling columnarity of pyramidal cells in the human cerebral cortex. Australian & New Zealand Journal of Statistics, 63(1): 33–54, March 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cascos:2021:DOS

[2350] Ignacio Cascos, Qiyu Li, and Ilya Molchanov. Depth and outliers for samples of sets and random sets distributions. Australian & New Zealand Journal of Statistics, 63(1):55–82, March 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Diggle:2021:CIP

[2351] Peter J. Diggle. Conditional intensity: a powerful tool for modelling and analysing point process data. Australian & New Zealand Journal of Statistics, 63(1):83–92, March 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Shaw:2021:GIR

[2352] Thomas Shaw, Jesper Møller, and Rasmus Plenge Waagepetersen. Globally intensity-reweighted estimators for K — and pair correlation functions. Australian & New Zealand Journal of Statistics, 63(1):93–118, March 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Choiruddin:2021:ICI

[2353] Achmad Choiruddin, Jean-François Coeurjolly, and Rasmus Waagepetersen. Information criteria for inhomogeneous spatial point processes. Australian & New Zealand Journal of Statistics, 63(1):119–143, March 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Renner:2021:WES

[2354] Ian W. Renner, David I. Warton, and Francis K. C. Hui. What is the effective sample size of a spatial point process? Australian & New Zealand Journal of Statistics, 63(1):144–158, March 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

vanLieshout:2021:IAA

[2355] M. N. M. van Lieshout. Infill asymptotics for adaptive kernel estimators of spatial intensity. Australian & New Zealand Journal of Statistics, 63 (1):159–181, March 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Cressie:2021:FSP

[2356] Noel Cressie. A few statistical principles for data science. Australian & New Zealand Journal of Statistics, 63(1):182–200, March 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gill:2021:AKT

[2357] R. D. Gill. Anna Karenina and the two envelopes problem. Australian & New Zealand Journal of Statistics, 63(1):201–218, March 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2021:IIb

[2358] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 63(2):i–iv, June 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hui:2021:SPM

[2359] Francis K. C. Hui and Howard D. Bondell. A shared parameter mixture model for longitudinal income data with missing responses and zero rounding. Australian & New Zealand Journal of Statistics, 63(2):221– 240, June 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hyndman:2021:FOA

[2360] Rob J. Hyndman, Yijun Zeng, and Han Lin Shang. Forecasting the oldage dependency ratio to determine a sustainable pension age. Australian & New Zealand Journal of Statistics, 63(2):241–256, June 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Shi:2021:NMM

[2361] Yanlin Shi. New moderation methods of higher school certificate assessments: a case study of the New South Wales practice. Australian & New Zealand Journal of Statistics, 63(2):257–283, June 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Verbyla:2021:MTG

[2362] A. P. Verbyla, J. De Faveri, D. M. Deery, and G. J. Rebetzke. Modelling temporal genetic and spatio-temporal residual effects for highthroughput phenotyping data. Australian & New Zealand Journal of Statistics, 63(2):284–308, June 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Dehghan:2021:NPD

[2363] Sakineh Dehghan and Mohammad Reza Faridrohani. Non-parametric depth-based tests for the multivariate location problem. Australian & New Zealand Journal of Statistics, 63(2):309–330, June 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Dhar:2021:DBG

[2364] Subhra Sankar Dhar and Ujjwal Das. On distance based goodness of fit tests for missing data when missing occurs at random. Australian & New Zealand Journal of Statistics, 63(2):331–356, June 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ejaz:2021:ARA

[2365] Muhammad Ejaz, Chaitanya Joshi, and Stephen Joe. Adversarial risk analysis for first-price sealed-bid auctions. Australian & New Zealand Journal of Statistics, 63(2):357–376, June 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ozturk:2021:MBI

[2366] Omer Ozturk and Konul Bayramoglu Kavlak. Model-based inference using judgement post-stratified samples in finite populations. Australian & New Zealand Journal of Statistics, 63(2):377–393, June 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Long:2021:BDR

[2367] Yuqi Long and Xingzhong Xu. Bayesian decision rules to classification problems. Australian & New Zealand Journal of Statistics, 63(2):394– 415, June 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2021:IIc

[2368] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 63(3):i–iv, September 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Rakshit:2021:VSU

[2369] Suman Rakshit, Greg McSwiggan, Gopalan Nair, and Adrian Baddeley. Variable selection using penalised likelihoods for point patterns on a linear network. Australian & New Zealand Journal of Statistics, 63 (3):417–454, September 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Williams:2021:EDP

[2370] E. R. Williams, C. G. Forde, J. Imaki, and K. Oelkers. Experimental design in practice: The importance of blocking and treatment structures. Australian & New Zealand Journal of Statistics, 63(3):455–467, September 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Spencer:2021:AAA

[2371] Simon E. F. Spencer. Accelerating adaptation in the adaptive Metropolis–Hastings random walk algorithm. Australian & New Zealand Journal of Statistics, 63(3):468–484, September 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Nitithumbundit:2021:EAE

[2372] Thanakorn Nitithumbundit and Jennifer S. K. Chan. ECM algorithm for estimating vector ARMA model with variance gamma distribution and possible unbounded density. Australian & New Zealand Journal of Statistics, 63(3):485–516, September 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Maestrini:2021:IGW

[2373] Luca Maestrini and Matt P. Wand. The Inverse G-Wishart distribution and variational message passing. Australian & New Zealand Journal of Statistics, 63(3):517–541, September 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Arbel:2021:BBN

[2374] J. Arbel, G. Kon Kam King, A. Lijoi, L. Nieto-Barajas, and I. Prünster. BNPdensity: Bayesian nonparametric mixture modelling in R. Australian & New Zealand Journal of Statistics, 63(3):542–564, September 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2021:IId

[2375] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 63(4):i–iv, December 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chandrakar:2021:MVC

[2376] Pranjal Chandrakar and Shubhabrata Das. Measuring the values of cricket players. Australian & New Zealand Journal of Statistics, 63(4): 565–578, December 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Liu:2021:PIG

[2377] Pengyi Liu, Guo-Liang Tian, Kam Chuen Yuen, Chi Zhang, and Man-Lai Tang. Proportional inverse Gaussian distribution: a new tool for analysing continuous proportional data. Australian & New Zealand Journal of Statistics, 63(4):579–605, December 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Novo:2021:FEA

[2378] Silvia Novo, Philippe Vieu, and Germán Aneiros. Fast and efficient algorithms for sparse semiparametric bifunctional regression. Australian & New Zealand Journal of Statistics, 63(4):606–638, December 2021. CO-DEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Pirikahu:2021:BCI

[2379] Sarah Pirikahu, Geoffrey Jones, and Martin L. Hazelton. Bayesian credible intervals for population attributable risk from case-control, cohort and cross-sectional studies. Australian & New Zealand Journal of Statistics, 63(4):639–657, December 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Jiang:2021:VSD

[2380] Rong Jiang, Yexun Peng, and Yufei Deng. Variable selection and debiased estimation for single-index expectile model. Australian & New Zealand Journal of Statistics, 63(4):658–673, December 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ando:2021:OSM

[2381] Shuji Ando. Odds-symmetry model for cumulative probabilities and decomposition of a conditional symmetry model in square contingency tables. Australian & New Zealand Journal of Statistics, 63(4):674–684, December 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Barreto-Souza:2021:BRB

[2382] Wagner Barreto-Souza, Vinícius D. Mayrink, and Alexandre B. Simas. Bessel regression and bbreg package to analyse bounded data. Australian & New Zealand Journal of Statistics, 63(4):685–706, December 2021. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2022:IIa

[2383] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 64(1):i–iv, March 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Huijser:2022:PAI

[2384] David Huijser, Jesse Goodman, and Brendon J. Brewer. Properties of the affine-invariant ensemble sampler’s ‘stretch move’ in high dimensions. Australian & New Zealand Journal of Statistics, 64(1):1–26, March 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Ma:2022:EEP

[2385] Yaolan Ma and Bo Wei. Efficient estimation of partially linear tail index models using B-splines. Australian & New Zealand Journal of Statistics, 64(1):27–44, March 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Liang:2022:FDR

[2386] Beiting Liang, Taoxuan Gao, Defa Bai, and Guochang Wang. Functional dimension reduction based on fuzzy partition and transformation. Australian & New Zealand Journal of Statistics, 64(1):45–66, March 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Rayner:2022:STG

[2387] J. C. W. Rayner, Paul Rippon, Thomas Suesse, and Olivier Thas. Smooth tests of goodness of fit for the distributional assumption of regression models. Australian & New Zealand Journal of Statistics, 64(1): 67–85, March 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Teimouri:2022:MRP

[2388] Mahdi Teimouri and Saralees Nadarajah. MPS: an R package for modelling shifted families of distributions. Australian & New Zealand Journal of Statistics, 64(1):86–108, March 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2022:E

[2389] Anonymous. Erratum. Australian & New Zealand Journal of Statistics, 64(1):109, March 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2022:IIb

[2390] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 64(2):i–iv, June 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Nguyen:2022:FGM

[2391] Hien Nguyen, Sharon Lee, and Florence Forbes. A Festschrift for Geoff McLachlan. Australian & New Zealand Journal of Statistics, 64(2):111– 116, June 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Tomarchio:2022:MSC

[2392] Salvatore D. Tomarchio, Salvatore Ingrassia, and Volodymyr Melnykov. Modelling students’ career indicators via mixtures of parsimonious matrix-normal distributions. Australian & New Zealand Journal of Statistics, 64(2):117–132, June 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hui:2022:SDR

[2393] F. K. C. Hui and L. H. Nghiem. Sufficient dimension reduction for clustered data via finite mixture modelling. Australian & New Zealand Journal of Statistics, 64(2):133–157, June 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Scrucca:2022:MCP

[2394] Luca Scrucca. Modal clustering on PPGMMGA projection subspace. Australian & New Zealand Journal of Statistics, 64(2):158–170, June 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Durand:2022:BNP

[2395] J.-B. Durand, F. Forbes, C. D. Phan, L. Truong, H. D. Nguyen, and F. Dama. Bayesian non-parametric spatial prior for traffic crash risk mapping: a case study of Victoria, Australia. Australian & New Zealand Journal of Statistics, 64(2):171–204, June 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Greve:2022:SPN

[2396] Jan Greve, Bettina Grün, Gertraud Malsiner-Walli, and Sylvia Frühwirth-Schnatter. Spying on the prior of the number of data clusters and the partition distribution in Bayesian cluster analysis. Australian & New Zealand Journal of Statistics, 64(2):205–229, June 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Hennig:2022:AAD

[2397] Christian Hennig and Pietro Coretto. An adequacy approach for deciding the number of clusters for OTRIMLE robust Gaussian mixture-based clustering. Australian & New Zealand Journal of Statistics, 64(2):230– 254, June 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Nguyen:2022:GIF

[2398] Hien Duy Nguyen and Florence Forbes. Global implicit function theorems and the online expectation-maximisation algorithm. Australian & New Zealand Journal of Statistics, 64(2):255–281, June 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Stewart:2022:DBS

[2399] Michael I. Stewart. Detection boundary for a sparse gamma scale mixture model. Australian & New Zealand Journal of Statistics, 64(2):282–296, June 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Arief:2022:VPL

[2400] Vivi N. Arief, Ian H. DeLacy, Thomas Payne, and Kaye E. Basford. Visualising the pattern of long-term genotype performance by leveraging a genomic prediction model. Australian & New Zealand Journal of Statistics, 64(2):297–312, June 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Zhang:2022:BHM

[2401] Huizi Zhang, Ben Swallow, and Mayetri Gupta. Bayesian hierarchical mixture models for detecting non-normal clusters applied to noisy genomic and environmental datasets. Australian & New Zealand Journal of Statistics, 64(2):313–337, June 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2022:IIc

[2402] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 64(3):i–iv, September 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Smith:2022:RSS

[2403] Connor Smith, Boris Guennewig, and Samuel Muller. Robust subtractive stability measures for fast and exhaustive feature importance ranking and selection in generalised linear models. Australian & New Zealand Journal of Statistics, 64(3):339–355, September 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Mukherjee:2022:MKT

[2404] Amitava Mukherjee and Hidetoshi Murakami. Multivariate Kruskal– Wallis tests based on principal component score and latent source of independent component analysis. Australian & New Zealand Journal of Statistics, 64(3):356–380, September 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Piladaeng:2022:PPP

[2405] Janjira Piladaeng, S. Ejaz Ahmed, and Supranee Lisawadi. Penalised, post-pretest, and post-shrinkage strategies in nonlinear growth models. Australian & New Zealand Journal of Statistics, 64(3):381–405, September 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Marmolejo-Ramos:2022:PPD

[2406] Fernando Marmolejo-Ramos, Raydonal Ospina, and Freddy Hernández-Barajas. The place of probability distributions in statistical learning. A commented book review of ‘Distributions for modeling location, scale, and shape using GAMLSS in R’ by Rigby et al. (2021). Australian & New Zealand Journal of Statistics, 64(3):406–412, September 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2022:IId

[2407] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 64(4):i–iv, December 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Gerhard:2022:RGM

[2408] Daniel Gerhard and Elena Moltchanova. A Richards growth model to predict fruit weight. Australian & New Zealand Journal of Statistics, 64 (4):413–421, December 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Chattopadhyay:2022:MCC

[2409] Bhargab Chattopadhyay and Swarnali Banerjee. Minimum costcompression risk in principal component analysis. Australian & New Zealand Journal of Statistics, 64(4):422–441, December 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Huang:2022:PEV

[2410] Xin Huang, Han Lin Shang, and David Pitt. Permutation entropy and its variants for measuring temporal dependence. Australian & New Zealand Journal of Statistics, 64(4):442–477, December 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Qian:2022:NMI

[2411] Lianyong Qian and Fukang Zhu. A new minification integer-valued autoregressive process driven by explanatory variables. Australian & New Zealand Journal of Statistics, 64(4):478–494, December 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Sefidkar:2022:SAE

[2412] Reyhane Sefidkar, Mahmoud Torabi, and Amir Kavousi. Small area estimation under a semi-parametric covariate measured with error. Australian & New Zealand Journal of Statistics, 64(4):495–515, December 2022. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Anonymous:2023:II

[2413] Anonymous. Issue information. Australian & New Zealand Journal of Statistics, 65(1):i–iv, March 2023. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Giaimo:2023:TCA

[2414] Stefano Giaimo. On two conjectures about perturbations of the stochastic growth rate. Australian & New Zealand Journal of Statistics, 65(1):1– 13, March 2023. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Mei:2023:VSH

[2415] Xiaoling Mei, Bin Peng, and Huanjun Zhu. Variable selection in heterogeneous panel data models with cross-sectional dependence. Australian & New Zealand Journal of Statistics, 65(1):14–34, March 2023. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

OGorman:2023:MRW

[2416] T. W. O’Gorman. A method to reduce the width of confidence intervals by using a normal scores transformation. Australian & New Zealand Journal of Statistics, 65(1):35–48, March 2023. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).

Xue:2023:NMA

[2417] Xiangjie Xue and Yong Wang. A nonparametric mixture approach to density and null proportion estimation in large-scale multiple comparison problems. Australian & New Zealand Journal of Statistics, 65(1):49– 75, March 2023. CODEN ???? ISSN 1369-1473 (print), 1467-842X (electronic).